A Bibliography of Publications of the USENIX Association: 1980–1989

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

06 February 2024 Version 2.16

Title word cross-reference

γ [Gra88c].

* [DC85, JN88]. */ [DC85].

-GLA [Gra88c, Gra88a, Gra88b]. -like [Ros87b]. -or- [Ros88, Spe88]. -v [Pik83].

/* [DC85]. /etc/passwd [Hoo83]. /rdb [Man83a, Man83b].

1 [Kol86, Tor83a, Tor83b, Wil83b, Wil83c]. 1-2-3 [TK88a, TK88b]. 1/2 [Kri84a, Kri84b]. 1/2-inch [Kri84a, Kri84b]. 11 [Kar83]. 11/44 [Tuo82]. 11/780 [GM82a, GM82b, Tuo82]. 11s [See83].

12/84 [Wei85a, Wei85b]. 16-bit [Bis83]. 1988 [Til88].

2 [Bak89, Bam85a, Bam85b, Car87c, DO85, HK86a, Pow84]. 2-D [Bam85a, Bam85b]. 2-inch [Kri84a, Kri84b]. 20 [Hat82].

3 [Ale87a, Ale87b, KK89b, KK89c, LOR88a, LOR88b, TF89a, TF89b, Zuc83b, Zuc83c]. 3-D [KK89b, KK89c, LOR88a, LOR88b, TF89a, TF89b]. 3-Dimensional [Ale87a, Ale87b]. 3/5 [Sch83c]. 32/27 [Ble83]. 370 [Eyk86]. 3D [BPM87, SGS89]. 3DWorks [SGS89].

4.1BSD [CQ83a, CB83]. 4.1C [CQ83b]. 4.2bsd [All83c, BH86, CKM85, Eva83, JTUB85, Joy82a, Joy82b, JLMM82, KM83, Kri84a, Kri84b, Lef82, LKM84, Len86b, LS83b, MK85a, McK82b, Mos82, O’D83a, OTW85, PBL86, WG84]. 4.3BSD [BL89, Gou86, HS89, KM86, MK85b, MK88, MKB89, Ste89b, Tre88]. 4/95 [Poz83]. 44 [Tuo82]. 4BSD [JLSG84, Lan84]. 4GL [Sto88].

64 [JS89a]. 644-2332 [Lan86]. 68000/Unix [Gar86].

80 [Cox83]. 8x [Ros87d].

9 [Pre88]. 9660/HSG [KGTM89].

Abstract [LC87b]. Abstractions [SA86a, SA86b]. Abstracts [USE88c]. Academic [Mid87]. Accelerate [Sch89]. Access [CJ88, Chr89a, Chr89b, FA88, HP89a, HP89b, Kra88a, LG88, MC85a, MC85b, Man87b, Mog89, OLJ^{+}88, Ort88, Ric85, Win88, Lil88, MC85a]. Accessing [GBM87a, GBM87b]. Account [Abb87, FT89, PMD88, TH87]. Accounting [Eat88, GP88, McK88b, WH83]. Accounts [Bis87c]. Accuracy [SF86]. Acorn [RSSW89]. ACSNET [DLKE84, Lau85]. ACSS [Mid87]. Active [MN85a, MN85b]. Activities [McC88]. Activity [BG88a, Hae89, Isl83a]. ad [TK88a, TK88b]. Ada [CC86, CKK87, EST86, Fis86a, Fis86b, Gar86, Gro84, Lam83, Nyb86, SR85a, SR85b, Sch86a, Shu89]. Adaptable [AQ84]. Adaptation [Bey88]. adapted [Yos85]. Adapting [Rob84b]. Adaptive [Cic88]. Adding [HJAW88, NY88, DJM86]. Address [Che87b]. Addresses [HP85, HB86]. Addressing [All83c, Lau85, Sal89a]. Administering [Zad89]. Administration [Abb87, Car88d, Con88, Cyg88, Dat88, FS89, HH88, Hun88a, Joi87, Jon88, KGL89, Lil88, Ond89, Smi87b, Ste84, USE88f, USE89e, Van88, Woh88]. Administrative [Hei87, PL89, RW86a].

Administrator [EVS88a, EVS88b]. Administrators [USE87d]. Advanced [Tur87]. Adventures [KM84b, KM84a, Mor88d]. Advice [BH86]. Advisor [RU88a, RU88b]. Affecting [SM89]. AFQL [CS86]. Afraid [LK82]. after [JN88, O’D83a]. AFUU [DB88]. Again [PS82]. Age [McD87, Red85, RW86b, RW86a, Tay86]. Aid [Kle87, LFN^{+}89a, LFN^{+}89b]. Aide [LW89b, LW89a]. Aided [Tho85a, Tho85b, Wat88c]. Aids [RW86a, SM84]. AIX [CLH^{+}89, TGB^{+}89]. Algorithm [BK87, SW84a, SW84b]. Algorithms [MT89]. Aliased [Coh87]. aliasing [PG88c]. All-Natural [Har85a, Har85b]. Allegro [LQC87]. Allocation [LB89a, LB89b, ZH88, Cor82]. Allocator [Cap88, MK88, Smi89]. Alpha 1 [Tho85a, Tho85b]. Alphanumeric [Bam85a, Bam85b]. Alternative [DLKE84, DO85]. Alternatives [FW89]. Always [Tay88, Til89]. Amdahl [Wal82a, Wal82b, Wal82c]. Amsterdam [KTv83]. Analysis [BL89, Jac83, Leb87, Mil87, Per82a, SD87]. Analyzers [Gra88c, Gra88a, Gra88b]. Analyzing [Kor89a, Kor89b]. Anarchy [TS87]. Andrew [BERS88a, Mor88b, USE88d, BERS88b, Cyg88, Hec88, How88, Kaz88, Man87a, Mor88c, PHS^{+}88a, PHS^{+}88b]. Angus [War82]. Animated [JHRR86]. Animating [Duf85]. Animation [BK87, Hal85b, LOR88a, LOR88b, LM83a, LM83b, Pet87, SGS89]. Annotation [CM86a, CM86b]. ANSI [Gro88, Bol88]. ANSI/ISO [Bol88]. Answers [JLMM82, Red85]. Anti [Coh87]. Anti-Aliased [Coh87]. AOS [WM82]. AOS/VS [WM82]. APL [Yao83]. Apollo [SSWW83]. Application [Bis88a, Hor82a, KI82, Mar84, Mee85, SM89, Sto85, UTC84]. Applications [BDWW89, Bil86, BWH87, Che89, DLM^{+}87, Gri85, HS82, Jam88, Kep85, Nor84, Rei89, SW82, Ste88a, Sun88, TBS87, Sto88]. Applying [VL88a, VL88b]. Approach [HCE88, KAH83a, Mor88e, Oti88, Son88a, Son88b, Van87, Pre85]. Approximate [Nac88a, Nac88b]. April [USE87d, USE89a, USE89j]. Arachne [DR89]. ARCADE [CDT89a, CDT89b]. Architected [Sal82]. Architectural [DNQ^{+}83, RFH^{+}86a, RFH^{+}86b, Wat88a]. Architecture [CE89a, CE89b, Dew87, DLM^{+}87, Fel84, GMS87, Hal85b, HQZ^{+}87, Kle86, Ral88a, Ral88b, Req85, Tra85, TGB^{+}89, Eyk86]. Architectures [Kle89]. Area [DGM82, Wat88b]. Aren’t [Nac86]. ARIEL [HK83]. Arithmetic [Mor88d]. Arizona [USE87f]. Army [Mas83a, Ton87]. Array [Ber85, Col83, HCN85, JTUB85, DJM86]. array-processor [DJM86]. arrays [Koe88]. Arsenal [Per83]. ART [GGSW88a, GGSW88b]. asm [Loc87]. Asmodeus [EVS88b, EVS88a]. Aspects [FKN85a, FKN85b, Kel89, NHR84]. Assessment [SGH^{+}89, Zho87]. Assist [BK85]. Assistance [MN85a, MN85b]. Assisted [Ges86, Ivi84a, Ivi84b]. Association [UE88, Sof83, Sof84, Usr82, USE85c, USE86e, USE87g]. Associative [Koe88]. AT&T [DeJ86a, DeJ86b, EST86, Isl83b, Jun85a, Jun85b]. Atari [Col88, GD89]. Athena [Abb87, Dav89b, Gee88a, Gee88b, Get84, RGL88, Tre88, Vas87a, Vas87b]. Atlanta [USE86c, USE86a]. ATLAS [Hor82a]. Attaching [Col83, Eng88]. Attacks [Duf89b]. Audio [Haw89a, Haw89b]. Auditing [Bis88b]. August [USE88g, USE88b]. Austin [USE89e]. Australia [Hau83]. Australian [DLKE84]. Authentication [NS88, SNS88]. Authoring [GBM87a, GBM87b]. Automated [BC89, FT89, TH87, Vas87a, Vas87b]. Automatic [AN88, Gra87, Gro87, Koe84, Nac86, O’B85, Sig87]. Automating [Egg89, PL89]. Automation [Poz83]. Automounter [CL89]. Autonomous [MR88d, MR88e]. Autumn [USE87b]. Availability [Ado89]. Avalon [DHKW87]. Avalon/C [DHKW87]. Avionics [Mar84]. awk [Tut82].

B [Kno87, MP84, Sta87a, Sta87b, Tho86]. B-Level [Kno87]. B-spline [Sta87a, Sta87b]. B-Splines [Tho86]. B1 [ST89a, ST89b, SM89]. Back [AGHR89a, AGHR89b, Uit87a, Uit87b]. Backend [Duf82]. Backup [AN88, Har88b, Hec88, Hom87, Hum88a, Par88b, Poe87, YKK89, Zwi88a, Zwi88b]. Backup/Restore [YKK89]. Backups [MR89]. Bad [Dye82a, Dye82b, SW84a, SW84b]. Bad-Block [SW84a, SW84b]. Bad-Sector [Dye82a, Dye82b]. Balancing [Ber86, Cab86, JS87]. Baltimore [USE89f, USE89b]. Band [Rag89a, Rag89b]. Base [Ben82, GBM87a, GBM87b, Hae83, Hoo83, Isa83a, Lev83, Man83a, Man83b, Ohk84, Wai82]. Based [AM85a, AM85b, BNB87, CE89a, CE89b, Che82, CM89a, CM89b, CP84a, CP84b, DHKW87, Gan88a, Gan88b, GGSW88a, GGSW88b, HQZ^{+}87, HP89a, HP89b, JHRR86, JH86, KTS^{+}86a, KTS^{+}86b, KT88a, KT88b, Kle85, Kle87, Lew86a, Lew86b, LZ82, Man87b, Neu86a, Neu86b, Roc89a, Son88a, Son88b, STT86a, STT86b, Tra85, Wam83a, Wam83b, WRM^{+}89, BL88a, BL88b, Epp89, HK86a, KLP88a, KLP88b, LM88a, LB89a, LB89b, LOR88a, LOR88b, NLR84a, NLR84b, PG87, Sam87, McD84a, McD84b, Per87a, PY84a, PY84b, Wat88b, WH83, HK83, HK86b, KK89a, LM88b, Mog89, SR85a, SR85b, Sti83]. Basic [Ric85, Bak89]. Basis [Fis86c, Fis86d, FH88]. Batching [Har88a]. Battle [TRG^{+}87]. BBN [Dye82a, Dye82b, WG84]. Bcc [Ken83]. Be [Col87b, Pik88, Ros88, TK88a, TK88b]. bed [Ada83a, Ada83b]. Been [TA82]. Behaved [Mur88d, Mur88e]. Behind [RP84]. Bell [Car88a, Car88b, LK82, Pre88]. Bellcore [Con88]. Benchmark [Ban83, Mil86, SCW89a, SCW89b, Dro84]. Benchmarking [DB88, Dro82, Tuo82]. Benchmarks [BWH87, Dro84, Sax85a, TH82]. Benchwarmers [Dro82]. Benefits [HCE87]. Berkeley [Akh87, CKM85, Car88c, GZ84a, GZ84b, O’D83a, TPRZ84, Tre88, ZDS85]. Better [KV85, Pla89, TK88a, TK88b]. Between [Roc89b, Roc89c, SL88]. Beyond [BW88a, SH85, Ste88b]. BIBFIND [Moy83a, Moy83b]. Bibliographic [Moy83a, Moy83b]. Big [Miy86, PG88a, PG88b]. Binary [Ada83a, Ada83b, RP84]. BIND [BD86]. Bit [Ker83, Ban82, Bis83]. bitblt [Loc87]. Bitmap [BEHW86a, BEHW86b, BWH87, FL82, Pik82]. bitmapped [Col88]. Blits [Pik84a, Pik84b]. Block [Rob84a, SW84a, SW84b]. BNF [Mil87]. Board [Lou82, SM84]. Book [Don88, Don89a, Gro88, Lea89, Sal89a, Sal89b]. Bösendorfer [Haw86b]. Boston [Usr82, USE82a, USE82b]. Bourne [KR85]. Boxes [Gle89]. Break [Spa89]. Break-Ins [Spa89]. breaks [Rob87]. Bricks [SA88]. Bridges [UTC84]. Bringing [RW86b, RW86a, TS87]. Brings [AGHR89a, AGHR89b]. BRL [Dyk87, MSM88c, MSM88a, MSM88b, MSM88d]. BRL/USNA [MSM88c, MSM88a, MSM88b, MSM88d]. Broadcast [McK88a]. Broadcasting [Wei84b]. Brother [PG88a, PG88b]. Browser [RRS87, WK83a, WK83b]. Brushes [Str87a]. BSD [AW89, BE89, CQ83b, Fer85, Pla89]. Buddy [LB89a, LB89b]. Budget [Ton87]. Buffer [CKM85, Pea80, TBJW88]. Bug [Hop89, Tut83]. Bugs [Fil85]. Build [McI87]. Building [ACK89, BK85, LSC^{+}88, MLRC88, Mur88d, Mur88e, SA88, USE89k, UTC84, WS83a, WS83b, WRM^{+}89]. Bulletin [Lou82]. BUMP [MSM88c, MSM88a, MSM88b, MSM88d]. Bus [Cer83, Cer82]. Business [HS82, Wil82a, Wil82b]. Business-Oriented [Wil82a, Wil82b]. Butterfly [SC88]. Button [Pit89].

C

[BG88b, Lea88, Lea89, Ros87b, Tie88, AB85, AG88a, AG88b, BBT83, Bak89, BD87a, Bol88, Boy84, Bre88a, Bre88b, Bro87, Bru88, Cap88, Car87c, Che89, Con87, Cox82, Cox83, DHKW87, Dew87, DO85, DG87, Don89b, Dye82a, Dye82b, Ecc88, FKT83, Feu84, Feu85, Fis86a, Fis86b, FJ82, Fri87, Fuh87a, Fuh87b, GS87, GR85, GMW86, Gor87a, Gor87b, Gro87, Gro88, Hae89, Hop87c, Joh88a, Joh88b, KLP88a, KLP88b, KM85, Ken83, Kir87, Koe88, Kol86, Kre83, Kri86, LR84, LM88c, LS88, Lon82a, Lon82b, MLRC88, Man87b, Mey82a, Mur88d, Mur88e, O’R88, Oti88, Pow83, Raf87, Raf88a, Raf88b, RRS87, Ree82a, Ree82b, Ric87, dR85, RS87, Ros87c, RK88, Sch82, Sch83b, Sch88b, Sho87, Ste83a, Ste83b, Ste85, Sto88, Str85a, Str85b, Str87b, Str87c, SS87, Str88a, Str88b]. C [Str88c, Str88d, Str88e, Str88f, Str89a, Str89b, Str89c, Til83, Tri87, USE87a, USE88a, USE88k, Wal87, WS84, WM82, Woo83a, ZH88]. C* [Ros87b, RS87]. C-1 [Kol86]. C2 [HJAW88]. CA [USE83a, Ass83a, Ass83b, USE85d, USE86d, USE88f, USE89c, USE89i]. Cache [BLMY87, Che87b]. Caching [Kaz88]. CAD [Dyk87, HQZ^{+}87]. CAIS [Fis86c, Fis86d, GBH86]. Cake [Som88]. California [USE88h, USE89g]. Call [Kar83, Leb87, McK83c, McK83d, Rod86].

Callout [BL88a, BL88b]. Calls [CLH^{+}89, MMTW88]. Cambridge [USE87c]. Camphor [Kaz85b, Kaz85a]. Can [Gra88c, Gra88a, Gra88b, Les88a, O’D83c, Biv87, Ros88, Feu84]. Canada [Sof83]. Can’t [DC85]. Capabilities [Ale87a, Ale87b]. Capability [Erl88, Kle85, Tra85]. Capacity [RWNA87, Wat88c, Wat88b]. Capital [Wil83a]. CAPS [LFN^{+}89a, LFN^{+}89b]. Care [Ham87, HB86]. CAS [Fun87]. Case [BG88b, CKM85, Car87a, Car87b, Fen87, Ham87, KL87a, KL87b, Tri87, HM89]. Cases [DR83b]. CASPER [BC88]. cat [Pik83]. Catalog [Les88c]. CCSLAND [Bro89a]. CD [KGTM89]. CD-ROM [KGTM89]. CDA [SM84]. CDC [LR84]. Center [BB87, Woh88]. Central [Ond89]. Centralized [Har87a, Har87b]. Centuries [Les87]. Century [Til87a, Til87b]. Certifiable [ST89a, ST89b]. Challenge [Hos83]. Change [BED^{+}85]. Changes [Lon82b, Tre88]. Changing [BED^{+}85, Mas88]. CHAOS [GGSW88a, GGSW88b]. CHAOS-ART [GGSW88a, GGSW88b]. Character [Orr83, Roc89a]. Character-Based [Roc89a]. Character-Oriented [Orr83]. Characterization [Miy88]. Charge [McK88b, Uit87a, Uit87b]. Check [Sax85b, Sax85c]. Checkers [Ros82]. Checking [Ken83, Cor82]. Chips [LS85]. Choices [KM83]. CHORUS [RAA^{+}88]. Circuit [SM84]. Circuits [Les87]. Circular [Mur83]. City [Sof84, USE84a, USE84b]. CLAM [CCM87]. Class [Cog87, Fuh87a, Fuh87b, Gan88a, Gan88b, Gor87a, Gor87b, LS88, MLRC88, RRS87]. Class-Based [Gan88a, Gan88b]. Classes [Den89, Kir87, SS87]. Cleaning [TS87]. Client [LA89a, LA89b, Ros88]. Client-Server [LA89a, LA89b]. Clients [LR89]. Clinic [San89]. Clinical [KTS^{+}86a, KTS^{+}86b]. Clipboard [Nic85].

Clocks [Miy88]. Clone [Tan87a]. Cloned [Yos85]. Cloning [Hun88b, Hun88c, Sto87]. Closures [Bre88b]. Clouds [WRM^{+}89]. CMS [TH86]. Co [Bar88a, Bar88b, SS87, USE86b, USE88a, USE88k]. Co-Resident [Bar88a, Bar88b]. Co-routine [SS87]. Coast [USJ83]. COBOL [CM83]. Code [AQ84, BBT83, Bou89, Kal82, KM85, Lor88c, Ree82a, Ree82b, Spe88, Tan87a, Les88a]. Coding [LFN^{+}89a, LFN^{+}89b, Mey82a]. Coexistence [JLSG84]. Cogito [BH86]. cognitive [van86]. Collection [Cap88]. Collector [WHM89]. Color [Ges86, Hal85a, McD84a, McD84b]. Colorado [USE86e]. Columbia [USE87g]. Combinatorial [Sta87a, Sta87b]. Combining [Gol88, SV83]. Coming [Hos83, McK82a]. Command [Bee84a, Bee84b, Der83, Per85]. Commands [HS89, HA84]. Commercial [Dun88, Kin83, SW82, UTC84]. Commercialization [TY82]. Committee [Lyc83b, Lyc83a]. Common [Gol88, Ecc88]. Communicating [Don88]. Communication [Gil86, Phi84, PR85, PA89, Rag89a, Rag89b, Sch89, SL88, Ste88a, Tay86]. Communications [Joy82a, Kee88, Lef82, Req85, Rob84b]. Community [Cle83a, Cle83b]. Compare [Mil86]. Comparing [Mur88c]. Comparison [Bih88, CQ83a, GBH86, Kle89]. Comparisons [Nac88a, Nac88b]. Compatibility [Sch83c, See83]. Compatible [FH88, HMP83a, HMP83b, Len86b, Zuc83b, Zuc83c]. Competition [Kat83]. Compilation [EST86]. Compiler [CM83, Cox82, Dew87, Gri85, KR85, KTv83, Kle89, Kri86, LR84, Pow83, Rei87a, Rei87b, Ros87c, RSW83, WM82, Woo83a]. Compilers [Miy88, TvK83, ZS83]. Compound [Ros82]. Computation [DW88, Pet87, Ros87b]. Computational [GPF^{+}86a, GPF^{+}86b]. Computations [LN88]. Computer [BED^{+}85, BO83a, BO83b, Che82, Cog87, CH83a, CH83b, Ges86, Haw89a, Haw89b, Ivi84a, Ivi84b, LG88, Les88d, LM83a, LM83b, Min82a, Min82b, Miy88, O’B82, Pet87, SB88a, SB88b, Spa89, Ste84, Sul87, Tay86, Tho85a, Tho85b, USE85d, USE86d, USE87c, USE89i, Wat88c, Wil88, Don88, KL82a, Poz83]. Computer-Aided [Tho85a, Tho85b]. Computers [FT83, HZ89a, HZ89b, Lin84, SSWW83, SF86]. Computing [UE88, Bry88, DLM^{+}87, HZ89a, HZ89b, HSY88, Mid87, RRSZ89, TWM86a, TWM86b, Ass88a, Ass88b, Ass88c, Ass88d, Ass89a, Ass89b, Ass89c, Ass89d]. Concentrix [Tes86b, Tes86a]. Concept [FT83]. Concerns [LAKS88]. Concise [FJ82]. Concurrency [Gro87, Kat82c]. Concurrent [HMP83b, OLJ^{+}88, Ort88, Pik89, GR85, HMP83a]. Conference [Sof83, Sof84, Til88, Usr82, USE82a, USE82b, USE83a, Ass83a, USE83b, Ass83b, USE84a, USE84b, USE84c, USE85c, USE85a, USE85b, USE86e, USE86a, USE86b, USE87b, USE87f, USE87g, USE88a, USE88e, USE88h, Ass88f, USE88j, USE88k, USE89f, USE89g, USE89c, USE89b, USE85e, USE86c, USE88d]. Conferencing [RRSZ89, STT86a, STT86b]. Configuration [Bla89, DF89, Ful89b, Gle89, RWNA87, Sch86b, Sch86c, Van87, ML88]. Configured [Van88]. Configuring [HZ89a, HZ89b]. Connect [RCB83]. Connecting [Mil84a]. Connection [Bla83a, Del87, KNI88]. Considerations [EGL86, Mye86, Pat83, Rei89, Woh88]. Considered [Pik83]. Consistent [Har87d, Hil89]. Console [Lin88a]. Constrained [Per87a]. Constructed [BEHW86a, BEHW86b]. Constructing [KL87a, KL87b, UTC84]. Construction [CM83]. Consulting [CAG89a, CAG89b, Man87a]. Contention [Wil82a, Wil82b]. Contiguous [Zuc83a].

Control

[Bih88, Bou89, Bru88, CJ88, FA88, HP89c, Hum89, Kal82, Kat82c, Kra88a, Len86b, Lib85a, Lib85b, MR88d, MR88e, Sch86a, Sho87, TRG^{+}87, Wil82a, Wil82b, Yos85]. Controlled [Nac88a, Nac88b]. Controller [Cic88, GZ84a, GZ84b]. Controlling [Don88, Mil89, Tri89]. Controls [Elz84, LG88, Mog89]. Convenience [JS87]. Conventional [DV89]. Conversation [CP84a, CP84b, Mye86]. Conversation-Based [CP84a, CP84b]. Conversion [Ges86]. Converting [WG84]. Cookbook [Hag83]. Cooperating [BEHW86a, BEHW86b]. Copy [NO88a, NO88b, SM88a, SM88b]. Copy-on-Write [NO88a, NO88b, SM88a, SM88b]. Core [DC85]. Coroutines [dR85]. Corrections [DGM82]. Correctness [Jus89]. Cost [JHRR86, Tho85c, Col88]. Counterpoint [Pea88]. Coupled [BFS89, Inm85]. cpp [Loc87]. CPU [FKN85a, FKN85b, Ree81]. CRACK [RU88a, RU88b]. Crash [McK82a]. Cray [AO86, Eng88, Fou88, HK86a, HK86b, Par88a]. Creating [Smi87a]. Creation [Hef82]. Criteria [Swa83]. Criticalness [BSR88]. Cron [Har87c]. Cross [HCE87]. Cross-Module [HCE87]. crunchers [DJM86]. Cscope [Ste85]. CSL [Per82a]. CSNET [BO83a, BO83b, OL84a, OL84b, O’B85, Rei83]. Ctrace [Ste83b, Ste83a]. CTSS [AO86, Bro88]. CTSS/POSIX [Bro88]. Cult [Col87a]. Current [Bla83a, KL82a, KL82b]. Curses [Hor82b, Nyb86]. Custom [Gri85]. CYBER [LR84]. Cycle [Lit87]. Cypress [CN88].

D [Bam85a, Bam85b, KK89b, KK89c, LOR88a, LOR88b, TF89a, TF89b, Ber86]. Daemon [BC88, EVS88a, EVS88b, Fed88a, Fed88b, Len87]. Daemons [Jon88]. Dallas [Til88, USE85c, USE85b, USE88j]. Darkly [DR86]. DARPA [BD86]. Darth [Ree82b, Ree82a]. DASH [AF87]. Data [All87, BKT89, Ben82, Bis88a, Bre88a, Con87, Gri89, Hae85a, Hae85b, Hae86a, Hae86b, Hae83, Hoo83, Isa83a, Jac83, JN88, KBT89, Kal82, Lee87b, Lev83, Man83a, Man83b, Nic85, Par87, Per82a, PG87, Pyn82a, Pyn82b, Rob84b, Ros87b, RS87, Ton87, Wai82, Dro84, GS87, Han82, Hen83, McL83, WM82]. Data-Flow [Hae86a, Hae86b, Hae85a, Hae85b]. Data-Object [BKT89]. Data-Parallel [Ros87b]. Database [Bla89, BC89, Che89, CS86, DW88, Duf82, Haw85, KL82a, Mey82b, Son88a, Son88b, Sti83, Ton87, War83, WK83a, WK83b, Sto88]. Dataflow [BS85a, BS85b]. Datagram [Mog89]. day [Ing87, RR85]. DB [War83]. DBMS [Kin83, TBS87]. Dbxtool [AM85b, AM85a]. DC [USE84c, USE87g, Ass88e]. Deactivation [FT89]. Deadline [BSR88]. Deadlock [Pea80]. Debugger [AM85a, AM85b, FKT83, PG87, Ste83a, Ste83b, Zim85]. Debugger-based [PG87]. Debugging [BV88, BC84, Dav89a, HR85a, HR85b, Kat89, O’R88, VM84]. DEC [Kri84a, Kri84b]. December [USE85d]. Decentralized [Shu89]. Decisions [Mor88b, Mor88c]. DECNET [JLSG84, Mur88c]. Decreasing [Len86a]. Defined [Bol88]. Definition [Pyn82a, Pyn82b, Sta87a, Sta87b]. Definitions [Cra83, Mil87]. Delegating [DM88]. Deletion [FT89]. Delivering [Den83]. Demand [Jun85a, Jun85b, Mil84b]. Denver [USE86e, USE86b, USE88a, USE88k]. Department [GM82a, GM82b, Ond89, Les88d]. Dependent [JHRR86, LGZ88]. Description [Ada83a, Ada83b, Bas81, Cla87a, Cla87b, FJ82, Hef82, Tes82]. Design [AF87, Baa88, BLK87, BWS87a, BWS87b, Big85, BL89, CRJ87, Che87c, Cic88, Col84c, ELS88a, ESS89, FL82, LQC87, LC87a, MdM88, MK88, Mer82, Mye86, Rei89, SGK^{+}85, Sch88a, SREC88, Sen87, SM84, TM82, Tho85a, Tho85b, VL88a, VL88b, War83]. Designers [War82]. Detection [BK88a, BK88b]. Develop [Wal87]. Developing [BDWW89, CKK87, DLM^{+}87, FN83, Neu86a, Neu86b]. Development [ABB^{+}86a, ABB^{+}86b, Bih88, Che87a, CB83, ESS89, Gro82, Hop87c, KR85, Lor88c, Mas83b, Mas83c, NM83a, NM83b, Ohk84, Per87a, Per83, RC83, SW82, Tut83, War84, KNN88a, KNN88b, STV87]. Device [Alb84, Gou85, Hid83, KM82, MVB84, VM84, Wat83]. Devices [WO88]. DG [Kel89]. DG/UX [Kel89]. Diagram [MW84]. Dial [LP89]. Dial-up [LP89]. Dialectic [Rit87]. Dialogue [SE88]. Diamond [CFLT87]. DIBOL [ASS85]. DIBOLIX [ASS85]. Dictionary [Bra89, Hae83]. Did [LS83b]. Diego [USE83a, Ass83a, Ass83b, USE89g, USE89c]. Different [Koe87a]. Differentiation [Mar83]. Digital [Kin86, RC83]. Dimensional [Ale87a, Ale87b]. Direct [Ric85]. Directional [Hop87a, Hop87b]. Directions [Str87c, Tag83, Wat88a]. Directly [Eng88, Mey88]. Directory [Sal89a]. Discipline [TS87]. Discretionary [CJ88, LG88]. Discuss [RRSZ89]. Disk [GRS88a, GRS88b, Har88b, JSW87, KM83, LEG88, Ste89b, Van87, YKK89, Zwi89]. Diskless [ACF^{+}86, CGFCKT88]. Disks [Eng88, Van87]. Dispatch [Len86a]. Display [Bam85a, Bam85b, JHRR86, Lew86a, Lew86b, PG87, SSNU87]. Displays [McG86a, McG86b]. Distributed [AF87, And88a, BKT89, BV88, Bar88a, Bar88b, BD86, BP84, CM86a, CM86b, CDT89a, CDT89b, DRK^{+}89, DLM^{+}87, Duc89, EBFH85a, EBFH85b, Ful89a, Gos86a, Gos86b, GZ84a, GZ84b, HSY88, HLW84, Hom87, JH86, JC89a, JC89b, KBT89, Ker88, LW89a, LW89b, Lee89, LSC^{+}88, McG85, MF89, Muu87, Pet87, RRSZ89, RU88a, RU88b, RK89, Rei89, RAA^{+}88, SJL^{+}87, Sen87, Sha89, Shu89, Smi89, Son88a, Son88b, Spe87, Sun88, STT86a, STT86b, Tal89a, Tal89b, TM82, Tay88, TR84, TWM86a, TWM86b, USE89d, USE89k, Wam83a, Wam83b, Wat88a, WRM^{+}89, STV87]. Distributing [AGHR89a, AGHR89b, BFS89]. Distribution [Bro85, DGM82, Kim87, Koe84, Lad88a, Lad88b, Mar83, Nac86, Rod87, Sig87, TP86]. District [USE87g]. DITROFF [BD87b]. Diverse [JH86]. Do [Bou89, Dro84, Har88b, Hoo83, JN88, O’D83a]. Document [CM86a, CM86b, vH87]. Documentation [YT83, MdM88]. Documenting [SH85]. Does [Ree82a, Ree82b]. Domain [Hor84a, Hor84b, Lau85, Par86, TPRZ84]. Domains [PL89]. done [RR85]. Don’t [O’D87c]. Doomed [Gre82a, Gre82b]. Downtime [Har87c]. Draft [Bol88]. DRAGONMAIL [CP84a, CP84b]. Drawing [Coh87]. DREGS [BCL^{+}87]. Drive [Ste89b]. Driven [Bla83b, Bla83c, Hae83, Les83, Ney83a, Ney83b]. driver [Lau81a]. Drivers [Alb84, Gou85, Gur88, MVB84, VM84, Wat83]. Dual [GM82a, GM82b]. Dublin [USE87b]. Dump [DC85, Pla89, PK88]. Dumping [Haw88a, Haw88b, Vas87a, Vas87b]. Dumps [Jaf87]. DUNE [PA89]. Duplex [Ste86]. duplicating [Hun88b, Hun88c]. Duplication [Hal87]. Durra [BDWW89]. DV [CR89]. Dynamic [HF89, KGL89, Par88a, RKPP88, SW84a, SW84b, GS87]. Dynamically [GM89, ROS87a]. Dynamics [GPF^{+}86a, GPF^{+}86b, PF84a, PF84b, Wil87a, Wil88].

Early [FT83]. Ease [Sch86c, Sch86b]. Eastman [Les88d]. Easy [Hop87a, Hop87b]. ED [Mok88]. Eddie [Lan86]. Edit [SV83]. Editing [Ale87a, Ale87b, BPM87, PG87, SV83]. Edition [HH86, PR85]. Editor [Ada83a, Ada83b, CFLT87, MW84, Rug82a, Rug82b, WK83a, WK83b, GS87, MD87, Sal89b]. EE [GM82a, GM82b]. Eedie [Lan86]. Effectively [Car88c, Nov83]. Effects [KM83, SM88a, SM88b]. Efficiency [Mur88c]. Efficient [BS85a, BS85b, FKV89, KLB89, PA89]. Effort [Lyc84, Mei84]. Eighth [HH86, PR85]. Electronic [Alt87, Bro85, HP85, Kim87, O’D83b, RRSZ89, Sal89a, SREC88, STT86a, STT86b, Tay88, van86]. Elements [Gle89]. Eliminate [Dro82]. Elmer [Pre82a, Pre82b, Tuo82]. Elmer’s [Car82a, Car82b]. Embedded [Isa83a, Shu89]. Embedding [NMP82]. Empirical [MFS89]. Emulate [LS83b]. Emulation [PS82, SSWW83]. Emulator [Cap82a, Cap82b]. Encryption [Bis88a]. End [Den83, GPF^{+}86a, GPF^{+}86b]. End-User [Den83]. Enforcing [Mok88]. Engine [Che87c]. Engineering [Dav89b, Ela83, Les88d, MKB89, Mor88e, AG88a, AG88b]. Engines [Inm85]. England [USE88e]. English [Bra89]. Enhanced [BW88b]. Enhancement [LN88]. Enhancements [Cal83, DGM82, Goo84, HS89, Hid83, Kol86, MK85b, RW86a, Tut83]. Enhancing [AW89, BBT83, Hir83]. Enough [Nac86, Hoo83]. Entities [NS88]. Entry [Cle83a, Cle83b]. Environment [Ado89, Ale87a, Ale87b, AN88, BDWW89, Boy84, Bry88, CMM88, Col83, Col84b, CD85, DF89, DLM^{+}87, FS89, FHW88, Gen86, Gin88, GM89, Gou85, Hae85a, Hae85b, Hae86a, Hae86b, Har87d, HSY88, HCN85, Hom87, HM89, vH87, JTUB85, JH86, KLP88a, KLP88b, Kaz85a, Kaz85b, Kre83, KI82, Lam83, LW89a, LW89b, Lib85a, Lib85b, LQC87, LSC^{+}88, MLS88, McI87, Mer82, MR89, Mur88a, Mur88b, Nor88, Pat83, PL89, RRSZ89, RC83, Rod87, Sen87, Smi87a, TG86, Tay88, Tho85a, Tho85b, Zem83, ZP89, STV87]. Environments [Bre83, CM86a, CM86b, Har88a, DJM86, HOG88, KNN88a, KNN88b]. Equities [LSC^{+}88]. Eradication [Hop89]. ergonomic [van86]. Error [AF86, Gra87, RW86a]. Establishing [Fer85]. etc [Hoo83, Kod82a, Kod82b]. Ethernet [Fos83a, Fos83b, Sku88]. EtherTIP [Fos83a, Fos83b]. Ethics [Spa89]. Euclid [HMP83a, HMP83b]. EUNET [RCB83]. EUNICE [Wil83d]. Europe [McK83a, McK83b]. European [Ass89e, WB85a, WB85b]. EUUG [USE87b, USE88e, Ass89f]. Evaluation [McD87, Pod82]. Every [RP84]. Everyone [Wil87a]. Everything [Bal83, LK82, Tan87b]. Evolution [Fed83, Gin88, Str87b, Str89a, Weh83]. Evolving [MS89]. Examination [Ste85]. Example [Mac83a, Mac83b]. Excelan [NB84, Smi87b]. Exception [AB85, Eyk88, Mil88]. Exceptions [HCE88]. Executable [KT88a, KT88b]. Execution [McK83c, McK83d, Shu89]. Exercise [Hei87]. Expandable [NLR84a, NLR84b]. Expansion [Bol88, Hei87]. Experience [Bil86, CM89a, CM89b, DRK^{+}89, Dou89, Duf89a, HJAW88, Hop87c, JLSG84, KBT89, MBBP89, Ros87c, Ste84, Tuo83, WRM^{+}89, MdM88]. Experiences [Bec84, BD86, BCL^{+}87, Bro85, CB83, CM83, Fou88, FT83, GB89, Nic89, PA89, TBS87, USE89k]. Experiment [Lan86, Sun89]. Experimental [Fos88, HK83, HC88, NHR84, SD87, Zho87]. Experiments [IvW87, Les88c]. Expert [BH86, BK88a, BK88b, PG88a, PG88b]. Exploitation [Gro87]. Exptools [Ste84]. Extended [CFA85, Eyk86, RS87]. Extending [BP88a, BP88b, FA88, Man87b, Raf87, Sho87]. Extensible [Gos86a, Gos86b, Kaz85a, Kaz85b, RLML86a, RLML86b, ROS87a, SE88, SREC88, Str85a, Str85b]. Extension [TF89a, TF89b, Zuc83b, Zuc83c].

Extensions [Cla88, CCF89, DHKW87, GGSW88a, GGSW88b, LH84a, LH84b, Mil88, WO88, GS87].

F77 [MC82]. Face [PP85, Cla87a, Cla87b]. Faces [Kin86]. Facilities [GBH86, Gil86, RW86b]. Facility [BM87, BK84, Har87c, MF83, Rit88, Str85a, Str85b, Nic85]. Factors [MT89, SM89]. Fall [Ass88a, Ass89a]. Family [GB89, LZ82]. Farms [Har88b]. Fast [Bis88a, Gra87, Loc87, Pea88, Sun88, Woo83b]. Faster [BE89, PK88]. Fault [LA89a, LA89b]. Fault-tolerant [LA89a, LA89b]. Fe [USE87a]. Features [DJ88, HJAW88, MH88, Ton87]. February [USE89g]. Federal [BB87]. Feeding [HB86]. Feel [Car86, Feu84]. Fetters [MR88a, MR88b]. Fewer [MR88a, MR88b]. Fifth [Som88, USE88c, Ass88e, USE89i]. Figures [BD87b]. File [Ama88, BLMY87, BG88a, Bas81, BP88a, BP88b, BEHW86a, BEHW86b, Bry83a, Bry83b, CFA85, CR89, Cot87, Cyg88, FKV89, Gou86, HLW84, HP89a, HP89b, HH86, How88, Hug86, Hum88a, JN88, KGTM89, Kaz88, Kle86, Koe87a, Koe87b, KK89b, KK89c, LEG88, LL83a, LL83b, MLRC88, McK82b, Nac86, Par88b, Pat83, Pea88, Pos88, RKH86, SJL^{+}87, Sax85b, Sax85c, SCW89a, SCW89b, Spe89, TR84, WDL87, WLS^{+}85, Weh83, Wei84a, WO88, Wil82a, Wil82b, ZDS85, Lil88, Yos85]. Files [Bis88b, Kil84, Lio88, Mey88, Nac88a, Nac88b, Per82b, TRY^{+}87, Woo83b, Lil88]. Filesystem [AF86, MR89, SGK^{+}85]. Filesystems [Haw88a, Haw88b]. Finding [Woo83b]. Fine [MP89a, MP89b, Sau88]. Fine-Grain [MP89a, MP89b]. First [TA82, Tut82, USE88g]. Fish [MBS86a, MBS86b]. Fit [LS85]. Fixes [Tut83]. FL [USE89d]. Flags [Til89]. Flamingo [SA86b, SA86a]. Flavors [Ecc88]. Flexibility [Car82a, Car82b]. Flexible [ABD^{+}89, Chr89a, Chr89b, CS86, Har88b, Mog89]. Flight [PJ89a, PJ89b]. Floating [Gri85, Sch88b, SF86]. flocks [Ing87]. Flow [Hae86a, Hae86b, Hae85a, Hae85b]. Fluid [GPF^{+}86a, GPF^{+}86b]. Focus [WK83a, WK83b]. Focus/USE [WK83a, WK83b]. Folding [Hum85]. Font [Big85, RSSW89, WDL87]. Force [Les83]. fork [SM88a, SM88b]. Format [LL83a, LL83b, WDL87, Cal83]. Formats [Raf87]. Formatted [Raf88a, Raf88b]. Formatter [KV89]. Formatting [Lon82a]. Forms [All87, Joh87, Pyn82a, Pyn82b]. Fort [USE89d]. FORTRAN [GMW86, Ros87d]. Forwarding [O’B85, SW84a, SW84b]. Foundation [ABB^{+}86a, ABB^{+}86b]. Four [Koe87a, Kri86]. Fourth [Fow85, USE87c]. FP [MK84a, MK84b]. FP-Shell [MK84a, MK84b]. Framework [BLK87, CKK87, DS88, Nic89, OMI86]. Francisco [USE88h, Ass88f]. French [Bey88]. Friendly [BC88, Zem83]. friends [DJM86]. Frontend [Hae83]. fsck [BE89]. Full [Lee87b, Ste86, Kat82a, Kat82b]. full-screen [Kat82a, Kat82b]. Full-Text [Lee87b]. Fully [Van88]. Fun [Akk87]. Functional [Bih88, MK85b, Tur87]. Functions [Inm85, SF86]. Future [Bla83a, KL82b, Mas87, Mas88, WB85a, WB85b, Bro87].

G [MLS88]. GA [USE86a]. Galadriel [Lew86b, Lew86a]. gamma [Gra88a, Gra88b]. Garbage [Cap88, WHM89]. Gated [Fed88a, Fed88b]. Gateways [Mog89]. General [CS86, CH83a, CH83b, Han82, Isa83a, LL83a, LL83b, MH88, MK88, WM82, Hen83, McL83]. General-Purpose [LL83a, LL83b]. Generalizations [Ros87c]. Generalized [WDL87]. Generated [Hae83]. Generation [Bam85a, Bam85b, Cha87a, Fow85, Lan86, McG86a, McG86b, Som88]. Generations [Kri86]. Generator [Gra88c, Gra88a, Gra88b, Per85]. Generic [Ale87a, Ale87b, RKH86]. GENIX [Nef83]. geographers [Biv87]. Geometric [Sta87a, Sta87b, Tho85a, Tho85b]. Georgia [USE86c]. Geritol [LNSZ85]. get [O’D83a, RR85]. Getting [Dav89a, Wil83a]. GFS [Koe87a]. Gigabytes [Har88b]. Give [BH86]. GKS [Ric87]. GLA [Gra88c, Gra88a, Gra88b]. Glass [DR86]. GLO [Neu86a, Neu86b]. Global [GMW86, Pow83]. GNU [Lea88, Tie88]. going [O’D83a]. Gone [Alt87]. Good [Feu84]. GOSIP [Bak89]. Got [LNSZ85]. GOTHIC [CMM88]. GOTHIX [Ker88]. Gould [Ble83, FT83]. government [Bak89]. GPIO [SREC88]. gprof [McK83c, McK83d]. GRAB [Les88b]. Grain [MP89a, MP89b]. Granularity [Sau88]. Graph [McK83c, McK83d]. Graphical [BWS87a, BWS87b, CCM87, Che87a, Con87, JA88, Mye86, PG87, Roc89a, SC88]. Graphics [Ale87a, Ale87b, And88a, BB83, BS85a, BS85b, BWH87, Cog87, Dan83a, Dan83b, GPF^{+}86a, GPF^{+}86b, GW86a, GW86b, Hae85a, Hae85b, Hid83, Leb87, McD84a, McD84b, McG85, PJ89a, PJ89b, Pik82, Sha83, Sul87, SSNU87, TG88, Tho85c, USE85d, USE86d, USE87c, USE89i, VL88a, VL88b, Wil88]. Graphs [Leb87]. Grep [Hum88b]. Group [Hag83, KL82a, Ond89, Sof83, Sof84, Usr82, Ass89e, DGM82, USJ83]. Groups [Car88c]. Grouse [Wil87b]. Grows [OL84a, OL84b]. gsck [Sax85b, Sax85c]. Guaranteed [RT83]. Guardians [DM88]. Guest [Rit88]. Guide [Dat88, Dav89a, Sax85b, Sax85c]. Guidelines [Hay88]. Gurus [Kol85].

Hacct [Uit87b, Uit87a]. HACKMAN [SB88a, SB88b]. Hairy [Str87a]. Hanasemasu [JK86]. Handler [AB85]. Handling [Bis83, Dye82a, Dye82b, Eyk88, Mil88, RW86a]. HandS [KNN88a, KNN88b]. Happen [DC85]. Hard [BSR88, Gro87, LLS88, Ros88]. Hardware [Car82a, Car82b, McG86a, McG86b, War84]. Hardware/I [Car82a, Car82b]. Hardware/I-O [Car82b, Car82a]. Harmful [Pik83]. Health [Ham87]. Heap [BL88a, BL88b]. Heap-based [BL88a, BL88b]. Hello [Ros88]. Help [Lio88, BK84]. HEMS [Par88c]. Henderson [Daw82]. Heritage [RP84]. Hesiod [Dye88]. Heterogeneous [BDWW89, CDT89a, CDT89b, FS89, Har88a, HH88, KGL89, MR89, PMD88, PL89, SS88]. Heuristics [Ste89b]. Hewlett [Cle83a, Cle83b]. Hideous [PW85]. Hierarchic [Tra85]. Hierarchical [Lib85a, Lib85b, Lor88c, Wal86a, Wal86b]. Hierarchies [Bry83a, Bry83b]. Hierarchy [FKV89, HP89a, HP89b, MLRC88]. High [Ado89, Bam85a, Bam85b, CCF89, Dan83a, Dan83b, GPF^{+}86a, GPF^{+}86b, Kol86, Min82a, Min82b, PBL86, Ren88a, Ren88b, Ren88c, Wil83b, Wil83c, Pos88]. High-End [GPF^{+}86a, GPF^{+}86b]. High-Performance [CCF89, GPF^{+}86a, GPF^{+}86b, Min82a, Min82b, Bam85a, Bam85b, PBL86]. High-speed [Ren88a, Ren88b, Ren88c]. Highly [ELS88b, Tay88]. Hire [SA88]. History [Fel84, Ker83, Pet83, SV83, Tan84]. HITAC [KAH83a]. hoc [TK88a, TK88b]. Holes [SB88a, SB88b]. Home [Opp89a, Opp89b]. HoneyDanBer [RW86b, RW86a]. Hopkins [KTS^{+}86a, KTS^{+}86b]. Horses [LG88]. HOSE [SC88]. Hospital [KTS^{+}86a, KTS^{+}86b, KTS^{+}86a]. Host [HSHK84]. Hosts [Hil89, McK88a]. HP [KGTM89, Sto87]. HP-SDD [Sto87]. HP-UX [KGTM89]. HP9000 [Lin84]. HPC [Kat89]. HPC/VORX [Kat89]. HSG [KGTM89]. HUB [O’D87a, O’D87b]. hybrid [Sto88]. Hygiene [Ste89a]. HYPERchannel [Wat88b]. HYPERchannel-Based [Wat88b]. Hypercube [CM89b, CM89a]. Hypertext [Bro89b, Nic89, Wal87].

I-O [Car82b, Car82a]. I/O

[DP83a, Haw89b, Orr83, Raf87, Raf88b, RLML86a, Rob84a, Str85b, vMM88, vM88, DP83b, Haw89a, Raf88a, RLML86b, Str85a]. IAFORM [Pyn82a, Pyn82b]. iAPX286 [Bar83]. IBM [EGL86, Eng88, Tan87a, Wil83b, Wil83c]. Icon [Gri89]. Ideas [CJ88]. Identifying [LAKS88]. Idle [Lit87]. IEC [Dun89]. IEEE [USE88c]. II [ELS88a, HP89a, HP89b, KAH83b]. III [CQ83a, Dav89a, LK82, LZ82, USE89e, Zuc83b, Zuc83c]. ILMON [BM87]. I’m [Lio88]. Image [Bee86, Cog87, Coh87, Gom85, Kin86, Sau88]. Imbalances [McK88a]. Immovable [Les83]. Impact [CKM85]. Impersonal [Tay86]. Implementation [ASS85, AK88, Baa88, BKT89, BL88a, BL88b, Bar83, Bas81, Bis88a, CE89a, CE89b, Cic88, Col84a, CFA85, DF84, Fis86c, Fis86d, Hat82, Hen83, HCE87, HOG88, JC89a, JC89b, Jun85a, Jun85b, Kar83, Kep85, Ker84, KLB89, KM87, LP89, LW89a, LW89b, Len86b, Lin84, LC87a, MP84, Mor88a, NHR84, Par88c, PBL86, RK89, ROS87a, dR85, SGK^{+}85, SW88, SCC86a, SCC86b, War83, Wil83b, Wil83c, YSF89, Cor82, Daw82]. Implemented [Gou86]. Implementing [HRO82a, BD86, DRK^{+}89, Get86, HRO82b, Hil89, Inm85, KT88a, KT88b, Nyb86, OTW85, Par87, Ros87c, SLM89]. Implementors [DGM82, USJ83]. Implets [CN88]. Implications [DNQ^{+}83]. importation [Egg89]. Imposing [Orr83]. Improved [Pea83]. Improvements [MK85b]. Improving [Gri85, Jus89, LKM84]. inch [Kri84a, Kri84b]. Include [Raf87]. Incorporating [ESS89, Kra88a]. Incremental [Hum88a, Par88b]. Independent [Hid83, KM84b, KM84a, KAH83a, Mac83a, Mac83b]. Indexes [Les88b]. Indices [Zho87]. indirect [Lau81a]. Infinite [Sch88b]. Influence [Cab86]. Informal [Par86]. Information [HK83, Hoo83, KTS^{+}86a, KTS^{+}86b, Lee87b, Red85, RW86b, RW86a, Wai82, WS83a, WS83b, Wat88c]. Informix [Kin83]. Inheritance [Str89b]. Inhibit [LG88]. inner [GS87]. Inpatient [KTS^{+}86a, KTS^{+}86b]. Input [Tuo83]. Ins [Spa89, RP84]. Insecure [NS88]. Installation [BH86, BC89, TP86, USE87d, USE88f, USE89e]. Installations [Hay88, Lee87a]. Instruction [Kle89, Mor88e]. Instrumentation [O’R88]. Instruments [Kod82a, Kod82b]. Integral [HCN85]. Integrated [BW89, Boy84, BPM87, Cog87]. Integrating [Lee89]. Integration [BSR88, URK85, RGDP88]. Integrity [Ful89a, Ton87]. Intel [Bar83, Lev83]. Intelligent [BFGK89, BNB87, FL82, Spe87]. Interactive [Hae85a, Hae85b, Hae86a, Hae86b, HK83, Hop89, Jac83, KAH83b, Lev83, Mye86, Phi84, Ste85, WS83a, WS83b, MD87]. Interactivity [Hos84]. Interchange [Nic85]. Interconnecting [SS88]. Interface [AW89, BEHW86a, BEHW86b, BNB87, BLSS83, CS86, ESS89, Fos83a, Fos83b, Gan86a, Gan86b, Gol88, GBM87a, GBM87b, Har85a, Har85b, IvW87, Isa83a, MN85a, MN85b, MS89, Old88a, Old88b, Oti88, Per82b, Per83, Roc89a, Rug83, SA86a, SA86b, TRY^{+}87, Biv87, Kat82a, Kat82b]. Interfaces [BLK87, CCM87, KNN88a, KNN88b, SA88]. Interfacing [Duf82]. Intermediate [KM85]. Internet [BD86, Hed89, Mur88c, Sku88, TPRZ84]. Interpolation [Seq86]. Interpret [MC85a, MC85b]. Interpretation [Hef82]. Interpreter [Feu85, KLP88a, KLP88b, RK88].

Interpreter-based [KLP88a, KLP88b]. Interpreting [Sax85a]. Interpretor [BQd86]. Interprocess [Joy82a, PR85, PA89, Ste88a]. InterViews [LC87a]. Interwindow [Gil86]. Introducing [McL84]. Introduction [Ger82, Hed89, Rob89]. Intruder [HE88]. Intrusion [BK88a, BK88b]. Inventing [Spe88, Hir83]. Inverted [Les88b]. Invoking [MMTW88]. IP [FFH^{+}86, Fer85, JLSG84, KM86, LP89, NB84, WG84]. IP/TCP [KM86]. IPC [BWH87]. Ireland [USE87b]. Iris [Gan88b, RHH85a, RHH85b, Gan88a]. Iron [Miy86]. Irresistible [Les83]. IS/1 [Tor83a, Tor83b]. IS/3 [Zuc83b, Zuc83c]. ISAM [Wil82a, Wil82b]. isn’t [Hoo83]. ISO [Bol88, Dun89, HOG88, KGTM89]. ISO-9660 [KGTM89]. ISO-9660/HSG [KGTM89]. ISO/IEC [Dun89]. Isolation [HE88]. Issues [AF87, Fer82, Kaz88, NSB85, SLM89, Ton87]. Iterative [Mer82]. ITTDCD [Hid83]. ix [SW88].

January [USE84c, USE85c, USE86e, USE87g, USE89g]. Japanese [JK86]. Job [KK89a, Len86b]. Johns [KTS^{+}86a, KTS^{+}86b]. Joint [Usr82]. journal [UE88]. JTC1 [Dun89]. JTC1/SC22/WG15 [Dun89]. Judicial [BB87]. July [Sof83, Usr82]. June [Sof84, USE85e, USE86c, USE87f, USE88h, USE89f]. JUNET [Mur88a, Mur88b]. just [Hoo83].

Kanji [JK86]. keep [PS89]. Keeping [Ing87]. Kerberos [SNS88]. KERMIT [Col84a]. Kernel [ABB^{+}86a, ABB^{+}86b, CMM88, Eyk88, FKN85a, FKN85b, FH88, GGSW88a, GGSW88b, HC88, HMP83a, HMP83b, Kel89, Kor89a, Kor89b, LJ84, LB89a, LB89b, Len86a, Lin84, LS83a, McK88b, MK88, MMTW88, NY88, NLR84a, NLR84b, PMI88, Req85, SH89, TRG^{+}87, War84, Zim85].

Keystroke [WK83a, WK83b]. Kit [TvK83, KTv83]. Know [Bal83, LK82]. Knowing [Feu84]. Knowledge [GBM87a, GBM87b, HQZ^{+}87, Ohk84, Sam87]. Knowledge-based [Sam87]. Kodak [Les88d]. KSH [Kor83a, Kor83b].

Labeling [FW89]. Laboratories [Ham87]. Laboratory [Sul87]. Labs [Pre88]. Lake [Sof84, USE84a, USE84b]. lambda [CB83]. LAN [HOG88]. Land [MR88d, MR88e]. Language [Bec84, Bey87, BQd86, Cla87a, Cla87b, CH83a, CH83b, Cox82, Cox83, DJ88, FKT83, Feu85, Gri89, HZ89a, HZ89b, HMM^{+}88a, HMM^{+}88b, Hor82a, Isa83a, Jen83, Joh87, KLP88a, KLP88b, KT88a, KT88b, Koe85, KM87, Kor83a, Kor83b, MP84, Mil88, Ros87b, RS87, Sch82, Sch86b, Sch86c, Ter87, Tut82]. Large [AF87, Bec84, Bil86, Bis83, Bob88, ELS88a, GP88, Har88b, Hun88a, Lee87a, Mar84, PMD88, PF84a, PF84b, PL89, Sim88, STA86, TP86, USE87d, USE88f, USE89e, Wat88c, ZP89]. Large-Scale [ELS88a, GP88]. Latency [Len86a]. Latent [Fil85]. Later [Lyc85]. Lauderdale [USE89d]. Layered [Lin84]. Lazy [Dat88, Epp89, LB89a, LB89b]. Leading [Kra83]. Learn [O’D83c]. Learning [WHM89, Sal89b]. Lengths [Zho87]. Lessons [Bra89]. Level [Bre88a, Jac84a, Jac84b, Lib85a, Lib85b, ST89a, ST89b, SM89, WO88, Epp89, MR88c, Kno87]. Levels [Isa83b, Lin88b]. Leverage [Mas87]. LEVI [MN85a, MN85b]. lex [CM83, Pax84, Jac87]. Lexical [Bre88b, Gra88c, Gra88a, Gra88b]. Liability [NSB85]. libc [Kuc89]. libg [Lea88]. Libraries [Arn86, DF84, GLDW87, Sun89]. Library [Fuh87a, Fuh87b, Gan88a, Gan88b, Gor87a, Gor87b, Lea88, Les88c, Oti88, Sch86a, Sch88b, SSNU87, Tri89]. License [FHW88, RP84]. Licensing [Isl83a, Isl83b, Mos82, OLJ^{+}88, Ort88]. Life [LNSZ85, PB84, Tan87b]. light [Rob87].

Lightweight

[GM89, Kep85, O’D87a, O’D87b]. like [Kor89a, Ros87b, MLRC88, Bee84a, Bee84b, Kor89b, LM88a, LM88b]. Limited [War82]. Limiting [MT89]. LINCS [Req85]. Linda [KBT89]. Line [AW89, Coh87, CAG89a, CAG89b, Der83, Mas83b, Mas83c, Per85, ZP89, Kle87, NM83a, NM83b, Tal89a, Tal89b]. lines [Lau81a]. link [GS87]. Linkage [Str88b, Str88c, Str88d, Str88e, Str88f]. Links [Fer85, Gle89]. Lint [Kor89a, Kor89b]. Lint-like [Kor89a, Kor89b]. LINUS [Kra83]. LIPs [Ohk84]. LISP [DM83, Ecc88, Tri87, ZH88]. List [Lew86a, Lew86b, SV83, Wat83]. List-Based [Lew86a, Lew86b]. Lists [FA88, Kra88a]. Lived [Koe87a]. Livermore [Phi84]. Load [Ber86, Cab86, McK88a, Zho87, Zuc83a]. Loader [Tim85]. Local [GB83, Koe87a, Sal82, Wam83a, Wam83b, Wat88b]. LOCK [SW88]. LOCK/ix [SW88]. Locking [Bas81, GM89]. Locus [BP84]. Logic [KT88a, KT88b]. Logic-Based [KT88a, KT88b]. Logical [Kat82a, Kat82b]. Login [Con88, Lee87a]. LOGIX [Kat82c]. Logo [Har83]. London [USE88e]. Lookalike [All83a, All83b]. Lookaside [TBJW88]. Loosely [BFS89, Inm85]. Loosely-Coupled [BFS89]. Loosing [Ban82]. Losing [Lio88]. Lot [LNSZ85]. Lots [Til82]. Louisiana [USE89j]. Low [JHRR86, Les88b, Tho85c, WK83a, WK83b, Col88]. Ltd. [Ree82a, Ree82b]. Lucasfilm [HL85a, HL85b, Ree82a, Ree82b]. Lucasfilms [Law83a, Law83b]. LWP [SS88].

M [KAH83a]. M-series [KAH83a]. M4 [MAB83]. MA [USE82a, USE82b, USE87c]. Mach [ABB^{+}86a, ABB^{+}86b, BL89, TRG^{+}87, TRY^{+}87]. Mach/4.3BSD [BL89]. Machine [BWP85, BDWW89, BLK87, DG87, Gen86, Hum89, Joi87, KAH83a, KNN88a, KNN88b, Daw82, Del87, KNI88]. Machines [Bis88b, Duf82, Dye82a, Dye82b, KAH83a, Ond89, WJ82]. Macintosh [Fri87, Seq89, SGS89]. MacMix [Fre85]. Macro [Bol88]. Madness [Tut83]. Mail [All83c, Alt87, Bis87b, CP84a, CP84b, DeJ86a, DeJ86b, HH86, HP85, IvW87, Ker84, Kim87, Man87a, O’B85, O’D83b, Par86, Sal89a, Sve83, Tay88, Wei84b, PG88c, Pre85]. Mailer [OK85a, OK85b]. Mainframe [HP89a, HP89b]. Mainframes [Ste86]. Maintaining [Har87d, Kal82, MAB83]. Maintenance [Bry83a, Bry83b, Kim87, TP86, Van88]. Maintenance/Distribution [Kim87]. Maitre [Ber86]. Major [DR86]. Make [Baa88, Bak89, Kin86, MAB83, Mor88b, Mor88c, DJM86, Hum87, Som88, Hir83, Fow85, Nov83]. Makealiases [PG88c]. Makefiles [MAB83]. Makeup [KV89]. Making [Kuc89, Lin88a, Nac88a, Nac88b, Sim88, TvK83]. Malloc [KV85]. Man [BLK87, Cla87a, Cla87b, Dat88, KNN88a, KNN88b, SH85]. Man-Machine [KNN88a, KNN88b]. Manage [Nic89]. Management [Akh87, Ben82, Bla89, BLSS83, CKM85, CS86, CR89, Cot87, DF89, ELS88a, ELS88b, Fen87, FHW88, Ful89b, Gee88a, Gee88b, Gle89, HSY88, HP89a, HP89b, HM89, KM86, Kiv84, LJ84, Lee87a, Lee87b, Lor88c, Man83a, Man83b, MF83, Mok88, Nef83, PL89, RGL88, SA86a, SA86b, SM88a, SM88b, Sti83, Tes86a, Tes86b, Tra85, USE89a, Wai82, War83, Wat88c, Wil89, YSF89, Zwi89, ML88, USE89j]. Manager [Bel88a, Hae86a, Hae86b, Jac84a, Jac84b, Lew86a, Lew86b, Mil84b, RHH85a, RHH85b, Wil82a, Wil82b, Bel88b, RSSW89]. Managers [McG85]. Managing [EST86, Per87a, Sle87, Woz82]. Mandatory [FW89]. manipulate [Yos85]. Manipulating [LC87b]. Manipulation [Les88c, Mil87]. Manual [NM83a, NM83b, Don89a]. Manufacturing [Dix82, SM84]. Many [HZ89a, HZ89b, Hil89, TH83a, TH83b]. Mapped [Mey88, TRY^{+}87]. Mapping [KSH83, KSH84]. March [Mas83a]. Market [Den83, FKN85a, FKN85b]. Marketplace [Tan87b]. Markets [Ela83, Kat83]. Maryland [USE89f, Cot87, TW84]. MASCOT [Jac82]. Mass [AN88, HP89a, HP89b]. Mass-Storage [AN88]. Massively [EGL86]. Matches [Car82a, Car82b]. Math [SF86]. Maturation [Fun87]. May [Ass88e, USE89h]. MD [USE89b]. Me [Bey87]. Measurement [Isa83b, RSV85]. Measurements [JSW87, STA87c]. Measures [Les88d]. Measuring [BG88a, LKM84]. Mechanism [Eyk88, Kle85]. Media [Mye86, BERS88a, BERS88b]. Medical [GBM87a, GBM87b, Ham87]. Meet [BL88a, BL88b]. Meeting [Dun89, Hos83]. Meets [Joh88a, Joh88b, Les83]. Members [LS88]. Memory [Akh87, BC84, Cap88, Cla88, ELS88a, FBS89, GMS87, HP89c, JC89a, JC89b, LJ84, LB89a, LB89b, MK85a, MK88, Mil84b, MF89, Mor88a, Muu87, Nef83, Par88a, RK89, SM88a, SM88b, SCC86a, SCC86b, TRY^{+}87, Til82, ZH88]. Menu [Bla83b, Bla83c, Hef82, Ney83a, Ney83b]. Menu-Driven [Bla83b, Bla83c, Ney83a, Ney83b]. MENUNIX [Per82b]. Menus [Hop87a, Hop87b]. Merging [Pik82]. Message [BERS88a, BERS88b, PY84a, PY84b, Son88a, Son88b]. Message-Based [Son88a, Son88b]. Messages [Wil87b, RR85]. Messaging [DN87]. Mether [MF89]. Method [Ric85, Yos85]. Methodology [STA87c]. Methods [Cox82, Cox83]. Mex [RHH85b, RHH85a]. MGR [Uhl87]. MH.5 [RR85]. Micro [Ben82, Bis83]. Microcomputer [PY84a, PY84b]. Microcomputers [Jun85a, Jun85b, Per87a, Rob84a, Sti83]. Microfabrication [Seq89]. Microprocessor [Lyc85]. Micros [ACF^{+}86, Sch83a]. Microworlds [BPM87]. MIDI [Haw86a, Haw87a]. Migration [AK88, Dou89, MS88, MSM88c, MSM88a, MSM88b, MSM88d]. Mine [Nor84]. Mini [Miy88]. Mini-supercomputer [Miy88]. Minimal [DR83b]. Minimalist [HP89c]. Minis [Sch83a]. MINIX [GD89, Tan87a]. MIPS [PY84a, PY84b]. Miranda [Tur87, Tur88]. Miro [HMM^{+}88b, HMM^{+}88a]. Miscellaneous [Bak89]. Mistress [KL82b]. Mixed [Bec84]. Mixed-Language [Bec84]. Mixing [Fre85]. Mk [Hum87]. Mkuser [PS89]. MLS [FW89]. Mm4 [MAB83]. MMDFII [Kin84]. Mode [LS83a, War84]. Model [BKT89, Bam85a, Bam85b, BS85a, BS85b, BO87, BWH87, Cla89, FA88, Kno87, LA89a, LA89b]. Modeling [JHRR86, LOR88a, LOR88b, RWNA87, Seq85, Sta87a, Sta87b]. Modelling [Bru88, Con87, Wal86a, Wal86b]. Modems [Sle87]. Modest [Hos84]. Modifications [McK88b, Ree81]. Modifying [ZP89]. Modix [SH89]. Modula [DO85, Pow84]. Modula-2 [DO85, Pow84]. Modular [Boy84, Seq85]. Modularity [KLB89]. Module [HCE87]. Modules [Zuc83a]. Monitor [BFGK89, HSYY89, Spe89]. Monitoring [BM87, Bob88, HE88, JA88, Kat89, Mey85]. Monotonic [LN88]. Monterey [USE85d, USE86d, USE88f, USE89i]. MOS [BP86]. MOSIX [BW89]. Motel [Hum88a]. Mountain [DGM82]. Mounts [Sau88]. Mouse [Fre85]. Movie [JHRR86]. MP [Eng88, Par88a]. MS [URK85]. MS-DOS [URK85]. MSG [Mer82]. MSS [HP89a, HP89b]. MSS-II [HP89a, HP89b]. Multi [BEHW86a, BEHW86b, BERS88a, BERS88b, Cra87a, Cra87b, Fed88a, Fed88b, Har87a, Har87b, HR85a, HR85b, JS89a, JS89b, Rei89, ST89a, ST89b, Tes86a, Tes86b, FKN85a, FKN85b]. Multi-CPU [FKN85a, FKN85b]. Multi-Level [ST89a, ST89b]. Multi-media [BERS88a, BERS88b]. Multi-platform [Rei89]. Multi-process [HR85a, HR85b]. Multi-Processor [JS89a, JS89b, Tes86a, Tes86b]. Multi-Representation [BEHW86a, BEHW86b]. Multi-Routing [Fed88a, Fed88b]. Multi-System [Har87a, Har87b]. Multi-User [Cra87a, Cra87b]. Multicasting [Per87b]. Multilevel [MR88a, MR88b, MR88c]. Multimedia [CFLT87]. Multiple [HRO82a, HRO82b, Hun88a, Jac86a, Jac86b, Jaf87, KM82, Kle86, KM87, Lib87, Mil89, Pat83, Str89b]. Multiple-Process [HRO82a, HRO82b]. Multiplexed [KAH83b]. multiplexing [Lau81a]. Multiplexor [Pik84a, Pik84b]. Multiprocessor [BB84, BW89, BK85, BFS89, BC84, CRJ87, EBFH85a, EBFH85b, GB89, Gou85, HC88, HCN85, JC88, Kel89, LLM87, Mil86, PBL86, RSV85, SLM89, TBJW88, USE89d, USE89k, vMM88, vM88]. Multiprocessors [ELS88a]. Multitasking [Fou88, Rei89]. Multithreaded [Kor89a, Kor89b, MS89]. Multiuser [Mil86]. Multivariable [Cic88]. MUSH [Elz84]. Music [Fre85, Haw86a, Haw87b, Kee88, Lan86, Fox87, Fox87]. Musings [Spa89]. MUSK [Cra87a, Cra87b]. Mutation [CM89a, CM89b]. MX [Lau81a]. My [Bey87, Lio88].

Name [BD86, Dye88, PW85, TPRZ84]. Names [Par86]. NAPS [BG88b]. Nation [PP85]. National [Hag83, Kod82a, Kod82b, NMS83, SJ83]. Native [Ter87]. Natural [Har85a, Har85b]. Navy [Gro82, SSNU87]. NBS [Che82]. Need [Col87b, FKN85a, FKN85b]. Needs [BL88a, BL88b]. Negotiating [JLSG84]. NERECO [STV87]. Nervous [McK88a].

Nest [BSY88b, BSY88a]. Nested [Duc89, Epp89]. Net [Ker84, Lau81c]. Netdump [Haw88b, Haw88a]. NETIX [Wam83a, Wam83b]. Netnews [Wei85a, Wei85b, NSB85, TH86, Wei84b]. Network [ACF^{+}86, BSY88a, BSY88b, BM87, Bis88b, BO83a, BO83b, BP84, DeJ86a, DeJ86b, DLM^{+}87, Fen87, Fer85, FHW88, GB83, GZ84a, GZ84b, HS89, HH88, Hor83, Joi87, KM86, KGL89, KSH83, KSH84, Lef82, Mil84a, Muu87, Nac86, NS88, PMD88, PG88a, PG88b, Phi84, Pre88, Rod87, SGK^{+}85, SCW89a, SCW89b, Sim88, SNS88, Sun89, TGB^{+}89, WLS^{+}85, Wat88b, Wei84a, Wei84b, Yam88, YSF89, Bak89, Lil88, Pre85, Gou86]. Networked [Pic83, Rei89, ZP89]. Networking [CKM85, Mul87, OMI86, Req85, RW86b, TG86, Wam83a, Wam83b, Ren88a, Ren88b, Ren88c]. Networks [BDWW89, MS88, Sal82, Sch89, Ste84, Sal89a]. Newcastle [Bla83a]. News [Col87b, Dun88, GRS88a, GRS88b, Hor83, Per87b, Pre82a, Pre82b, Den86, Den89, Opp89a, Opp89b, Rob87, Sch88a]. Newsletter [Ass89f, Ass89e]. Next [Cha87a, McG86a, McG86b, Haw89a, Haw89b]. Next-Generation [McG86a, McG86b]. NFS [CGFCKT88, Jus89, Lil88, RWFC86]. NFSSTONE [SCW89a, SCW89b]. NIAL [Jen83]. NIDX [BK88a, BK88b]. night [Ing87]. Nihongo [JK86]. NIX [Cap82a, Cap82b]. NLS [Ort88]. NM [USE87a]. Non [Pea83]. Non-Paged [Pea83]. Notesfiles [Ess85]. Notice [Erl88]. Noticeable [Kra83]. Notification [DEF^{+}88]. Notifier [Eva86]. NOTREACHED [DC85]. NOVA [Han82]. November [USE86d, USE87a, USE88f, USE89i]. Novice [Smi87a]. NRS [Bro87]. NS16000 [ZS83]. NS16032 [SJ83]. Number [McK88b, DJM86]. number-crunchers [DJM86]. Numeric [BB83]. Numerical [Ste88a]. NUnix [Tes82]. NYU [EGL86].

o [JK86, Car82a, Car82b, DP83a, DP83b, Haw89a, Haw89b, Orr83, Raf87, Raf88a, Raf88b, RLML86a, RLML86b, Rob84a, Str85a, Str85b, vMM88, vM88]. Object [Ale87a, Ale87b, AQ84, BKT89, Bih88, Bla89, Car87a, Car87b, CKK87, Cox82, DRK^{+}89, Den86, Fuh87a, Fuh87b, GGSW88a, GGSW88b, Gor87a, Gor87b, JC89a, JC89b, LW89a, LW89b, Les83, LL83a, LL83b, NLR84a, NLR84b, O’D87a, O’D87b, SE88, Sha89, SGH^{+}89, SA86a, SA86b, Str87d, Str87e, VL88a, VL88b, Wal86a, Wal86b, WRM^{+}89]. Object-Based [GGSW88a, GGSW88b, WRM^{+}89, NLR84a, NLR84b]. Object-File [LL83a, LL83b]. Object-Oriented [Ale87a, Ale87b, Bih88, Bla89, Car87a, Car87b, CKK87, Fuh87a, Fuh87b, Gor87a, Gor87b, JC89a, JC89b, LW89a, LW89b, Sha89, SGH^{+}89, SA86a, SA86b, Str87d, Str87e, VL88a, VL88b]. Objective [Cox83]. Objects [DV89, SL88, SREC88, Sta87a, Sta87b]. Observations [Miy88, TBS87]. October [USE87c, USE88a, USE88k, USE89d]. ODEs [Ste88a]. OEM [Mee85]. offer [Biv87]. Offering [Guf83, San83]. Offerings [Isl83b]. Office [Che82, Ney83a, Ney83b, van86]. Offs [Kri84a, Kri84b]. OFS [Ama88]. Ohio [Woh88, Zwi88a, Zwi88b]. OLC [CAG89a, CAG89b]. Old [LNSZ85]. On-Line [CAG89a, CAG89b, Mas83b, Mas83c, Kle87, NM83a, NM83b, Tal89a, Tal89b]. On-Screen [Pyn82a, Pyn82b]. Once [LEG88, Spe88]. One [HZ89a, HZ89b, Lib87]. Online [Cla89, Les88c]. Only [Spe88]. Ontario [Sof83, USE83b]. ONYX [Cor82]. Op [Chr89b, Chr89a]. Open [CCM87, SE88, SNS88, TM82]. Openness [Dav88]. Operating [Bar88a, Bar88b, Bec84, CRJ87, CDT89a, CDT89b, Col84c, Dan83a, Dan83b, DV89, DJ88, DP83a, DP83b, EBFH85a, EBFH85b, Gie83, GB89, GD89, HK86a, HK86b, JC88, KNI88, KK89a, Kiv84, Kra88a, LM88a, LM88b, Lee89, LH84a, LH84b, MC85a, MC85b, MH88, McG85, MFS89, RL88, Ral88a, Ral88b, RKPP88, RGDP88, RAA^{+}88, Sal88, Sam87, SLM89, Sha89, SGH^{+}89, Shu89, ST89a, ST89b, Sve83, TM82, Tes86a, Tes86b, USE88c, Ass88e, Ups82, UJ83, Wam83a, Wam83b, WRM^{+}89, WJ82, Don88]. Operations [SM88a, SM88b]. Operators [Win88]. Optical [Ama88, GRS88a, GRS88b, Kiv84, LEG88, YKK89]. Optimization [Rob84a]. Optimizations [HCE87]. Optimizer [AQ84, GMW86, KM85, Tim85]. Optimizing [Gri85, LR84, Mey82b, Pow83]. Option [Per85]. Options [Kri84a, Kri84b]. OPUS [Bak89]. ORE [Jam88]. Oregon [USE85e]. Organization [Isa83b]. Orientation [Duf85, Man87b]. Oriented [Ale87a, Ale87b, Bih88, Bla89, Car87a, Car87b, CKK87, Cox82, DRK^{+}89, Den86, Fuh87a, Fuh87b, Gor87a, Gor87b, JC89a, JC89b, Lad88a, Lad88b, LW89a, LW89b, Orr83, Pik84a, Pik84b, Sha89, SGH^{+}89, SA86a, SA86b, Str87d, Str87e, Sun88, VL88a, VL88b, Wil82a, Wil82b]. Original [AGHR89a, AGHR89b]. Orleans [USE89a, USE89j]. OS/2 [Bak89, Car87c]. OSI [Bak89, FFH^{+}86]. OSx [Bot84]. Othello [Col80]. Other [Tay86, Elz84]. our [LS83b]. Out-Of-Band [Rag89a, Rag89b]. Outline [RSSW89]. Output [KV89]. Overhead [Les88b]. Overview [CC86, Dav89c, Dyk87, GR85, How88, Joy82b, Ker88, OCD^{+}87, PHS^{+}88a, PHS^{+}88b, Per85, RFH^{+}86a, RFH^{+}86b, Sch88a, Ter87, Tur88, WLS^{+}85, Wat88a]. Oxford [Bra89].

PA [USE87d, USE88i, USE89h]. Pacific [Car88a, Car88b]. Package [BPM87, Dan83a, Dan83b, Der83, Dyk87, Ell85, Hor82b, Lad88a, Lad88b, Mar84, NB84, Pyn82a, Pyn82b, WH83, ZDS85]. Packaged [Hos84]. Packages [Mac83a, Mac83b]. Packard’s [Cle83a, Cle83b]. Packet [NY88]. Packets [Les87]. Page [KV89]. Paged [Jun85a, Jun85b, Pea83]. Pages [SH85]. Paging [Mil84b]. Panel [Tri89]. papers [USE88d]. Papillon [Che87a]. Paradigm [KBT89]. Paradigms [LGZ88]. Parallel [Baa88, BO87, Bob88, Bre88a, Bry88, DG87, EGL86, ELS88b, Gle89, JTUB85, Kuc89, LM88a, LM88b, LFN^{+}89a, LFN^{+}89b, Muu87, Ros87b, RS87, SC88]. Parallelization [BL89]. Parameterized [Str88a, Str89c]. Parlez [WB85a, WB85b]. Parlez-Vous [WB85a, WB85b]. PARPC [MBBP89]. Parser [Bee84a, Bee84b, Gra87, HP85, Per85]. Part [PJ89a, PJ89b, RW86b, RW86a]. Part-Task [PJ89a, PJ89b]. Partial [Col89, DW88, Kno87]. PARX [LM88a, LM88b]. Pascal [GMW86]. PASM [LFN^{+}89a, LFN^{+}89b]. passwd [Hoo83, Lil88]. Password [And88b, Cot87, DS88, Hun88a, RU88a, RU88b]. Past [Mas87, Mas88, WB85a, WB85b]. Patch [BC89]. Patchwork [BS85b, BS85a]. Path [Fos88]. PATHALIAS [HB86]. Pathname [MC85a, MC85b]. Paths [Bre83]. Pattern [Cog87]. PC [SJL^{+}87, Tan87a]. PCTE [KL87a, KL87b]. PDP [Kar83, See83, Tuo82]. PDP-11 [Kar83, Tuo82]. PDP-11/44 [Tuo82]. PDP-11s [See83]. Peace [Mul87]. Peaceful [JLSG84]. Ped [Rug82b, Rug82a]. People [McL84]. perform [Yos85]. Performance [BFGK89, Bob88, BL89, CKM85, CCF89, Dan83a, Dan83b, Dro84, Fed83, Fer82, Geo82a, Geo82b, GPF^{+}86a, GPF^{+}86b, Goo84, Gri85, JA88, Jus89, KM86, Kat89, Kol86, KM83, Lan84, Leb87, LLM87, LKM84, MT89, McD87, MK85b, Mey85, Min82a, Min82b, Miy88, MC82, Per87a, Pos88, RW86b, RSV85, Sam87, Sch83a, SCW89a, SCW89b, SD87, STA87c, Wil83b, Wil83c, Bam85a, Bam85b, PBL86]. Performance-Constrained [Per87a]. Periodic [Mok88]. Perkin [Car82a, Car82b, Pre82a, Pre82b, Tuo82]. Perkin-Elmer [Pre82a, Pre82b, Tuo82]. Perkin-Elmer’s [Car82a, Car82b]. Permissions [MC85a, MC85b]. Person [Yos85]. Personal [Big85, Haw85, O’B82, Sha83]. Personalizing [Tay86]. Perspective [Fai86, Fun87, Lee87b, WB85a, WB85b]. Perspectives [SGH^{+}89]. PEX [TF89a, TF89b]. Phi [Cap82a, Cap82b]. PHIGS [Bru88]. Philadelphia [USE87d]. Philosophy [BL89]. Phoenix [USE87f]. Physical [HP89c]. Pi [Car87b, Car86, Car87a]. Pictorial [Mye86]. Pie [Hop87a, Hop87b]. Pipeline [Gom85]. Pitfalls [DNQ^{+}83, Don89b]. Pittsburgh [USE88i, USE89h]. Place [TA82]. Plan [Pre88]. Planning [Cha87a, Nor84, RWNA87, Wat88c]. Plans [Bla83a, KL82b]. Plasm [MBS86b, MBS86a]. Platform [CDT89a, CDT89b, Rei89]. Plexus [Pic83]. Plotting [And82a, And82b, Mac83a, Mac83b, Spe87]. PLP [HOG88]. PMON [JA88]. PMothra [CM89a, CM89b]. Point [Fer85, Gri85, Sch88b, SF86]. Point-to-Point [Fer85]. Pointers [LS88]. Policies [Lor88c]. Policy [FW89]. Political [Tan84]. Pollster [CM86b, CM86a]. Polyhedra [Hum85]. Poor [Cla87a, Cla87b]. Port [Ble83, GD89, PF84a, PF84b]. Portability [Fil85, O’D83c, Roc89b, Roc89c, RWFC86, Sch82, SM89, Til83, Ups82]. Portable [Bee84a, Bee84b, HMP83a, HMP83b, KM85, Kri86, LR84, LZ82, Rug82a, Rug82b, RSW83, Ste83a, Ste83b, Sve83, TvK83, Tim85].

Porters [DNQ^{+}83]. Porting [ACF^{+}86, CB83, Ecc88, Eyk86, Haw86b, Jun85a, Jun85b, MK85a, O’B82, VB83, WJ82]. Portland [USE85e, USE85a, USE88g, USE88b]. Ports [GB83, Sle87]. Positioning [Ste89b]. POSIX [Bro88, Dun89, Old88a, Old88b]. Possible [Str87c]. Postman [Tay88]. Postprocessing [KV89]. POSTSCRIPT [BD87b, BQd86, Hop89]. pounds [Dro84]. Powerful [WH83]. Practical [CQ83a]. Practice [LM88c]. Pre [Cox82]. Pre-Compiler [Cox82]. Precision [Sch88b]. predictable [TK88a, TK88b]. Preemption [Len86a]. Preparation [vH87]. Preprocessor [BD87b, Fox87, Man87b]. Present [Mas87, Mas88]. Presentation [BBT83, War82]. Presenting [Sau88]. Pricing [Isl83a]. Prime [WJ82]. Primer [Joy82a]. Primitives [LN88]. Principles [Big85]. Printed [SM84]. Printer [ZP89]. Printing [Gol88, JH86, Spe87, Fox87]. Printing/Plotting [Spe87]. Priv [Hei87]. Privileges [Elz84, Win88]. Problem [Har87a, Har87b, Tie88]. Problems [Get86, LLM87]. Procedural [Seq86]. Proceedings [Sof83, Usr82, USE82a, USE82b, USE83a, Ass83a, USE83b, Ass83b, USE84a, USE84b, USE84c, USE85c, USE85d, USE85a, USE85b, USE86d, USE86a, USE86b, USE87a, USE87b, USE87c, USE87d, USE87e, USE87f, USE88a, USE88e, Ass88e, USE88f, USE88g, USE88h, Ass88f, USE88j, USE88b, USE88i, USE89d, USE89e, USE89f, USE89g, USE89a, USE89c, USE89b, USE89h, Sof84, USE85e, USE86c, USE86e, USE87g, USE88k, USE89j]. Process [HRO82a, AK88, BO87, Dou89, ELS88b, HRO82b, Hun88b, Hun88c, KK89a, Len86a, Lib87, MS88, Mey85, HR85a, HR85b, RR85]. Processes [ABD^{+}89, BEHW86a, BEHW86b, BFS89, Kep85, Kil84, LGZ88, Hun88b, Hun88c].

Processing [Ado89, ACK89, BB83, Bar88a, Bar88b, Ber85, CCF89, Cla89, HCN85, HSYY89, JTUB85, KM87, LFN^{+}89a, LFN^{+}89b, RT83, Tal89a, Tal89b, USE89h, Zem83, vMM88]. Processor [Col83, DP83a, DP83b, GM82a, GM82b, Gri85, Jac86a, Jac86b, JS89a, JS89b, Lev83, Pat83, Tes86a, Tes86b, Bro87, DJM86]. Processors [Kri84a, Kri84b]. Product [Isl83b, Les88d, NMP82]. Production [Gom85]. Productivity [Pan88]. Professional [Bak89]. Profiler [McK83c, McK83d, ZH88]. Program [Bis87a, Bob88, Bry83a, Bry83b, Che89, Cra87a, Cra87b, LC87b, MC85a, MC85b, Nac88a, Nac88b, PK88, Ros88, Sax85b, Sax85c, Sch86a, Ste85]. Programdb [Kal82]. Programmers [Gra88c, Gra88a, Gra88b]. Programming [BS85a, BS85b, BO87, Boy84, Bre88a, Car87a, Car87b, Cox82, Cox83, DR86, DR89, Den86, DHKW87, DO85, DJ88, Don89a, Gin88, Gri89, Hae86a, Hae86b, Isa83a, Jam88, KLP88a, KLP88b, Kaz85a, Kaz85b, Koe85, Kor83a, Kor83b, Kre83, KI82, Lam83, LQC87, MP84, Pow84, Roc89a, RS87, SC88, SS87, Str87d, Str87e, Tur87, Tut82, Ros87d, Gro88, Lea89]. Programs [Ban83, Bih88, Bis83, DC85, Gor87a, Gor87b, Ken83, Kra88b, Kuc89, LNSZ85, Lib87, MS89, Muu87, Neu86a, Neu86b, O’R88, Per82a, Per82b, dR83a, Sch82, Ste83a, Ste83b, Wil87b, YTS88, ZH88, TH83a, TH83b]. Progress [Wei85a, Wei85b]. Project [AF87, BG88b, Bro88, Eat88, Gee88b, GP88, Jen83, Mas83a, MSM88c, MSM88a, MSM88b, MSM88d, OCD^{+}87, PF84a, PF84b, SHH85, USE88d, Vas87b, Yos85, Abb87, Bra89, Dav89b, Gee88a, Get84, Ste89a, Vas87a, Wei85c]. Projections [Dro84]. Projects [Che82, Mil89]. Prolog [CLH^{+}89]. Prompts [Wil87b]. Proposal [Col80, Gil86, HSHK84, Hos84]. Proposals [O’D83b]. Proposed [HA84, Lon82b, USJ83]. Protection [FA88, Kle85]. Protocol [Che87c, Fed88a, Fed88b, GM89, NB84, Par87, SL88, Sku88, Ste88a, Sun88, YSF89]. Protocols [Duc89, FFH^{+}86, Hed89, Mur88c, OTW85, Sch89, Yam88]. Prototext [Gur88]. Prototype [CP84a, CP84b, Fos88, MN85a, MN85b, RWNA87]. Prototyping [BSY88a, BSY88b, Sha89, Son88a, Son88b, Sto88]. Pseudo [WO88]. Psfig [BD87b]. PSIBER [Hop89]. Psyche [SLM89]. Psychology [Les83]. Purdue [GM82a, GM82b]. Purdue/EE [GM82a, GM82b]. Purpose [CS86, CH83a, CH83b, Isa83a, LL83a, LL83b, MH88, MK88, Pet83, Yam88]. Putting [BPM87]. PWB [KAH83b]. PWB/II [KAH83b].

QDP [And82a, And82b]. Quaternion [Duf85]. Queries [Mey82b]. Questions [JLMM82]. Queue [Zho87]. Queuing [KM82]. Quick [And82a, And82b]. QuickPak [Bak89]. Quotas [Zwi89].

Radio [NY88]. RAID [KLB89]. Rapid [HP89a, HP89b, BC89, WS83a, WS83b]. RASH [HP89a, HP89b]. Raster [Coh87, Dan83a, Dan83b, PBT86]. Ratfor [BB83, Col84a, Mar84, Nor84, Gro82]. ratfor-T [Gro82]. Ray [Del87, Muu87, Sta87a, Sta87b]. Ray-Tracing [Muu87, Sta87a, Sta87b]. rdb [Man83a, Man83b]. RDBMS [Ton87]. RDOS [Hen83]. RDP [Par87]. Re [AG88a, AG88b, Hir83]. Re-engineering [AG88a, AG88b]. Re-inventing [Hir83]. Readers [Ivi84a, Ivi84b]. Reading [Ivi84a, Ivi84b]. Real [HRO82a, All83a, All83b, BFGK89, BV88, BL88a, BL88b, Bar88a, Bar88b, BK88a, BK88b, BSR88, Bla83b, Bla83c, Cic88, DC85, DW88, DJ88, Geo82a, Geo82b, GB89, GW86a, GW86b, HRO82b, Hor82a, Isa82a, Isa82b, Jam88, JS89a, JS89b, LLS88, LH84a, LH84b, MR88d, MR88e, NMP82, Oti88, Pan88, PB84, RL88, Ral88a, Ral88b, RT83, RGDP88, SB88a, SB88b, Sho87, STT86a, STT86b, TBS87, USE88c, Ass88e, Wat88a, RR85]. Real-Time [HRO82a, All83a, All83b, BFGK89, BV88, BL88a, BL88b, Bar88a, Bar88b, BK88a, BK88b, BSR88, Bla83b, Bla83c, Cic88, DW88, Geo82a, Geo82b, GB89, GW86a, GW86b, HRO82b, Isa82a, Isa82b, Jam88, JS89a, JS89b, LLS88, MR88d, MR88e, NMP82, Oti88, PB84, RL88, Ral88a, Ral88b, RT83, Sho87, USE88c, Ass88e, Wat88a, DJ88, LH84a, LH84b, STT86a, STT86b]. Really [Gro87, Ros88]. Realtime [Len86a, Sen87]. Receiving [Dun88]. Recipe [Fer85]. Recognition [Cog87]. Recollections [Tan86]. Record [Orr83]. Record-Oriented [Orr83]. Recovery [AF86, Gra87, KK89a]. Reduction [Kin86, Par88a]. refer [Tut83]. Referee [Col80]. Reference [Don89a, PF84a, PF84b, Tim85]. Refinement [LN88]. Reflections [Boh86, van86]. Registration [PMD88]. Registry [HSHK84]. Regular [Hum85]. REGULUS [All83a, All83b]. Related [Ral88a, Ral88b]. Relating [Dro84]. Relational [CS86, Man83a, Man83b, War83]. Relationships [Mur88d, Mur88e]. Relative [HB86]. Release [Ban83, Dav89b, Koe87b, MKB89, CGFCKT88, Dav89c, Wil89]. Reliability [MFS89]. Reliable [Ker84, Sim88, Par87]. Reloadable [Alb84]. Remote [AF86, BLMY87, BC89, Hug86, Lit87, Zad89]. Remote-File [BLMY87]. REMRT [Muu87]. Render [Pit89]. Rendering [Seq85]. Replacement [Akh87]. Replicated [Koe87b]. Report [Dun89, Ker84, Lyc83b, Rei83, SF86, Wei85a, Wei85b]. Representation [BEHW86a, BEHW86b]. reprint [TH83b]. Reproducing [dR83a]. Require [Win88]. Requirements [BSR88, Miy86, Par88a]. Rescuing [JN88]. Research [BO83a, BO83b, KL82a, Law83a, Law83b, LM88c, NSB85, Sti83, Sul87]. Resident [Bar88a, Bar88b]. Resource [BW88a, BW88b, Elz84, GM89, GW86a, GW86b, Hal87, Smi89, Zho87]. Resources [ABD^{+}89]. Response [JSW87, SM88a, SM88b]. Responsibility [Hom87]. Restore [Poe87, YKK89]. Restoring [Jaf87]. Restricted [Chr89a, Chr89b]. Results [JHRR86, STA87c, Tuo82]. Retained [SE88]. Retaining [Kra88b]. Retrieval [Lee87b, Moy83a, Moy83b, Pyn82a, Pyn82b]. Review [BWH87, Don88, Don89a, Gro88, Kin84, Lea89, Sal89a, Sal89b]. Revision [Sch86a, Yos85]. Revisited [AA85, Fer82, Koe87a, Nac86]. Revolution [AGHR89a, AGHR89b]. RFS [BLMY87, Cha87b, RFH^{+}86a, RFH^{+}86b]. RIACS [Bis87b]. Rich [Yos85]. Ring [Sal82]. Rings [Tay88]. RISC [HCE88]. Robot [Bih88]. Robot-Control [Bih88]. Rocky [DGM82]. Rogue [TA82]. Roles [Mas88]. ROM [KGTM89]. Roomful [Woz82]. Root [Win88]. Rotation [Coh87]. Rounder [Hir83]. routine [SS87]. Routines [Hop89]. Routing [Fed88a, Fed88b, Par86]. RP3 [Bry88, EGL86]. RPC [Rei87a, Rei87b, SS88, Ste88a, Tie88]. RPC/LWP [SS88]. RPCC [Rei87a, Rei87b]. RPCODE [Ste88a]. RT [Muu87, SJL^{+}87]. Rtools [HSY88]. Rule [Man87b]. Running [Bec84, KM83, LS83a, Tur87]. Runtime [HCE88, Ken83, SR85a, SR85b].

Saber [KLP88a, KLP88b]. Saber-C

[KLP88a, KLP88b]. safe

[Str88b, Str88c, Str88d, Str88e, Str88f].

SAGE [Sal88]. sailing [Til89]. Salt [Sof84, USE84a, USE84b]. Sample [MBS86a, MBS86b]. San [USE83a, Ass83a, Ass83b, USE88h, Ass88f, USE89g, USE89c]. Santa [USE87a]. Satellite [Wei84b, Wei85a, Wei85b]. Say [Jac87]. SC22 [Dun89]. Scale [ELS88a, GP88, Hun88a]. Scaling [BP86]. Scanning [Der83]. Scattered [Hal85a, Tho86]. Scheduler [Boh86, Jac86a, Jac86b, STA87c, TK88a, TK88b]. Schedulers [Pea83]. Scheduling [Lau81b, LGZ88, LLS88, MP89a, MP89b, Mok88, STA86]. Scheme [RK88]. Science [BO83a, BO83b, Pan88]. Scientific [Gri85, HS82]. scores [Fox87]. Screen [Cla87a, Cla87b, Pyn82a, Pyn82b, SV83, Kat82a, Kat82b]. SDD [Sto87]. Search [Bre83, KV85]. Searching [Les88c]. SECD [Daw82]. Second [USE85d]. secret [Les88a]. Sector [Dye82a, Dye82b]. Secure [CJL^{+}89, Kra88b, Lin88a, ST89a, ST89b, TG86]. Securing [YTS88]. Security [And88b, Car88a, Car88b, CLH^{+}89, DM88, Duf89b, FH88, HJAW88, HMM^{+}88a, HMM^{+}88b, JS87, Les88d, LAKS88, Lin88b, MR88a, MR88b, NHR84, RW86b, SB88a, SB88b, Sku88, Spe89, USE88g, USE88b, Kra83]. SEL [FT83]. Selected [SF86, USE88d]. Selecting [Ben82]. Selection [DS88, Hop87a, Hop87b]. Self [dR83a, YTS88]. Self-Reproducing [dR83a]. Self-Securing [YTS88]. Sema [Kor89a, Kor89b]. Semaphore [Kor89a, Kor89b]. Semi [PF84a, PF84b]. Semi-Large [PF84a, PF84b]. Semiconductor [SJ83]. Sendmail [AA85, Sch86c, Sch86b]. Sensitivity [SD87]. Separate [EST86, Mul87, vMM88]. September [USE88i, USE89e]. Serial [AW89, Sle87]. Series [Lin84, Wil83b, Wil83c, KAH83a]. Series/1 [Wil83b, Wil83c]. SERIX [Wil83b, Wil83c]. Servant [EVS88a, EVS88b]. Server [BD86, Dye88, Epp89, FBS89, HF89, Jus89, LA89a, LA89b, SCW89a, SCW89b, TPRZ84, Ste88a]. Server-based [Epp89]. Servers [Lit87]. Service [DeJ86a, DeJ86b, DEF^{+}88, Gee88a, Gee88b, Ker84, RGL88, SNS88]. Services [Lee89, Mid87, PG88a, PG88b, SJL^{+}87]. Session [Bel88a, Bel88b, Wil89]. Set [FN83, Kle89, Mok88, SS87]. SETOPT [Per85]. setting [Hag83]. Setuid [Bis87a]. SGID [Win88]. SGML [vH87]. Shape [Hop89]. Share [Lau81b]. Shared [ABD^{+}89, Arn86, BKT89, BC84, DF84, ELS88a, FBS89, GLDW87, Lib85a, Lib85b, MF89, Muu87, RK89, Sun89, TRY^{+}87]. Sharing [AO86, BW88a, BW88b, Bis87c, GW86a, GW86b, Ney83a, Ney83b]. Shell [CC86, Col89, KR85, Kor83a, Kor83b, MLS88, SV83, MK84a, MK84b]. Should [Bou89, Ess85, Pik88]. Shouldn’t [TA82]. si [Feu85]. Simple [MF83, Mog89, Par88b, Ros88, Rug83, WS84, YSF89]. simpler [Pre85]. Simulation [BSY88a, BSY88b, Mar84, PJ89a, PJ89b, RC83, WS84]. Simulator [JC88]. Single [Bot84, MR88c, Sau88, DJM86]. Single-level [MR88c]. single-user [DJM86]. Site [Dun88, HZ89a, HZ89b, HZ89a]. Sites [Zad89]. Sketch [Cra87a, Cra87b]. Sky [Kat84]. slave [DJM86]. Slow [Col87b, MT89]. Slow-start [MT89]. SM [OK85a, OK85b]. Small [Dun88, Kin86, OK85a, OK85b, Ree81]. Smalltalk [Cox82, Cox83]. Smalltalk-80 [Cox83]. smooth [Til89]. SMP [Sin88]. SmScript [BQd86]. Snocone [Koe85]. Softshell [Kat82a, Kat82b]. Software [AG88a, AG88b, BV88, BB87, BWS87a, BWS87b, Bih88, Bla89, Bre83, Bro85, CKM85, CKK87, Che87a, Che82, DM83, DR83b, ESS89, FHW88, Gro82, Hal85b, Han82, Har87d, HSY88, Hat82, Haw86a, Haw87a, Haw87b, Hen83, HM89, Hop87c, Hum89, Jac83, Kee88, Koe84, Koe87b, Lad88a, Lad88b, Leb87, Les83, Les88d, Lou82, Mas83a, McG85, Mei84, Mil89, Mor88e, Nac86, NM83a, NM83b, Nic89, O’D83c, PF84a, PF84b, Per87a, Per83, Rod87, Sof83, Sof84, Sch83b, Sig87, Ste84, Sti83, Sve83, TP86, Usr82, USE82b, USE83b, Ass83b, USE84b, USE88c, Ass88e, USE89a, Wam83a, Wam83b, War84, Egg89, KNN88a, KNN88b, ML88, MdM88, STV87, USE89j]. Sol [Gie83]. Solid [Sta87a, Sta87b, Mas83b, Mas83c]. solve [Ste88a]. Solving [LLM87, Tie88]. Some [Bal83, Miy88, dR83a, SF86, Spa89, TH82, Tuo83, RR85]. SOS [SGH^{+}89]. Source [BBT83, Bou89, CR89, FKT83, Fil85, Kal82, Lor88c, Tan87a, TS87, Les88a]. Sources [Wat83]. Space [Hop89, Yam88, Zwi89]. Spaces [Akk87]. Spacing [Akk88]. Speaks [JK86]. Special [Yam88]. Specific [PF84a, PF84b]. Specification [KT88a, KT88b]. Specifying [HMM^{+}88a, HMM^{+}88b, LGZ88]. Speech [Tuo83]. Speed [Ban82, Ren88a, Ren88b, Ren88c]. Spell [Bey88]. Spelling [Ros82]. Spellings [Ros82]. SPIFF [Nac88a, Nac88b]. Spline [Seq86, Sta87a, Sta87b]. Splines [Duf85, Tho86]. Spring [Ass88b, USE88e, Ass89b, Ass89f, Ass89e]. Sprite [Dou89, NO88a, NO88b, OCD^{+}87, WO88]. SPUDS [Lad88b, Lad88a]. spy [Spe89]. SQL [Mey82b]. Squads [HF89]. ST [Col88, GD89]. Stable [Ell85]. Standard [Bis88a, Bol88, Gre82a, Gre82b, HA84, Kee88, Old88a, Old88b, RW87, Til84, Bak89]. Standards [Car88d, Cra83, Hae89, Hay88, Isa83b, Lyc83b, Lyc83a, Lyc84, Mac83a, Mac83b, McC88, Mei84, Mey82a, Pet83, Sha83, Swa83, USJ83]. Stargate [Wei85c]. start [MT89]. Started [Dav89a]. State [TWM86a, TWM86b, Zwi88a, Zwi88b]. State-wide [TWM86a, TWM86b]. Stateful [AF86]. Stateless [GM89]. Static [Par88a].

Status [Bla83a, Ker84, KL82b, Lyc83b, Rei83, SHH85]. STD [Cer82, Cer83]. Steal [Spe88]. Stellix [TG88]. Sticky [Ban82]. still [RR85]. Stone [McD87]. Storage [AN88, Ell85, HP89a, HP89b, Kiv84, Les88b]. Storms [McK88a]. straight [PS89]. Straightforward [PBL86]. Strategies [Cab86, Mas83a]. Strategy [Sin88]. Stream [Raf87]. STREAMS [CE89a, CE89b, KM87, ROS87a, Rag89a, Rag89b]. STREAMS-Based [CE89a, CE89b]. Strengthening [LG88]. StrongBox [YTS88]. Structure [KBT89, TM82]. Structured [Big85, VL88a, VL88b]. Structures [BED^{+}85, Cog87, Gri89, PG87]. Stub [Rei87a, Rei87b]. Student [TH87]. Studies [ESS89]. Study [BG88b, CKM85, Car87a, Car87b, Fen87, Ham87, Kin86, KL87a, KL87b, MFS89, SB88a, SB88b, Tri87]. Style [Kir87, Mey82a, Pik83, SS87, Ros87d]. Subjects [Ral88a, Ral88b]. Subscription [Lad88a, Lad88b]. Subscription-Oriented [Lad88a, Lad88b]. Substrate [O’D87a, O’D87b]. Subsystem [AO86, KM83, MK85a]. successor [Hum87]. Suggested [Lin88b]. Suggestions [Mac83a, Mac83b]. SUID [Kra88b, Win88]. Suitable [Kuc89]. Suited [Min82a, Min82b]. Summer [LS83b, Sof83, Sof84, USE82a, USE82b, USE83b, USE84a, USE84b, USE85e, USE85a, USE86c, USE86a, USE87f, Ass88c, USE88h, Ass88f, Ass89c, USE89f, USE89b]. SUN [Bec82, Dav89a, Lau81c, AM85a, AM85b, GMW86, Kle86, LS83b, Rei87a, Rei87b, SGK^{+}85, TG86, WLS^{+}85]. SUNDEW [Gos86a, Gos86b]. SunNet [Cha85]. SunOS [Cha87b, GLDW87, GMS87, Gin88, MF89, Mor88a]. SunView [GBM87a, GBM87b]. Super [Ben82, Bis83]. Super-micro [Bis83]. Supercomputer [And88a, Miy88, TG88, Woh88]. Supercomputers [LM88a, LM88b, Ren88a, Ren88b, Ren88c, USE88i]. Supermicro [Gri85]. Superminis [Bot84]. Superuser [Chr89a, Chr89b, HCC^{+}87]. Support [BWS87a, BWS87b, Che87a, KGTM89, Lam83, LW89a, LW89b, Man87b, MH88, MS89, Req85, RT83, San83, SE88, Ste86, YTS88]. Supported [Nor88]. Supporting [DV89, Shu89]. Survey [HS82]. survive [Les88a]. SVID [Fis86c, Fis86d]. Switching [PY84a, PY84b]. Sydney [Lau81c]. Symbol [Kal82]. Symbolic [AM85a, AM85b]. Symmetric [HC88]. Symunix [ELS88a]. Synchronization [BB87, Kaz88, RL88, TBJW88]. Synchronized [GM89]. Syntactic [FJ82]. Syntax [HA84, LC87b]. Synthesis [Bee86, PMI88]. SYSTANT [Boy84]. System [HRO82a, AK88, Ama88, ACK89, Arn86, BB84, BG88a, BH86, BFGK89, Bar88a, Bar88b, BK88a, BK88b, BK85, Bec84, Ben82, BK87, BP88a, BP88b, Bis87b, Bla83b, Bla83c, BEHW86a, BEHW86b, BWH87, BERS88a, BERS88b, Bot84, Bou89, BC84, Bro89b, BK84, BP84, BLSS83, CM86a, CM86b, CCM87, CRJ87, Car82a, Car82b, CQ83a, CQ83b, CB83, CFA85, Col84c, CP84a, CP84b, CAG89a, CAG89b, CR89, Cra83, CLH^{+}89, Cyg88, Dat88, Dav89c, DRK^{+}89, Del87, DV89, DLM^{+}87, DO85, DJ88, Duf89b, DP83a, DP83b, EST86, EVS88a, EVS88b, EBFH85a, EBFH85b, Fed83, Fel84, FFH^{+}86, Fil85, FS89, FW89, Ful89b, Gie83, GBH86, Gle89, GPF^{+}86a, GPF^{+}86b, Gos86a, Gos86b, GP88, GD89, GZ84a, GZ84b, HZ89a, HZ89b, HK83, HQZ^{+}87, Har87a, Har87b, Har88a, Har88b, HRO82b, HLW84, Hef82, HA84]. System [HP89a, HP89b, Hil89, HH86, HK86a, HK86b, HH88, Hom87, Hop87c, Hum88a, Ivi84a, Ivi84b, JS89a, JS89b, JHRR86, JC88, Jon88, KTS^{+}86a, KTS^{+}86b, KNI88, KGTM89, Kar83, Ker88, KK89a, KM82, Kiv84, Kle86, Koe87b, KM87, KAH83b, KK89b, KK89c, Kra88a, Lad88a, Lad88b, Lam83, LM88a, LM88b, Lan84, LEG88, LLM87, Lee89, LB89a, LB89b, Les88c, Lev83, LH84a, LH84b, LFN^{+}89a, LFN^{+}89b, LSC^{+}88, MC85a, MC85b, Man83a, Man83b, MF83, MH88, McG85, MS89, McK82b, Mer82, Mey85, Mil84a, MFS89, MF89, Min82a, Min82b, MMTW88, Moy83a, Moy83b, NM83a, NM83b, Ney83a, Ney83b, Ohk84, OMI86, Orr83, Par88b, PMD88, Pat83, PF84a, PF84b, Pea88, Pik89, PL89, PG87, Pow84, PR85, PY84a, PY84b, RRSZ89, Ral88a, Ral88b, RC83, RLML86a, RLML86b, Rod86, RKPP88, Sal88, Sam87]. System [SW82, SJL^{+}87, Sau88, Sax85b, Sax85c, SR85a, SR85b, SLM89, Seq85, Sha89, SGH^{+}89, Sho87, Shu89, SM84, ST89a, ST89b, Spe89, Sta87a, Sta87b, SD87, Sto87, STT86a, STT86b, Sve83, Tag83, TM82, TH87, Tes82, Tes86a, Tes86b, Tho85a, Tho85b, TBJW88, TR84, TWM86a, TWM86b, Tur87, USE87d, Uhl87, Uit87a, Uit87b, UJ83, Van88, WLS^{+}85, Wam83a, Wam83b, War83, Wat88c, Weh83, WRM^{+}89, Woh88, YKK89, Yos85, ZDS85, Zuc83b, Zuc83c, ZP89, vM88, Hun88b, Hun88c, PG88c, SS88, WJ82, AO86, Bal83, CGFCKT88, Cyg88, Eyk86, Goo84, Gou86, Guf83, Hec88, How88, Hug86, Kaz88, Len86b, LK82, LZ82, Mil84b, Phi84, Pos88, RKH86, RGL88, RGDP88, San83, Sch83c, Ter87, Tim85, Ton87, TW84, Ups82, Wei84a, WO88, Wil89]. System/370 [Eyk86]. System/Three [Lev83]. Systematic [SB88a, SB88b]. Systems [All83c, AF87, ACK89, AQ84, BSR88, Bob88, BFS89, Bru88, CJL^{+}89, Car88d, Che82, CS86, Cla89, CDT89a, CDT89b, CN88, Dan83a, Dan83b, DW88, Don88, Duf89a, EGL86, ELS88b, Get86, Gol88, GB89, Hal85b, Hun88a, Isa82a, Isa82b, JN88, Kaz85a, Kaz85b, Koe87a, KL82a, KM83, LP89, LW89a, LW89b, LZ82, MLRC88, Man87b, Mar84, Mas83b, Mas83c, McG85, MVB84, Mid87, Mil86, Nac86, O’D87c, Oti88, Pea83, Pik88, Poe87, RL88, RWNA87, RW86b, RW86a, Rob89, Roc89a, Roc89b, Roc89c, RAA^{+}88, SS88, Sen87, Smi89, SC88, SNS88, STA86, TGB^{+}89, USE88c, Ass88a, Ass88b, Ass88c, Ass88d, Ass88e, USE88f, Ass89a, Ass89b, Ass89c, Ass89d, USE89d, Ass89e, USE89e, USE89k, War82, WS83a, WS83b, Wat88a, WH83, Woz82, YT83, UE88, Hag83, TH83a, TH83b, SSNU87].

T [Gro82]. table [MD87]. Tables [Hae83]. Tabstar [Wai82]. Tactical [SSNU87]. Tactics [Mas83a]. Take [Mor88b, Mor88c]. Taking [McD87]. Tales [Tay86]. Talking [Tuo83]. Talks [Bey87]. Tape [Jaf87, Kri84a, Kri84b]. Targeting [Gar86]. Task [Mok88, Oti88, PJ89a, PJ89b, RL88, Sho87]. Tasking [BP84]. Tasks [PL89, Shu89]. TCB [SW88]. TCP [FFH^{+}86, Fer85, JLSG84, KM86, MT89, NB84, WG84]. TCP/IP [FFH^{+}86, Fer85, JLSG84, NB84, WG84]. Teaching [BD87a, Tut82]. Technical [FKN85a, FKN85b, Kin84, Til88, USE88d]. technique [Cor82]. Techniques [Mok88, VM84]. Technologies [Mas88]. Technology [Che82, Les83]. Telephone [Red85]. Tell [O’D87c, LK82]. TEMPO [GZ84a, GZ84b]. Terminal [FL82, Fos83a, Fos83b, Lee89, Mac83a, Mac83b, MR88c, Pik84a, Pik84b, Col88]. Terminal-Independent [Mac83a, Mac83b]. Terminals [Ney83a, Ney83b]. Terminfo [Hor82b]. Test [DR83b, Hor82a]. Testing [CM89a, CM89b, CLH^{+}89, Dix82, Wat88b]. Texas [USE85c]. Text [Gur88, KV89, Lee87b, Pik84a, Pik84b, Wal86a, Wal86b, Zem83]. Text-Oriented [Pik84a, Pik84b]. Them [BED^{+}85, Ess85]. There [TA82, O’D83a]. Third [USE86d]. Thoughts [Hal85a, Har85a, Har85b, Tho86]. Threads [BW88a, CHS89, TRG^{+}87]. Three [Lev83]. Tightly [Inm85]. Tilde [CD85].

Time [HRO82a, All83a, All83b, AO86, BFGK89, BV88, BL88a, BL88b, Bar88a, Bar88b, BK88a, BK88b, BSR88, Bla83b, Bla83c, Cic88, DW88, Geo82a, Geo82b, GB89, GW86a, GW86b, GZ84a, GZ84b, HRO82b, Hor82a, Hum89, Isa82a, Isa82b, Jam88, JS89a, JS89b, JSW87, JHRR86, Kat82c, LGZ88, LLS88, MR88d, MR88e, Ney83a, Ney83b, NMP82, Oti88, PB84, RL88, Ral88a, Ral88b, RT83, RGDP88, Sho87, SM88a, SM88b, USE88c, Ass88e, Wat88a, DJ88, LH84a, LH84b, STT86a, STT86b]. Time-sharing [Ney83a, Ney83b]. Timesharing [MF83]. Tinfoil [ACF^{+}86]. Together [BPM87]. tolerant [LA89a, LA89b]. Tool [And82a, And82b, BSY88a, BSY88b, Cha87a, CM89a, CM89b, Chr89a, Chr89b, Haw88a, Haw88b, JA88, Kor89a, Kor89b, LOR88a, LOR88b, Neu86a, Neu86b, Pax84, RWNA87, RSV85, TvK83, Tuo82, WS83a, WS83b, MdM88]. Tool-based [LOR88a, LOR88b]. Toolkit [MA88, RW87, Roc89b, Roc89c, SA88, Wal87, WS84, ML88, PHS^{+}88a, PHS^{+}88b, PBT86, RGDP88]. Tools [And88a, BB83, Bre83, Che87a, DM83, DR83b, FS89, Gro82, Han82, HSY88, Hat82, Hau83, Haw85, Hen83, Jac83, Kee88, KL87a, KL87b, Lor88c, Lou82, Mei84, O’D83c, Per83, Sch86a, Sch83a, Sch83b, SREC88, Sve83, TP86, Usr82, USE82b, USE83b, Ass83b, USE84b, UJ83, Yos85]. Top [Epp89]. Top-level [Epp89]. TOPS [Bee84a, Bee84b, Hat82]. TOPS-20 [Hat82]. TOPS-20-like [Bee84a, Bee84b]. Toronto [Sof83, USE83b, KL82a]. Tour [Col89, Par86, See89]. TOWER [CE89a, CE89b]. TRACE [Cla88]. TRACE/UNIX [Cla88]. Tracer [Rod86]. Tracing [Del87, Muu87, ZDS85, Sta87a, Sta87b]. Tracking [Har87a, Har87b]. Trade [Kri84a, Kri84b]. Trade-Offs [Kri84a, Kri84b]. Trading [LSC^{+}88]. Training [Hos83, Hos84]. Transaction [Ado89, ACK89, CCF89, Cla89, DHKW87, Epp89, Ful89a, HSYY89, LLS88, LA89a, LA89b, MH88, Sun88, Tal89a, Tal89b, USE89h]. Transaction-Based [DHKW87]. Transactions [Duc89]. Transcending [PL89]. Transition [MW84]. Transitioning [Nor88]. Translation [Gil86, LC87b, TBJW88]. Transmission [NSB85]. Transparent [DF84, Pik88, URK85, Hug86]. Transputer [LM88a, LM88b, Sch89]. Transputer-based [LM88a, LM88b]. Traps [Don89b]. Tree [Yos85]. Trees [CD85, Gle89, LC87b]. Trials [O’D83b]. Tribulations [O’D83b]. Tricks [Dav89b]. TROFF [Akk87, KM84b, KM84a, Akk88, Fox87, LNSZ85]. Trojan [LG88]. TRS [DP83a, DP83b]. TRS-XENIX [DP83a, DP83b]. True [Jac87]. Trusted [Fos88, SM89, TGB^{+}89]. Tty [Bel88a, Bel88b, Lau81a, CE89a, CE89b, KM87]. Tumult [JS89b, JS89a]. Tumult-64 [JS89b, JS89a]. tunable [Jac86a, Jac86b]. Tuning [Jac87, Sam87, Woh88]. Tunis [EBFH85b, FH88, HMP83a, HMP83b, EBFH85a]. Tunnels [UTC84]. Tuples [Kat82c]. Turing [CH83a, CH83b]. Turn [Til87a, Til87b]. Turning [Lit87]. Tutorial [Til83, Wil88]. Tutorials [Kle87]. Twice [Tay88]. Two [Bih88, GS87, Gol88, Les87, TH83a, TH83b]. TX [Til88, USE85b, USE88j, USE89e]. Type [Mur88d, Mur88e, Str88b, Str88c, Str88d, Str88e, Str88f]. Type-safe [Str88b, Str88c, Str88d, Str88e, Str88f]. Types [Kle86, Str88a, Str89c]. Typesetter [KM84b, KM84a]. Typesetter-Independent [KM84b, KM84a].

UBOAT [Kle87]. UCSD [LM83a, LM83b, Mer82]. UFOS [BFGK89]. UIMS [SE88]. Ultracomputer [EGL86].

ULTRIX [CHS89, Sin88, HC88, NY88]. Uncle [KL87a, KL87b]. Undergraduate [Mor88e]. UNICOS [Eat88, Woh88, Fou88, Rit88]. UNICUBIX [Seq86]. Unification [Dav88]. Unified [Roc89a]. UNIFLEX [Pod82]. Uniform [Rug83]. UniForum [USE84c]. Unison [RGDP88]. Units [LJ84]. Universal [Gur88]. UNIverse [Tan87b]. University [KL82a, Sto85, Cot87]. UNIX [CKM85, EST86, JLSG84, Kra83, HRO82a, ABB^{+}86a, ABB^{+}86b, Ado89, ASS85, Alb84, All83a, All83b, All87, AW89, Ama88, And88a, And88b, ACK89, AQ84, ACF^{+}86, AGHR89a, AGHR89b, Arn86, AO86, BB84, BG88a, BH86, BFGK89, Ban82, BP86, BW89, BM87, Bar88a, Bar88b, Bar83, BW88a, BW88b, BK85, Bec84, Bee86, BED^{+}85, Ber85, BP88a, BP88b, Bey87, BE89, Bis88b, Biv87, Bla83b, Bla83c, Ble83, BEHW86a, BEHW86b, Boh86, BFS89, BQd86, BNB87, Bot84, Boy84, Bro89b, Bry83a, Bry83b, BK84, BP84, Cap82a, Cap82b, Car88a, Car88b, CJL^{+}89, Car82a, Car82b, Cer82, Cer83, CMM88, CQ83a, CQ83b, CB83, Che87b, CCD^{+}87, Cla88, CS86, CCF89, Cle83a, Cle83b, Col83, CFA85, Col84b, Col84c, Col87a, Col89, CD85, CN88, CAG89a, CAG89b, CH83a, CH83b, Cra83, Dan83a, Dan83b]. UNIX [Das88, Dat88, DM88, Dav89c, Daw82, Den83, DNQ^{+}83, Dix82, DN87, DJM86, Duf89a, Duf89b, Duf82, EGL86, ELS88b, Ela83, Erl88, Eyk86, Eyk88, Fed83, Fed88a, Fed88b, Fel84, FFH^{+}86, FA88, Fer85, Fil85, FKN85a, FKN85b, Fis86a, Fis86b, FN83, FT83, Fun87, FH88, Ger82, Get86, GBH86, Gom85, GP88, Gre82a, Gre82b, GZ84a, GZ84b, Hae89, Hag83, HK83, Ham87, HS82, HQZ^{+}87, Har83, Haw85, HL85a, HL85b, Haw86a, Haw87a, HRO82b, HLW84, HCC^{+}87, HJAW88, HA84, HP89a, HP89b, HCN85, HH86, HP89c, HMP83a, HMP83b, Hoo83, Hor82a, Hos83, Hos84, Hun88b, Hun88c, IvW87, Isl83b, Jac84a, Jac84b, Jac87, JTUB85, Joh87, JH86, JN88, Jun85a, Jun85b, JK86, KNI88, Kar83, Kat83, Kat82a, Kat82b, Kep85, Kin83, KK89a, Kiv84, Kle86, Kod82a, Kod82b, KM87, KAH83a]. UNIX [Kor89a, Kor89b, Kra88a, Kra88b, KM83, Kri84a, Kri84b, KI82, LJ84, Lam83, LM88a, LM88b, Lan84, LP89, LEG88, Law83a, Law83b, LLM87, Lee89, Len87, Les88a, Les88c, Lib87, Lin84, LH84a, LH84b, LSC^{+}88, LS83a, Lyc85, MLRC88, MC85a, MC85b, MS88, MF83, MK85a, MH88, Mas87, Mas88, McC88, McD84a, McD84b, MS89, McK82a, MK88, McL83, McL84, Mer82, Mil84a, Mil86, Min82a, Min82b, MMTW88, Miy86, Mor88d, Mul87, Mur83, NM83a, NM83b, NMS83, Ney83a, Ney83b, NLR84a, NLR84b, NMP82, O’B82, O’D83a, O’D83c, O’D83b, O’D87d, Pan88, Par88c, PS82, PMD88, Pat83, PF84a, PF84b, Pea80, Pea83, Per87a, Per82a, Per82b, PJ89a, PJ89b, Pic83, Pik82, Pik83, Pla89, Poe87, PK88, PB84, Pos88, Pow84, Poz83, PY84a, PY84b, RWNA87, RC83, RP84, Ree82a]. UNIX [Ree82b, Req85, Ric85, RW86b, RW86a, Rob84a, Rob84b, Rod86, RKPP88, RT83, Rug83, Sam87, SW82, San89, Sax85a, Sax85b, Sax85c, SW88, SR85a, SR85b, SH85, SSWW83, SJ83, Sku88, ST89a, ST89b, Smi83, SM88a, SM88b, Smi89, SM89, SD87, Ste88b, Sti83, STA86, STA87c, STT86a, STT86b, Tag83, Tal89a, Tal89b, TH83a, TH83b, TvK83, Tan87a, Tan84, Tan87b, TG88, TY82, Tho85a, Tho85b, TS87, Til84, Til87a, Til87b, TBS87, Tra85, Tre88, TWM86a, TWM86b, Tuo82, Tuo83, Tur87, USE88b, USE88i, Ass89e, USE89h, Uhl87, UTC84, URK85, van86, VB83, Wal82a, Wal82b, Wal82c, Wam83a, Wam83b, Wat83, Weh83, WJ82, Wil83b, Wil83c, Wil82a, Wil82b, WH83, WB85a, WB85b, Woz82, Yao83, YT83, YSF89, Yos85, ZDS85, Zim85, Zuc83b, Zuc83a, Zuc83c, vMM88]. UNIX [vM88, AK88, AN88, Bil86, Car88c, Con88, Gar86, Gen86, Har88a, Haw86b, Haw87b, Hay88, HK86a, HK86b, Hum88a, Joi87, Kle85, Kle87, Kno87, Kol86, Lau81c, Lit87, Mog89, PR85, Ree81, Rit87, Sha89, Spe89, SCC86a, SCC86b, TRG^{+}87, TRY^{+}87]. UNIX-Based [McD84a, McD84b, Per87a, PY84a, PY84b, WH83, HK83, KK89a, SR85a, SR85b, Sti83, HK86b, Mog89, HK86a]. UNIX-Like [MLRC88, LM88a, LM88b]. UNIX/Prime [WJ82]. Unknown [NS88]. unofficially [Tut83]. Unorthodox [Mor88e]. Untrusted [NS88]. Upas [Pre85]. Update [Car88d, Hae89, Hen83, Lad88a, Lad88b, McC88, Mei84]. Uptime [Hal87]. USA [Sof84, USE84c, USE85c, USE85e, USE86c, USE86e, USE87f, USE87g, USE88g, USE88h, USE88j, USE89f, USE89g]. Usability [ESS89]. Usage [HS82, Kor89a, Kor89b]. Use [DP83a, DP83b, Ess85, Gra88c, Gra88a, Gra88b, Hum89, Kal82, Kuc89, Til82, WK83a, WK83b]. used [LFN^{+}89a, LFN^{+}89b]. USENET [Fai86, GRS88a, GRS88b, Kat84, Hor83, HSHK84]. USENIX [Sof83, Sof84, SHH85, Til88, Usr82, Bak89]. USENIX/Software [USE82b, USE83b, Ass83b, USE84b]. User [Abb87, BLSS83, CCM87, Cra87a, Cra87b, DM88, Den83, Gan86a, Gan86b, Gol88, GBM87a, GBM87b, Har85a, Har85b, Hoo83, IvW87, Jac84a, Jac84b, Jac86a, Jac86b, Lib85a, Lib85b, LS83a, Miy86, PMD88, Per83, RW86a, Rug83, SA86a, SA86b, SA88, Ass89e, War84, WO88, Yam88, YT83, Biv87, DJM86]. User-Interface [GBM87a, GBM87b, Per83]. User-Interfaces [SA88]. User-Level [Jac84a, Jac84b, Lib85a, Lib85b, WO88]. User-Mode [War84]. User-Space [Yam88]. User-tunable [Jac86a, Jac86b]. usernames [PS89]. Users [Hag83, Les83, Nor88, Smi87a, Usr82]. Using [And88a, Car88c, CKK87, CM83, CLH^{+}89, DRK^{+}89, FT83, FH88, Gri85, HM89, Hop87c, Leb87, Les88c, MA88, Nic89, Nov83, Per87b, Pow84, Ros87d, Sax85b, Sax85c, Sch89, SE88, Wal87, AG88a, AG88b, Jac83, Par86]. USNA [MSM88c, MSM88a, MSM88b, MSM88d]. UT [USE84a, USE84b]. Utah [Sof84, PBT86]. UTek [McI87]. Utilities [MFS89, Bak89]. Utilization [Kle89]. UTMOST [Ney83a, Ney83b]. UTS [Wal82a, Wal82b, Wal82c]. UUCP [DLKE84, HSHK84, Ker84, KSH83, KSH84, NHR84, Per87b, RW86b, RW86a, SHH85]. UUCP/Usenet [HSHK84]. Uwm [Gan86b, Gan86a]. UX [KGTM89, Kel89].

V [CGFCKT88, Dav89c, KM87, Wil89, Arn86, BFGK89, Bal83, CQ83b, FW89, Goo84, Guf83, Lan84, Len86b, Mil84b, OMI86, San83, Tim85, Pik83]. V.3 [Ste88b]. V/MLS [FW89]. Vacation [LS83b]. Vader [Ree82a, Ree82b]. Validation [FN83]. Variable [ABD^{+}89]. Variables [Lib85a, Lib85b]. Variant [Ros82]. VAX [Cap82a, Cap82b, DJM86, GM82a, GM82b, Kri84a, Kri84b, LS83b, Tor83a, Tor83b, Tuo82, KM83]. VAX/VMS [Cap82a, Cap82b, Tor83a, Tor83b]. VAX11 [See83]. VCHK [Bry83a, Bry83b]. VDM [CKK87]. Vehicle [MR88d, MR88e]. Vendor [War82]. Venture [Wil83a]. Version [FKN85a, FKN85b, KL82b, RW87, Som88, Sch83c, Wei84a]. versus [Sch83a, Tri87]. Very [AF87, Bis83, Wat88c]. vfork [Kar83]. vi [Sal89b]. via [Man87a, Man87b, Wei84b, Wei85a, Wei85b]. Vice [Koe87b]. Video [HK83, JHRR86]. Videotape [Ger82]. View [Ama88, O’D87d]. Viewing [BPM87]. Viral [Duf89b]. Virology [McI89]. Virtual [BWP85, Che87b, Cla88, Fos83a, Fos83b, GB83, Gen86, GMS87, JC89a, JC89b, KAH83a, MK85a, Mil84b, Mor88a, Nef83, Roc89b, Roc89c, Sve83, SCC86a, SCC86b, Tri89, UJ83, Van87, Lau81a, Ups82].

Virtual-Memory [SCC86a, SCC86b]. Virtues [AGHR89a, AGHR89b]. Viruses [Duf89a, LG88]. Visual [DR89, HMM^{+}88a, HMM^{+}88b]. Visualization [Sul87]. Visualized [WHM89]. Visualizing [LR89]. VLIW [CCD^{+}87]. VLSI [BK85, CB83]. VM [TH86]. VM/CMS [TH86]. VMS [Cap82a, Cap82b, JLSG84, Tor83a, Tor83b, Woo83a]. VMUNIX [Fer82]. Vnodes [Kle86]. VORX [Kat89]. VOS [DM83, DR83b, Gro84, Pax84]. Vous [WB85a, WB85b]. vs [Bih88, Les83, Les87, WM82].

wa [JK86]. Walker [FKV89]. want [O’D83a]. Wanted [Bal83, LK82]. Wars [Hum88b]. Was [TA82, LK82]. Washington [USE84c, USE87g, Ass88e]. watch [Ing87]. Watchdogs [BP88a, BP88b]. Watermark [LB89a, LB89b]. Watermark-based [LB89a, LB89b]. WEBDMS [USE89k]. Weight [ABD^{+}89]. Well [Mur88d, Mur88e]. Well-Behaved [Mur88d, Mur88e]. West [USJ83]. WG15 [Dun89]. Wheel [Hir83, Spe88]. Where [McK83a, McK83b, O’D83a, TA82]. Whither [Kol85]. Who [Red85]. wide [TWM86a, TWM86b]. Widget [MA88]. Widgets [SA88]. Wild [Alt87]. Will [BED^{+}85]. Window [AM85a, AM85b, BNB87, CJL^{+}89, Ful89b, Gan88a, Gan88b, Get86, Gos86a, Gos86b, Jac84a, Jac84b, Lew86a, Lew86b, MF83, McG85, Neu86a, Neu86b, O’D87c, Pik88, Pik89, RHH85a, RHH85b, Rob89, Roc89a, Roc89b, Roc89c, Tes82, TW84, Tra85, Uhl87, Rob87]. Window-Based [AM85a, AM85b, Neu86a, Neu86b]. Windowed [McG86a, McG86b]. Windowing [Rei89, ST89a, ST89b]. Windows [DR86, KTS^{+}86a, Col84b, Eva83, Gan86a, Gan86b, HL85a, HL85b, KTS^{+}86b, MR88c, Opp89a, Opp89b, TF89a, TF89b].

WINDX [Col84b]. Winter [USE83a, Ass83a, Ass83b, USE85c, USE85b, USE86e, USE86b, USE87g, USE88d, Ass88d, Ass89d, USE89g, USE89c]. Wire [Lan86]. Within [MMTW88]. Without [McK88b, Pan88, Zwi89, HCC^{+}87, Mil88]. Word [Les88c]. Words [Ros82, Tri89]. work [RR85]. Workbench [Ivi84a, Ivi84b, Smi83, Tor83a, Tor83b]. Workload [Cab86]. Works [Kee88, Lau81b]. Workshop [USE85d, USE86d, USE87a, USE87c, USE87d, USE88c, Ass88e, USE88f, USE88g, USE88b, USE88i, USE89d, USE89e, USE89a, USE89h, USE89k, USE89i, USE89j]. Workspace [SE88]. Workstation [Bec82, Big85, Hay88, KTS^{+}86a, KTS^{+}86b, Leb87, LOR88a, LOR88b, LS83b, McD84a, McD84b, PJ89a, PJ89b, Tho85c]. Workstation-Based [KTS^{+}86a, KTS^{+}86b]. Workstations [AM85a, AM85b, CGFCKT88, DR86, GW86a, GW86b, Lit87, NS88, Sha83, Tre88, Van88]. Worksteps [Ral88a, Ral88b]. World [Das88, O’D83c, O’D87d, Ros88, TBS87]. Worm [See89]. Worth [Dun88]. Write [Bis87a, LEG88, Len87, MA88, NO88a, NO88b, Ros88, SM88a, SM88b]. Write-Once [LEG88]. Writer [Smi83]. Writing [Kir87, MVB84, YT83]. Written [Feu84, HMP83a, HMP83b]. WYSISYG [Wal87]. WYSIWYG [MD87].

X [Eng88, Ful89b, Gan86a, Gan86b, MA88, Par88a, RW87, SA88, TF89a, TF89b]. X-MP [Eng88, Par88a]. X.25 [HOG88, Mil84a]. X.400 [DN87]. X11 [LR89, Opp89a, Opp89b, Ros88, Sch88a]. X11/NeWS [Opp89a, Opp89b, Sch88a]. XENIX [DP83a, DP83b, MVB84, RSW83, VM84]. XINU [BWP85]. Xlib [Don89a]. XNS [OTW85]. XVT [Roc89b, Roc89c].

YABS [Sim89]. yacc [CM83, Joh88a, Joh88b]. Yackos [HF89]. Years [Lyc85]. yonder [Rob87]. You’re [Red85]. Yunikkusu [JK86].

Z [Dan83a, Dan83b]. Z80 [DP83a, DP83b]. Zephyr [DEF^{+}88].

References

Allman:1985:SR

[AA85] Eric Allman and Miriam Amos. Sendmail revisited. In USENIX Association [USE85e], pages 547–555. LCCN QA76.8.U65 U8 1985.

Allman:1985:EHC

[AB85] Eric Allman and David Been. An exception handler for C. In USENIX Association [USE85e], pages 25–45. LCCN QA76.8.U65 U8 1985.

Accetta:1986:MAN

[ABB^{+}86a] Mike Accetta, Robert Baron, William Bolosky, David Golub, Richard Rashid, Avadis Tevanian, and Michael Young. Mach: A new kernel foundation for UNIX development. In USENIX Association [USE86c], pages 93–112.

Accetta:1986:MNK

[ABB^{+}86b] Mike Accetta, Robert Baron, William Bolosky, David Golub, Richard Rashid, Avadis Tevanian, and Michael Young. Mach: A new kernel foundation for UNIX development. In USENIX [USE86a], pages 93– 112.

Abbate:1987:UAA

[Abb87]  Janet Abbate.  User ac-

count administration at Project Athena. In USENIX Association [USE87d], pages 28– ?? ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

Aral:1989:VWP

[ABD^{+}89] Ziya Aral, James Bloom, Thomas Doeppner, Ilya Gertner, Alan Langerman, and Greg Schaffer. Variable weight processes with flexible shared resources. In USENIX Association [USE89g], pages 405–412.

Appelbe:1986:PUN

[ACF^{+}86] W. Appelbe, D. Coleman, A. Fratkin, J. Hutchison, and W. J. Savitch. Porting UNIX to a network of diskless micros or UNIX on tinfoil. In USENIX Association [USE86c], pages 486–498.

Andrade:1989:BTP

[ACK89] Juan M. Andrade, Mark T.

Carges, and Kurt R. Kovach. Building a transaction processing system on UNIX systems. In USENIX Association [USE89h], pages 13–22. ISBN ???? LCCN ????

Adamson:1983:DBB

[Ada83a] A. Adamson. Description of the ‘bed’ binary editor. EUUG Newsletter, 3(4):12–23, Winter 1983. CODEN EONLE8. ISSN 1011-4211.

Adamson:1983:DBE

[Ada83b] A. Adamson. Description of the ‘bed’ binary editor. EUUG Newsletter, 3(4):12–23, Winter 1983. CODEN EONLE8. ISSN 1011-4211.

Adolph:1989:HAU

[Ado89] W. Stephen Adolph. High availability in a UNIX transaction processing environment. In USENIX Association [USE89h], pages 23–32. ISBN ???? LCCN ????

Atlas:1986:ERS

[AF86] Alan Atlas and Perry Flinn. Error recovery in a stateful remote filesystem. In USENIX Association [USE86c], pages 355–366.

Anderson:1987:DPD

[AF87] David P. Anderson and Domenico Ferrari. The DASH project: Design issues for very large distributed systems. ;login: the USENIX Association newsletter, 12(2):13–14, March/April 1987. CODEN LOGNEM. ISSN 1044-6397.

Anderson:1988:SRE

[AG88a] Bruce Anderson and Sanjiv Gossain. Software reengineering using C++. In USENIX Association [USE88e], pages 213–218. ISBN ???? LCCN ????

Anderson:1988:SRU

[AG88b] Bruce Anderson and Sanjiv Gossain. Software reengineering using C++. In USENIX Association [USE88e], pages 213–218. ISBN ???? LCCN ????

Armand:1989:RDU

[AGHR89a] Francois Armand, Michel Gien, Frederic Herrmann, and Marc Rozier. Revolution 89 or “distributing UNIX brings it back to its original virtues”. In USE-NIX Association [USE89d], pages 153–174. ISBN ???? LCCN ????

Armand:1989:RUB

[AGHR89b] Francois Armand, Michel Gien, Frederic Herrmann, and Marc Rozier. Revolution 89 or “distributing UNIX brings it back to its original virtues”. In USE-NIX Association [USE89d], pages 153–174. ISBN ???? LCCN ????

Alonso:1988:PMI

[AK88] Rafael Alonso and Kriton Kyrimis. A process migration implementation for a Unix system. In USENIX Association [USE88j], pages 365–372. ISBN ???? LCCN ????

Akhtar:1987:RBM

[Akh87] Pervaze Akhtar. A replacement for Berkeley memory management. In USENIX Association [USE87f], pages 69–79.

Akkerhuis:1987:FST

[Akk87] Jaap Akkerhuis. Fun with spaces in TROFF. EUUG Newsletter, 7(4):63–67, Winter 1987. CODEN EONLE8. ISSN 1011-4211.

Akkerhuis:1988:ST

[Akk88] Jaap Akkerhuis. Spacing out on Troff. ;login: the USENIX Association newsletter, 13(1): 9–12, January/February 1988. CODEN LOGNEM. ISSN 1044-6397.

Alborough:1984:RUD

[Alb84] Thomas Alborough. Reloadable UNIX device drivers. In USENIX Association [USE84c], pages 135–144. ISBN none. LCCN QA76.8.U65 U55 1984.

Alecci:1987:GOG

[Ale87a] Donald V. Alecci. Generic object-oriented 3-Dimensional graphics environment with editing capabilities. In USE-NIX Association [USE87c], pages 102–?? ISBN ???? LCCN ???? Abstract only.

Alecci:1987:GOO

[Ale87b] Donald V. Alecci. Generic object-oriented 3-Dimensional graphics environment with editing capabilities. In USE-NIX Association [USE87c], pages 102–?? ISBN ???? LCCN ???? Abstract only.

Allen:1983:RRT

[All83a] Bill Allen. REGULUS, a realtime UNIX lookalike. In USE-NIX [USE83a], pages 268–?? Abstract only.

Allen:1983:RRU

[All83b] Bill Allen. REGULUS, a realtime UNIX lookalike. In Association [Ass83a], pages 268–?? Abstract only.

Allman:1983:MSA

[All83c] Eric Allman. Mail systems and addressing in 4.2bsd. In Association [Ass83a], pages 53–62.

Allman:1987:UDF

[All87]  Eric Allman.  UNIX: The

data forms. In USENIX Association [USE87g], pages 9– 15. LCCN QA 76.76 O63 U84 1987. Cover title: Conference proceedings. Spine title: Winter 1987 USENIX Association Conference proceedings, Washington, DC.

Alter:1987:EMG

[Alt87] Diane Alter. Electronic mail gone wild. In USENIX Association [USE87d], pages 24– 25. ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

Adams:1985:DAW

[AM85a] Evan Adams and Steven S.

Muchnick. Dbxtool — A window-based symbolic debugger for Sun workstations. In USENIX Association [USE85e], pages 213–227. LCCN QA76.8.U65 U8 1985.

Adams:1985:DWB

[AM85b] Evan Adams and Steven S.

Muchnick. Dbxtool — A window-based symbolic debugger for Sun workstations. In USENIX [USE85a], pages 213– 227.

Amaral:1988:OOV

[Ama88] Paolo Amaral. OFS — an optical view of a UNIX file system. In USENIX Association [USE88e], pages 203–211. ISBN ???? LCCN ????

Arnold:1988:AUB

[AN88] Edward R. Arnold and Marc E. Nelson. Automatic Unix backup in a mass-storage environment. In USENIX Association [USE88j], pages 131–136. ISBN ???? LCCN ????

Anderson:1982:QAQ

[And82a] Kenneth R. Anderson. QDP — A quick plotting tool. In Usr Group [Usr82], pages 323–341.

Anderson:1982:QQP

[And82b] Kenneth R. Anderson. QDP — A quick plotting tool. In USE-NIX [USE82b], pages 323–341.

Anderson:1988:DSG

[And88a] H. Stephen Anderson. Distributed supercomputer graphics using UNIX tools. In USENIX Association [USE88i], pages 25–32. ISBN ???? LCCN ????

Anderson:1988:UPS

[And88b] Lon E. Anderson. UNIX password security. In USENIX Association [USE88g], pages 5– 7. LCCN QA76.8.U65 U55 1988(1)-1990(2)//.

Auerbach:1986:USC

[AO86] Karl Auerbach and Robin O’Neill. A UNIX subsystem on the Cray Time Sharing System (CTSS). In USENIX Association [USE86e], pages 211–218.

Appelbe:1984:AOC

[AQ84] Bill Appelbe and Bob Querido.

An adaptable object code optimizer for UNIX systems. In Software Tools Users Group [Sof84], pages 111–118. LCCN QA76.8.U65 U83 1984.

Arnold:1986:SLU

[Arn86] James Q. Arnold. Shared libraries on UNIX system V. In USENIX Association [USE86c], pages 395–404.

USENIX:1983:UCPb

[Ass83a] USENIX Association, editor.

USENIX Conference Proceedings, Winter, 1983. San Diego, CA. USENIX, Berkeley, CA, USA, Winter 1983.

USENIX:1983:USTb

[Ass83b] USENIX Association, editor.

USENIX/Software Tools Conference Proceedings, Winter, 1983. San Diego, CA. Software Tools User Group, Berkeley, CA, USA, Winter 1983.

Aitken:1985:DID

[ASS85] Gary Aitken, Christine Scott, and Kenneth Scott. DIBOLIX — an implementation of DI-BOL under UNIX. In USENIX Association [USE85c], pages 30–33.

USENIX:1988:CSF

[Ass88a] USENIX Association, editor. Computing Systems, Fall, 1988. USENIX, Berkeley, CA, USA, Fall 1988.

USENIX:1988:CSSa

[Ass88b] USENIX Association, editor. Computing Systems, Spring, 1988. USENIX, Berkeley, CA, USA, Spring 1988.

USENIX:1988:CSSb

[Ass88c] USENIX Association, editor. Computing Systems, Summer, 1988. USENIX, Berkeley, CA, USA, Summer 1988.

USENIX:1988:CSW

[Ass88d] USENIX Association, editor. Computing Systems, Winter, 1988. USENIX, Berkeley, CA, USA, Winter 1988.

USENIX:1988:FRT

[Ass88e] USENIX Association, editor. Fifth Real-Time Software and Operating Systems Workshop Proceedings, May 12–13, 1988. Washington, DC. USENIX, Berkeley, CA, USA, May 12– 13, 1988.

USENIX:1988:UCPa

[Ass88f] USENIX Association, editor. USENIX Conference Proceedings, Summer, 1988. San Francisco. USENIX, Berkeley, CA, USA, Summer 1988.

USENIX:1989:CSF

[Ass89a] USENIX Association, editor. Computing Systems, Fall, 1989. USENIX Association, Berkeley, CA, USA, Fall 1989.

USENIX:1989:CSSa

[Ass89b] USENIX Association, editor.

Computing Systems, Spring, 1989. USENIX Association, Berkeley, CA, USA, Spring 1989.

USENIX:1989:CSSb

[Ass89c] USENIX Association, editor.

Computing Systems, Summer, 1989. USENIX Association, Berkeley, CA, USA, Summer 1989.

USENIX:1989:CSW

[Ass89d] USENIX Association, editor.

Computing Systems, Winter, 1989. USENIX, Berkeley, CA, USA, Winter 1989.

USENIX:1989:EUS

[Ass89e] USENIX Association, editor. European UNIX Systems User Group Newsletter, Spring, 1989. EUUG, Berkeley, CA, USA, Spring 1989. CODEN EONLE8. ISSN 1011-4211.

USENIX:1989:ENS

[Ass89f] USENIX Association, editor.

EUUG Newsletter, Spring, 1989. European UNIX Systems User Group, Berkeley, CA, USA, Spring 1989.

Almada:1989:EBU

[AW89] Alfredo Almada and David H.

Williams. Enhancing the 4.3 BSD UNIX serial line interface. ;login: the USENIX Association newsletter, 14(1):6–29, January/February 1989. CO-DEN LOGNEM. ISSN 1044-6397.

Baalbergen:1988:DIP

[Baa88] Erik H. Baalbergen. Design and implementation of parallel make. In Association [Ass88b], pages 135–158.

Baker:1989:MOM

[Bak89] M. Steven Baker. Miscellaneous: OPUS make 5.1 (DOS and OS/2); QuickPak professional (BASIC utilities); GOSIP — the government’s OSI network standard; Usenix and C++. Programmer’s Journal, 7(1):92–??, January 1, 1989. ISSN 0747-5861.

Balter:1983:EYW

[Bal83] Jim Balter. Everything you wanted to know about System V, and then some. In Software Tools Users Group [Sof83], pages 263–?? LCCN QA76.8.U65 U74 1983. Abstract only.

Bame:1985:HMA

[Bam85a] Paul Bame. A high-performance model for 2-D alphanumeric display generation. In USENIX Association [USE85c], pages 65–68.

Bame:1985:HPM

[Bam85b] Paul Bame. A high-performance model for 2-D alphanumeric display generation. In USENIX [USE85b], pages 65–68.

Banahan:1982:LSB

[Ban82] Mike Banahan. The loosing of the sticky bit, or how to speed up your UNIX. EUUG Newsletter, 2(4):125–128, Winter 1982. CODEN EONLE8. ISSN 1011-4211.

Banahan:1983:BPR

[Ban83] Mike Banahan. Benchmark programs release ‘A’. EUUG Newsletter, 3(3):20–23, Autumn 1983. CODEN EONLE8. ISSN 1011-4211.

Barrett:1983:IUI

[Bar83] P. L. Barrett. An implementation of UNIX for the Intel iAPX286. In Software Tools Users Group [Sof83], pages 79– ?? LCCN QA76.8.U65 U74 1983. Abstract only.

Barr:1988:COS

[Bar88a] John R. Barr. Co-resident operating system: UNIX and realtime distributed processing. In Association [Ass88e], pages 47– 53.

Barr:1988:CRO

[Bar88b] John R. Barr. Co-resident operating system: UNIX and realtime distributed processing. In Association [Ass88e], pages 47– 53.

Bass:1981:IDF

[Bas81] John L. Bass. Implementation description for file locking. ;login: the USENIX Association newsletter, 6(2):11–32, February 1981. CODEN LOGNEM. ISSN 1044-6397.

Bader:1983:NRT

[BB83] Morris Bader and William Allan Bader. New Ratfor tools for numeric and graphics processing. In Software Tools Users Group [Sof83], pages 411–417. LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Bach:1984:MUS

[BB84] Maurice J. Bach and Steven J. Buroff. A multiprocessor UNIX system. In Software Tools Users Group [Sof84], pages 174–177. LCCN QA76.8.U65 U83 1984.

Becker-Berlin:1987:SSF

[BB87] Julie Becker-Berlin. Software synchronization at the Federal Judicial Center. In USENIX Association [USE87d], pages 12–13. ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

Baecker:1983:EPC

[BBT83] Ronald Baecker, Paul Breslin, and Trevor Thompson. On enhancing the presentation of C source code. In Software Tools Users Group [Sof83], pages 17– ?? LCCN QA76.8.U65 U74 1983. Abstract only.

Britten:1984:MDS

[BC84] Chet Britten and Paul Chen. Multiprocessor debugging on a shared memory system. In USENIX Association [USE84c], pages 33–38. ISBN none. LCCN QA76.8.U65 U55 1984.

Barkley:1988:CFD

[BC88] Ronald E. Barkley and Danny Chen. CASPER the friendly daemon. In Association [Ass88f], pages 251–260.

Brand:1989:RRA

[BC89] Russell Brand and D. Brent Chapman. RAPID: Remote automated patch installation database. In USENIX Association [USE89j], pages 77–84.

Bricker:1987:ED

[BCL^{+}87] Allan Bricker, Morgan Clark, Tad Lebeck, Barton P. Miller, and Peter Wu. Experiences with DREGS. In USENIX Association [USE87f], pages 471– 481.

Bloom:1986:EIB

[BD86] James M. Bloom and Kevin J.

Dunlap. Experiences implementing BIND, A distributed name server for the DARPA Internet. In USENIX Association [USE86c], pages 172–181.

Bar-David:1987:TC

[BD87a]  Tsvi Bar-David.  Teaching

C++. In USENIX Association [USE87a], pages 232–237. ISBN ???? LCCN ????

Batchelder:1987:PDP

[BD87b] Ned Batchelder and Trevor Darrell. Psfig — A DITROFF preprocessor for POSTSCRIPT figures. In USENIX Association [USE87f], pages 31–42.

Barbacci:1989:DAH

[BDWW89] Mario R. Barbacci, Dennis L. Doubleday, Charles B. Weinstock, and Jeannette M. Wing. Developing applications for heterogeneous machine networks: The durra environment. In Association [Ass89d], pages 7–35.

Bina:1989:FFB

[BE89] Eric J. Bina and Perry A. Emrath. A faster fsck for BSD UNIX. In USENIX Association [USE89g], pages 173–185.

Bechtolsheim:1982:SW

[Bec82] Andreas Bechtolsheim. The SUN workstation. In Usr Group [Usr82], pages 61–?? Abstract only.

Becker:1984:ELM

[Bec84] Richard A. Becker. Experiences with a large mixedlanguage system running under the UNIX operating system. In Software Tools Users Group [Sof84], pages 326–331. LCCN QA76.8.U65 U83 1984.

Bell:1985:CSC

[BED^{+}85] C. Gordon Bell, Steve Emmerich, Ivor Durham, Daniel P. Siewiorek, and Andrew Wilson. Computer structures are changing: Will UNIX change with them? In USENIX Association [USE85e], pages 1–4. LCCN QA76.8.U65 U8 1985.

Beebe:1984:PTC

[Bee84a] Nelson Beebe. A portable TOPS-20-like command parser. In Software Tools Users Group [Sof84], pages 354–?? LCCN QA76.8.U65 U83 1984. Abstract only.

Beebe:1984:PTL

[Bee84b] Nelson Beebe. A portable TOPS-20-like command parser. In USENIX [USE84b], pages 354–?? Abstract only.

Beeker:1986:ISU

[Bee86] Etienne Beeker. Image synthesis with UNIX. EUUG Newsletter, 6(1):4–9, Spring 1986. CO-DEN EONLE8. ISSN 1011-4211.

Blewett:1986:MBI

[BEHW86a] C. D. Blewett, J. T. Edmark, J. I. Helfman, and M. Wish. A multi-representation, bitmap interface to the UNIX file system constructed from cooperating processes. In USENIX Association [USE86d], pages 41– 48. ISBN ???? LCCN ????

Blewett:1986:MRB

[BEHW86b] C. D. Blewett, J. T. Edmark, J. I. Helfman, and M. Wish. A multi-representation, bitmap interface to the UNIX file system constructed from cooperating processes. In USENIX Association [USE86d], pages 41– 48. ISBN ???? LCCN ????

Bellovin:1988:STM

[Bel88a] S. M. Bellovin. The “session tty” manager. In USE-NIX Association [USE88j], pages 339–354. ISBN ???? LCCN ???? FTP - research.att.com:/dist/sessext.ps.Z; local - sessext.ps.

Bellovin:1988:TM

[Bel88b] S. M. Bellovin. The “Session Tty” Manager. In Association [Ass88f], pages 339–354. URL ftp://ftp.

uu.net/doc/security/att/
smb/sessext.ps;   ftp://
research.bell-labs.com/dist/
sessext.ps.Z. FTP - re-

search.att.com:/dist/sessext.ps.Z; local - sessext.ps.

Bender:1982:SDB

[Ben82] Mike Bender. Selecting a data base management system for a super micro. In Usr Group [Usr82], pages 257–268.

Berens:1985:APU

[Ber85] Peter H. Berens. Array processing under UNIX. In USENIX Association [USE85e], pages 175–181. LCCN QA76.8.U65 U8 1985.

Bershad:1986:LBM

[Ber86] Brian Bershad. Load balancing with Maitre d’. ;login: the USENIX Association newsletter, 11(1):32–45, January/February 1986. CODEN LOGNEM. ISSN 1044-6397.

Borenstein:1988:MMM

[BERS88a] Nathaniel Borenstein, Craig Everhart, Jonathan Rosenberg, and Adam Stoller. A multimedia message system for andrew. In Association [Ass88f], pages 37–42.

Borenstein:1988:MMS

[BERS88b] Nathaniel Borenstein, Craig Everhart, Jonathan Rosenberg, and Adam Stoller. A multimedia message system for Andrew. In USENIX Association [USE88j], pages 37–42. ISBN ???? LCCN ????

Beyls:1987:NUT

[Bey87] Pascal Beyls. Now UNIX talks to me in my language. In USENIX Association [USE87f], pages 307–321.

Beyls:1988:ASF

[Bey88] Pascal Beyls. An adaptation of spell to French. EUUG Newsletter, 8(1):11–14, Spring 1988. CODEN EONLE8. ISSN 1011-4211.

Ballance:1989:UIR

[BFGK89] Charles Ballance, Sean Fleming, Jay Goldberg, and Nelly Karasik. UFOS: An intelligent real-time performance monitor for UNIX System V. In USENIX Association [USE89f], pages 277–286. LCCN QA 76.76 O63 U83 1989.

Bonomi:1989:DPL

[BFS89] F. Bonomi, P. J. Fleming, and P. D. Steinberg. Distributing processes in looselycoupled UNIX multiprocessor systems. In USENIX Association [USE89f], pages 61–72. LCCN QA 76.76 O63 U83 1989.

Bach:1988:MFS

[BG88a] Maurice J. Bach and Ron Gomes. Measuring file system activity in the UNIX system. In USENIX Association [USE88e], pages 43–52. ISBN ???? LCCN ????

Berman:1988:NCP

[BG88b] C. Berman and R. Gur. NAPS — A C++ project case study. In USENIX Association [USE88k], pages 137–152.

Bahill:1986:CES

[BH86] A. Terry Bahill and Pat Harris. Cogito, an expert system to give installation advice for UNIX 4.2BSD. ;login: the USENIX Association newsletter, 11(5):4–8, September/October 1986. CODEN LOGNEM. ISSN 1044-6397.

Bigelow:1985:PSF

[Big85] Charles Bigelow. Principles of structured font design for the personal workstation. ;login: the USENIX Association newsletter, 10(4):65–83, October/November 1985. CODEN LOGNEM. ISSN 1044-6397.

Bihari:1988:FVO

[Bih88] Thomas E. Bihari. Functional vs. object-oriented development of robot-control software (A comparison of two robot-control programs). In Association [Ass88e], pages 80–81.

Bilyeu:1986:ELA

[Bil86] Bob Bilyeu. Experience with large applications on Unix. In USENIX Association [USE86e], pages 110–?? Title listed only, no paper or abstract.

Bishop:1983:HVL

[Bis83] Mitch Bishop. Handling very large programs on a 16-bit super-micro. In Association [Ass83a], pages 41–47.

Bishop:1987:HWS

[Bis87a] Matt Bishop. How to write a Setuid program. ;login: the USENIX Association newsletter, 12(1):5–11, January/February 1987. CO-DEN LOGNEM. ISSN 1044-

6397. URL http://nob.cs.
ucdavis.edu/~bishop/papers/
1987-setuid/.

Bishop:1987:RMS

[Bis87b] Matt Bishop. The RIACS mail system. ;login: the USENIX Association newsletter, 12(3): 3–26, May/June 1987. CODEN LOGNEM. ISSN 1044-6397.

Bishop:1987:SA

[Bis87c] Matt Bishop. Sharing accounts. In USENIX Association [USE87d], pages 36–?? ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

Bishop:1988:AFD

[Bis88a] Matt Bishop. An application of a fast data encryption standard implementation. In Association [Ass88c], pages 221–254.

Bishop:1988:AFN

[Bis88b] Matthew A. Bishop. Auditing files on a network of UNIX machines. In USENIX Association [USE88g], pages 51– 52. LCCN QA76.8.U65 U55 1988(1)-1990(2)//. Abstract only.

Bivand:1987:UIG

[Biv87] Roger Bivand. A user interface for geographers — what can UNIX offer? In USENIX Association [USE87b], pages 183– 190. ISBN ???? LCCN ????

Butler:1984:USH

[BK84] T. W. Butler and L. A. Kennedy. The UNIX system help facility. In Software Tools Users Group [Sof84], pages 253–257. LCCN QA76.8.U65 U83 1984.

Beck:1985:VAB

[BK85] Bob Beck and Bob Kasten. VLSI assist in building a multiprocessor UNIX system. In USENIX Association [USE85e], pages 255–275. LCCN QA76.8.U65 U8 1985.

Bentley:1987:SAA

[BK87] Jon L. Bentley and Brian W. Kernighan. A system for algorithm animation. In USENIX Association [USE87c], pages 13–23. ISBN ???? LCCN ????

Bauer:1988:NAR

[BK88a] David S. Bauer and Michael E.

Koblentz. NIDX — A real-time intrusion detection expert system. In Association [Ass88f], pages 261–273.

Bauer:1988:NRT

[BK88b] David S. Bauer and Michael E.

Koblentz. NIDX — A real-time intrusion detection expert system. In USENIX Association [USE88j], pages 261–273. ISBN ???? LCCN ????

Bal:1989:DIS

[BKT89] Henri E. Bal, M. Frans Kaashoek, and Andrew S. Tanenbaum. A distributed implementation of the shared data-object model. In USENIX Association [USE89d], pages 1–19. ISBN ???? LCCN ????

Barkley:1988:HBC

[BL88a] Ronald E. Barkley and T. Paul Lee. A heap-based callout implementation to meet real-time needs. In USENIX Association [USE88j], pages 213–222. ISBN ???? LCCN ????

Barkley:1988:HCI

[BL88b] Ronald E. Barkley and T. Paul Lee. A heap-based callout implementation to meet real-time needs. In Association [Ass88f], pages 213–222.

Boykin:1989:PMD

[BL89] Joseph Boykin and Alan Langerman. The parallelization of Mach/4.3BSD: Design philosophy and performance analysis. In USENIX Association [USE89d], pages 105–126. ISBN ???? LCCN ????

Black:1983:NCC

[Bla83a] J. P. Black. The Newcastle Connection: Current status and future plans. In Software Tools Users Group [Sof83], pages 377–382. LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Blackett:1983:MDR

[Bla83b] Kent Blackett. A menu-driven real-time UNIX system. In USENIX [USE83a], pages 279– 283.

Blackett:1983:MRU

[Bla83c] Kent Blackett. A menu-driven real-time UNIX system. In Association [Ass83a], pages 279– 283.

Black:1989:SCM

[Bla89] Eric Black. Software configuration management with an object-oriented database. In USENIX Association [USE89g], pages 257–272.

Blevins:1983:PUG

[Ble83] Jack Blevins. The port of UNIX to the Gould 32/27. In Association [Ass83a], pages 273–278.

Beaudouin-Lafon:1987:FMM

[BLK87] Michel Beaudouin-Lafon and Solange Karsenty. A framework for man machine interfaces design. In USENIX Association [USE87b], pages 1–10. ISBN ???? LCCN ????

Bach:1987:RFC

[BLMY87] M. J. Bach, M. W. Luppi, A. S.

Melamed, and K. Yueh. A remote-file cache for RFS. In USENIX Association [USE87f], pages 273–279.

Buxton:1983:UIM

[BLSS83] W. Buxton, M. R. Lamb, D. Sherman, and K. C. Smith. A user interface management system. In Software Tools Users Group [Sof83], pages 177–?? LCCN QA76.8.U65 U74 1983. Abstract only.

Barnett:1987:IUN

[BM87] Lewis Barnett and Michael K.

Malloy. ILMON: A UNIX network monitoring facility. In USENIX Association [USE87g], pages 133–144. LCCN QA 76.76 O63 U84 1987. Cover title: Conference proceedings. Spine title: Winter 1987 USE-NIX Association Conference proceedings, Washington, DC.

Borthwick:1987:IWB

[BNB87] Stuart Borthwick, John R.

Nicol, and Gordon S. Blair. An intelligent, window based interface to UNIX. In USENIX Association [USE87b], pages 225– 241. ISBN ???? LCCN ????

Breeden:1983:CAC

[BO83a] Laura Breeden and Mike O’Brien. CSNET: A computer science research network. In Software Tools Users Group [Sof83], pages 371–376. LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Breeden:1983:CCS

[BO83b] Laura Breeden and Mike O’Brien. CSNET: A computer science research network. In Association [Ass83a], pages 371–376.

Beck:1987:PPP

[BO87] Bob Beck and Dave Olien. A parallel programming process model. In USENIX Association [USE87g], pages 83– 102. LCCN QA 76.76 O63 U84 1987. Cover title: Conference proceedings. Spine title: Winter 1987 USENIX Association Conference proceedings, Washington, DC.

Bobey:1988:MPP

[Bob88]  Kenneth Bobey.  Monitor-

ing program performance on large parallel systems. In USENIX Association [USE88i], pages 43–49. ISBN ???? LCCN ????

Bohannon:1986:RUS

[Boh86] Thomas A. Bohannon. Reflections on A UNIX scheduler. ;login: the USENIX Association newsletter, 11(4):39– 45, July/August 1986. CODEN LOGNEM. ISSN 1044-6397.

Boldyreff:1988:MED

[Bol88]  Cornelia Boldyreff.  Macro

expansion as defined by the ANSI/ISO C draft standard. EUUG Newsletter, 8(2):36– 39, Summer 1988. CODEN EONLE8. ISSN 1011-4211.

Bott:1984:OTS

[Bot84] Ross Bott. OSx: Towards a single UNIX system for superminis. In USENIX Association [USE84c], pages 145–167. ISBN none. LCCN QA76.8.U65 U55 1984.

Bourne:1989:WSC

[Bou89]  Steven R. Bourne.  What

a source code control system should do. In USENIX Association [USE89j], page ?? Listed in contents only, no abstract, no paper.

Boyd:1984:SIP

[Boy84] Stowe Boyd. SYSTANT: An integrated programming environment for modular C under UNIX. In Software Tools Users Group [Sof84], pages 141–150. LCCN QA76.8.U65 U83 1984.

Butterfield:1984:NTL

[BP84] David A. Butterfield and Gerald J. Popek. Network tasking in the locus distributed UNIX system. In Software Tools Users Group [Sof84], pages 62–71. LCCN QA76.8.U65 U83 1984.

Barak:1986:MSU

[BP86] Amnon Barak and On G. Paradise. MOS — scaling up UNIX. In USENIX Association [USE86c], pages 414–418.

Bershad:1988:WEUa

[BP88a] Brian N. Bershad and C. Brian Pinkerton. Watchdogs — extending the UNIX file system. In USENIX Association [USE88j], pages 267–275. ISBN ???? LCCN ????

Bershad:1988:WEUb

[BP88b] Brian N. Bershad and C. Brian Pinkerton. Watchdogs — extending the UNIX file system. In Association [Ass88b], pages 169–188.

Brett:1987:PIA

[BPM87] Cliff Brett, Steve Pieper, and David Meltzer. Putting it all together: An integrated package for viewing and editing 3D microworlds. In USENIX Association [USE87c], pages 2–12. ISBN ???? LCCN ????

Borghi:1986:SIP

[BQd86] Bruno Borghi, Stephane Querel, and Daniel deRauglaudre. SmScript: An interpretor for the PostScript language under UNIX. In USENIX Association [USE86c], pages 284–293.

Bray:1989:LNO

[Bra89] Tim Bray. Lessons of the New Oxford English Dictionary Project. In USENIX Association [USE89g], pages 187–199. This dictionary is coded with SGML markup.

Breckon:1983:ESP

[Bre83]  Theresa Breckon.  Envi-

ronments and search paths for the software tools. In Software Tools Users Group [Sof83], pages 419–424. LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Breuel:1988:DLP

[Bre88a] Thomas M. Breuel. Data level parallel programming in C++. In USENIX Association [USE88k], pages 153–167.

Breuel:1988:LCC

[Bre88b] Thomas M. Breuel. Lexical closures for C++. In USENIX Association [USE88k], pages 293– 304.

Brooks:1985:EES

[Bro85] Catherine A. Brooks. Experiences with electronic software distribution. In USENIX Association [USE85e], pages 433– 436. LCCN QA76.8.U65 U8 1985.

Brooks:1987:NPC

[Bro87] Piete Brooks. An NRS processor in C and the future. EUUG Newsletter, 7(1):65–67, 1987. CODEN EONLE8. ISSN 1011-4211.

Brown:1988:CPP

[Bro88] Jonathan Brown. The CTSS/ POSIX project. In USENIX Association [USE88i], pages 33–?? ISBN ???? LCCN ???? Abstract only.

Bronson:1989:C

[Bro89a] Nathaniel R. Bronson, III. CC-SLAND. In USENIX [USE89a], pages 87–94.

Brown:1989:HSU

[Bro89b] P. J. Brown. A hypertext system for UNIX. In Association [Ass89d], pages 37–53.

Bruck:1988:MCS

[Bru88] Dag M. Bruck. Modelling of control systems with C++ and PHIGS. In USENIX Association [USE88k], pages 183–192.

Bryan:1983:VAM

[Bry83a] Scott Bryan. VCHK — A maintenance program for UNIX file hierarchies. In Software Tools Users Group [Sof83], pages 41– 44. LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Bryan:1983:VMP

[Bry83b] Scott Bryan. VCHK — A maintenance program for UNIX file hierarchies. In Association [Ass83a], pages 41–44.

Bryant:1988:RPC

[Bry88] Ray Bryant. The RP3 parallel computing environment. In USENIX Association [USE88i], pages 69–92. ISBN ???? LCCN ????

Barzel:1985:PAD

[BS85a] Ronen Barzel and David Salesin. Patchwork: A dataflow model for efficient graphics programming. ;login: the USE-NIX Association newsletter, 10 (4):43–53, October/November 1985. CODEN LOGNEM. ISSN 1044-6397.

Barzel:1985:PDM

[BS85b] Ronen Barzel and David Salesin. Patchwork: A dataflow model for efficient graphics programming. ;login: the USE-NIX Association newsletter, 10 (4):43–53, October/November 1985. CODEN LOGNEM. ISSN 1044-6397.

Biyabani:1988:IDC

[BSR88] Sara Biyabani, John A. Stankovic, and Krithi Ramamritham. The integration of deadline and criticalness requirements in hard real-time systems. In Association [Ass88e], pages 12–17.

Bacon:1988:NAN

[BSY88a] David F. Bacon, Jed Schwartz, and Yechiam Yemini. Nest: A network simulation and prototyping tool. In USENIX Association [USE88j], pages 71–77. ISBN ???? LCCN ????

Bacon:1988:NNS

[BSY88b] David F. Bacon, Jed Schwartz, and Yechiam Yemini. Nest: A network simulation and prototyping tool. In Association [Ass88f], pages 71–77.

Banda:1988:DDR

[BV88] Venu P. Banda and Richard A. Volz. Debugging distributed real-time software. In Association [Ass88e], pages 66–70.

Barton:1988:BTR

[BW88a] J. M. Barton and J. C. Wagner. Beyond threads: Resource sharing in UNIX. In USENIX Association [USE88j], pages 259–266. ISBN ???? LCCN ????

Barton:1988:ERS

[BW88b] J. M. Barton and J. C. Wagner. Enhanced resource sharing in UNIX. In Association [Ass88b], pages 111–133.

Barak:1989:MIM

[BW89] Amnon Barak and Richard Wheeler. MOSIX: An integrated multiprocessor UNIX. In USENIX Association [USE89g], pages 101–112.

Blewett:1987:NIS

[BWH87] C. Douglas Blewett, Myron (Mike) Wish, and Jonathan I. Helfman. A new IPC system for bitmap graphics applications: Review, model, and benchmarks. In USENIX Association [USE87f], pages 159– 184.

Bachrach:1985:XVM

[BWP85] Jonathan Bachrach, John Wallerius, and Jehan-Francois Paris. A XINU virtual machine. In USENIX Association [USE85e], pages 348–355. LCCN QA76.8.U65 U8 1985.

Beer:1987:DSG

[BWS87a] Stephen Beer, Ray Welland, and Ian Sommerville. DES — support for the graphical design of software. In USENIX Association [USE87b], pages 261–273. ISBN ???? LCCN ????

Beer:1987:SGD

[BWS87b] Stephen Beer, Ray Welland, and Ian Sommerville. DES — support for the graphical design of software. In USENIX Association [USE87b], pages 261–273. ISBN ???? LCCN ????

Cabrera:1986:IWL

[Cab86] Luis-Felipe Cabrera. The influence of workload on load balancing strategies. In USENIX Association [USE86c], pages 446–458.

Coppeto:1989:OLC

[CAG89a] Thomas J. Coppeto, Beth L.

Anderson, and Daniel E. Geer, Jr. OLC: An online consulting system for UNIX. In USENIX [USE89b], pages 83–94. FTP - aeneas.mit.edu:/pub/usenix/olc.PS; local - olc.ps.

Coppeto:1989:OOC

[CAG89b] Thomas J. Coppeto, Beth L.

Anderson, and Daniel E. Geer, Jr. OLC: An online consulting system for UNIX. In USENIX Association [USE89f], pages 83–94. LCCN QA 76.76 O63 U83 1989.

                                        URL ftp://aeneas.mit.edu/
                                        pub/usenix/olc.PS;   ftp:
                                        //ftp.uu.net/networking/
                                        athena/usenix/olc.PS.Z. FTP

- aeneas.mit.edu:/pub/usenix/olc.PS; local - olc.ps.

Calland:1983:EF

[Cal83] Bob Calland. Enhancements to format. In Software Tools Users Group [Sof83], pages 7– ?? LCCN QA76.8.U65 U74 1983. Abstract only.

Caplinger:1982:NUE

[Cap82a] Michael Caplinger. [Phi]NIX: A UNIX emulator for VAX/VMS. In Usr Group [Usr82], pages 249–255.

Caplinger:1982:PUE

[Cap82b] Michael Caplinger. [Phi]NIX: A UNIX emulator for VAX/ VMS. In USENIX [USE82a], pages 249–255.

Caplinger:1988:MAG

[Cap88] Michael Caplinger. A memory allocator with garbage collection for C. In USENIX Association [USE88j], pages 325–330. ISBN ???? LCCN ????

Carter:1982:PEH

[Car82a] Joel R. Carter. Perkin-Elmer’s hardware/I-O system: Flexibility that matches UNIX. In USENIX [USE82a], pages 108– ?? Abstract only.

Carter:1982:PHI

[Car82b] Joel R. Carter. Perkin-Elmer’s hardware/I-O system: Flexibility that matches UNIX. In Usr Group [Usr82], pages 108– ?? Abstract only.

Cargill:1986:FP

[Car86] T. A. Cargill. The feel of Pi. In USENIX Association [USE86e], pages 62–71.

Cargill:1987:PAC

[Car87a] T. A. Cargill. Pi: A case study in object-oriented programming. In USENIX Association [USE87a], pages 282–303. ISBN ???? LCCN ????

Cargill:1987:PCS

[Car87b] T. A. Cargill. Pi: A case study in object-oriented programming. In USENIX Association [USE87a], pages 282–303. ISBN ???? LCCN ????

Carolan:1987:C

[Car87c]  John Carolan.  C++ for

OS/2. In USENIX Association [USE87a], pages 47–65. ISBN ???? LCCN ????

Carlin:1988:USA

[Car88a] Jerry M. Carlin. UNIX security at Pacific Bell. In USENIX Association [USE88g], pages 86– 87. LCCN QA76.8.U65 U55 1988(1)-1990(2)//. Abstract only.

Carlin:1988:USP

[Car88b] Jerry M. Carlin. UNIX security at Pacific Bell. In USENIX [USE88b], pages 86–87. Abstract only.

Carson:1988:UGE

[Car88c] Scott D. Carson. Using groups effectively in Berkeley Unix. In USENIX Association [USE88j], pages 171–173. ISBN ???? LCCN ????

Carter:1988:USA

[Car88d]  Steve Carter.  Update on

systems administration standards. In USENIX Association [USE88f], pages 49–?? ISBN ???? LCCN ????

Chen:1983:EPU

[CB83] Paul Chen and Chet Britten. Experiences in porting 4.1BSD UNIX to the [lambda]750 VLSI development system. In Association [Ass83a], pages 132–?? Abstract only.

Campbell:1986:OAS

[CC86] Lisa M. Campbell and Mark D. Campbell. An overview of the Ada [1] shell. In USENIX Association [USE86e], pages 302– 313.

Clancy:1987:UV

[CCD^{+}87] Patrick Clancy, Benjamin F. Cutler, J. Christopher Dodd, Douglas W. Gilmore, Robert P. Nix, John J. O’Donnell, and Christopher P. Ryland. UNIX on a VLIW. In USENIX Association [USE87f], pages 225– 241.

Clay:1989:UEH

[CCF89] Larry Clay, George Copeland, and Mike Franklin.

UNIX extensions for highperformance transaction processing. In USENIX Association [USE89h], pages 73–80. ISBN ???? LCCN ????

Call:1987:COS

[CCM87] Lisa A. Call, David L. Cohrs, and Barton P. Miller. CLAM — an open system for graphical user interfaces. In USENIX Association [USE87a], pages 305– 326. ISBN ???? LCCN ????

Comer:1985:TTU

[CD85] Douglas Comer and Ralph E.

Droms. Tilde trees in the UNIX environment. In USENIX Association [USE85c], pages 23–29.

Cohn:1989:AAP

[CDT89a] David L. Cohn, William P. Delaney, and Karen M. Tracey. ARCADE: A platform for heterogeneous distributed operating systems. In USENIX Association [USE89d], pages 373– 390. ISBN ???? LCCN ????

Cohn:1989:APH

[CDT89b] David L. Cohn, William P. Delaney, and Karen M. Tracey. ARCADE: A platform for heterogeneous distributed operating systems. In USENIX Association [USE89d], pages 373– 390. ISBN ???? LCCN ????

Campbell:1989:TSB

[CE89a] Mark. D. Campbell and Tracy R. Edmonds. TOWER STREAMS-Based TTY: Architecture and implementation. In USENIX [USE89b], pages 15– 27.

Campbell:1989:TST

[CE89b] Mark. D. Campbell and Tracy R. Edmonds. TOWER STREAMS-Based TTY: Architecture and implementation. In USENIX Association [USE89f], pages 15–27. LCCN QA 76.76 O63 U83 1989.

Cerofolini:1982:USB

[Cer82] Luigi Cerofolini. UNIX for the STD bus. EUUG Newsletter, 2(4):49–51, Winter 1982. CO-DEN EONLE8. ISSN 1011-4211.

Cerofolini:1983:USB

[Cer83] Luigi Cerofolini. UNIX for the STD bus. In Association [Ass83a], pages 185–?? Abstract only.

Cole:1985:IEF

[CFA85] Clement T. Cole, Perry B. Flinn, and Alan B. Atlas. An implementation of an extended file system for UNIX. In USENIX Association [USE85e], pages 131–149. LCCN QA76.8.U65 U8 1985.

Crowley:1987:DME

[CFLT87] Terrence Crowley, Harry Forsdick, Matt Landau, and Virginia Travers. The Diamond multimedia editor. In USENIX Association [USE87f], pages 1– 18.

Cranmer-Gordon:1988:SVR

[CGFCKT88] Robert Cranmer-Gordon, Bill Fraser-Campbell, Mike Kelly, and Peter Tyrell. System V Release 3, diskless workstations and NFS. In USENIX Association [USE88e], pages 301–310. ISBN ???? LCCN ????

Cordy:1983:TAN

[CH83a] J. R. Cordy and R. C. Holt.

Turing: A new general purpose computer language under UNIX. In Software Tools Users Group [Sof83], pages 249–254. LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Cordy:1983:TNG

[CH83b] J. R. Cordy and R. C. Holt.

Turing: A new general purpose computer language under UNIX. In Association [Ass83a], pages 249–254.

Chang:1985:S

[Cha85] JoMei Chang. SunNet. In USENIX Association [USE85e], pages 71–78. LCCN QA76.8.U65 U8 1985.

Chahley:1987:NGP

[Cha87a] Richard Chahley. Next generation planning tool. In USENIX Association [USE87d], pages 19–?? ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

Chartock:1987:RS

[Cha87b] Howard Chartock. RFS in SunOS. In USENIX Association [USE87f], pages 281–290.

Chernick:1982:NPS

[Che82] Mike Chernick. NBS projects on software technology and computer based office systems. In Usr Group [Usr82], pages 342–?? Abstract only.

Chedgey:1987:PST

[Che87a] Chris Chedgey. Papillon — support tools for the development of graphical software. In USENIX Association [USE87b], pages 59–71. ISBN ???? LCCN ????

Cheng:1987:VAC

[Che87b] Ray Cheng. Virtual address cache in UNIX. In USENIX Association [USE87f], pages 217– 224.

Chesson:1987:PED

[Che87c] Greg Chesson. Protocol engine design. In USENIX Association [USE87f], pages 209–215.

Chen:1989:CPD

[Che89] Yih-Farn Chen. The C program database and its applications. In USENIX Association [USE89f], pages 157–171. LCCN QA 76.76 O63 U83 1989.

Christiansen:1989:OAF

[Chr89a] Tom Christiansen. Op: A flexible tool for restricted superuser access. In USENIX Association [USE89e], pages 89–94. ISBN ???? LCCN ????

Christiansen:1989:OFT

[Chr89b] Tom Christiansen. Op: A flexible tool for restricted superuser access. In USENIX Association [USE89e], pages 89–94. ISBN ???? LCCN ????

Conde:1989:UT

[CHS89] Daniel S. Conde, Felix S. Hsu, and Ursula Sinkewicz. ULTRIX threads. In USENIX Association [USE89f], pages 257–268. LCCN QA 76.76 O63 U83 1989.

Ciccarella:1988:DIR

[Cic88] Gianfranco Ciccarella. Design and implementation of a real-time multivariable adaptive controller. In Association [Ass88e], pages 82–86.

Carson:1988:NID

[CJ88] Mark E. Carson and Wen-Der Jiang. New ideas in discretionary access control. In USENIX Association [USE88g], pages 35– 37. LCCN QA76.8.U65 U55 1988(1)-1990(2)//.

Carson:1989:SWS

[CJL^{+}89] Mark E. Carson, Wen-Der Jiang, Jeremy G. Liang, Gary L. Luckenbaugh, and Debra H. Yakov. Secure window systems for UNIX. In USENIX Association [USE89g], pages 441–455. An architecture for a CMW based on Trusted XENIX and a text-based windowing system. Also mentions some X related issues.

Chedgey:1987:DAS

[CKK87] Chris Chedgey, Seamus Kearney, and Hans-Jurgen Kugler. Developing Ada software using VDM in an object-oriented framework. In USENIX Association [USE87b], pages 41–58. ISBN ???? LCCN ????

Cabrera:1985:IBM

[CKM85] Luis Felipe Cabrera, Michael J. Karels, and David Mosher. The impact of buffer management on networking software performance in Berkeley UNIX 4.2BSD: A case study. In USENIX Association [USE85e], pages 507–518. LCCN QA76.8.U65 U8 1985.

Callaghan:1989:A

[CL89] Brent Callaghan and Tom Lyon. The Automounter. In USENIX Association [USE89g], pages 43–51.

Clanton:1987:FAP

[Cla87a] Chuck Clanton. FACE: A poor man’s screen description language. In USENIX Association [USE87c], pages 101–?? ISBN ???? LCCN ???? Abstract only.

Clanton:1987:FPM

[Cla87b] Chuck Clanton. FACE: A poor man’s screen description language. In USENIX Association [USE87c], pages 101–?? ISBN ???? LCCN ???? Abstract only.

Clancy:1988:VME

[Cla88] Patrick Clancy. Virtual memory extensions in TRACE/ UNIX. In USENIX Association [USE88i], pages 137–150. ISBN ???? LCCN ????

Claybrook:1989:TMO

[Cla89] Bill Claybrook. A transaction model for online transaction processing systems. In USE-NIX Association [USE89h], pages 33–44. ISBN ???? LCCN ????

Clegg:1983:HEU

[Cle83a] Frederick W. Clegg. Hewlett– Packard’s entry into the UNIX community. In Association [Ass83a], pages 119–131.

Clegg:1983:HPE

[Cle83b] Frederick W. Clegg. Hewlett– Packard’s entry into the UNIX community. In USENIX [USE83a], pages 119–131.

Cugini:1989:STA

[CLH^{+}89] Janet A. Cugini, Shau-Ping Lo, Matthew S. Hecht, Chii-Ren Tsai, Virgil D. Gligor, Radhakrishna Aditham, and T. John Wei. Security testing of AIX system calls using Prolog. In USENIX Association [USE89f], pages 223–237. LCCN QA 76.76 O63 U83 1989.

Conant:1983:CCC

[CM83] Robert E. Conant and Herbert G. Mayer. COBOL compiler construction experiences using lex and yacc. In Association [Ass83a], pages 69–98.

Cabrera:1986:PAD

[CM86a] Luis-Felipe Cabrera and Eric Mowat. Pollster: A document annotation system for distributed environments. In USENIX Association [USE86c], pages 142–158.

Cabrera:1986:PDA

[CM86b] Luis-Felipe Cabrera and Eric Mowat. Pollster: A document annotation system for distributed environments. In USENIX [USE86a], pages 142– 158.

Choi:1989:EPA

[CM89a] ByoungJu Choi and Aditya P. Mathur. Experience with PMothra: A tool for mutation based testing on a Hypercube. In USENIX Association [USE89d], pages 237–253. ISBN ???? LCCN ????

Choi:1989:EPT

[CM89b] ByoungJu Choi and Aditya P. Mathur. Experience with PMothra: A tool for mutation based testing on a hypercube. In USENIX Association [USE89d], pages 237–253. ISBN ???? LCCN ????

Certen:1988:UEG

[CMM88] Pascale Le Certen, Beatrice Michel, and Gilles Muller. A UNIX environment for the GOTHIC kernel. In USENIX Association [USE88e], pages 219–229. ISBN ???? LCCN ????

Comer:1988:USC

[CN88] Douglas Comer and Thomas Narten. UNIX systems as Cypress implets. In USENIX Association [USE88j], pages 55– 62. ISBN ???? LCCN ????

Coggins:1987:ICS

[Cog87] James M. Coggins. Integrated class structures for image pattern recognition and computer graphics. In USENIX Association [USE87a], pages 240–245. ISBN ???? LCCN ????

Cohen:1987:RIR

[Coh87] Ephraim Cohen. Raster image rotation and anti-aliased line drawing. In USENIX Association [USE87c], pages 38–48. ISBN ???? LCCN ????

Collyer:1980:POR

[Col80] Geoff Collyer. A proposal for an Othello referee. ;login: the USENIX Association newsletter, 5(6):10–13, August 1980. CODEN LOGNEM. ISSN 1044-6397.

Cole:1983:AAP

[Col83] Clement T. Cole. Attaching an array processor in the UNIX environment. In Software Tools Users Group [Sof83], pages 135–?? LCCN QA76.8.U65 U74 1983. Abstract only.

Cole:1984:RIK

[Col84a] Allen Cole. A Ratfor implementation of KERMIT. In Software Tools Users Group [Sof84], pages 355–367. LCCN QA76.8.U65 U83 1984.

Collins:1984:WWU

[Col84b] Peter E. Collins. WINDX — Windows for the UNIX environment. In Software Tools Users Group [Sof84], pages 159–165. LCCN QA76.8.U65 U83 1984.

Collinson:1984:DUO

[Col84c] Peter Collinson. On the design of the UNIX operating system. ;login: the USENIX Association newsletter, 9(3):12–13, July 1984. CODEN LOGNEM. ISSN 1044-6397.

Collinson:1987:UC

[Col87a] Peter Collinson. UNIX: The cult. In USENIX Association [USE87g], pages 22–28. LCCN QA 76.76 O63 U84 1987. Cover title: Conference proceedings. Spine title: Winter 1987 USENIX Association Conference proceedings, Washington, DC.

Collyer:1987:NNS

[Col87b] Geoff Collyer. News need not be slow. In USENIX Association [USE87g], pages 181– 190. LCCN QA 76.76 O63 U84 1987. Cover title: Conference proceedings. Spine title: Winter 1987 USENIX Association Conference proceedings, Washington, DC.

Collinson:1988:LCB

[Col88] Peter Collinson. A low cost bitmapped terminal on the Atari ST. In Association [Ass88f], pages 355–364.

Collyer:1989:PTT

[Col89]  Geoff Collyer.  A par-

tial tour through the UNIX shell. In USENIX Association [USE89g], pages 343–353.

Conrad:1987:MGD

[Con87] Al Conrad. Modelling graphical data with C++. In USENIX Association [USE87a], pages 238–239. ISBN ???? LCCN ????

Connelly:1988:ULA

[Con88] Wayne C. Connelly. Unix login administration at Bellcore. In USENIX Association [USE88f], pages 13–15. ISBN ???? LCCN ????

Cornah:1982:OIA

[Cor82] Tony Cornah. An ONYX implementation of an allocation checking technique. EUUG Newsletter, 2(4):129–131, Winter 1982. CODEN EONLE8. ISSN 1011-4211.

Cottrell:1987:PFM

[Cot87] Pete Cottrell. Password file management at the University of Maryland. In USENIX Association [USE87d], pages 32– 33. ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

Cox:1982:OOP

[Cox82] Brad J. Cox. The object oriented pre-compiler: Programming Smalltalk 80 methods in C language. In Usr Group [Usr82], pages 44–?? Abstract only.

Cox:1983:OCP

[Cox83]  Brad J. Cox.  Objective

C: Programming Smalltalk-80 methods in C language. In Software Tools Users Group [Sof83], pages 236–?? LCCN QA76.8.U65 U74 1983. Abstract only.

Comer:1984:DAP

[CP84a] Douglas E. Comer and Larry L. Peterson. DRAGONMAIL: A prototype conversation-based mail system. In Software Tools Users Group [Sof84], pages 42– 51. LCCN QA76.8.U65 U83 1984.

Comer:1984:DPC

[CP84b] Douglas E. Comer and Larry L. Peterson. DRAGONMAIL: A prototype conversation-based mail system. In USENIX [USE84a], pages 42–51.

Chambers:1983:USI

[CQ83a] John Chambers and John Quarterman. UNIX System III and 4.1BSD; a practical comparison. In Association [Ass83a], pages 25–38.

Chambers:1983:USV

[CQ83b] John Chambers and John Quarterman. UNIX system V and 4.1C BSD. In Software Tools Users Group [Sof83], pages 265–291. LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Costantinidis:1989:DSS

[CR89] Peter Costantinidis, Jr. and Hamish Reid. The DV system of source file management. In USENIX Association [USE89j], pages 29–38.

Cragun:1983:USD

[Cra83] Don Cragun. UNIX system definitions and standards. In Association [Ass83a], pages 112– ?? Abstract only.

Crampton:1987:MMS

[Cra87a]  Chris Crampton.  MUSK

— a multi-user sketch program. In USENIX Association [USE87b], pages 17–29. ISBN ???? LCCN ????

Crampton:1987:MMU

[Cra87b]  Chris Crampton.  MUSK

— a multi-user sketch program. In USENIX Association [USE87b], pages 17–29. ISBN ???? LCCN ????

Campbell:1987:DMO

[CRJ87] Roy Campbell, Vincent Russo, and Gary Johnston. The design of a multiprocessor operating system. In USENIX Association [USE87a], pages 109–125. ISBN ???? LCCN ????

Clark:1986:AAF

[CS86] Paul Clark and Andrew Simms. AFQL — A flexible, general purpose interface to relational database management systems under UNIX. EUUG Newsletter, 6(3):7–25, Winter 1986. CODEN EONLE8. ISSN 1011-4211.

Cyganik:1988:SAA

[Cyg88] Marybeth Schultz Cyganik. System administration in the Andrew File System. In USENIX Association [USE88f], pages 67–69. ISBN ???? LCCN ????

Daniel:1983:ZAH

[Dan83a] Steve Daniel. Z — A high performance raster graphics package for UNIX operating systems. In Software Tools Users Group [Sof83], pages 135–?? LCCN QA76.8.U65 U74 1983. Abstract only.

Daniel:1983:ZHP

[Dan83b] Steve Daniel. Z — A high performance raster graphics package for UNIX operating systems. In Association [Ass83a], pages 135–?? Abstract only.

Das:1988:UAW

[Das88] Sunil K. Das. UNIX around the world. In USENIX Association [USE88e], pages 1–6. ISBN ???? LCCN ????

Datdeva:1988:LMG

[Dat88] Bjorn Datdeva. Lazy man’s guide to UNIX system administration. In USENIX Association [USE88f], pages 25–?? ISBN ???? LCCN ???? Abstract only.

Davis:1988:UO

[Dav88] Janet Davis. Unification and openness. EUUG Newsletter, 8 (2):69–72, Summer 1988. CO-DEN EONLE8. ISSN 1011-4211.

Davey:1989:GSD

[Dav89a]  David F. Davey.  Getting

started and debugging guide for SUN III. Australian UNIX systems User Group Newsletter, 10 (1):17–38, February 1989.

Davis:1989:PAR

[Dav89b] Don Davis. Project Athena’s release engineering tricks. In USENIX Association [USE89j], pages 101–106.

Davis:1989:OUS

[Dav89c] Janet Davis. Overview of UNIX System V Release 4.0. EUUG Newsletter, 9(1):74–78, Spring 1989. CODEN EONLE8. ISSN 1011-4211.

Dawson:1982:IHS

[Daw82] Mark Dawson. An implementation of Henderson’s SECD machine under UNIX. EUUG Newsletter, 2(4):33–39, Winter 1982. CODEN EONLE8. ISSN 1011-4211.

Doduc:1988:BA

[DB88] Nhuan Doduc and Christophe Binot. Benchmarking in the AFUU. EUUG Newsletter, 8 (1):15–18, Spring 1988. CO-DEN EONLE8. ISSN 1011-4211.

Darwin:1985:CHN

[DC85] Ian Darwin and Geoff Collyer. Can’t happen or /* NOTREACHED */ or real programs dump core. In USENIX Association [USE85c], pages 136–151.

DellaFera:1988:ZNS

[DEF^{+}88] C. Anthony DellaFera, Mark W. Eichin, Robert S. French, David C. Jedlinsky, John T. Kohl, and William E. Sommerfeld. The Zephyr notification service. In USENIX Association [USE88j], pages 213–219. ISBN ???? LCCN ????

DeJager:1986:AMS

[DeJ86a] Dale S. DeJager. The AT&T mail service and network. In USENIX [USE86a], pages 377– 390.

DeJager:1986:ATM

[DeJ86b] Dale S. DeJager. The AT&T mail service and network. In USENIX Association [USE86c], pages 377–390.

Delaney:1987:RTC

[Del87] Hubert C. Delaney. Ray tracing on the Connection Machine system. In USENIX Association [USE87c], pages 37–?? ISBN ???? LCCN ???? Abstract only.

Denny:1983:DUE

[Den83] Michael Denny. Delivering UNIX to the end-user market. In Association [Ass83a], pages 311–312. Abstract and summary.

Densmore:1986:OOP

[Den86] Owen M. Densmore. Object oriented programming in NeWS. In USENIX Association [USE86d], pages 117–135. ISBN ???? LCCN ????

Densmore:1989:NC

[Den89]  Owen Densmore.  NeWS

classes. In USENIX Association [USE89i], pages 37–68. ISBN ???? LCCN ????

Deroo:1983:CLS

[Der83] Jerry J. Deroo. A command line scanning package. In Software Tools Users Group [Sof83], pages 387–392. LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Dewhurst:1987:ACC

[Dew87] S. C. Dewhurst. The architecture of a C++ compiler. In USENIX Association [USE87a], pages 35–45. ISBN ???? LCCN ????

Downing:1984:TIS

[DF84] Curtis B. Downing and Frank Farance. Transparent implementation of shared libraries. In USENIX Association [USE84c], pages 209–222.

ISBN none. LCCN QA76.8.U65 U55 1984.

Dart:1989:CME

[DF89] Susan A. Dart and Peter Feiler. Configuration management of an environment. In USENIX Association [USE89j], pages 85–?? Abstract only.

Doeppner:1987:CPM

[DG87] Thomas W. Doeppner, Jr. and Alan J. Gebele. C++ on a parallel machine. In USENIX Association [USE87a], pages 95– 107. ISBN ???? LCCN ????

Domenico:1982:RMA

[DGM82] Ben Domenico, Ken Garnett, and Bill Meine. Rocky Mountain Area Implementors Group — corrections and enhancements for the distribution. In Usr Group [Usr82], pages 342– 354.

Detlefs:1987:ACC

[DHKW87] David Detlefs, Maurice Herlihy, Karen Kietzke, and Jeanette Wing. Avalon/C++: C++ extensions for transaction-based programming. In USENIX Association [USE87a], pages 451– 459. ISBN ???? LCCN ????

Dixon:1982:UMT

[Dix82] Jack Dixon. UNIX and manufacturing testing. In Usr Group [Usr82], pages 177–?? Abstract only.

Donner:1988:LOS

[DJ88] Marc D. Donner and David H. Jameson. Language and operating system features for realtime programming. In Association [Ass88d], pages 33–62.

Dudek:1986:HMF

[DJM86] Gregory Dudek, Michael Jenkin, and Howard Marcus. How to make friends with numbercrunchers: adding single-user array-processor slave environments to VAX UNIX. In USENIX Association [USE86c], pages 200–208.

Dick-Lauder:1984:AAA

[DLKE84] Piers Dick-Lauder, R. J. Kummerfeld, and Robert Elz. AC-SNET — the Australian alternative to UUCP. In Software Tools Users Group [Sof84], pages 11–17. LCCN QA76.8.U65 U83 1984.

Dineen:1987:NCA

[DLM^{+}87] Terence H. Dineen, Paul J.

Leach, Nathaniel W. Mishkin, Joseph N. Pato, and Geoffrey L. Wyant. The network computing architecture and system: An environment for developing distributed applications. In USENIX Association [USE87f], pages 385–398.

Dolan:1983:LST

[DM83] Charlie Dolan and Dave Martin. LISP for the software tools VOS. In Association [Ass83b], pages 15–?? Listing only, no abstract or paper available.

Davida:1988:UGD

[DM88] George I. Davida and Brian J. Matt. UNIX guardians: Delegating security to the user. In USENIX Association [USE88g], pages 14– 23. LCCN QA76.8.U65 U55 1988(1)-1990(2)//.

Draskoy:1987:XMU

[DN87] Andrew Draskoy and Gerald Neufeld. X.400 messaging on UNIX. In USENIX Association [USE87f], pages 111–115.

Dickey:1983:AIU

[DNQ^{+}83] Matt Dickey, Greg Noel, Bob Querido, Bill Appelbe, and Jim McGinness. Architectural implications of UNIX (or pitfalls for UNIX porters!). In Association [Ass83a], pages 307–?? Abstract only.

Djavaheri:1985:MAC

[DO85] Morris Djavaheri and Stan Osborne. Modula-2 — an alternative to C for system programming. In USENIX Association [USE85c], pages 34–42.

Donner:1988:BRO

[Don88] Marc D. Donner. Book review: Operating Systems: Communicating with and Controlling the Computer. ;login: the USENIX Association newsletter, 13(6):13–??, November/ December 1988. CODEN LOGNEM. ISSN 1044-6397.

Donner:1989:BRX

[Don89a] Marc D. Donner. Book review: Xlib Programming Manual and Xlib Reference Manual. ;login: the USENIX Association newsletter, 14(4):16– ??, July/August 1989. CODEN LOGNEM. ISSN 1044-6397.

Donner:1989:CTP

[Don89b] Marc D. Donner. C traps and pitfalls. ;login: the USENIX Association newsletter, 14(2): 9–??, March/April 1989. CO-DEN LOGNEM. ISSN 1044-6397.

Douglis:1989:EPM

[Dou89]  Fred Douglis.  Experience

with process migration in Sprite. In USENIX Association [USE89d], pages 59–72. ISBN ???? LCCN ????

Dunietz:1983:UZO

[DP83a] Jerry Dunietz and Robert Powell. The use of the Z80 I/O processor by the TRS-XENIX operating system. In Software Tools Users Group [Sof83], pages 101–108. LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Dunietz:1983:UZP

[DP83b] Jerry Dunietz and Robert Powell. The use of the Z80 I/O processor by the TRS-XENIX operating system. In Association [Ass83a], pages 101–108.

Ridder:1983:SSP

[dR83a] Theo de Ridder. Some selfreproducing programs. EUUG Newsletter, 3(4):9–11, Winter 1983. CODEN EONLE8. ISSN 1011-4211.

Domenico:1983:MTC

[DR83b] Ben Domenico and Russell K. Rew. Minimal test cases for the VOS software tools. In Software Tools Users Group [Sof83], pages 393–409. LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Ridder:1985:YAI

[dR85] Theo de Ridder. Yet another implementation of coroutines for C. EUUG Newsletter, 5(2): 1–8, Summer 1985. CODEN EONLE8. ISSN 1011-4211.

Daniel:1986:PWM

[DR86] Stephen Daniel and C. Durward Rogers. Programming with windows on the major workstations or through a glass darkly. In USENIX Association [USE86c], pages 441–445.

Danskin:1989:VPA

[DR89] John Danskin and Sally N. Rosenthal. Visual programming with Arachne. In USE-NIX Association [USE89i], pages 69–82. ISBN ???? LCCN ????

DeCouchant:1989:EIU

[DRK^{+}89] Dominique DeCouchant, Michel Riveill, Sacha Krakowiak, Chris Horn, Edward Finn, and Neville Harris. Experience with implementing and using an object oriented distributed system. In USENIX Association [USE89d], pages 301–310. ISBN ???? LCCN ????

Dronek:1982:BEB

[Dro82] Eugene F. Dronek. Benchmarking to eliminate the benchwarmers. In Usr Group [Usr82], pages 235–?? Abstract only.

Dronek:1984:RBP

[Dro84] Gene Dronek. Relating benchmarks to performance projections, or what do you do with 20 pounds of benchmark data? In Software Tools Users Group [Sof84], pages Addendum 12– 25. LCCN QA76.8.U65 U83 1984. Abstract p. 227, paper in Addendum pp. 12–25.

DeAlvare:1988:FPS

[DS88] Ana Maria De Alvare and E. Eugene Schultz, Jr. A framework for password selection. In USENIX Association [USE88g], pages 8– 9. LCCN QA76.8.U65 U55 1988(1)-1990(2)//. Abstract only.

Duchamp:1989:PDN

[Duc89] Dan Duchamp. Protocols for distributed and nested transactions. In USENIX Association [USE89h], pages 45–53. ISBN ???? LCCN ????

Duffy:1982:IUB

[Duf82] Michael E. Duffy. Interfacing UNIX to backend database machines. In Usr Group [Usr82], pages 19–28.

Duff:1985:QSA

[Duf85]  Tom Duff.  Quaternion

splines for animating orientation. In USENIX Association [USE85d], pages 54–62. ISBN ???? LCCN ????

Duff:1989:EVU

[Duf89a] Tom Duff. Experience with viruses on UNIX systems. In Association [Ass89b], pages 155–171.

Duff:1989:VAU

[Duf89b] Tom Duff. Viral attacks on UNIX system security. In USE-NIX Association [USE89g], pages 165–171.

Dunlop:1988:RNA

[Dun88] Dominic Dunlop. Receiving news at a small commercial site: Is it worth it? EUUG Newsletter, 8(2):66–68, Summer 1988. CODEN EONLE8. ISSN 1011-4211.

Dunlop:1989:RII

[Dun89] Dominic Dunlop. Report on ISO/IEC JTC1/SC22/WG15 (POSIX) meeting. ;login: the USENIX Association newsletter, 14(6):25–??, November/ December 1989. CODEN LOGNEM. ISSN 1044-6397.

Dewan:1989:SOC

[DV89] Prasun Dewan and Eric Vasilik. Supporting objects in a conventional operating system. In USENIX Association [USE89g], pages 273–285.

Davidson:1988:PCR

[DW88] Susan B. Davidson and Aaron Watters. Partial computation in real-time database systems. In Association [Ass88e], pages 117–121.

Dyer:1982:BHB

[Dye82a] Steve Dyer. Bad-sector handling on the BBN C machines. In Usr Group [Usr82], pages 69–78.

Dyer:1982:BSH

[Dye82b] Steve Dyer. Bad-sector handling on the BBN C machines. In USENIX [USE82a], pages 69–78.

Dyer:1988:HNS

[Dye88] Stephen P. Dyer. The Hesiod name server. In USENIX Association [USE88j], pages 183– 189. ISBN ???? LCCN ????

Dykstra:1987:BCP

[Dyk87]  Phillip C. Dykstra.  The

BRL CAD package — an overview. In USENIX Association [USE87c], pages 73–80. ISBN ???? LCCN ????

Eaton:1988:PAU

[Eat88] Charles K. Eaton. Project accounting on UNICOS. In USENIX Association [USE88j], pages 163–169. ISBN ???? LCCN ????

Ewens:1985:TAD

[EBFH85a] P. Ewens, D. R. Blythe, M. Funkenhauser, and R. C. Holt. Tunis: A distributed multiprocessor operating system. In USENIX Association [USE85e], pages 247–254. LCCN QA76.8.U65 U8 1985.

Ewens:1985:TDM

[EBFH85b] P. Ewens, D. R. Blythe, M. Funkenhauser, and R. C. Holt. Tunis: A distributed multiprocessor operating system. In USENIX [USE85a], pages 247–254.

Eccles:1988:PCL

[Ecc88] Joseph Eccles. Porting from Common Lisp with Flavors to C++. In USENIX Association [USE88k], pages 31–40.

Eggert:1989:AIS

[Egg89] Paul R. Eggert. Automating the importation of software. In USENIX Association [USE89j], pages 115–118.

Edler:1986:CMP

[EGL86] Jan Edler, Allan Gottlieb, and Jim Lipkis. Considerations for massively parallel UNIX systems on the NYU Ultracomputer and IBM RP3. In USENIX Association [USE86e], pages 193–210.

Elahian:1983:NUM

[Ela83] Camran Elahian. New UNIX markets in engineering. In Association [Ass83a], pages 313– ?? Abstract only.

Ellis:1985:SSP

[Ell85] Bruce Ellis. A stable storage package. In USENIX Association [USE85e], pages 209–212. LCCN QA76.8.U65 U8 1985.

Edler:1988:MMS

[ELS88a] Jan Edler, Jim Lipkis, and Edith Schonberg. Memory management in Symunix II: A design for large-scale shared memory multiprocessors. In USENIX Association [USE88i], pages 151–168. ISBN ???? LCCN ????

Edler:1988:PMH

[ELS88b] Jan Edler, Jim Lipkis, and Edith Schonberg. Process management for highly parallel UNIX systems. In USENIX Association [USE88i], pages 1–17. ISBN ???? LCCN ????

Elz:1984:RCP

[Elz84] Robert Elz. Resource controls, privileges, and other MUSH. In Software Tools Users Group [Sof84], pages 183–191. LCCN QA76.8.U65 U83 1984.

Engert:1988:AID

[Eng88] Douglas E. Engert. Attaching IBM disks directly to a Cray X-MP. In USENIX Association [USE88i], pages 227–229. ISBN ???? LCCN ????

Eppinger:1989:NTL

[Epp89]  Jeffrey L. Eppinger.  The

nested top-level lazy serverbased transaction. In USENIX Association [USE89h], pages 81–82. ISBN ???? LCCN ????

Erlinger:1988:NCU

[Erl88] Michael A. Erlinger. A notice capability for UNIX. In USENIX Association [USE88f], pages 21–22. ISBN ???? LCCN ????

Essick:1985:NWY

[Ess85] Raymond B. Essick, IV. Notesfiles: Why you should use them. In USENIX [USE85b], pages 195–200.

Ehrlich:1989:IUS

[ESS89] Kate Ehrlich, Barbara Stanley, and Tim Shea. Incorporating usability studies and interface design into software development. In USENIX Association [USE89f], pages 133–145. LCCN QA 76.76 O63 U83 1989.

Elsesser:1986:MSC

[EST86] G. W. Elsesser, M. S. Safran, and T. Tieger. Managing separate compilation in AT&T’s UNIX Ada system. In USENIX Association [USE86e], pages 252–260.

Evans:1983:W

[Eva83] Steven R. Evans. Windows with 4.2BSD. In Association [Ass83a], pages 260–?? Abstract only.

Evans:1986:N

[Eva86] Steve Evans. The Notifier. In USENIX Association [USE86c], pages 344–354.

Epstein:1988:AAD

[EVS88a] Mark E. Epstein, Curt Vandetta, and John Sechrest. Asmodeus — A daemon servant for the system administrator. In Association [Ass88f], pages 377–391.

Epstein:1988:ADS

[EVS88b] Mark E. Epstein, Curt Vandetta, and John Sechrest. Asmodeus — A daemon servant for the system administrator. In USENIX Association [USE88j], pages 377–391. ISBN ???? LCCN ????

Eykholt:1986:PUS

[Eyk86] Joseph R. Eykholt. Porting UNIX to the System/370 Extended Architecture. In USENIX Association [USE86e], pages 157–164.

Eykholt:1988:NEH

[Eyk88] Joseph R. Eykholt. A new exception handling mechanism for the UNIX kernel. In USENIX Association [USE88j], pages 291–295. ISBN ???? LCCN ????

Fernandez:1988:EUP

[FA88] Gary Fernandez and Larry Allen. Extending the UNIX protection model with access control lists. In Association [Ass88f], pages 119–132.

Fair:1986:PU

[Fai86] Erik E. Fair. A perspective on the USENET. ;login: the USENIX Association newsletter, 11(1):46–52, January/February 1986. CODEN LOGNEM. ISSN 1044-6397.

Forin:1989:SMS

[FBS89] Alessandro Forin, Joseph Barrera, and Richard Sanzi. The shared memory server. In USE-NIX Association [USE89g], pages 229–243.

Feder:1983:EUS

[Fed83] Jerome Feder. Evolution of UNIX system performance. In Association [Ass83a], pages 110–111. Abstract only.

Fedor:1988:GAM

[Fed88a] Mark S. Fedor. Gated: A multirouting protocol daemon for UNIX. In Association [Ass88f], pages 365–376.

Fedor:1988:GMR

[Fed88b] Mark S. Fedor. Gated: A multi-routing protocol daemon for UNIX. In USENIX Association [USE88j], pages 365–376. ISBN ???? LCCN ????

Feldman:1984:AHU

[Fel84] Stuart I. Feldman. An architecture history of the UNIX system. In Software Tools Users Group [Sof84], pages xi–xvi. LCCN QA76.8.U65 U83 1984.

Fenlon:1987:CSN

[Fen87] M. K. Fenlon. A case study of network management. In USE-NIX Association [USE87d], pages 2–3. ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

Ferrin:1982:PIV

[Fer82] Thomas E. Ferrin. Performance issues of VMUNIX revisited. ;login: the USENIX Association newsletter, 7(5):3– 6, November 1982. CODEN LOGNEM. ISSN 1044-6397.

Ferrin:1985:REP

[Fer85] Thomas Ferrin. A recipe for establishing point-to-point TCP/ IP network links with 4.2 BSD UNIX. In USENIX Association [USE85e], pages 113–117. LCCN QA76.8.U65 U8 1985.

Feuer:1984:YCF

[Feu84] Alan R. Feuer. You CAN feel good knowing it is written in C. In USENIX Association [USE84c], pages 57–64. ISBN none. LCCN QA76.8.U65 U55 1984.

Feuer:1985:SIC

[Feu85] Alan R. Feuer. si — an interpreter for the C language. In USENIX Association [USE85e], pages 47–55. LCCN QA76.8.U65 U8 1985.

Fenart:1986:OTI

[FFH^{+}86] Jean Marc Fenart, Marc Fievet, Christian Huitema, Bernard Martin, Annie Remille, and Guy Vaysseix. OSI and TCP/ IP protocols on a UNIX system. In USENIX Association [USE86c], pages 46–58.

Funkenhauser:1988:UTU

[FH88] M. J. Funkenhauser and R. C.

Holt. Using TUNIS, A UNIX compatible kernel, as a basis for security. In USENIX Association [USE88g], pages 70– 77. LCCN QA76.8.U65 U55 1988(1)-1990(2)//.

French:1988:SLM

[FHW88] Andrew H. French, Antoinette F. Hershey, and Edward J. Wilkens. Software license management in a network environment. In USENIX Association [USE88g], pages 68– 69. LCCN QA76.8.U65 U55 1988(1)-1990(2)//. Abstract only.

Filipski:1985:LSB

[Fil85] Alan Filipski. Latent source bugs and UNIX system portability. In USENIX Association [USE85c], pages 125–130.

Fischer:1986:ACU

[Fis86a] Herman Fischer. Ada, “C”, and UNIX. In USENIX [USE86b], pages 225–240.

Fischer:1986:AU

[Fis86b] Herman Fischer. Ada, “C”, and UNIX. In USENIX Association [USE86e], pages 225–240.

Fischer:1986:SAB

[Fis86c] Herman Fischer. SVID as A basis for CAIS implementation. In USENIX Association [USE86e], pages 294–301.

Fischer:1986:SBC

[Fis86d] Herman Fischer. SVID as a basis for CAIS implementation. In USENIX [USE86b], pages 294– 301.

Fitzhorn:1982:CTC

[FJ82] Patrick A. Fitzhorn and Gearold R. Johnson. C: Toward a concise syntactic description. EUUG Newsletter, 2 (4):54–69, Winter 1982. CO-DEN EONLE8. ISSN 1011-4211.

Finger:1985:MCV

[FKN85a] Eric J. Finger, Mihcael M.

Krueger, and Alan F. Nugent. A Multi-CPU version of the UNIX kernel — technical aspects and market need. In USE-NIX [USE85b], pages 11–22.

Finger:1985:MVU

[FKN85b] Eric J. Finger, Mihcael M.

Krueger, and Alan F. Nugent. A Multi-CPU version of the UNIX kernel — technical aspects and market need. In USENIX Association [USE85c], pages 11–22.

Farley:1983:CSL

[FKT83] Michael Farley, Paul Kunkel, and Trevor Thompson. A C source language debugger. In Software Tools Users Group [Sof83], pages 31–40. LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Fowler:1989:EFH

[FKV89] Glenn S. Fowler, David G.

Korn, and K.-Phong Vo. An efficient file hierarchy walker. In USENIX Association [USE89f], pages 173–188. LCCN QA 76.76 O63 U83 1989.

Fortier:1982:DIB

[FL82] Richard Fortier and Anthony Lake. Design of an intelligent bitmap terminal. In Usr Group [Usr82], pages 51–60.

Fostel:1983:DUV

[FN83] Gary Fostel and Alison Naylor. Developing a UNIX validation set. In Software Tools Users Group [Sof83], pages 329–339. LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Foster:1983:EAV

[Fos83a] Dick Foster. EtherTIP — A virtual terminal interface to Ethernet. In Software Tools Users Group [Sof83], pages 311–?? LCCN QA76.8.U65 U74 1983. Abstract only.

Foster:1983:EVT

[Fos83b] Dick Foster. EtherTIP — A virtual terminal interface to Ethernet. In Association [Ass83a], pages 311–?? Abstract only.

Foster:1988:ETP

[Fos88]  Brian Foster.  An exper-

imental trusted path prototype. In USENIX Association [USE88g], pages 53– 56. LCCN QA76.8.U65 U55 1988(1)-1990(2)//.

Fouts:1988:MUU

[Fou88] Martin Fouts. Multitasking under UniCos: Experiences with the Cray 2. In USENIX Association [USE88i], pages 121–131. ISBN ???? LCCN ????

Fowler:1985:FGM

[Fow85] Glenn S. Fowler. The fourth generation Make. In USENIX Association [USE85e], pages 159–174. LCCN QA76.8.U65 U8 1985.

Foxley:1987:MTP

[Fox87] Eric Foxley. Music: A Troff preprocessor for printing music scores. EUUG Newsletter, 7(2): 7–23, 1987. CODEN EONLE8. ISSN 1011-4211.

Freed:1985:MMM

[Fre85] Adrian Freed. MacMix: Mixing music with a mouse. In USE-NIX Association [USE85d], pages 23–37. ISBN ???? LCCN ????

Friedenbach:1987:CM

[Fri87] Ken Friedenbach. C++ on the Macintosh. In USENIX Association [USE87a], pages 67–76. ISBN ???? LCCN ????

Finkel:1989:TSA

[FS89] Raphael Finkel and Brian Sturgill. Tools for system administration in a heterogeneous environment. In USENIX Association [USE89e], pages 15–29. ISBN ???? LCCN ???? n.

Franke:1983:EEU

[FT83] Deborah L. Franke and Thomas R. Truscott. Early experiences using UNIX on the Gould SEL concept computers. In Software Tools Users Group [Sof83], pages 341–347. LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Foree:1989:AAD

[FT89] Don Foree and Margaret Tiano. Automated account deactivation and deletion. In USENIX Association [USE89e], pages 31–33. ISBN ???? LCCN ???? n.

Fuhrman:1987:OCL

[Fuh87a] Ken Fuhrman. Object-oriented class library for C++. In USE-NIX Association [USE87a], pages 209–231. ISBN ???? LCCN ????

Fuhrman:1987:OOC

[Fuh87b] Ken Fuhrman. Object-oriented class library for C++. In USE-NIX Association [USE87a], pages 209–231. ISBN ???? LCCN ????

Fulton:1989:DTI

[Ful89a]  Dorm Fulton.  Distributed

transaction integrity. In USE-NIX Association [USE89h], pages 55–61. ISBN ???? LCCN ????

Fulton:1989:CMX

[Ful89b] Jim Fulton. Configuration management in the X window system. In USENIX Association [USE89j], pages 107–113.

Funk:1987:CPM

[Fun87] Susan A. Funk. CAS perspective on the maturation of UNIX. In USENIX Association [USE87f], pages 95–104.

Flink:1989:SVM

[FW89] Charles W. Flink, II and Jonathan D. Weiss. System V/MLS labeling and mandatory policy alternatives. In USENIX [USE89c], pages 413– 427.

Gancarz:1986:UAU

[Gan86a] Michael Gancarz. Uwm: A user interface for X Windows. In USENIX Association [USE86c], pages 429–440.

Gancarz:1986:UUI

[Gan86b] Michael Gancarz. Uwm: A user interface for X Windows. In USENIX [USE86a], pages 429– 440.

Gansner:1988:IAC

[Gan88a] E. R. Gansner. Iris: A classbased window library. In USE-NIX Association [USE88k], pages 283–292.

Gansner:1988:ICB

[Gan88b] E. R. Gansner. Iris: A classbased window library. In USE-NIX [USE88a], pages 283–292.

Gart:1986:TAU

[Gar86] Mitchell Gart. Targeting Ada to 68000/Unix. In USENIX Association [USE86c], pages 261– 274.

Gafke:1983:LNV

[GB83] Gary Gafke and Eric Bergan. Local network with virtual ports. In Software Tools Users Group [Sof83], pages 293–?? LCCN QA76.8.U65 U74 1983. Abstract only.

Gopinath:1989:EFM

[GB89] Prabha Gopinath and Thomas Bihari. Experiences with a family of multiprocessor real-time operating systems. In USENIX Association [USE89d], pages 205–225. ISBN ???? LCCN ????

Gill:1986:CUC

[GBH86] Helen Gill, Rebecca Bowerman, and Chuck Howell. A comparison of UNIX and CAIS system facilities. In USENIX Association [USE86e], pages 275–293.

Groundwater:1987:SUA

[GBM87a] Neil P. Groundwater, Neil Bodick, and Andre Marquis. A SunView user-interface for authoring and accessing a medical knowledge base. In USENIX Association [USE87b], pages 93–104. ISBN ???? LCCN ????

Groundwater:1987:SUI

[GBM87b] Neil P. Groundwater, Neil Bodick, and Andre Marquis. A SunView user-interface for authoring and accessing a medical knowledge base. In USENIX Association [USE87b], pages 93–104. ISBN ???? LCCN ????

Gull:1989:PMO

[GD89] Aarron Gull and Sunil K. Das.

A port of the MINIX operating system to the Atari ST. In Association [Ass89e], pages 2–14. CODEN EONLE8. ISSN 1011-4211.

Geer:1988:SMA

[Gee88a] Daniel E. Geer, Jr. Service management at Project Athena. In USENIX Association [USE88f], pages 71–?? ISBN ???? LCCN ????

Geer:1988:SMP

[Gee88b] Daniel E. Geer, Jr. Service management at project Athena. In USENIX Association [USE88f], pages 71–?? ISBN ???? LCCN ????

Genter:1986:UVM

[Gen86] Robert E. (Rick) Genter. Unix as a virtual machine environment. In USENIX Association [USE86c], pages 475–485.

George:1982:RP

[Geo82a] Johann George. Real-time performance. In Usr Group [Usr82], pages 15–?? Abstract only.

George:1982:RTP

[Geo82b] Johann George. Real-time performance. In USENIX [USE82a], pages 15–?? Abstract only.

Gerkin:1982:IUV

[Ger82] Fred Gerkin. Introduction to UNIX — videotape. In Usr Group [Usr82], pages 108–?? Abstract only.

Geshwind:1986:CAC

[Ges86] David M. Geshwind. Computer assisted color conversion. In USENIX Association [USE86d], pages 137–144. ISBN ???? LCCN ????

Gettys:1984:PA

[Get84] James Gettys. Project Athena. In Software Tools Users Group [Sof84], pages 72–77. LCCN QA76.8.U65 U83 1984.

Gettys:1986:PIW

[Get86] James Gettys. Problems implementing window systems in UNIX. In USENIX Association [USE86e], pages 89–97.

Gheith:1988:CCA

[GGSW88a] Ahmed Gheith, Prabha Gopinath, Karsten Schwan, and Peter Wiley. CHAOS and CHAOS-ART — extensions to an object-based kernel. In Association [Ass88e], pages 32–41.

Gheith:1988:CCE

[GGSW88b] Ahmed Gheith, Prabha Gopinath, Karsten Schwan, and Peter Wiley. CHAOS and CHAOS-ART — extensions to an object-based kernel. In Association [Ass88e], pages 32–41.

Gien:1983:SOS

[Gie83] Michel Gien. The Sol operating system. In Software Tools Users Group [Sof83], pages 75– 78. LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Gill:1986:PIC

[Gil86] Daniel P. Gill. A proposal for interwindow communication and translation facilities. In USENIX Association [USE86e], pages 79–88.

Gingell:1988:ESP

[Gin88] Robert A. Gingell. Evolution of the SunOS programming environment. In USENIX Association [USE88e], pages 255–269. ISBN ???? LCCN ????

Gingell:1987:SLS

[GLDW87] Robert A. Gingell, Meng Lee, Xuong T. Dang, and Mary S. Weeks. Shared libraries in SunOS. In USENIX Association [USE87f], pages 131–145.

Glew:1989:BLP

[Gle89] Andy Glew. Boxes, links, and parallel trees: Elements of a configuration management system. In USENIX Association [USE89j], pages 17–28.

Goble:1982:WNA

[GM82a] George Goble and Michael H.

Marsh. What’s new at Purdue/EE department — A dual processor VAX 11/780. In Usr Group [Usr82], pages 113–138.

Goble:1982:WNP

[GM82b] George Goble and Michael H.

Marsh. What’s new at Purdue/ EE department — A dual processor VAX 11/780. In USE-NIX [USE82a], pages 113–138.

Gloor:1989:DSL

[GM89] Peter Gloor and Rudolf Marty. Dynamically synchronized locking — a lightweight locking protocol for resource locking in a stateless environment. In USE-NIX Association [USE89g], pages 13–27.

Gingell:1987:VMA

[GMS87] Robert A. Gingell, Joseph P. Moran, and William A. Shannon. Virtual memory architecture in SunOS. In USENIX Association [USE87f], pages 81– 94.

Ghodssi:1986:GOS

[GMW86] Vida Ghodssi, Steven S. Muchnick, and Alex Wu. Global optimizer for Sun FORTRAN, C & Pascal. In USENIX Association [USE86c], pages 318–334.

Goldberg:1988:CTP

[Gol88] Dave Goldberg. Combining two printing systems under a common user interface. In USENIX Association [USE88f], pages 29–31. ISBN ???? LCCN ????

Gomez:1985:UIP

[Gom85] Julian E. Gomez. A UNIX image production pipeline. ;login: the USENIX Association newsletter, 10(4):40–42, October/November 1985. CODEN LOGNEM. ISSN 1044-6397.

Goodwin:1984:SVP

[Goo84] Ken Goodwin. System V performance enhancements. ;login: the USENIX Association newsletter, 9(3):6–11, July 1984. CODEN LOGNEM. ISSN 1044-6397.

Gorlen:1987:OCL

[Gor87a] Keith E. Gorlen. An objectoriented class library for C++ programs. In USENIX Association [USE87a], pages 181–207. ISBN ???? LCCN ????

Gorlen:1987:OOC

[Gor87b] Keith E. Gorlen. An objectoriented class library for C++ programs. In USENIX Association [USE87a], pages 181–207. ISBN ???? LCCN ????

Gosling:1986:SAD

[Gos86a] James Gosling. SUNDEW: A distributed and extensible window system. In USENIX Association [USE86e], pages 98–103.

Gosling:1986:SDE

[Gos86b] James Gosling. SUNDEW:

A distributed and extensible window system. In USENIX [USE86b], pages 98–103.

Gould:1985:DDM

[Gou85] Ed Gould. Device drivers in a multiprocessor environment. In USENIX Association [USE85e], pages 357–360. LCCN QA76.8.U65 U8 1985.

Gould:1986:NFS

[Gou86]  Ed Gould.  The Network

File System implemented on 4.3BSD. In USENIX Association [USE86c], pages 294–298.

Gray:1988:PAL

[GP88] W. H. Gray and A. K. Powers. Project accounting on a large-scale UNIX system. In USENIX Association [USE88f], pages 7–12. ISBN ???? LCCN ????

Gomez:1986:HEH

[GPF^{+}86a] Julian E. Gomez, Frank Preston, Steve Fine, Tony Hasegawa, Bock Lee, and Blaine Walker. A high-end high-performance graphics system for computational fluid dynamics. In USENIX Association [USE86d], pages 13–14. ISBN ???? LCCN ????

Gomez:1986:HHG

[GPF^{+}86b] Julian E. Gomez, Frank Preston, Steve Fine, Tony Hasegawa, Bock Lee, and Blaine Walker. A high-end high-performance graphics system for computational fluid dynamics. In USENIX Association [USE86d], pages 13–14. ISBN ???? LCCN ????

Gehani:1985:CCO

[GR85] N. H. Gehani and W. D. Roome. Concurrent C — an overview. In USENIX Association [USE85c], pages 43–50.

Gray:1987:AER

[Gra87] Robert W. Gray. Automatic error recovery in a fast parser. In USENIX Association [USE87f], pages 337–346.

Gray:1988:GAG

[Gra88a] Robert W. Gray. [gamma]- GLA — A generator for lexical analyzers that programmers can use. In Association [Ass88f], pages 147–160.

Gray:1988:GGG

[Gra88b] Robert W. Gray. [gamma]- GLA — A generator for lexical analyzers that programmers can use. In USENIX Association [USE88j], pages 147–160. ISBN ???? LCCN ????

Gray:1988:AGL

[Gra88c] Robert W. Gray. γ-GLA— A generator for lexical analyzers that programmers can use. In USENIX Association [USE88h], pages 147–160.

Greenberg:1982:IUS

[Gre82a] Robert B. Greenberg. Is UNIX as a standard doomed? In Usr Group [Usr82], pages 280–281. Abstract only.

Greenberg:1982:USD

[Gre82b] Robert B. Greenberg. Is UNIX as a standard doomed? In USENIX [USE82a], pages 280– 281. Abstract only.

Gridley:1985:IPS

[Gri85] Curt Gridley. Improving the performance of scientific applications on a supermicro using a custom floating point processor and an optimizing compiler. In USENIX Association [USE85e], pages 597–610. LCCN QA76.8.U65 U8 1985.

Griswold:1989:DSI

[Gri89] Ralph E. Griswold. Data structures in the Icon programming language. In Association [Ass89a], pages 339–365.

Groundwater:1982:NSD

[Gro82] Neil P. Groundwater. Navy software development with ratfor-T and software tools. In Usr Group [Usr82], pages 342– ?? Abstract only.

Groundwater:1984:AYA

[Gro84] Neil Groundwater. Ada? Yet another VOS? In Software Tools Users Group [Sof84], pages 353–?? LCCN QA76.8.U65 U83 1984. Abstract only.

Grob:1987:AEC

[Gro87] Lori S. Grob. Automatic exploitation of concurrency in C: Is it really so hard? In USENIX Association [USE87b], pages 209–223. ISBN ???? LCCN ????

Gronke:1988:BRP

[Gro88] Ed Gronke. Book review: Programming in ANSI C. ;login: the USENIX Association newsletter, 13(6):12–??, November/ December 1988. CODEN LOGNEM. ISSN 1044-6397.

Garribo:1988:UNO

[GRS88a] A. Garribo, L. Regoli, and G. Succi. UNO: USENET news on optical disk. In USENIX Association [USE88e], pages 97– 103. ISBN ???? LCCN ????

Garribo:1988:UUN

[GRS88b] A. Garribo, L. Regoli, and G. Succi. UNO: USENET news on optical disk. In USENIX Association [USE88e], pages 97– 103. ISBN ???? LCCN ????

Gautron:1987:TEC

[GS87] Philippe Gautron and Marc Shapiro. Two extensions to C++: A dynamic link editor and inner data. In USENIX Association [USE87a], pages 23– 32. ISBN ???? LCCN ????

Guffy:1983:SVO

[Guf83] W. R. Guffy. System V offering.

In Association [Ass83a], pages 48–?? Abstract only.

Guruprasad:1988:PUT

[Gur88] V. Guruprasad. Prototext:

Universal text drivers. In Association [Ass88f], pages 331–338.

Grossman:1986:RRS

[GW86a] Mark S. Grossman and Glen E.

Williams. Real-time resource sharing for graphics workstations. In USENIX Association [USE86e], pages 23–33.

Grossman:1986:RTR

[GW86b] Mark S. Grossman and Glen E.

Williams. Real-time resource sharing for graphics workstations. In USENIX [USE86b], pages 23–33.

Gusella:1984:TAN

[GZ84a] Riccardo Gusella and Stefano Zatti. TEMPO — A network time controller for a distributed Berkeley UNIX system. In Software Tools Users Group [Sof84], pages 78–85. LCCN QA76.8.U65 U83 1984.

Gusella:1984:TNT

[GZ84b] Riccardo Gusella and Stefano Zatti. TEMPO — A network time controller for a distributed Berkeley UNIX system. In USENIX [USE84a], pages 78– 85.

Hemenway:1984:PSS

[HA84] Kathleen Hemenway and Helene Armitage. Proposed syntax standard for UNIX system commands. In USENIX Association [USE84c], pages 89–98. ISBN none. LCCN QA76.8.U65 U55 1984.

Haenlin:1983:DBF

[Hae83] Edward Haenlin. A data base frontend, driven by tables generated from a data dictionary. In Software Tools Users Group [Sof83], pages 313–320. LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Haeberli:1985:DEI

[Hae85a] Paul Haeberli. A data-flow environment for interactive graphics. In USENIX Association [USE85d], pages 1–12. ISBN ???? LCCN ????

Haeberli:1985:DFE

[Hae85b] Paul Haeberli. A data-flow environment for interactive graphics. In USENIX Association [USE85d], pages 1–12. ISBN ???? LCCN ????

Haeberli:1986:DFM

[Hae86a] Paul E. Haeberli. A dataflow manager for an interactive programming environment. In USENIX [USE86a], pages 419– 428.

Haeberli:1986:DMI

[Hae86b] Paul E. Haeberli. A dataflow manager for an interactive programming environment. In USENIX Association [USE86c], pages 419–428.

Haemer:1989:UUC

[Hae89] Jeffrey S. Haemer. An update on UNIX and C standards activity. ;login: the USENIX Association newsletter, 14(6):31– 45, November/December 1989. CODEN LOGNEM. ISSN 1044-6397.

Hagen:1983:CSN

[Hag83] Teus Hagen. Cookbook for setting up a national UNIX systems users group. EUUG Newsletter, 3(2):7–10, Summer 1983. CODEN EONLE8. ISSN 1011-4211.

Hall:1985:STC

[Hal85a] Roy Hall. Scattered thoughts on color. In USENIX Association [USE85d], pages 63–75. ISBN ???? LCCN ????

Hall:1985:SAA

[Hal85b] Roy Hall. Software architecture for animation systems. ;login: the USENIX Association newsletter, 10(4):35–39, October/November 1985. CODEN LOGNEM. ISSN 1044-6397.

Hall:1987:RDU

[Hal87] Pennie Hall. Resource duplication for 100% uptime. In USE-NIX Association [USE87d], pages 43–?? ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

Hamburger:1987:UHC

[Ham87] Sheldon Hamburger. UNIX in health care: Medical laboratories — A case study. In USE-NIX Association [USE87g], pages 192–195. LCCN QA 76.76 O63 U84 1987. Cover title: Conference proceedings. Spine title: Winter 1987 USE-NIX Association Conference proceedings, Washington, DC.

Hanshew:1982:STD

[Han82] Jon Hanshew. The software tools on the Data General NOVA. In Usr Group [Usr82], pages 285–289.

Harvey:1983:UL

[Har83] Brian Harvey. UNIX logo. In Association [Ass83a], pages 145–150.

Harris:1985:TAN

[Har85a] Marion O. Harris. Thoughts on an all-natural user interface. In USENIX [USE85a], pages 343– 347.

Harris:1985:TAU

[Har85b] Marion O. Harris. Thoughts on an all-natural user interface. In USENIX Association [USE85e], pages 343–347. LCCN QA76.8.U65 U8 1985.

Harkness:1987:CMP

[Har87a] Ken Harkness. A centralized multi-system problem tracking system. In USENIX Association [USE87d], pages 40– ?? ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

Harkness:1987:CMS

[Har87b] Ken Harkness. A centralized multi-system problem tracking system. In USENIX Association [USE87d], pages 40– ?? ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

Harkness:1987:CFD

[Har87c] Ken Harkness. A Cron facility for downtime. In USENIX Association [USE87d], pages 41– 42. ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

Harrison:1987:MCS

[Har87d] Helen E. Harrison. Maintaining a consistent software environment. In USENIX Association [USE87d], pages 16–17. ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

Harrison:1988:BSH

[Har88a] Helen E. Harrison. A batching system for heterogeneous Unix environments. In USENIX Association [USE88f], pages 23– 24. ISBN ???? LCCN ????

Harrison:1988:FBS

[Har88b] Helen E. Harrison. A flexible backup system for large disk farms or what to do with 20 gigabytes. In USENIX Association [USE88f], pages 33–34. ISBN ???? LCCN ????

Hathaway:1982:STT

[Hat82] Steve Hathaway. Software tools TOPS-20 implementation. In Usr Group [Usr82], pages 291– 301.

Hausman:1983:TA

[Hau83] Paul Hausman. Tools in Australia. In Association [Ass83b], pages 14–?? Abstract only.

Hawley:1985:UTP

[Haw85] Michael J. Hawley. UNIX tools for a personal database. In USENIX Association [USE85e], pages 333–341. LCCN QA76.8.U65 U8 1985.

Hawley:1986:MMS

[Haw86a] Michael Hawley. MIDI music software for UNIX. In USENIX Association [USE86c], pages 1– 12.

Hawley:1986:PUB

[Haw86b] Michael Hawley. Porting Unix to the Bösendorfer. In USENIX Association [USE86d], pages 83–90. ISBN ???? LCCN ????

Hawley:1987:MMSa

[Haw87a] Michael Hawley. More MIDI software for UNIX. In USENIX Association [USE87b], pages 201–208. ISBN ???? LCCN ????

Hawley:1987:MMSb

[Haw87b] Michael Hawley. More music software for Unix. In USENIX Association [USE87c], pages 1– ?? ISBN ???? LCCN ???? Abstract only.

Hawley:1988:NAT

[Haw88a] D. Ryan Hawley. Netdump: A tool for dumping filesystems. In USENIX Association [USE88f], pages 27–?? ISBN ???? LCCN ???? man page only.

Hawley:1988:NTD

[Haw88b] D. Ryan Hawley. Netdump: A tool for dumping filesystems. In USENIX Association [USE88f], pages 27–?? ISBN ???? LCCN ???? man page only.

Hawley:1989:ANC

[Haw89a] Michael Hawley. Audio I/O with the NeXT computer. ;login: the USENIX Association newsletter, 14(6):19– 23, November/December 1989. CODEN LOGNEM. ISSN 1044-6397.

Hawley:1989:AON

[Haw89b] Michael Hawley. Audio I/O with the NeXT computer. ;login: the USENIX Association newsletter, 14(6):19– 23, November/December 1989. CODEN LOGNEM. ISSN 1044-6397.

Hayes:1988:SGU

[Hay88] James Hayes. Standards and guidelines for Unix workstation installations. In USENIX Association [USE88f], pages 51–61. ISBN ???? LCCN ????

Honeyman:1986:PCF

[HB86] Peter Honeyman and Steven M. Bellovin. PATHALIAS or the care and feeding of relative addresses. In USENIX Association [USE86c], pages 126–141.

Hamilton:1988:ESM

[HC88] Graham Hamilton and Daniel S. Conde. An experimental symmetric multiprocessor Ultrix kernel. In USENIX Association [USE88j], pages 283–290. ISBN ???? LCCN ????

Hecht:1987:US

[HCC^{+}87] M. S. Hecht, M. E. Carson, C. S. Chandersekaran, R. S. Chapman, L. J. Dotterrer, V. D. Gligor, W. D. Jiang, A. Johri, G. L. Luckenbaugh, and N. Vasudevan. UNIX without the superuser. In USENIX Association [USE87f], pages 243–256.

Himelstein:1987:CMO

[HCE87] Mark I. Himelstein, Fred C.

Chow, and Kevin Enderby. Cross-module optimizations: Its implementation and benefits. In USENIX Association [USE87f], pages 347–356.

Himelstein:1988:RAR

[HCE88] Mark Himelstein, Steven Correll, and Kevin Enderby. A RISC approach to runtime exceptions. In Association [Ass88f], pages 239–249.

Hewson:1985:IAP

[HCN85] Denise Hewson, Gregory Cullen, and Alan Nugent. Integral array processing in a multiprocessor UNIX environment. In USENIX Association [USE85e], pages 183–187. LCCN QA76.8.U65 U8 1985.

Hansen:1988:IIM

[HE88] Stephen Hansen and Michael Eldredge. Intruder isolation and monitoring. In USENIX Association [USE88g], pages 63–64. LCCN QA76.8.U65 U55 1988(1)-1990(2)//. Abstract only.

Hecht:1988:ABS

[Hec88]  Stephen Hecht.  Andrew

Backup System. In USENIX Association [USE88f], pages 35–38. ISBN ???? LCCN ????

Hedrick:1989:IIP

[Hed89] Charles L. Hedrick. Introduction to the Internet protocols.

Australian UNIX systems User Group Newsletter, 10(1):66–96, February 1989.

Heffler:1982:DMC

[Hef82] Michael J. Heffler. Description of a menu creation and interpretation system. In Usr Group [Usr82], pages 235–?? Abstract only.

Heilman:1987:PEA

[Hei87] Eric Heilman. Priv: An exercise in administrative expansion. In USENIX Association [USE87d], pages 38–39. ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

Henshew:1983:UST

[Hen83] John Henshew. Update on software tools implementation — Data General’s RDOS. In Association [Ass83b], pages 14–?? Abstract only.

Hensgen:1989:DSS

[HF89] Debra Hensgen and Raphael Finkel. Dynamic server squads in Yackos. In USENIX Association [USE89d], pages 73–89. ISBN ???? LCCN ????

Hitz:1986:MFS

[HH86] David Hitz and Peter Honeyman. A mail file system for Eighth Edition UNIX. In USENIX Association [USE86c], pages 391–394.

Hofkin:1988:SAH

[HH88] Bob Hofkin and W. Terry Hardgrave. System administration in a heterogeneous network. In USENIX Association [USE88j], pages 119–123. ISBN ???? LCCN ????

Hidley:1983:DIG

[Hid83] Greg Hidley. Device independent graphics enhancements at ITTDCD. In Association [Ass83a], pages 247–250.

Hillery:1989:ICS

[Hil89] Nathan H. Hillery. Implementing a consistent system over many hosts. In USENIX Association [USE89e], pages 69–73. ISBN ???? LCCN ????

Hirgelt:1983:EMR

[Hir83]  Edward Hirgelt.  Enhanc-

ing MAKE or re-inventing a rounder wheel. In Software Tools Users Group [Sof83], pages 45–58. LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Hecht:1988:EAC

[HJAW88] Matthew S. Hecht, Abhai Johri, Radhakrishna Aditham, and T. John Wei. Experience adding C2 security features to UNIX. In Association [Ass88f], pages 133–146.

Haight:1983:AEU

[HK83] R. C. Haight and D. B. Knudsen. ARIEL: An experimental UNIX-based interactive video information system. In Association [Ass83a], pages 167–168. Abstract only.

Hoel:1986:UBO

[HK86a] Timothy W. Hoel and Bruce A. Keller. A Unix-based operating system for the Cray 2. In USE-NIX [USE86b], pages 219–224.

Hoel:1986:UOS

[HK86b] Timothy W. Hoel and Bruce A. Keller. A Unix-based operating system for the Cray 2. In USENIX Association [USE86e], pages 219–224.

Hawley:1985:WUA

[HL85a] Michael J. Hawley and Samuel J. Leffler. Windows for UNIX at Lucasfilm. In USENIX Association [USE85e], pages 393–406. LCCN QA76.8.U65 U8 1985.

Hawley:1985:WUL

[HL85b] Michael J. Hawley and Samuel J. Leffler. Windows for UNIX at Lucasfilm. In USENIX [USE85a], pages 393–406.

Hecht:1984:DFS

[HLW84] Matthew S. Hecht, John R. Levine, and Justin C. Walker. A distributed file system for UNIX. In USENIX Association [USE84c], pages 23–32. ISBN none. LCCN QA76.8.U65 U55 1984.

Honda:1989:SMU

[HM89] Masahiro Honda and Terrence Miller. Software management using a CASE environment. In USENIX Association [USE89j], pages 11–16.

Heydon:1988:MAV

[HMM^{+}88a] C. A. Heydon, M. W. Maimone, A. F. Moorman, J. D. Tygar, and J. M. Wing. Miro: A visual language for specifying security. In USENIX Association [USE88g], pages 49– ?? LCCN QA76.8.U65 U55 1988(1)-1990(2)//. Abstract only.

Heydon:1988:MVL

[HMM^{+}88b] C. A. Heydon, M. W. Maimone, A. F. Moorman, J. D. Tygar, and J. M. Wing. Miro: A visual language for specifying security. In USENIX [USE88b], pages 49–?? Abstract only.

Holt:1983:TAP

[HMP83a] R. C. Holt, M. P. Mendel, and S. G. Perelgut. TUNIS: A portable, UNIX compatible kernel written in Concurrent Euclid. In Software Tools Users Group [Sof83], pages 61– 74. LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Holt:1983:TPU

[HMP83b] R. C. Holt, M. P. Mendel, and S. G. Perelgut. TUNIS: A portable, UNIX compatible kernel written in concurrent Euclid. In Association [Ass83a], pages 61–74.

Hussain:1988:IXP

[HOG88] S. A. Hussain, J. Olnes, and T. Grimstad. Implementation of X.25 PLP in ISO 8802 LAN environments. In USENIX Association [USE88e], pages 89– 95. ISBN ???? LCCN ????

Hommel:1987:SBD

[Hom87] Carlton B. Hommel. System backup in a distributed responsibility environment. In USE-NIX Association [USE87d], pages 8–?? ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

Hoover:1983:UID

[Hoo83] Clyde W. Hoover. A user information data base for UNIX (what to do when /etc/passwd just isn’t enough). In Software Tools Users Group [Sof83], pages 121–134. LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Hopkins:1987:DSE

[Hop87a] Don Hopkins. Directional selection is easy as pie menus! In USENIX Association [USE87c], pages 103–?? ISBN ???? LCCN ???? Abstract only.

Hopkins:1987:DSI

[Hop87b] Don Hopkins. Directional selection is easy as pie menus! In USENIX Association [USE87c], pages 103–?? ISBN ???? LCCN ???? Abstract only.

Hopkins:1987:EUC

[Hop87c] William E. Hopkins. Experience in using C++ for software system development. In USE-NIX Association [USE87a], pages 327–344. ISBN ???? LCCN ????

Hopkins:1989:SPS

[Hop89] Don Hopkins. The shape of PSIBER space: PostScript interactive bug eradication routines. In USENIX Association [USE89i], pages 129–147. ISBN ???? LCCN ????

Horbal:1982:ATL

[Hor82a] Mark T. Horbal. ATLAS test language — A real time application under UNIX. In Usr Group [Usr82], pages 165–176.

Horton:1982:NCT

[Hor82b] Mark Horton. The new Curses and Terminfo package. In Usr Group [Usr82], pages 79–91.

Horton:1983:UNN

[Hor83] Mark R. Horton. Usenet: The network news. ;login: the USE-NIX Association newsletter, 8 (3):10–13, June 1983. CODEN LOGNEM. ISSN 1044-6397.

Horton:1984:WD

[Hor84a] Mark R. Horton. What is a domain? In USENIX [USE84a], pages 368–372.

Horton:1984:WID

[Hor84b] Mark R. Horton. What is a domain? In Software Tools Users Group [Sof84], pages 368–372. LCCN QA76.8.U65 U83 1984.

Hosler:1983:MCU

[Hos83] Jay R. Hosler. Meeting the coming UNIX training challenge. In Association [Ass83a], pages 177–184.

Hosler:1984:IPU

[Hos84]  Jay Hosler.  Interactivity

in packaged UNIX training: A modest proposal. In Software Tools Users Group [Sof84], pages 346–349. LCCN QA76.8.U65 U83 1984.

Howard:1988:OAF

[How88] John H. Howard. On overview of the Andrew File System. In USENIX Association [USE88j], pages 23–26. ISBN ???? LCCN ????

Honeyman:1985:PEM

[HP85] Peter Honeyman and Pat E. Parseghian. A parser for electronic mail addresses. In USENIX Association [USE85c], pages 184–190.

Henderson:1989:MIR

[HP89a] Robert L. Henderson and Alan Poston. MSS-II and RASH — A mainframe UNIX based mass storage system with a rapid access storage hierarchy file management system. In USENIX [USE89c], pages 65–84.

Henderson:1989:MRA

[HP89b] Robert L. Henderson and Alan Poston. MSS-II and RASH — A mainframe UNIX based mass storage system with a rapid access storage hierarchy file management system. In USENIX Association [USE89g], pages 65–84.

Holderbaugh:1989:MPM

[HP89c] Mark C. Holderbaugh and Scott E. Preece. Minimalist physical memory control in UNIX. In USENIX Association [USE89g], pages 245–256.

Hanrot:1987:KBC

[HQZ^{+}87] S. Hanrot, P. Quintrand, J. Zoller, E. Chouraqui, P. Dugerdil, P. Francois, and M. Ricard. A knowledge based CAD system in architecture on UNIX. In USENIX Association [USE87b], pages 169–181. ISBN ???? LCCN ????

Himelstein:1985:MD

[HR85a] Mark Himelstein and Peter Rowell. Multi-process debugging. In USENIX Association [USE85e], pages 155–158. LCCN QA76.8.U65 U8 1985.

Himelstein:1985:MPD

[HR85b] Mark Himelstein and Peter Rowell. Multi-process debugging. In USENIX [USE85a], pages 155–158.

AVHays:1982:IMP

[HRO82a] A. V. Hays, Jr., B. J. Richmond, and L. M. Optican. Implementing a multiple-process real-time system under UNIX. In USENIX [USE82a], pages 15–?? Abstract only.

Hays:1982:IMR

[HRO82b] A. V. Hays, Jr., B. J. Richmond, and L. M. Optican. Implementing a multiple-process real-time system under UNIX. In Usr Group [Usr82], pages 15–?? Abstract only.

Hanley:1982:SUU

[HS82] James R. Hanley and Jeffry A. Scott. A survey of UNIX usage in scientific and business applications. In Usr Group [Usr82], pages 281–?? Abstract only.

Harrison:1989:ENC

[HS89] Helen E. Harrison and Tim Seaver. Enhancements to 4.3BSD network commands. In USENIX Association [USE89e], pages 49–51. ISBN ???? LCCN ????

Horton:1984:PUU

[HSHK84] Mark R. Horton, Karen Summers-Horton, and Berry Kercheval. Proposal for a UUCP/Usenet registry host. In Software Tools Users Group [Sof84], pages 373–?? LCCN QA76.8.U65 U83 1984. Abstract only.

Harrison:1988:RTS

[HSY88] Helen E. Harrison, Stephen P. Schaefer, and Terry S. Yoo. Rtools: Tools for software management in a distributed computing environment. In Association [Ass88f], pages 85–93.

Hydar:1989:TPM

[HSYY89] Dan Hydar, David Scott, Felix Yang, and Morris Yeh. A transaction processing monitor. In USENIX Association [USE89h], pages 1–6. ISBN ???? LCCN ????

Hughes:1986:TRF

[Hug86] Ronald P. Hughes. The Transparent Remote File System. In USENIX Association [USE86c], pages 306–317.

Hume:1985:FRP

[Hum85] Andrew Hume. Folding regular polyhedra. In USENIX Association [USE85d], pages 76–81. ISBN ???? LCCN ????

Hume:1987:MSM

[Hum87] Andrew Hume. Mk: a successor to make. In USENIX Association [USE87f], pages 445–457.

Hume:1988:FMI

[Hum88a] Andrew Hume. The file motel — an incremental backup system for Unix. In Association [Ass88f], pages 61–72.

Hume:1988:GW

[Hum88b] Andrew Hume. Grep wars. In USENIX Association [USE88e], pages 237–245. ISBN ???? LCCN ????

Hume:1989:UTM

[Hum89] Andrew G. Hume. The use of a time machine to control software. In USENIX Association [USE89j], pages 119–124.

Hunter:1988:PAM

[Hun88a] Bruce H. Hunter. Password administration for multiple large scale systems. In USENIX Association [USE88f], pages 1–?? ISBN ???? LCCN ????

Hunter:1988:PCA

[Hun88b] Chad Hunter. Process cloning: A system for duplicating UNIX processes. In USENIX Association [USE88j], pages 373–379. ISBN ???? LCCN ????

Hunter:1988:PCS

[Hun88c] Chad Hunter. Process cloning: A system for duplicating UNIX processes. In Association [Ass88f], pages 373–379.

Hagemark:1989:SAL

[HZ89a] Bent Hagemark and Kenneth Zadeck. Site: A language and system for configuring many computers as one computing site. In USENIX Association [USE89e], pages 1–14. ISBN ???? LCCN ????

Hagemark:1989:SLS

[HZ89b] Bent Hagemark and Kenneth Zadeck. Site: A language and system for configuring many computers as one computing site. In USENIX Association [USE89e], pages 1–14. ISBN ???? LCCN ????

Ingham:1987:KWF

[Ing87] Kenneth Ingham. Keeping watch over the flocks by night (and day). In USENIX Association [USE87f], pages 105–110.

Inman:1985:ILC

[Inm85]  Jack Inman.  Implement-

ing loosely coupled functions on tightly coupled engines. In USENIX Association [USE85e], pages 277–298. LCCN QA76.8.U65 U8 1985.

Isaak:1982:RS

[Isa82a] Jim Isaak. Real-time systems.

In Usr Group [Usr82], pages 15–?? Abstract only.

Isaak:1982:RTS

[Isa82b] Jim Isaak. Real-time systems.

In USENIX [USE82a], pages 15–?? Abstract only.

Isaacson:1983:GPP

[Isa83a]  Joel Isaacson.  A general

purpose programming language with an embedded data base interface. In Software Tools Users Group [Sof83], pages 247–?? LCCN QA76.8.U65 U74 1983. Abstract only.

Isaak:1983:SOL

[Isa83b] Jim Isaak. Standards organization: Levels and measurement. In Association [Ass83a], pages 348–349. Abstract only.

Isley:1983:LAP

[Isl83a] Larry K. Isley. Licensing activity and pricing. In Association [Ass83a], pages 49–50. Abstract only.

Isley:1983:ULN

[Isl83b] Larry K. Isley. UNIX licensing and new AT&T product offerings. In Software Tools Users Group [Sof83], pages 3– ?? LCCN QA76.8.U65 U74 1983. Abstract only.

Ivie:1984:RWA

[Ivi84a] Evan L. Ivie. The readers workbench — A system for computer assisted reading. In Software Tools Users Group [Sof84], pages 270–279. LCCN QA76.8.U65 U83 1984.

Ivie:1984:RWS

[Ivi84b] Evan L. Ivie. The readers workbench — A system for computer assisted reading. In USE-NIX [USE84a], pages 270–279.

Innocent:1987:EUI

[IvW87] Peter R. Innocent, Gerrit C. van der Veer, and Yvonne Waern. Experiments with the user interface for UNIX mail. In USENIX Association [USE87b], pages 73–91. ISBN ???? LCCN ????

Jatkowski:1988:PGP

[JA88] Paul Jatkowski and Mike Akre. PMON: Graphical performance monitoring tool. In USENIX Association [USE88j], pages 111–118. ISBN ???? LCCN ????

Jackson:1982:M

[Jac82] Ken Jackson. MASCOT. In Usr Group [Usr82], pages 93– 106.

Jacobson:1983:IDA

[Jac83]  Van Jacobson.  Interactive

data analysis using the software tools. In Association [Ass83b], pages 8–?? Abstract only.

Jacob:1984:ULW

[Jac84a] Robert J. K. Jacob. User-level window manager for UNIX. In USENIX Association [USE84c], pages 123–134. ISBN none. LCCN QA76.8.U65 U55 1984.

Jacob:1984:UWM

[Jac84b] Robert J. K. Jacob. User-level window manager for UNIX. In USENIX Association [USE84c], pages 123–134. ISBN none. LCCN QA76.8.U65 U55 1984.

Jacobs:1986:UMP

[Jac86a] Herb Jacobs. A user-tunable multiple processor scheduler. In USENIX Association [USE86e], pages 183–191.

Jacobs:1986:UTM

[Jac86b] Herb Jacobs. A user-tunable multiple processor scheduler. In USENIX [USE86b], pages 183– 191.

Jacobson:1987:TUL

[Jac87]  Van Jacobson.  Tuning

UNIX Lex or it’s NOT true what they say about Lex. In USENIX Association [USE87g], pages 163–164. LCCN QA 76.76 O63 U84 1987.

URL https://www.tuhs.org/
Archive/Documentation/Papers/
TuningUnixLex_Jacobson_USENIX_
Winter_1987_pp163_164.pdf.

Abstract only.

Jaffee:1987:RMT

[Jaf87] Harris Jaffee. Restoring from multiple tape dumps. In USE-NIX Association [USE87d], pages 9–?? ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

Jameson:1988:OPR

[Jam88] David H. Jameson. ORE: Programming real-time applications. In Association [Ass88e], pages 23–26.

Johnston:1988:MOS

[JC88] Gary M. Johnston and Roy H. Campbell. A multiprocessor operating system simulator. In USENIX Association [USE88k], pages 169–182.

Johnston:1989:OID

[JC89a] Gary M. Johnston and Roy H. Campbell. An object-oriented implementation of distributed virtual memory. In USENIX Association [USE89d], pages 39–57. ISBN ???? LCCN ????

Johnston:1989:OOI

[JC89b] Gary M. Johnston and Roy H. Campbell. An object-oriented implementation of distributed virtual memory. In USENIX Association [USE89d], pages 39–57. ISBN ???? LCCN ????

Jenkins:1983:NLP

[Jen83] M. A. Jenkins. The NIAL language project. In Association [Ass83a], pages 331–332.

Johnston:1986:UBD

[JH86] William E. Johnston and Dennis E. Hall. UNIX based distributed printing in a diverse environment. In USENIX Association [USE86c], pages 514– 528.

Johnston:1986:LCV

[JHRR86] William E. Johnston, Dennis E. Hall, Fritz Renema, and David Robertson. A low cost, video based, animated movie system for the display of time dependent modeling results. In USENIX Association [USE86d], pages 91–115. ISBN ???? LCCN ????

Jung:1986:KUY

[JK86] Robert S. Jung and Joseph T.

Kalash. Kanji UNIX: Yunikkusu wa Nihongo o Hanasemasu (UNIX speaks Japanese). In USENIX Association [USE86c], pages 209–222.

Joy:1982:QA

[JLMM82] Bill Joy, Sam Leffler, Kirk McKusick, and David Mosher. 4.2BSD questions and answers. In Usr Group [Usr82], pages 32–33.

Jacobson:1984:UTI

[JLSG84] Van Jacobson, Craig Leres, Joseph Sventek, and Wayne Graves. 4BSD UNIX TCP/ IP and VMS DECNET: Experience in negotiating a peaceful coexistence. In Software Tools Users Group [Sof84], pages 323–325. LCCN QA76.8.U65 U83 1984. Abstract only.

Joyce:1988:RDU

[JN88] Jim Joyce and Bob Nystrom. Rescuing data in UNIX file systems (what to do after rm *). In USENIX Association [USE88j], pages 331–334. ISBN ???? LCCN ????

Johnson:1987:ULF

[Joh87] Stephen C. Johnson. UNIX: The language forms. In USE-NIX Association [USE87g], pages 16–20. LCCN QA 76.76 O63 U84 1987. Cover title: Conference proceedings. Spine title: Winter 1987 USENIX Association Conference proceedings, Washington, DC.

Johnson:1988:YMCa

[Joh88a] Stephen C. Johnson. Yacc meets C++. In USENIX Association [USE88e], pages 53–57. ISBN ???? LCCN ????

Johnson:1988:YMCb

[Joh88b] Stephen C. Johnson. Yacc meets C++. In Association [Ass88b], pages 159–167.

Joiret:1987:AUM

[Joi87]  Denis Joiret.  Administra-

tion of a Unix machine network. In USENIX Association [USE87d], pages 1–?? ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

Jones:1988:SAD

[Jon88] Von Jones. System administration daemons. In USENIX Association [USE88j], pages 137– 143. ISBN ???? LCCN ????

Joy:1982:ICP

[Joy82a] Bill Joy. 4.2BSD interprocess communications primer. In Usr Group [Usr82], pages 30–31. Abstract only.

Joy:1982:O

[Joy82b] Bill Joy. 4.2BSD overview. In Usr Group [Usr82], pages 29– 30. Abstract only.

Jones:1987:BSC

[JS87] Von Jones and David Schrodel.

Balancing security and convenience. In USENIX Association [USE87d], pages 5–6. ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

Jansen:1989:TAR

[JS89a] Pierre G. Jansen and Gerard J. M. Smit. Tumult-64: A real-time multi-processor system. In USENIX Association [USE89d], pages 193–203. ISBN ???? LCCN ????

Jansen:1989:TRT

[JS89b] Pierre G. Jansen and Gerard J. M. Smit. Tumult-64: A real-time multi-processor system. In USENIX Association [USE89d], pages 193–203. ISBN ???? LCCN ????

Johnson:1987:DRT

[JSW87] Thomas D. Johnson, Jonathan M.

Smith, and Eric S. Wilson. Disk response time measurements. In USENIX Association [USE87g], pages 147– 162. LCCN QA 76.76 O63 U84 1987. Cover title: Conference proceedings. Spine title: Winter 1987 USENIX Association Conference proceedings, Washington, DC.

Jaenson:1985:PAP

[JTUB85] Richard Jaenson, Gregory Taylor, Cyrus Umrigar, and Alison Brown. A parallel array processing environment under 4.2BSD UNIX. In USENIX Association [USE85e], pages 195– 208. LCCN QA76.8.U65 U8 1985.

Jung:1985:PAD

[Jun85a] Robert S. Jung. Porting the AT&T demand paged UNIX implementation to microcomputers. In USENIX [USE85a], pages 361–372.

Jung:1985:PAT

[Jun85b] Robert S. Jung. Porting the AT&T demand paged UNIX implementation to microcomputers. In USENIX Association [USE85e], pages 361–372. LCCN QA76.8.U65 U8 1985.

Juszczak:1989:IPC

[Jus89] Chet Juszczak. Improving the performance and correctness of an NFS server. In USENIX Association [USE89g], pages 53– 63.

Kojima:1983:AMI

[KAH83a] Tomihiko Kojima, Hidehiko Akita, and Hisashi Hashimoto.

An approach to a machine independent UNIX — UNIX on HI-TAC M-series virtual machines. In Software Tools Users Group [Sof83], pages 363–369. LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Kojima:1983:MIS

[KAH83b] Tomihiko Kojima, Hidehiko Akita, and Hisashi Hashimoto. A multiplexed interactive system PWB/II. In Software Tools Users Group [Sof83], pages 355–362. LCCN QA76.8.U65 U74 1983. Sponsored by USE-NIX Association in cooperation with Software Tools Users Group.

Kalish:1982:PMS

[Kal82] Douglas I. Kalish. Programdb:

Maintaining symbol use data for source code control. In Usr Group [Usr82], pages 92–?? Abstract only.

Karels:1983:IVS

[Kar83] Michael Karels. An implementation of the vfork system call for PDP-11 UNIX. In Association [Ass83a], pages 40–?? Abstract only.

Katz:1982:LSA

[Kat82a] Fred M. Katz. The logical softshell: A full-screen interface to UNIX. In Usr Group [Usr82], pages 92–?? Abstract only.

Katz:1982:LSF

[Kat82b] Fred M. Katz. The logical softshell: A full-screen interface to UNIX. In USENIX [USE82a], pages 92–?? Abstract only.

Katz:1982:TTC

[Kat82c] Fred M. Katz. Time and tuples: Concurrency control in LOGIX. In Usr Group [Usr82], pages 29–?? Abstract only.

Katsive:1983:UMC

[Kat83] Bob Katsive. UNIX markets and competition. In Association [Ass83a], pages 308–309. Abstract only.

Katz:1984:US

[Kat84] Lou Katz. USENET in the sky. ;login: the USENIX Association newsletter, 9(6):8– 10, December 1984. CODEN LOGNEM. ISSN 1044-6397.

Katseff:1989:DPM

[Kat89] Howard P. Katseff. Debugging and performance monitoring in HPC/VORX. In USENIX Association [USE89d], pages 255– 267. ISBN ???? LCCN ????

Kazar:1985:CAP

[Kaz85a] Michael Leon Kazar. Camphor: A programming environment for extensible systems. In USENIX Association [USE85e], pages 107–112. LCCN QA76.8.U65 U8 1985.

Kazar:1985:CPE

[Kaz85b] Michael Leon Kazar. Camphor: A programming environment for extensible systems. In USENIX [USE85a], pages 107– 112.

Kazar:1988:SCI

[Kaz88] Michael Leon Kazar. Synchronization and caching issues in the Andrew File System. In USENIX Association [USE88j], pages 27–36. ISBN ???? LCCN ????

Kaashoek:1989:EDD

[KBT89] M. Frans Kaashoek, Henri E.

Bal, and Andrew S. Tanenbaum. Experience with the distributed data structure paradigm in Linda. In USENIX Association [USE89d], pages 175–191. ISBN ???? LCCN ????

Keeffe:1988:STM

[Kee88] David Keeffe. Software tools for music or communications standard works! In USENIX Association [USE88e], pages 149– 156. ISBN ???? LCCN ????

Kelley:1989:MAD

[Kel89] Michael H. Kelley. Multiprocessor aspects of the DG/UX kernel. In USENIX Association [USE89g], pages 85–99.

Kendall:1983:BRC

[Ken83] Samuel C. Kendall. Bcc: Runtime checking for C programs. In Software Tools Users Group [Sof83], pages 5–16. LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Kepecs:1985:LPU

[Kep85] Jonathan Kepecs. Lightweight processes for UNIX implementation and applications. In USENIX Association [USE85e], pages 299–308. LCCN QA76.8.U65 U8 1985.

Kernighan:1983:BH

[Ker83] Brian W. Kernighan. A bit of history. In Software Tools Users Group [Sof83], pages 385–?? LCCN QA76.8.U65 U74 1983. Abstract only.

Kercheval:1984:RMS

[Ker84] Berry Kercheval. A reliable mail service for the UUCP net: Implementation status report. In Software Tools Users Group [Sof84], pages 374–?? LCCN QA76.8.U65 U83 1984. Abstract only.

Kermarrec:1988:OGD

[Ker88] Alain Kermarrec. An overview of the GOTHIX distributed system. In USENIX Association [USE88e], pages 69–78. ISBN ???? LCCN ????

Kint:1989:ADH

[KGL89] Richard W. Kint, Charles V. Gale, and Andrew B. Liwen. Administration of a dynamic heterogeneous network. In USENIX Association [USE89e], pages 59–67. ISBN ???? LCCN ????

Kao:1989:SIH

[KGTM89] Ping-Hui Kao, Bill Gates, Bruce Thompson, and Dale McCluskey. Support of the ISO-9660/HSG CD-ROM file system in HP-UX. In USENIX Association [USE89f], pages 189–202. LCCN QA 76.76 O63 U83 1989.

Kurihara:1982:APE

[KI82] Masatoshi Kurihara and Yukio Ikadai. Application programming environment on UNIX. In Usr Group [Usr82], pages 178– ?? Abstract only.

Killian:1984:PF

[Kil84] T. J. Killian. Processes as files.

In Software Tools Users Group [Sof84], pages 203–207. LCCN QA76.8.U65 U83 1984.

Kim:1987:EMM

[Kim87] Yoon W. Kim. Electronic mail maintenance/distribution. In USENIX Association [USE87d], pages 27–?? ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

King:1983:ICD

[Kin83] Laura L. King. The Informix commercial DBMS for UNIX. In Association [Ass83a], pages 245–?? Abstract only.

Kingston:1984:MTR

[Kin84] Douglas P. Kingston, III.

MMDFII: A technical review. In USENIX [USE84a], pages 32–41.

Kingston:1986:SDI

[Kin86] Doug Kingston. A study in digital image reduction or how to make small faces. EUUG Newsletter, 6(3):31–39, Winter 1986. CODEN EONLE8. ISSN 1011-4211.

Kirslis:1987:SWC

[Kir87] Peter A. Kirslis. A style for writing C++ classes. In USE-NIX Association [USE87a], pages 147–148. ISBN ???? LCCN ???? Extended Abstract only.

Kivolowitz:1984:OSM

[Kiv84] Perry S. Kivolowitz. Optical storage management under the UNIX operating system. In Software Tools Users Group [Sof84], pages 297–311. LCCN QA76.8.U65 U83 1984.

Kingsbury:1989:JPR

[KK89a] Brent A. Kingsbury and John T. Kline. Job and process recovery in a UNIX-based operating system. In USENIX Association [USE89g], pages 355– 364.

Korn:1989:DFS

[KK89b] David G. Korn and Eduardo Krell. The 3-D file system. In USENIX [USE89b], pages 147– 156.

Korn:1989:FS

[KK89c] David G. Korn and Eduardo Krell. The 3-D file system. In USENIX Association [USE89f], pages 147–156. LCCN QA 76.76 O63 U83 1989.

Kornatowski:1982:CDR

[KL82a] John Z. Kornatowski and Ivor Ladd. Current database research at the Computer Systems Research Group, University of Toronto. In Usr Group [Usr82], pages 150–?? Abstract only.

Kornatowski:1982:CSM

[KL82b] John Z. Kornatowski and Ivor Ladd. Current status of Mistress (version 2) and future plans. In Usr Group [Usr82], pages 198–?? Abstract only.

Kugler:1987:UAC

[KL87a] Hans-Jurgen Kugler and Barry Lynch. Uncle — A case study in constructing tools for the PCTE. In USENIX Association [USE87b], pages 123–130. ISBN ???? LCCN ????

Kugler:1987:UCS

[KL87b] Hans-Jurgen Kugler and Barry Lynch. Uncle — A case study in constructing tools for the PCTE. In USENIX Association [USE87b], pages 123–130. ISBN ???? LCCN ????

Koelbel:1989:EIM

[KLB89] Charles Koelbel, Fady Lamaa, and Bharat Bhargava. Efficient implementation of modularity in RAID. In USENIX Association [USE89d], pages 127–143. ISBN ???? LCCN ????

Klein:1985:CBP

[Kle85]  Daniel Klein.  A capabil-

ity based protection mechanism under Unix. In USENIX Association [USE85c], pages 152– 159.

Kleiman:1986:VAM

[Kle86] S. R. Kleiman. Vnodes: An architecture for multiple file system types in Sun UNIX. In USENIX Association [USE86c], pages 238–247.

Klein:1987:UUB

[Kle87] Daniel V. Klein. UBOAT — A Unix based on-line aid to tutorials. In USENIX Association [USE87b], pages 31–39. ISBN ???? LCCN ????

Klein:1989:CCU

[Kle89] Daniel V. Klein. A comparison of compiler utilization of instruction set architectures. In USENIX Association [USE89g], pages 313–329.

Kaufer:1988:SCI

[KLP88a] Stephen Kaufer, Russell Lopez, and Sesha Pratap. Saber-C — an interpreter-based programming environment for the C language. In USENIX Association [USE88j], pages 161–171. ISBN ???? LCCN ????

Kaufer:1988:SIP

[KLP88b] Stephen Kaufer, Russell Lopez, and Sesha Pratap. Saber-C — an interpreter-based programming environment for the C language. In Association [Ass88f], pages 161–171.

Kingston:1982:MDQ

[KM82] Douglas P. Kingston, III and Michael J. Muuss. The multiple device queuing system. In USENIX [USE82a], pages 63– 67.

Kridle:1983:PED

[KM83] Bob Kridle and Kirk McKusick. Performance effects of disk subsystem choices for Vax systems running 4.2BSD UNIX. In Software Tools Users Group [Sof83], pages 155–169. LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Kahrs:1984:ATT

[KM84a] Mark Kahrs and Lee Moore.

Adventures with typesetterindependent TROFF. In Software Tools Users Group [Sof84], pages 258–269. LCCN QA76.8.U65 U83 1984.

Kahrs:1984:ATI

[KM84b] Mark Kahrs and Lee Moore.

Adventures with typesetterindependent TROFF. In USE-NIX [USE84a], pages 258–269.

Kelly:1985:PIC

[KM85] Thomas J. Kelly and Allen McIntosh. A portable intermediate code optimizer for C. In USENIX Association [USE85e], pages 577–589. LCCN QA76.8.U65 U8 1985.

Karels:1986:NPM

[KM86] Michael J. Karels and Marshall Kirk McKusick. Network performance and management with 4.3BSD and IP/TCP. In USENIX Association [USE86c], pages 182–188.

Kogure:1987:USV

[KM87] Hiromichi Kogure and Richard McGowan. A UNIX System V STREAMS TTY implementation for multiple language processing. In USENIX Association [USE87f], pages 323–336.

Kahle:1988:UCM

[KNI88] Brewster U. Kahle, William A. Nesheim, and Marshall Isman. UNIX and the Connection Machine operating system. In USENIX Association [USE88i], pages 93–107. ISBN ???? LCCN ????

Kuwana:1988:MIS

[KNN88a] Eiji Kuwana, Hironobu Nagano, and Yuzou Nakamura. Man-machine interfaces for software development environments (HandS). In USENIX Association [USE88j], pages 309–323. ISBN ???? LCCN ????

Kuwana:1988:MMI

[KNN88b] Eiji Kuwana, Hironobu Nagano, and Yuzou Nakamura. Man-machine interfaces for software development environments (HandS). In Association [Ass88f], pages 309–323.

Knowles:1987:PMB

[Kno87] Frank Knowles. A partial model for a B-Level Unix. In USENIX Association [USE87f], pages 257–271.

Kodosky:1982:UEA

[Kod82a] Jeffrey L. Kodosky. UNIX etc.

at National instruments. In Usr Group [Usr82], pages 141–149.

Kodosky:1982:UEN

[Kod82b] Jeffrey L. Kodosky. UNIX etc.

at National instruments. In USENIX [USE82a], pages 141– 149.

Koenig:1984:ASD

[Koe84] Andrew Koenig. Automatic software distribution. In Software Tools Users Group [Sof84], pages 312–322. LCCN QA76.8.U65 U83 1984.

Koenig:1985:SPL

[Koe85] Andrew Koenig. The Snocone programming language. In USENIX Association [USE85e], pages 87–106. LCCN QA76.8.U65 U8 1985.

Koehler:1987:GRH

[Koe87a] Matt Koehler. GFS revisited or how I lived with four different local file systems. In USENIX Association [USE87f], pages 291–305.

Koenigsberg:1987:RRS

[Koe87b] Christopher Koenigsberg. Release of replicated software in the vice file system. In USENIX Association [USE87d], pages 14–15. ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

Koenig:1988:AAC

[Koe88] Andrew Koenig. Associative arrays in C++. In Association [Ass88f], pages 173–186.

Kolstad:1985:WG

[Kol85] Rob Kolstad. Whither the gurus. In USENIX Association [USE85c], page ?? Title listed, no text or abstract.

Kolstad:1986:HPE

[Kol86] Bob Kolstad. High performance enhancements of C-1 Unix. In USENIX Association [USE86e], pages 192–?? Title listed only, no paper or abstract.

Korn:1983:KAS

[Kor83a] David Korn. KSH — A shell programming language. In Software Tools Users Group [Sof83], pages 191–202. LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Korn:1983:KSP

[Kor83b] David Korn. KSH — A shell programming language. In Association [Ass83a], pages 191– 202.

Korty:1989:SLL

[Kor89a]  Joseph A. Korty.  Sema:

a lint-like tool for analyzing semaphore usage in a multithreaded UNIX kernel. In USE-NIX [USE89c], pages 113–123.

Korty:1989:SLT

[Kor89b] Joseph A. Korty. Sema: a Lint-like tool for analyzing semaphore usage in a multithreaded UNIX kernel. In USE-NIX Association [USE89g], pages 113–123.

Kasten:1985:DCB

[KR85] Vince Kasten and Paul Ruel.

Development of a compiler for the Bourne shell. In USENIX Association [USE85c], pages 52–58.

Kramer:1983:LLN

[Kra83] Steven M. Kramer. LINUS (Leading Into Noticeable UNIX Security). In Association [Ass83a], pages 143–?? Abstract only.

Kramer:1988:IAC

[Kra88a] Steven M. Kramer. On incorporating access control lists into the UNIX operating system. In USENIX Association [USE88g], pages 38– 48. LCCN QA76.8.U65 U55 1988(1)-1990(2)//.

Kramer:1988:RSP

[Kra88b] Steven M. Kramer. Retaining SUID programs in a secure UNIX. In Association [Ass88f], pages 107–118.

Kretsch:1983:CPE

[Kre83] D. J. Kretsch. C programming environment. In Association [Ass83a], pages 111–?? Abstract only.

Kridle:1984:NIT

[Kri84a] Bob Kridle. New 1/2-inch tape options and trade-offs for 4.2BSD UNIX on DEC VAX processors. In USENIX Association [USE84c], pages 168–182. ISBN none. LCCN QA76.8.U65 U55 1984.

Kridle:1984:NTO

[Kri84b] Bob Kridle. New 1/2-inch tape options and trade-offs for 4.2BSD UNIX on DEC VAX processors. In USENIX Association [USE84c], pages 168–182. ISBN none. LCCN QA76.8.U65 U55 1984.

Kristol:1986:FGP

[Kri86] David M. Kristol. Four generations of portable C compiler. In USENIX Association [USE86c], pages 335–343.

Kolstad:1983:MUN

[KSH83] Rob Kolstad and Karen Summers-Horton. Mapping the UUCP network. EUUG Newsletter, 3(4):34–39, Winter 1983. CODEN EONLE8. ISSN 1011-4211.

Kolstad:1984:MUN

[KSH84] Rob Kolstad and Karen Summers-Horton. Mapping the UUCP network. In USENIX Association [USE84c], pages 251–258. ISBN none. LCCN QA76.8.U65 U55 1984.

Kirslis:1988:ILB

[KT88a] Peter A. Kirslis and Robert B.

Terwilliger. Implementing a logic-based executable specification language. In USENIX [USE88a], pages 211–225.

Kirslis:1988:ILE

[KT88b] Peter A. Kirslis and Robert B.

Terwilliger. Implementing a logic-based executable specification language. In USENIX Association [USE88k], pages 211–225.

Kahane:1986:WHA

[KTS^{+}86a] Stephen N. Kahane, Stephen G.

Tolchin, Marvin J. Schneider, Debra W. Richmond, Patrick Barta, Margaret K. Ardolino, and Howard S. Goldberg. Windows in the hospital or A workstation-based inpatient clinical information system in the Johns Hopkins Hospital. In USENIX Association [USE86e], pages 45–61.

Kahane:1986:WHW

[KTS^{+}86b] Stephen N. Kahane, Stephen G.

Tolchin, Marvin J. Schneider, Debra W. Richmond, Patrick Barta, Margaret K. Ardolino, and Howard S. Goldberg. Windows in the hospital or A workstation-based inpatient clinical information system in the Johns Hopkins hospital. In USENIX [USE86b], pages 45–61.

Keizer:1983:ACK

[KTv83] Ed Keizer, Andrew S. Tanenbaum, and Hans van Staveren. The Amsterdam Compiler Kit. EUUG Newsletter, 3(4):29– 33, Winter 1983. CODEN EONLE8. ISSN 1011-4211.

Kucera:1989:MLS

[Kuc89] Julie Kucera. Making libc suitable for use by parallel programs. In USENIX Association [USE89d], pages 145–152. ISBN ???? LCCN ????

Korn:1985:SBM

[KV85] David G. Korn and Kiem-Phong Vo. In search of a better malloc. In USENIX Association [USE85e], pages 489–506. LCCN QA76.8.U65 U8 1985.

Kernighan:1989:PMP

[KV89] Brian W. Kernighan and Christopher J. Van Wyk. Page makeup by postprocessing text formatter output. In Association [Ass89b], pages 103–132.

Lockwood:1989:FCT

[LA89a] Paul Lockwood and Divyakant Agrawal. A fault-tolerant client-server transaction model. In USENIX Association [USE89h], pages 63–71. ISBN ???? LCCN ????

Lockwood:1989:FTC

[LA89b] Paul Lockwood and Divyakant Agrawal. A fault-tolerant client-server transaction model. In USENIX Association [USE89h], pages 63–71. ISBN ???? LCCN ????

Ladipo:1988:SOS

[Lad88a] Ola Ladipo. A subscriptionoriented software package update distribution system (SPUDS). In USENIX Association [USE88f], pages 75–77. ISBN ???? LCCN ????

Ladipo:1988:SSP

[Lad88b] Ola Ladipo. A subscriptionoriented software package update distribution system (SPUDS). In USENIX Association [USE88f], pages 75–77. ISBN ???? LCCN ????

Lindsley:1988:ISC

[LAKS88] Rick Lindsley, Seth Alford, Richard Kurschner, and Roger Southwick. Identifying security concerns. In USENIX Association [USE88g], pages 82– 83. LCCN QA76.8.U65 U55 1988(1)-1990(2)//. Abstract only.

Lamb:1983:TUS

[Lam83] J. Eli Lamb. Towards a UNIX system Ada programming support environment. In Association [Ass83a], pages 143–?? Abstract only.

Lankford:1984:USV

[Lan84] Jeffrey P. Lankford. UNIX system V and 4BSD performance. In Software Tools Users Group [Sof84], pages 228–236. LCCN QA76.8.U65 U83 1984.

Langston:1986:EEW

[Lan86] Peter S. Langston. (201) 644-2332 or Eedie & Eddie on the wire — an experiment in music generation. In USENIX Association [USE86c], pages 13–27.

Lauder:1981:MID

[Lau81a] Piers Lauder. MX — an indirect driver for multiplexing virtual “tty” lines. ;login: the USENIX Association newsletter, 6(3):4–6, March 1981. CO-DEN LOGNEM. ISSN 1044-6397.

Lauder:1981:SSW

[Lau81b] Piers Lauder. Share scheduling works! ;login: the USE-NIX Association newsletter, 6 (3):10–13, March 1981. CO-DEN LOGNEM. ISSN 1044-6397.

Lauder:1981:SSU

[Lau81c] Piers Lauder. SUN — the Sydney Unix net. ;login: the USE-NIX Association newsletter, 6 (3):7–9, March 1981. CODEN LOGNEM. ISSN 1044-6397.

Lauder:1985:DAA

[Lau85] Piers Lauder. Domain addressing in ACSnet. ;login: the USE-NIX Association newsletter, 10 (3):5–8, August 1985. CODEN LOGNEM. ISSN 1044-6397.

Lawson:1983:URA

[Law83a] Jim Lawson. UNIX research at Lucasfilms. In Association [Ass83a], pages 167–?? Abstract only.

Lawson:1983:URL

[Law83b] Jim Lawson. UNIX research at Lucasfilms. In USENIX [USE83a], pages 167–?? Abstract only.

Lee:1989:WBL

[LB89a] T. Paul Lee and R. E. Barkley.

A watermark-based lazy buddy system for kernel memory allocation. In USENIX [USE89b], pages 1–13.

Lee:1989:WLB

[LB89b] T. Paul Lee and R. E. Barkley.

A watermark-based lazy buddy system for kernel memory allocation. In USENIX Association [USE89f], pages 1–13. LCCN QA 76.76 O63 U83 1989.

Linton:1987:DII

[LC87a] Mark A. Linton and Paul R.

Calder. The design and implementation of InterViews. In USENIX Association [USE87a], pages 256–267. ISBN ???? LCCN ????

Liu:1987:PTM

[LC87b] Xing Liu and Patrick Conley. Program translation by manipulating abstract syntax trees. In USENIX Association [USE87a], pages 345–360. ISBN ???? LCCN ????

Lea:1988:LGC

[Lea88] Douglas Lea. libg++, the GNU C++ library. In USENIX Association [USE88k], pages 243–

Leach:1989:BRP

[Lea89] George W. Leach. Book review: Programming in C++. ;login: the USENIX Association newsletter, 14(5):9–??, September/October 1989. CODEN LOGNEM. ISSN 1044-6397.

Leblang:1987:SPA

[Leb87] David B. Leblang. Software performance analysis using call graphs and workstation graphics. In USENIX Association [USE87g], pages 124–132. LCCN QA 76.76 O63 U84 1987. Cover title: Conference proceedings. Spine title: Winter 1987 USENIX Association Conference proceedings, Washington, DC.

Leeper:1987:LML

[Lee87a]  Evelyn C. Leeper.  Login

management for large installations. In USENIX Association [USE87d], pages 35–?? ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

Leese:1987:DMF

[Lee87b] Ken Leese. Data management: A full-text information retrieval perspective. In USENIX Association [USE87g], pages 191–?? LCCN QA 76.76 O63 U84 1987. Abstract only.

Lee:1989:IUT

[Lee89] Geoffrey M. Lee. Integrating UNIX terminal services into a distributed operating system. In USENIX Association [USE89g], pages 29–42.

Leffler:1982:NC

[Lef82] Sam Leffler. 4.2BSD network communications. In Usr Group [Usr82], pages 31–?? Abstract only.

Laskodi:1988:UFS

[LEG88] Terry Laskodi, Bob Eifrig, and Jason Gait. A UNIX file system for a write-once optical disk. In Association [Ass88f], pages 51– 60.

Lennert:1986:DRP

[Len86a] David C. Lennert. Decreasing realtime process dispatch latency through kernel preemption. In USENIX Association [USE86c], pages 405–414.

Lennert:1986:SVC

[Len86b] David C. Lennert. A System V compatible implementation of 4.2BSD job control. In USENIX Association [USE86c], pages 459–474.

Lennert:1987:HWU

[Len87] Dave Lennert. How to write a UNIX daemon. ;login: the USENIX Association newsletter, 12(4):17–23, July/August 1987. CODEN LOGNEM. ISSN 1044-6397.

Lesk:1983:TDS

[Les83]  Michael Lesk.  Technology

driven software vs. psychology of users: An irresistible force meets an immovable object. In Software Tools Users Group [Sof83], pages 2–?? LCCN QA76.8.U65 U74 1983. Abstract only.

Lesk:1987:PVC

[Les87] Michael Lesk. Packets vs. circuits, in two centuries. EUUG Newsletter, 7(2):3–5, 1987. CO-DEN EONLE8. ISSN 1011-4211.

Lesk:1988:CUS

[Les88a] Michael Lesk. Can UNIX survive secret source code? In Association [Ass88b], pages 189– 199.

Lesk:1988:GII

[Les88b] Michael Lesk. GRAB — inverted indexes with low storage overhead. In Association [Ass88c], pages 207–220.

Lesk:1988:WMO

[Les88c] Michael Lesk. Word manipulation in online catalog searching: Using the UNIX system for library experiments. In USENIX Association [USE88e], pages 135–147. ISBN ???? LCCN ????

Lester:1988:CSM

[Les88d] Ken Lester. Computer security measures at Eastman Kodak, Product Software Engineering Department. In USENIX Association [USE88g], pages 84– 85. LCCN QA76.8.U65 U55 1988(1)-1990(2)//. Abstract only.

Levine:1983:IST

[Lev83] John R. Levine. Interactive system/three and the Intel data base processor. In Association [Ass83a], pages 229–236.

Lewis:1986:GAD

[Lew86a] Bob Lewis. Galadriel: A display list-based window manager. In USENIX Association [USE86e], pages 1–10.

Lewis:1986:GDL

[Lew86b] Bob Lewis. Galadriel: A display list-based window manager. In USENIX [USE86b], pages 1–10.

Lumpp:1989:CAC

[LFN^{+}89a] James E. Lumpp, Jr., Samuel A.

Fineberg, Wayne G. Nation, Thomas L. Casavant, Edward C. Bronson, Howard Jay Siegel, Pierre H. Pero, Thomas Schwederski, and Dan C. Marinescu. CAPS — A coding aid used with the PASM parallel processing system. In USENIX Association [USE89d], pages 269–288. ISBN ???? LCCN ????

Lumpp:1989:CCA

[LFN^{+}89b] James E. Lumpp, Jr., Samuel A.

Fineberg, Wayne G. Nation, Thomas L. Casavant, Edward C. Bronson, Howard Jay Siegel, Pierre H. Pero, Thomas Schwederski, and Dan C. Marinescu. CAPS — A coding aid used with the PASM parallel processing system. In USENIX Association [USE89d], pages 269–288. ISBN ???? LCCN ????

Lai:1988:SDA

[LG88] Nick Lai and Terence E. Gray. Strengthening discretionary access controls to inhibit Trojan horses and computer viruses. In Association [Ass88f], pages 275–286.

Lee:1988:SSP

[LGZ88] I. Lee, R. Gerber, and A. Zwarico. Specifying scheduling paradigms for time dependent processes. In Association [Ass88e], pages 7–11.

Look:1984:REU

[LH84a] Byron Look and Gary Ho. Real-time extensions to the UNIX operating system. In USENIX Association [USE84c], pages 293–299. ISBN none. LCCN QA76.8.U65 U55 1984.

Look:1984:RTE

[LH84b] Byron Look and Gary Ho. Real-time extensions to the UNIX operating system. In USENIX Association [USE84c], pages 293–299. ISBN none. LCCN QA76.8.U65 U55 1984.

Libes:1985:ULS

[Lib85a] Don Libes. User-level shared variables (in a hierarchical control environment). In USENIX [USE85a], pages 317–324.

Libes:1985:USV

[Lib85b] Don Libes. User-level shared variables (in a hierarchical control environment). In USENIX Association [USE85e], pages 317–324. LCCN QA76.8.U65 U8 1985.

Libes:1987:MPO

[Lib87] Don Libes. Multiple programs in one UNIX process. ;login: the USENIX Association newsletter, 12(4):7–13, July/August 1987. CODEN LOGNEM. ISSN 1044-6397.

Lilly:1988:ANP

[Lil88] Deb Lilly. Administration of network passwd files and NFS file access. In USENIX Association [USE88f], pages 3–5. ISBN ???? LCCN ????

Lindberg:1984:LIU

[Lin84] Jeff Lindberg. A layered implementation of the UNIX kernel on the HP9000 series 500 computers. In USENIX Association [USE84c], pages 183–194. ISBN none. LCCN QA76.8.U65 U55 1984.

Lindsley:1988:MYC

[Lin88a] Rick Lindsley. Making your console secure. In USENIX Association [USE88g], pages 61– 62. LCCN QA76.8.U65 U55 1988(1)-1990(2)//. Abstract only.

Lindsley:1988:SLS

[Lin88b] Rick Lindsley. Suggested levels of security. In USENIX Association [USE88g], pages 78– 81. LCCN QA76.8.U65 U55 1988(1)-1990(2)//.

Lions:1988:HIL

[Lio88] John Lions. Help! I’m losing my files! In USENIX Association [USE88e], pages 23–28. ISBN ???? LCCN ????

Litzkow:1987:RUT

[Lit87] Michael J. Litzkow. Remote Unix — turning idle workstations into cycle servers. In USENIX Association [USE87f], pages 381–384.

Lai:1984:MMU

[LJ84] Clara S. Lai and Chris Peer Johnson. Memory management units and the UNIX kernel. In Software Tools Users Group [Sof84], pages 208–213. LCCN QA76.8.U65 U83 1984.

Lucas:1982:EYW

[LK82] Brian Lucas and Mark Kampe. Everything you wanted to know about System III but Bell was afraid to tell you. In Usr Group [Usr82], pages 68–?? Abstract only.

Leffler:1984:MIP

[LKM84] Sam Leffler, Mike Karels, and M. Kirk McKusick. Measuring and improving the performance of 4.2BSD. In Software Tools Users Group [Sof84], pages 237–252. LCCN QA76.8.U65 U83 1984.

Lucas:1983:GOF

[LL83a] Brian Lucas and Heinz Lycklama. A general-purpose object-file format. In Software Tools Users Group [Sof83], pages 119–?? LCCN QA76.8.U65 U74 1983. Abstract only.

Lucas:1983:GPO

[LL83b] Brian Lucas and Heinz Lycklama. A general-purpose object-file format. In Association [Ass83a], pages 119–?? Abstract only.

Lee:1987:SPP

[LLM87] T. P. Lee, M. W. Luppi, and R. E. Menninger. Solving performance problems on a multiprocessor UNIX system. In USENIX Association [USE87f], pages 399–405.

Liu:1988:SHR

[LLS88] Jane W. S. Liu, Kwei-Jay Lin, and X. Song. Scheduling hard real-time transaction. In Association [Ass88e], pages 112– 116.

Loomis:1983:CAA

[LM83a] Jeff Loomis and Phil Mercurio.

Computer animation at UCSD. In Association [Ass83a], pages 261–267.

Loomis:1983:CAU

[LM83b] Jeff Loomis and Phil Mercurio. Computer animation at UCSD. In USENIX [USE83a], pages 261–267.

Langue:1988:PUL

[LM88a] Y. Langue and T. Muntean.

PARX: A UNIX-like operating system for transputer-based parallel supercomputers. In USENIX Association [USE88i], pages 109–120. ISBN ???? LCCN ????

Langue:1988:PUO

[LM88b] Y. Langue and T. Muntean. PARX: A UNIX-like operating system for Transputer-based parallel supercomputers. In USENIX Association [USE88i], pages 109–120. ISBN ???? LCCN ????

Lippman:1988:CRP

[LM88c] S. B. Lippman and B. E. Moo. C++: From research to practice. In USENIX Association [USE88k], pages 123–135.

Lin:1988:REP

[LN88] Kwei-Jay Lin and Swaminathan Natarajan. Refinement and enhancement: Primitives for monotonic computations. In Association [Ass88e], pages 27– 31.

Lawson:1985:GOP

[LNSZ85] Robert P. Lawson, Avi Naiman, David Slocombe, and Mathew Zaleski. Geritol for old programs or Troff’s got a lot of life in it yet! In USENIX Association [USE85c], pages 165–169.

Locanthi:1987:FBA

[Loc87] Bart N. Locanthi. Fast bitblt() with asm() and cpp. In USE-NIX Association [USE87b], pages 243–259. ISBN ???? LCCN ????

Long:1982:FC

[Lon82a] Tim Long. Formatting C.

EUUG Newsletter, 2(4):73– 85, Winter 1982. CODEN EONLE8. ISSN 1011-4211.

Long:1982:PCC

[Lon82b] Tim Long. Proposed changes to C. EUUG Newsletter, 2(4): 70–72, Winter 1982. CODEN EONLE8. ISSN 1011-4211.

Leffler:1988:TBD

[LOR88a] Samuel J. Leffler, Eben F. Ostby, and William T. Reeves. A tool-based 3-D modeling and animation workstation. In USENIX Association [USE88e], pages 29–40. ISBN ???? LCCN ????

Leffler:1988:TMA

[LOR88b] Samuel J. Leffler, Eben F. Ostby, and William T. Reeves. A tool-based 3-D modeling and animation workstation. In USENIX Association [USE88e], pages 29–40. ISBN ???? LCCN ????

Lord:1988:TPH

[Lor88c] Thomas Lord. Tools and policies for the hierarchical management of source code development. In Association [Ass88f], pages 95–106.

Louden:1982:STB

[Lou82] Bill Louden. Software tools bulletin board. In Usr Group [Usr82], pages 283–?? Abstract only.

Lanzillo:1989:IDI

[LP89] Leo Lanzillo and Craig Partridge. Implementation of dial-up IP for UNIX systems. In USENIX Association [USE89g], pages 201–207.

Linton:1987:DAP

[LQC87] Mark A. Linton, Russell W. Quong, and Paul R. Calder. The design of the allegro programming environment. In USENIX Association [USE87a], pages 268–273. ISBN ???? LCCN ????

Lee:1984:OPC

[LR84] Kok-Weng Lee and Mario D. Ruggiero. An optimizing portable C compiler for the new CDC CYBER 180. In Software Tools Users Group [Sof84], pages 100–109. LCCN QA76.8.U65 U83 1984.

Lemke:1989:VXC

[LR89] David Lemke and David S. H. Rosenthal. Visualizing X11 clients. In USENIX Association [USE89g], pages 125–138.

Lutz:1983:RUK

[LS83a] Michael Lutz and Michael Shon. Running the UNIX kernel in user mode. In Software Tools Users Group [Sof83], pages 171–176. LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Lyon:1983:SWW

[LS83b] Tom Lyon and Bill Shannon.

4.2BSD on the Sun workstation (or what we did on our summer vacation) (or how to emulate a VAX on a 68000). In Association [Ass83a], pages 132– ?? Abstract only.

Lyon:1985:ACF

[LS85] Tom Lyon and Joseph Skudlarek. All the chips that fit. In USENIX Association [USE85e], pages 557–561. LCCN QA76.8.U65 U8 1985.

Lippman:1988:PCM

[LS88] S. B. Lippman and B. Stroustrup. Pointers to class members in C++. In USENIX Association [USE88k], pages 305–323.

Luppi:1988:BET

[LSC^{+}88] Mark Luppi, Mark Seiden, Joseph Collins, Daniel Fisher, Keith Iverson, Charles Marshall, Josef Sachs, and David Shaw. Building an equities trading system in a distributed UNIX environment. In USENIX Association [USE88j], pages 97–104. ISBN ???? LCCN ????

Lea:1989:IAA

[LW89a] Rodger Lea and Jonathan Walpole. The implementation of Aide: A support environment for distributed objectoriented systems. In USENIX Association [USE89d], pages 289–299. ISBN ???? LCCN ????

Lea:1989:IAS

[LW89b] Rodger Lea and Jonathan Walpole. The implementation of aide: A support environment for distributed object-oriented systems. In USENIX Association [USE89d], pages 289–299. ISBN ???? LCCN ????

Lycklama:1983:UGS

[Lyc83a] Heinz Lycklama. The /usr/ group standards committee. In Association [Ass83a], pages 335–347.

Lycklama:1983:SRU

[Lyc83b] Heinz Lycklama. Status report from the /usr/group standards committee. In Software Tools Users Group [Sof83], pages 353–?? LCCN QA76.8.U65 U74 1983. Abstract only.

Lycklama:1984:UGS

[Lyc84] Heinz Lycklama. /usr/group standards effort. In USENIX Association [USE84c], pages 231–244. ISBN none. LCCN QA76.8.U65 U55 1984.

Lycklama:1985:UMY

[Lyc85] Heinz Lycklama. UNIX on a microprocessor — 10 years later. In USENIX Association [USE85e], pages 5–16. LCCN QA76.8.U65 U8 1985.

Lycklama:1982:FPS

[LZ82] Heinz Lycklama and Steve Zucker. A family of portable systems based on System III. In Usr Group [Usr82], pages 198– ?? Abstract only.

McCormack:1988:UXT

[MA88] Joel McCormack and Paul Asente. Using the X toolkit or how to write a widget. In Association [Ass88f], pages 1–13.

URL ftp://aeneas.mit.edu/
pub/usenix/xtk.PS;   ftp:
//ftp.uu.net/networking/
athena/usenix/xtk.PS.Z. FTP

- aeneas.mit.edu:/pub/usenix/xtk.PS; local - xtk.ps.

McGowan:1983:MMM

[MAB83] Martin J. McGowan, III, William L. Anderson, and Allen H. Brumm. Mm4 — make with M4 for maintaining makefiles. In Association [Ass83a], pages 59–?? Abstract only.

Mackay:1983:TIP

[Mac83a]  Don Mackay.  Terminal-

independent plotting packages: An example and suggestions for standards. In USENIX [USE83a], pages 251–255.

Mackay:1983:TPP

[Mac83b]  Don Mackay.  Terminal-

independent plotting packages: An example and suggestions for standards. In Association [Ass83a], pages 251–255.

Manis:1983:RAR

[Man83a] Rod Manis. /rdb: A relational data base management system. In Association [Ass83a], pages 237–240.

Manis:1983:RRD

[Man83b] Rod Manis. /rdb: A relational data base management system. In USENIX [USE83a], pages 237–240.

Maniago:1987:CMA

[Man87a] Pierette Maniago. Consulting via mail at Andrew. In USE-NIX Association [USE87d], pages 22–23. ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

Mann:1987:PEC

[Man87b] Daniel Mann. A preprocessor extending C++ to support rule based systems via access orientation. EUUG Newsletter, 7(4): 12–35, Winter 1987. CODEN EONLE8. ISSN 1011-4211.

Martin:1983:DD

[Mar83] Marlene Martin. Distribution and differentiation. In Association [Ass83a], pages 313–?? Abstract only.

Martin:1984:ASP

[Mar84] Dave Martin. Avionics simulation package: A large systems application in Ratfor. In Software Tools Users Group [Sof84], pages 352–?? LCCN QA76.8.U65 U83 1984. Abstract only.

Mashey:1983:SAM

[Mas83a] John Mashey. Software army on the march — project strategies and tactics. In Association [Ass83a], pages 17–21.

Mashey:1983:SLS

[Mas83b] John R. Mashey. SOLID: for on-line systems development. In USENIX [USE83a], pages 333–334. Abstract only.

Mashey:1983:SOS

[Mas83c] John R. Mashey. SOLID: for on-line systems development. In Association [Ass83a], pages 333–334. Abstract only.

Mashey:1987:ULP

[Mas87]  John R. Mashey.  UNIX:

Leverage — past, present, future. In USENIX Association [USE87g], pages 1–8. LCCN QA 76.76 O63 U84 1987. Cover title: Conference proceedings. Spine title: Winter 1987 USE-NIX Association Conference proceedings, Washington, DC.

Mashey:1988:UPP

[Mas88] John R. Mashey. UNIX past, present, and future: Changing roles, changing technologies. In USENIX Association [USE88e], pages 7–13. ISBN ???? LCCN ????

Martin:1989:EP

[MBBP89] Bruce Martin, Charles Bergan, Walter Burkhard, and Jehan-Francois Paris. Experience with PARPC. In USENIX Association [USE89g], pages 1–12.

Myers:1986:PAF

[MBS86a] Rob Myers, Peter Broadwell, and Robin Schaufler. Plasm: A fish sample. In USENIX Association [USE86d], pages 37–39. ISBN ???? LCCN ????

Myers:1986:PFS

[MBS86b] Rob Myers, Peter Broadwell, and Robin Schaufler. Plasm: A fish sample. In USENIX Association [USE86d], pages 37–39. ISBN ???? LCCN ????

Mosher:1982:FP

[MC82] David A. Mosher and Robert P. Corbett. F77 performance. ;login: the USENIX Association newsletter, 7(3):9–14, June 1982. CODEN LOGNEM. ISSN 1044-6397.

Mahler:1985:AAP

[MC85a] Stephen J. Mahler and David A. Curry. Access — A program to interpret pathname access permissions for the UNIX operating system. In USENIX Association [USE85c], pages 59–64.

Mahler:1985:API

[MC85b] Stephen J. Mahler and David A. Curry. Access — A program to interpret pathname access permissions for the UNIX operating system. In USENIX [USE85b], pages 59–64.

McCarron:1988:UUS

[McC88] Shane P. McCarron. An update on UNIX standards activities. ;login: the USENIX Association newsletter, 13(5):18–22, September/October 1988. CO-DEN LOGNEM. ISSN 1044-6397.

McDowell:1984:UBC

[McD84a] Rex McDowell. A UNIX-Based color graphics workstation. In USENIX Association [USE84c], pages 115–122. ISBN none. LCCN QA76.8.U65 U55 1984.

McDowell:1984:UCG

[McD84b] Rex McDowell. A UNIX-Based color graphics workstation. In USENIX Association [USE84c], pages 115–122. ISBN none. LCCN QA76.8.U65 U55 1984.

McDonell:1987:TPE

[McD87] Ken J. McDonell. Taking performance evaluation out of the “stone” age. In USENIX Association [USE87f], pages 407– 417.

McGeady:1985:WMO

[McG85] S. McGeady. Window managers are operating systems: Software for a distributed graphics system. ;login: the USENIX Association newsletter, 10(4):23–34, October/ November 1985. CODEN LOGNEM. ISSN 1044-6397.

McGeady:1986:NGH

[McG86a] S. McGeady. Next-generation hardware for windowed displays. In USENIX [USE86b], pages 11–22.

McGeady:1986:NHW

[McG86b] S. McGeady. Next-generation hardware for windowed displays. In USENIX Association [USE86e], pages 11–22.

McIvor:1987:UBE

[McI87] Alan McIvor. UTek build environment. In USENIX Association [USE87f], pages 437–443.

McIlroy:1989:V

[McI89] M. Douglas McIlroy. Virology 101. In Association [Ass89b], pages 173–181.

McKee:1982:CUC

[McK82a] Roger McKee. The coming UNIX crash. In Usr Group [Usr82], pages 281–??

McKusick:1982:FS

[McK82b] Kirk McKusick. 4.2BSD file system. In Usr Group [Usr82], pages 31–?? Abstract only.

McKie:1983:WE

[McK83a] Jim McKie. Where is europe? In Association [Ass83a], pages 323–326.

McKie:1983:WIE

[McK83b] Jim McKie. Where is Europe? In Software Tools Users Group [Sof83], pages 323–326. LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

McKusick:1983:GAC

[McK83c] Marshall Kirk McKusick. gprof: A call graph execution profiler. In Software Tools Users Group [Sof83], pages 81–88. LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

McKusick:1983:GCG

[McK83d] Marshall Kirk McKusick. gprof:

A call graph execution profiler. In Association [Ass83a], pages 81–88.

McKenney:1988:BSN

[McK88a] Paul E. McKenney. Broadcast storms, nervous hosts, and load imbalances. ;login: the USENIX Association newsletter, 13(5):9–17, September/October 1988. CODEN LOGNEM. ISSN 1044-6397.

McKenney:1988:CNA

[McK88b] Paul E. McKenney. Charge number accounting without kernel modifications. ;login: the USENIX Association newsletter, 13(4):9–12, July/August 1988. CODEN LOGNEM. ISSN 1044-6397.

McLaren:1983:UDG

[McL83]  Wayne McLaren.  UNIX

à la Data General. In Software Tools Users Group [Sof83], pages 79–?? LCCN QA76.8.U65 U74 1983. Abstract only.

McLeod:1984:IPU

[McL84] Bubette McLeod. Introducing people to UNIX. In Software Tools Users Group [Sof84], pages 344–345. LCCN QA76.8.U65 U83 1984.

Murrel:1987:IWT

[MD87] S. L. Murrel and D. De Baer.

An interactive WYSIWYG table editor. In USENIX Association [USE87f], pages 19–29.

Manas:1988:DES

[MdM88] Jose A. Manas and Tomas de Miguel. Design of and experience with a software documentation tool. In USENIX Association [USE88e], pages 105– 115. ISBN ???? LCCN ????

Mee:1985:OA

[Mee85] Gary Mee. OEM application. In USENIX Association [USE85e], pages 188–190. LCCN QA76.8.U65 U8 1985.

Meine:1984:UST

[Mei84] Bill Meine. An update on the software tools standards effort. In Software Tools Users Group [Sof84], pages 352–?? LCCN QA76.8.U65 U83 1984. Abstract only.

Mercurio:1982:UMS

[Mer82] Philip J. Mercurio. The UCSD MSG system: Iterative design in the UNIX environment. In Usr Group [Usr82], pages 151– 163.

Meyer:1982:CSC

[Mey82a] Dennis F. Meyer. C style and coding standards. EUUG Newsletter, 2(4):86–106, Winter 1982. CODEN EONLE8. ISSN 1011-4211.

Meyer:1982:ODQ

[Mey82b] Dennis F. Meyer. Optimizing database queries in SQL. In Usr Group [Usr82], pages 16– 17. Abstract only.

Meyers:1985:MSP

[Mey85] William J. Meyers. Monitoring system and process performance. In USENIX Association [USE85c], pages 69–77.

Meyer:1988:DMF

[Mey88]  Andreas Meyer.  Directly

mapped files. In USENIX Association [USE88e], pages 231–

236. ISBN ???? LCCN ????

Mankins:1983:SWM

[MF83] David Mankins and Daniel Franklin. A simple window management facility for the UNIX timesharing system. In Software Tools Users Group [Sof83], pages 203–228. LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Minnich:1989:MSD

[MF89] Ronald G. Minnich and David J.

Farber. The Mether system: Distributed shared memory for SunOS 4.0. In USENIX Association [USE89f], pages 51–

60. LCCN QA 76.76 O63 U83 1989. URL ftp://

metropolis.super.org/pub/
mether/usenix.ps; local -
--   mether.ps.  FTP -

metropolis.super.org:/pub/mether/usenix.ps; local - mether.ps.

Miller:1989:ESR

[MFS89] Barton P. Miller, Lars Fredriksen, and Bryan So. An empirical study of the reliability of operating system utilities. In USENIX Association [USE89j], pages 59–76.

Marcie:1988:GPT

[MH88] S. G. Marcie and R. L. Holt. General purpose transaction support features for the UNIX operating system. In USENIX Association [USE88e], pages 179–184. ISBN ???? LCCN ????

Midden:1987:ACS

[Mid87] Marshall M. Midden. Academic Computing Services and Systems (ACSS). In USENIX Association [USE87d], pages 30– 31. ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

Miller:1984:CUS

[Mil84a] Joaquin Miller. Connecting a UNIX system to an X.25 network. In USENIX Association [USE84c], pages 47–56. ISBN none. LCCN QA76.8.U65 U55 1984.

Miller:1984:DPV

[Mil84b] Richard Miller. A demand paging virtual memory manager for System V. In Software Tools Users Group [Sof84], pages 178–182. LCCN QA76.8.U65 U83 1984.

Mills:1986:MMB

[Mil86] Philip M. Mills. A multiuser multiprocessor benchmark to compare UNIX systems. In USENIX Association [USE86c], pages 59–71.

Milne:1987:AMB

[Mil87] Allan C. Milne. The analysis and manipulation of BNF definitions. In USENIX Association [USE87b], pages 105–122. ISBN ???? LCCN ????

Miller:1988:EHL

[Mil88] William M. Miller. Exception handling without language extensions. In USENIX Association [USE88k], pages 327–341.

Miller:1989:CSM

[Mil89] Dale Miller. Controlling software for multiple projects. In USENIX Association [USE89j], pages 39–50.

Minter:1982:HCS

[Min82a]  Charles Minter.  A high-

performance computer system suited to UNIX. In Usr Group [Usr82], pages 107–?? Abstract only.

Minter:1982:HPC

[Min82b]  Charles Minter.  A high-

performance computer system suited to UNIX. In USENIX [USE82a], pages 107–?? Abstract only.

Miya:1986:URU

[Miy86] E. N. Miya. User requirements for UNIX on “big iron”. In USENIX Association [USE86e], pages 104–109.

Miya:1988:SOC

[Miy88] E. N. Miya. Some observations on computer performance characterization: Supercomputer and mini-supercomputer clocks and compilers. In USENIX Association [USE88i], pages 51– 66. ISBN ???? LCCN ????

Matthews:1984:F

[MK84a] Manton Matthews and Yogeesh Kamath. The FP-Shell. In Software Tools Users Group [Sof84], pages 133–140. LCCN QA76.8.U65 U83 1984.

Matthews:1984:FS

[MK84b] Manton Matthews and Yogeesh Kamath. The FP-Shell. In USENIX [USE84a], pages 133– 140.

Mankovich:1985:PUV

[MK85a] James E. Mankovich and Robert B. Kolstad. Porting the 4.2BSD UNIX virtual memory subsystem. In USENIX Association [USE85c], pages 4–10.

McKusick:1985:PIF

[MK85b] M. Kirk McKusick and Mike Karels. Performance improvements and functional enhancements in 4.3BSD. In USENIX Association [USE85e], pages 519–531. LCCN QA76.8.U65 U8 1985.

McKusick:1988:DGP

[MK88] Marshall Kirk McKusick and Michael J. Karels. Design of a general purpose memory allocator for the 4.3BSD UNIX kernel. In Association [Ass88f], pages 295–303.

McKusick:1989:RE

[MKB89] Marshall Kirk McKusick, Michael J.

Karels, and Keith Bostic. The release engineering of 4.3BSD. In USENIX Association [USE89j], pages 95–100.

Mahler:1988:TSC

[ML88] Axel Mahler and Andreas Lampen. A toolkit for software configuration management. In USENIX Association [USE88e], pages 185–202. ISBN ???? LCCN ????

Madany:1988:CCH

[MLRC88] Peter W. Madany, Douglas E.

Leyens, Vincent F. Russo, and Roy H. Campbell. A C++ class hierarchy for building UNIX-Like file systems. In USENIX Association [USE88k], pages 65–79.

Macklem:1988:GSE

[MLS88] Rick Macklem, Jim Linders, and Hugh Smith. G shell environment. In Association [Ass88f], pages 15–22.

Mitchell:1988:ISC

[MMTW88] Mike Mitchell, Kent Moat, Tom Truscott, and Bob Warren. Invoking system calls from within the UNIX kernel. In USENIX Association [USE88j], pages 277–282. ISBN ???? LCCN ????

Matthews:1985:LAP

[MN85a] Manton M. Matthews and Ted Nolan. LEVI: A prototype active assistance interface. In USENIX Association [USE85e], pages 325–331. LCCN QA76.8.U65 U8 1985.

Matthews:1985:LPA

[MN85b] Manton M. Matthews and Ted Nolan. LEVI: A prototype active assistance interface. In USENIX [USE85a], pages 325– 331.

Mogul:1989:SFD

[Mog89]  Jeffrey C. Mogul.  Simple

and flexible datagram access controls for Unix-based gateways. In USENIX Association [USE89f], pages 203–221. LCCN QA 76.76 O63 U83 1989.

Mok:1988:TMT

[Mok88] Aloysius K. Mok. Task management techniques for enforcing ED scheduling on periodic task set. In Association [Ass88e], pages 42–46.

Moran:1988:SVM

[Mor88a] Joseph P. Moran. SunOS virtual memory implementation. In USENIX Association [USE88e], pages 285–300. ISBN ???? LCCN ????

Morris:1988:MTD

[Mor88b] James H. Morris. “make or take” decisions in andrew. In Association [Ass88f], pages 1–8.

Morris:1988:TDA

[Mor88c] James H. Morris. “make or take” decisions in Andrew. In USENIX Association [USE88j], pages 1–8. ISBN ???? LCCN ????

Morris:1988:AUA

[Mor88d] Robert Morris. Adventures in UNIX arithmetic. In USENIX Association [USE88e], pages 157–159. ISBN ???? LCCN ????

Morris:1988:UAU

[Mor88e] Robert A. Morris. An unorthodox approach to undergraduate software engineering instruction. In Association [Ass88a], pages 405–419.

Mosher:1982:L

[Mos82] David Mosher. 4.2BSD licensing. In Usr Group [Usr82], pages 32–?? Abstract only.

Moyer:1983:BAB

[Moy83a] James A. Moyer. BIBFIND — A bibliographic retrieval system. In Association [Ass83a], pages 63–67.

Moyer:1983:BBR

[Moy83b] James A. Moyer. BIBFIND — A bibliographic retrieval system. In USENIX [USE83a], pages 63–67.

Meertens:1984:IBP

[MP84] Lambert Meertens and Steven Pemberton. An implementation of the B programming language. In USENIX Association [USE84c], pages 65–74. ISBN none. LCCN QA76.8.U65 U55 1984.

Massalin:1989:FGS

[MP89a] Henry Massalin and Calton Pu.

Fine-grain scheduling. In USE-NIX Association [USE89d], pages 91–104. ISBN ???? LCCN ????

Massalin:1989:FS

[MP89b] Henry Massalin and Calton Pu. Fine-grain scheduling. In USE-NIX Association [USE89d], pages 91–104. ISBN ???? LCCN ????

McIlroy:1988:MSFa

[MR88a] M. D. McIlroy and J. A. Reeds. Multilevel security with fewer fetters. In USENIX Association [USE88e], pages 117–122. ISBN ???? LCCN ????

McIlroy:1988:MSFb

[MR88b] M. D. McIlroy and J. A. Reeds. Multilevel security with fewer fetters. In USENIX Association [USE88g], pages 24– 31. LCCN QA76.8.U65 U55 1988(1)-1990(2)//.

McIlroy:1988:MWS

[MR88c] M. D. McIlroy and J. A. Reeds. Multilevel Windows on a single-level terminal. In USENIX Association [USE88g], pages 32– 34. LCCN QA76.8.U65 U55 1988(1)-1990(2)//. Describes a prototype of modifications to the Teletype 5620 and driving software to allow multi-level windows. Also explains some of the limitations of the method.

McSwain:1988:RCA

[MR88d] Jon McSwain and Tom Richardson. Real-time control of an autonomous land vehicle. In Association [Ass88e], pages 107– 111.

McSwain:1988:RTC

[MR88e] Jon McSwain and Tom Richardson. Real-time control of an autonomous land vehicle. In Association [Ass88e], pages 107– 111.

Montgomery:1989:FBH

[MR89] Ken Montgomery and Dan Reynolds. Filesystem backups in a heterogeneous environment. In USENIX Association [USE89e], pages 95–97. ISBN ???? LCCN ????

Mandelberg:1988:PMU

[MS88] K. I. Mandelberg and V. S.

Sunderam. Process migration in UNIX networks. In USENIX Association [USE88j], pages 357–363. ISBN ???? LCCN ????

McJones:1989:EUS

[MS89] Paul R. McJones and Garret F. Swart. Evolving the UNIX system interface to support multithreaded programs. In USENIX Association [USE89g], pages 393–404.

URL https://archive.org/
details/1989-conference-
proceedings-winter-usenixtech-
sd-full/page/393/mode/1up.

Muuss:1988:BBUb

[MSM88a] Michael John Muuss, Terry Slattery, and Donald F. Merritt. BUMP — the BRL/USNA migration project. In USENIX Association [USE88g], pages 183–214. LCCN QA76.8.U65 U55 1988(1)-1990(2)//.

Muuss:1988:BBUc

[MSM88b] Michael John Muuss, Terry Slattery, and Donald F. Merritt. BUMP — the BRL/ USNA migration project. In USENIX Association [USE88i], pages 183–214. ISBN ???? LCCN ????

Muuss:1988:BBUa

[MSM88c] Mike Muuss, Terry Slattery, and Don Merritt. BUMP — the BRL/USNA migration project. In USENIX Association [USE88f], pages 39–?? ISBN ???? LCCN ???? Abstract only.

Muuss:1988:BBUd

[MSM88d] Mike Muuss, Terry Slattery, and Don Merritt. BUMP — the BRL/USNA migration project. In USENIX Association [USE88f], pages 39–?? ISBN ???? LCCN ???? Abstract only.

Mankin:1989:LFP

[MT89] Allison Mankin and Kevin Thompson. Limiting factors in the performance of the slowstart TCP algorithms. In USE-NIX Association [USE89g], pages 219–228.

Mullen:1987:UNS

[Mul87]  John R. Mullen.  UNIX

and networking: A separate peace. In USENIX Association [USE87g], pages 21–?? LCCN QA 76.76 O63 U84 1987. Abstract only.

Murphy:1983:CU

[Mur83] Timothy Murphy. Circular UNIX. EUUG Newsletter, 3(3): 28–32, Autumn 1983. CODEN EONLE8. ISSN 1011-4211.

Murai:1988:JEa

[Mur88a] Jun Murai. The JUNET environment. EUUG Newsletter, 8 (2):3–12, Summer 1988. CO-DEN EONLE8. ISSN 1011-4211.

Murai:1988:JEb

[Mur88b] Jun Murai. The JUNET environment. In USENIX Association [USE88e], pages 41–?? ISBN ???? LCCN ???? Abstract only.

Murphy:1988:CEI

[Mur88c] R. L. Murphy. Comparing the efficiency of the Internet protocols to DECNET. In USENIX Association [USE88j], pages 105–110. ISBN ???? LCCN ????

Murray:1988:BWB

[Mur88d] R. B. Murray. Building wellbehaved type relationships in C++. In USENIX [USE88a], pages 19–30.

Murray:1988:BWT

[Mur88e] R. B. Murray. Building wellbehaved type relationships in C++. In USENIX Association [USE88k], pages 19–30.

Muuss:1987:RRS

[Muu87] Michael John Muuss. RT & REMRT: Shared memory parallel and network distributed ray-tracing programs. In USE-NIX Association [USE87c], pages 86–97. ISBN ???? LCCN ????

McNamara:1984:WDD

[MVB84] Jean McNamara, Paresh Vaish, and Richard N. Bryant. Writing device drivers for XENIX systems. In USENIX Association [USE84c], pages 195–208. ISBN none. LCCN QA76.8.U65 U55 1984.

Mills:1984:TDE

[MW84] Charles C. Mills and Anthony I. Wasserman. A transition diagram editor. In Software Tools Users Group [Sof84], pages 287–296. LCCN QA76.8.U65 U83 1984.

Myers:1986:PCD

[Mye86] Rob Myers. Pictorial conversation: Design considerations for interactive graphical media. In USENIX Association [USE86d], pages 17–35. ISBN ???? LCCN ????

Nachbar:1986:WNF

[Nac86] Daniel Nachbar. When network file systems aren’t enough: Automatic software distribution revisited. In USENIX Association [USE86c], pages 159–171.

Nachbar:1988:SAP

[Nac88a] Daniel Nachbar. SPIFF — A program for making controlled approximate comparisons of files. In Association [Ass88f], pages 73–84.

Nachbar:1988:SPM

[Nac88b] Daniel Nachbar. SPIFF — A program for making controlled approximate comparisons of files. In USENIX Association [USE88j], pages 73–84. ISBN ???? LCCN ????

Northlich:1984:ETI

[NB84] Bill Northlich and Bruce Borden. The Excelan TCP/IP protocol package. In USENIX Association [USE84c], pages 1–12. ISBN none. LCCN QA76.8.U65 U55 1984.

Neff:1983:VMM

[Nef83] Laura Neff. Virtual memory management in GENIX. In Software Tools Users Group [Sof83], pages 109–117. LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Neuendorffer:1986:GAT

[Neu86a] Thomas Neuendorffer. GLO — A tool for developing windowbased programs. In USENIX Association [USE86e], pages 34–44.

Neuendorffer:1986:GTD

[Neu86b] Thomas Neuendorffer. GLO — A tool for developing windowbased programs. In USENIX [USE86b], pages 34–44.

Neyer:1983:UTM

[Ney83a] James A. Neyer. UNIX timesharing menu-driven office system for terminals (UTMOST). In Association [Ass83a], pages 116–?? Abstract only.

Neyer:1983:UTS

[Ney83b] James A. Neyer. UNIX timesharing menu-driven office system for terminals (UTMOST). In USENIX [USE83a], pages 116–?? Abstract only.

Nowitz:1984:EIU

[NHR84] D. A. Nowitz, P. Honeyman, and B. E. Redman. Experimental implementation of UUCP — security aspects. In USENIX Association [USE84c], pages 245–250. ISBN none. LCCN QA76.8.U65 U55 1984.

Nicholson:1985:CDI

[Nic85] Robert T. Nicholson. The Clipboard Data Interchange Facility. In USENIX Association [USE85c], pages 131–135.

Nicklin:1989:EUH

[Nic89] Peter Nicklin. Experiences using a hypertext framework to manage software. In USENIX Association [USE89j], pages 125–136.

Nielsen:1984:EOB

[NLR84a] Erik Reeh Nielsen, Soren Lauesen, and Vilhelm Rosenqvist. An expandable object-based UNIX kernel. In USENIX [USE84a], pages 193–202.

Nielsen:1984:EOU

[NLR84b] Erik Reeh Nielsen, Soren Lauesen, and Vilhelm Rosenqvist. An expandable object-based UNIX kernel. In Software Tools Users Group [Sof84], pages 193–202. LCCN QA76.8.U65 U83 1984.

Nakamura:1983:LMS

[NM83a] Osamu Nakamura and Jun Murai. On-line manual system for software development on UNIX. In Association [Ass83a], pages 19–30.

Nakamura:1983:OMS

[NM83b] Osamu Nakamura and Jun Murai. On-line manual system for software development on UNIX. In Software Tools Users Group [Sof83], pages 19–30. LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Northlich:1982:EUP

[NMP82] William R. Northlich, Jr., T. D.

McCreery, and P. M. Powers. Embedding UNIX in a product (or, is ‘real-time’ real?). In Usr Group [Usr82], pages 1–14.

Neelands:1983:UN

[NMS83] Paul Neelands, Richard Miller, and Chris Sturgess. UNIX for the National 16032. In Association [Ass83a], pages 269–272. Extended abstract.

Nelson:1988:CS

[NO88a] Michael Nelson and John Ousterhout. Copy-on-write for Sprite. In Association [Ass88f], pages 187–201.

Nelson:1988:CWS

[NO88b] Michael Nelson and John Ousterhout. Copy-on-write for Sprite. In USENIX Association [USE88j], pages 187–201. ISBN ???? LCCN ????

Norred:1984:MPA

[Nor84] Michael Norred. Mine planning applications in Ratfor. In Software Tools Users Group [Sof84], pages 354–?? LCCN QA76.8.U65 U83 1984. Abstract only.

Norwood:1988:TUS

[Nor88] Earl W. Norwood, III. Transitioning users to a supported environment. In USENIX Association [USE88f], pages 45–46. ISBN ???? LCCN ????

Novak:1983:UME

[Nov83] Robert E. Novak. Using Make effectively. In Software Tools Users Group [Sof83], pages 59– ?? LCCN QA76.8.U65 U74 1983. Abstract only.

Neuman:1988:AUE

[NS88] B. Clifford Neuman and Jennifer G. Steiner. Authentication of unknown entities on an insecure network of untrusted workstations. In USENIX Association [USE88g], pages 10– 11. LCCN QA76.8.U65 U55 1988(1)-1990(2)//. Abstract only.

Nycum:1985:RLI

[NSB85] Susan Nycum, Gaston Snow, and Ely Bartlett. Research into liability issues in Netnews transmission. In USENIX Association [USE85c], page ?? Title listed, no text or abstract.

Neuman:1988:APR

[NY88] Clifford Neuman and Wayne Yamamoto. Adding packet radio to the Ultrix kernel. In USENIX Association [USE88j], pages 303–308. ISBN ???? LCCN ????

Nyberg:1986:ICA

[Nyb86] Karl A. Nyberg. Implementing Curses in Ada. In USENIX Association [USE86e], pages 314– 319.

OBrien:1982:PUP

[O’B82] Gregory J. O’Brien. Porting UNIX to a personal computer. In Usr Group [Usr82], pages 247–?? Abstract only.

OBrien:1985:AFM

[O’B85] Michael T. O’Brien. Automatic forwarding of mail in CSNET. In USENIX Association [USE85c], pages 191–194.

Ousterhout:1987:OSP

[OCD^{+}87] John Ousterhout, Andrew Cherenson, Fred Douglis, Michael Nelson, and Brent Welch. An overview of the Sprite project. ;login: the USE-NIX Association newsletter, 12 (1):13–17, January/February 1987. CODEN LOGNEM. ISSN 1044-6397.

ODell:1983:BUA

[O’D83a] Michael O’Dell. Berkeley UNIX after 4.2BSD: Where is it going and why do we want it to get there? In Software Tools Users Group [Sof83], pages 292–?? LCCN QA76.8.U65 U74 1983. Abstract only.

ODell:1983:UEM

[O’D83b] Michael D. O’Dell. UNIX and electronic mail: Trials, tribulations, and proposals. In Software Tools Users Group [Sof83], pages 327–?? LCCN QA76.8.U65 U74 1983. Abstract only.

ODell:1983:PUW

[O’D83c] Mike O’Dell. Portability in the UNIX world — what UNIX can learn from the software tools. In Association [Ass83a], pages 314–?? Abstract only.

ODell:1987:HAL

[O’D87a]  Michael D. O’Dell.  The

HUB: A lightweight object substrate. In USENIX Association [USE87b], pages 191–199. ISBN ???? LCCN ????

ODell:1987:HLO

[O’D87b]  Michael D. O’Dell.  The

HUB: A lightweight object substrate. In USENIX Association [USE87b], pages 191–199. ISBN ???? LCCN ????

ODell:1987:WTD

[O’D87c] Michael D. O’Dell. What they don’t tell you about window systems. In USENIX Association [USE87b], pages 11–16. ISBN ???? LCCN ????

ODell:1987:UWV

[O’D87d]  Mike O’Dell.  UNIX: The

world view. In USENIX Association [USE87g], pages 35– 45. LCCN QA 76.76 O63 U84 1987. Cover title: Conference proceedings. Spine title: Winter 1987 USENIX Association Conference proceedings, Washington, DC.

Ohkubo:1984:LKB

[Ohk84] Kiyoki Ohkubo. LIPs: Knowledge base development system. In Software Tools Users Group [Sof84], pages 151–158. LCCN QA76.8.U65 U83 1984.

Ostby:1985:SAS

[OK85a] Eben Ostby and Allan Kaplan. SM: A small mailer. In USENIX Association [USE85e], pages 539–546. LCCN QA76.8.U65 U8 1985.

Ostby:1985:SSM

[OK85b] Eben Ostby and Allan Kaplan.

SM: A small mailer. In USE-NIX [USE85a], pages 539–546.

OBrien:1984:CG

[OL84a] Michael T. O’Brien and Daniel B.

Long. CSNET grows up. In USENIX Association [USE84c], pages 13–16. ISBN none. LCCN QA76.8.U65 U55 1984.

OBrien:1984:CGU

[OL84b] Michael T. O’Brien and Daniel B. Long. CSNET grows up. In USENIX Association [USE84c], pages 13–16. ISBN none. LCCN QA76.8.U65 U55 1984.

Oldroyd:1988:PAS

[Old88a] Jim R. Oldroyd. POSIX — A standard interface. In USENIX Association [USE88e], pages 271–284. ISBN ???? LCCN ????

Oldroyd:1988:PSI

[Old88b] Jim R. Oldroyd. POSIX — A standard interface. In USENIX Association [USE88e], pages 271–284. ISBN ???? LCCN ????

Olson:1988:CAL

[OLJ^{+}88] S. Margaret Olson, Paul H. Levine, Stuart H. Jones, Stephanie Bodoff, and Stephen C. Bertrand. Concurrent access licensing. In Association [Ass88f], pages 287–294.

Olander:1986:FNS

[OMI86] David J. Olander, Gilbert J. McGrath, and Robert K. Israel. A framework for networking in system V. In USENIX Association [USE86c], pages 38–45.

Ondishko:1989:ADM

[Ond89]  Denise Ondishko.  Admin-

istration of department machines by a central group. In USENIX Association [USE89f], pages 73–82. LCCN QA 76.76 O63 U83 1989.

Opperman:1989:AHX

[Opp89a] Mark Opperman. At home with X11/NeWS Windows. In USENIX Association [USE89f], pages 119–131. LCCN QA 76.76 O63 U83 1989.

Opperman:1989:HXN

[Opp89b] Mark Opperman. At home with X11/NeWS Windows. In USE-NIX [USE89b], pages 119–131.

ORiordan:1988:DIC

[O’R88] Martin J. O’Riordan. Debugging and instrumentation of C++ programs. In USENIX Association [USE88k], pages 227–242.

Orr:1983:ICO

[Orr83]  Douglas Orr.  Imposing

character-oriented I/O on a record-oriented system. In Software Tools Users Group [Sof83], pages 418–?? LCCN QA76.8.U65 U74 1983. Abstract only.

Ortmeyer:1988:CAL

[Ort88] David Ortmeyer. Concurrent access licensing and NLS. In USENIX Association [USE88f], pages 73–74. ISBN ???? LCCN ????

Otillio:1988:CAR

[Oti88] Troy Otillio. C++ approach to real-time systems: Task interface library. In USENIX Association [USE88k], pages 257–

OToole:1985:IXP

[OTW85] James O’Toole, Chris Torek, and Mark Weiser. Implementing XNS protocols for 4.2BSD. In USENIX Association [USE85c], pages 90–97.

Pucci:1989:EEI

[PA89] Marc F. Pucci and J. L. Alberi. Experiences with efficient interprocess communication in DUNE. In USENIX Association [USE89d], pages 349–360. ISBN ???? LCCN ????

Panoff:1988:RPR

[Pan88] Robert M. Panoff. Real productivity for real science without real UNIX. In USENIX Association [USE88i], pages 35–?? ISBN ???? LCCN ???? Abstract only.

Partridge:1986:MRU

[Par86] Craig Partridge. Mail routing using domain names: An informal tour. In USENIX Association [USE86c], pages 366–376.

Partridge:1987:IRD

[Par87] Craig Partridge. Implementing the Reliable Data Protocol (RDP). In USENIX Association [USE87f], pages 367–379.

Pariser:1988:RSD

[Par88a] E. C. Pariser. Reduction of static and dynamic memory requirements on the Cray X-MP. In USENIX Association [USE88i], pages 169–182. ISBN ???? LCCN ????

Parseghian:1988:SIF

[Par88b] Patricia E. Parseghian. A simple incremental file backup system. In USENIX Association [USE88f], pages 41–42. ISBN ???? LCCN ????

Partridge:1988:UIH

[Par88c] Craig Partridge. A UNIX implementation of HEMS. In USENIX Association [USE88j], pages 89–96. ISBN ???? LCCN ????

Patriquin:1983:FSC

[Pat83] Ed Patriquin. File system considerations in a multiple processor UNIX environment. In Software Tools Users Group [Sof83], pages 118–?? LCCN QA76.8.U65 U74 1983. Abstract only.

Paxson:1984:LTV

[Pax84] Vern Paxson. A LEX tool for the VOS. In Software Tools Users Group [Sof84], pages 353–?? LCCN QA76.8.U65 U83 1984. Abstract only.

Polyak:1984:LUR

[PB84] Steven T. Polyak and Jeffrey S.

Barr. Life with UNIX in realtime. In USENIX Association [USE84c], pages 285–292. ISBN none. LCCN QA76.8.U65 U55 1984.

Probert:1986:SIH

[PBL86] Dave Probert, Jeff Berkowitz, and Mark Lucovsky. A straightforward implementation of 4.2BSD on a highperformance multiprocessor. In USENIX Association [USE86e], pages 141–156.

Peterson:1986:URT

[PBT86] John W. Peterson, Rod G. Bogart, and Spencer W. Thomas. The Utah Raster Toolkit. In USENIX Association [USE86d], pages 1–12. ISBN ???? LCCN ????

Peachey:1980:BDU

[Pea80] Darwyn Peachey. Buffer deadlock in UNIX. ;login: the USE-NIX Association newsletter, 5 (8):10–??, October 1980. CO-DEN LOGNEM. ISSN 1044-6397.

Peachey:1983:ISN

[Pea83] Darwyn Peachey. Improved schedulers for non-paged UNIX systems. In Association [Ass83a], pages 39–?? Abstract only.

Peacock:1988:CFF

[Pea88] J. Kent Peacock. The counterpoint fast file system. In USENIX Association [USE88j], pages 243–249. ISBN ???? LCCN ????

Perlman:1982:DAP

[Per82a] Gary Perlman. Data analysis programs on CSL UNIX. In Usr Group [Usr82], pages 211–224.

Perlman:1982:MIU

[Per82b] Gary Perlman. MENUNIX: An interface to UNIX files and programs. In Usr Group [Usr82], pages 225–234.

Perlman:1983:IAS

[Per83] Gary Perlman. The interface arsenal: Software tools for user-interface development. In Software Tools Users Group [Sof83], pages 177–?? LCCN QA76.8.U65 U74 1983. Abstract only.

Perlman:1985:OSC

[Per85] Gary Perlman. An overview of the SETOPT command line option parser generator. In USENIX Association [USE85c], pages 160–164.

Perkins:1987:MDP

[Per87a] Lawrence B. Perkins. Managing the development of performance-constrained UNIX-Based software on microcomputers. In USENIX Association [USE87g], pages 46–59. LCCN QA 76.76 O63 U84 1987. Cover title: Conference proceedings. Spine title: Winter 1987 USE-NIX Association Conference proceedings, Washington, DC.

Perry:1987:UNM

[Per87b] Rick Perry. Using news multicasting with UUCP. In USE-NIX Association [USE87d], pages 26–?? ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

Petersen:1983:HPS

[Pet83] Eric Petersen. The history and purpose of standards. In Association [Ass83a], pages 348–?? Brief description only.

Peterson:1987:DCC

[Pet87] John W. Peterson. Distributed computation for computer animation. In USENIX Association [USE87c], pages 24–36. ISBN ???? LCCN ????

Pawlowski:1984:DSL

[PF84a] Brian Pawlowski and Alan Filipski. The dynamics of a semilarge software project with specific reference to a UNIX system port. In USENIX [USE84a], pages 332–342.

Pawlowski:1984:DSS

[PF84b] Brian Pawlowski and Alan Filipski. The dynamics of a semilarge software project with specific reference to a UNIX system port. In Software Tools Users Group [Sof84], pages 332–342. LCCN QA76.8.U65 U83 1984.

Potrebic:1987:DBS

[PG87] Peter Potrebic and Phil Goldman. A debugger-based system for graphical display and editing of data structures. In USENIX Association [USE87f], pages 147–158.

Peacock:1988:BBA

[PG88a] Don Peacock and Mark Giuffrida. Big brother: A network services expert. In Association [Ass88f], pages 393–398.

Peacock:1988:BBN

[PG88b] Don Peacock and Mark Giuffrida. Big brother: A network services expert. In USENIX Association [USE88j], pages 393– 398. ISBN ???? LCCN ????

Phillips:1988:MMA

[PG88c] Gretchen Phillips and Don Gworek. Makealiases — a mail aliasing system. In USENIX Association [USE88f], pages 17–19. ISBN ???? LCCN ????

Phillips:1984:LIN

[Phi84]  Alex Phillips.  The Liv-

ermore Interactive Network Communication System. In Software Tools Users Group [Sof84], pages 98–99. LCCN QA76.8.U65 U83 1984. Abstract only.

Palay:1988:ATOa

[PHS^{+}88a] Andrew J. Palay, Wilfred J.

Hansen, Mark Sherman, Maria G. Wadlow, Thomas P. Neuendorffer, Zalman Stern, Miles Bader, and Thom Peters. The Andrew Toolkit — an overview. In USENIX Association [USE88j], pages 9–22. ISBN ???? LCCN ???? Abstract only.

Palay:1988:ATOb

[PHS^{+}88b] Andrew J. Palay, Wilfred J.

Hansen, Mark Sherman, Maria G. Wadlow, Thomas P. Neuendorffer, Zalman Stern, Miles Bader, and Thom Peters. The Andrew Toolkit — an overview. In USENIX Association [USE88e], pages 311–?? ISBN ???? LCCN ???? Abstract only.

Pickard:1983:PNU

[Pic83] Monte Pickard. The Plexus networked UNIX. In Association [Ass83a], pages 51–?? Abstract only.

Pike:1982:MBG

[Pik82] Rob Pike. Merging bitmap graphics and UNIX. In Usr Group [Usr82], pages 61–?? Abstract only.

Pike:1983:USC

[Pik83] Rob Pike. UNIX style, or cat -v considered harmful. In Software Tools Users Group [Sof83], pages 263–?? LCCN QA76.8.U65 U74 1983. Abstract only.

Pike:1984:TOT

[Pik84a] Rob Pike. A text-oriented terminal multiplexor for blits. In USENIX [USE84a], pages 173– ?? Abstract only.

Pike:1984:TTM

[Pik84b] Rob Pike. A text-oriented terminal multiplexor for blits. In Software Tools Users Group [Sof84], pages 173–?? LCCN QA76.8.U65 U83 1984. Abstract only.

Pike:1988:WSS

[Pik88] Rob Pike. Window systems should be transparent. In Association [Ass88c], pages 279–

Pike:1989:CWS

[Pik89] Rob Pike. A concurrent window system. In Association [Ass89b], pages 133–153.

Pittman:1989:RB

[Pit89] Jon H. Pittman. The render button. In USENIX Association [USE89i], pages 99–114. ISBN ???? LCCN ????

Phillipson:1989:PFS

[PJ89a] Steven H. Phillipson and Stefan Jeffers. Part-task flight simulation on a UNIX graphics workstation. In USENIX Association [USE89i], pages 115–128. ISBN ???? LCCN ????

Phillipson:1989:PTF

[PJ89b] Steven H. Phillipson and Stefan Jeffers. Part-task flight simulation on a UNIX graphics workstation. In USENIX Association [USE89i], pages 115–128. ISBN ???? LCCN ????

Polk:1988:FUD

[PK88] Jeff Polk and Rob Kolstad. A faster UNIX dump program. In USENIX Association [USE88j], pages 125–129. ISBN ???? LCCN ????

Pleasant:1989:TAD

[PL89] Melvin J. Pleasant, Jr. and Eliot Lear. Transcending administrative domains by automating system management tasks in a large heterogeneous environment. In USENIX Association [USE89j], pages 51–58.

Placeway:1989:BDB

[Pla89] Paul W. Placeway. A better dump for BSD UNIX. In USENIX Association [USE89e], pages 99–107. ISBN ???? LCCN ????

Pato:1988:UAR

[PMD88] Joseph N. Pato, Elizabeth Martin, and Betsy Davis. A user account registration system for a large (heterogeneous) UNIX network. In USENIX Association [USE88j], pages 155–161. ISBN ???? LCCN ????

Pu:1988:SK

[PMI88] Calton Pu, Henry Massalin, and John Ioannidis. The synthesis kernel. In Association [Ass88d], pages 11–32.

Podolski:1982:UE

[Pod82] Zdravko Podolski. UNIFLEX evaluation. EUUG Newsletter, 2(4):40–44, Winter 1982. CO-DEN EONLE8. ISSN 1011-4211.

Poepping:1987:BRU

[Poe87]  Mark Poepping.  Backup

and restore for UNIX systems. In USENIX Association [USE87d], pages 10–11. ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

Poston:1988:HPF

[Pos88] Alan Poston. A High Performance File System for UNIX. In USENIX Association [USE88i], pages 215–226. ISBN ???? LCCN ????

Powers:1983:GOC

[Pow83] George Powers. A global optimizing C compiler. In Association [Ass83a], pages 151–166.

Powell:1984:UMS

[Pow84]  Michael L. Powell.  Using

Modula-2 for system programming with UNIX. In Software Tools Users Group [Sof84], pages 119–132. LCCN QA76.8.U65 U83 1984.

Pozgaj:1983:UCA

[Poz83] Steve Pozgaj. UNIX for the Computer Automation 4/95. In Association [Ass83a], pages 307–?? Abstract only.

Pike:1985:FN

[PP85] Rob Pike and David L.

Presotto. Face the nation. In USENIX Association [USE85e], pages 81–86. LCCN QA76.8.U65 U8 1985.

Presotto:1985:ICE

[PR85] D. L. Presotto and D. M.

Ritchie. Interprocess communication in the Eighth Edition Unix system. In USENIX Association [USE85e], pages 309– 316. LCCN QA76.8.U65 U8 1985.

Preston:1982:NP

[Pre82a] David J. Preston. News from Perkin-Elmer. In Usr Group [Usr82], pages 150–?? Abstract only.

Preston:1982:NPE

[Pre82b] David J. Preston. News from Perkin-Elmer. In USENIX [USE82a], pages 150–?? Abstract only.

Presotto:1985:USA

[Pre85] David L. Presotto. Upas — a simpler approach to network mail. In USENIX Association [USE85e], pages 533–538. LCCN QA76.8.U65 U8 1985.

Presotto:1988:PBL

[Pre88] David Leo Presotto. Plan 9 from Bell Labs — the network. In USENIX Association [USE88e], pages 15–21. ISBN ???? LCCN ????

Patel:1982:UEA

[PS82] Sanand Patel and Richard Sniderman. UNIX emulation, again. In Usr Group [Usr82], pages 248–?? Abstract only.

Phillips:1989:MHW

[PS89] Gretchen Phillips and Ken Smith. Mkuser — or how we keep the usernames straight. In USENIX Association [USE89e], pages 35–39. ISBN ???? LCCN ????

Pike:1985:HN

[PW85] Rob Pike and P. J. Weinberger. The hideous name. In USENIX Association [USE85e], pages 563–568. LCCN QA76.8.U65 U8 1985.

Pyne:1984:MUB

[PY84a] T. Scott Pyne and Joseph S. D.

Yao. MIPS: A UNIX-Based microcomputer message switching system. In USENIX Association [USE84c], pages 99–114. ISBN none. LCCN QA76.8.U65 U55 1984.

Pyne:1984:MUM

[PY84b] T. Scott Pyne and Joseph S. D.

Yao. MIPS: A UNIX-Based microcomputer message switching system. In USENIX Association [USE84c], pages 99–114. ISBN none. LCCN QA76.8.U65 U55 1984.

Pyne:1982:IOD

[Pyn82a] T. Scott Pyne. IAFORM, an on-screen definition package for data retrieval forms. In Usr Group [Usr82], pages 280–?? Abstract only.

Pyne:1982:ISD

[Pyn82b] T. Scott Pyne. IAFORM, an on-screen definition package for data retrieval forms. In USE-NIX [USE82a], pages 280–?? Abstract only.

Rozier:1988:CDO

[RAA^{+}88] M. Rozier, V. Abrossimov, F. Armand, I. Boule, M. Gien, M. Guillemont, F. Herrmann, C. Kaiser, S. Langlois, P. Leonard, and W. Neuhauser. CHORUS distributed operating systems. In Association [Ass88a], pages 305–370.

Rafter:1987:ECS

[Raf87] Mark Rafter. Extending C++ stream I/O to include formats. In USENIX Association [USE87a], pages 149–157. ISBN ???? LCCN ????

Rafter:1988:FC

[Raf88a] Mark Rafter. Formatted I/O in C++. In USENIX Association [USE88e], pages 247–253. ISBN ???? LCCN ????

Rafter:1988:FOC

[Raf88b] Mark Rafter. Formatted I/O in C++. In USENIX Association [USE88e], pages 247–253. ISBN ???? LCCN ????

Rago:1989:BCS

[Rag89a] Stephen Rago. Out-of-band communication in STREAMS. In USENIX [USE89b], pages 29–37.

Rago:1989:OCS

[Rag89b] Stephen Rago. Out-of-band communication in STREAMS. In USENIX Association [USE89f], pages 29–37. LCCN QA 76.76 O63 U83 1989.

Ralya:1988:ROS

[Ral88a] Tom Ralya. Real-time operating system architecture: Worksteps and related subjects. In Association [Ass88e], pages 87– 106.

Ralya:1988:RTO

[Ral88b] Tom Ralya. Real-time operating system architecture: Worksteps and related subjects. In Association [Ass88e], pages 87– 106.

Raves:1983:DDS

[RC83] William Raves and James Cassidy. Development of a digital simulation system in a UNIX environment. In Association [Ass83a], pages 169–176.

Rosenthal:1983:HCE

[RCB83] David S. H. Rosenthal, John M.

Collins, and Piet Beertema. How to connect to EUNET. EUUG Newsletter, 3(2):1– 6, Summer 1983. CODEN EONLE8. ISSN 1011-4211.

Redman:1985:WAY

[Red85] Brian E. Redman. Who answers your telephone when you’re in the information age? In USENIX Association [USE85e], pages 569–576. LCCN QA76.8.U65 U8 1985.

Reek:1981:MUS

[Ree81]  Kenneth A. Reek.  Modi-

fications for Unix on small CPU’s. ;login: the USE-NIX Association newsletter, 6 (3):14–16, March 1981. CO-DEN LOGNEM. ISSN 1044-6397.

Reeves:1982:UAL

[Ree82a] Bill Reeves. UNIX at Lucasfilm Ltd. or does Darth Vader code in C? In Usr Group [Usr82], pages 29–?? Abstract only.

Reeves:1982:ULL

[Ree82b] Bill Reeves. UNIX at Lucasfilm Ltd. or does darth Vader code in C? In USENIX [USE82a], pages 29–?? Abstract only.

Reilley:1983:CSR

[Rei83] G. Brendan Reilley. CSNET status report. In Association [Ass83a], pages 51–?? Abstract only.

Reid:1987:RAS

[Rei87a] Irving Reid. RPCC — A stub compiler for Sun RPC. In USENIX Association [USE87f], pages 357–366.

Reid:1987:RSC

[Rei87b] Irving Reid. RPCC — A stub compiler for Sun RPC. In USENIX Association [USE87f], pages 357–366.

Reisman:1989:DCM

[Rei89] Ron Reisman. Design considerations for multitasking, windowing, networked, multiplatform, distributed applications. In USENIX Association [USE89i]. ISBN ???? LCCN ???? Presentation only, no paper.

Renwick:1988:HNSa

[Ren88a] John Renwick. High-speed networking with supercomputers. In USENIX Association [USE88g], pages 67– ?? LCCN QA76.8.U65 U55 1988(1)-1990(2)//. Abstract only.

Renwick:1988:HNSb

[Ren88b] John Renwick. High-speed networking with supercomputers. In USENIX Association [USE88i], pages 19–24. ISBN ???? LCCN ???? Abstract only.

Renwick:1988:HSN

[Ren88c] John Renwick. High-speed networking with supercomputers. In USENIX Association [USE88i], pages 67–?? ISBN ???? LCCN ???? Abstract only.

Requa:1985:UKN

[Req85] Joseph E. Requa. UNIX kernel networking support and the LINCS communications architecture. In USENIX Association [USE85c], pages 98–103.

Rifkin:1986:RAOa

[RFH^{+}86a] Andrew P. Rifkin, Michael P.

Forbes, Richard L. Hamilton, Michael Sabrio, Suryakanta Shah, and Kang Yueh. RFS architectural overview. EUUG Newsletter, 6(2):13–23, Spring 1986. CODEN EONLE8. ISSN 1011-4211.

Rifkin:1986:RAOb

[RFH^{+}86b] Andrew P. Rifkin, Michael P.

Forbes, Richard L. Hamilton, Michael Sabrio, Suryakanta Shah, and Kang Yueh. RFS architectural overview. In USENIX Association [USE86c], pages 248–259.

Rowe:1988:ITU

[RGDP88] P. Kim Rowe, D. Graham, A. Donenfeld, and B. Pagurek. The Integration Toolkit and the Unison Real Time Operating System. In USENIX Association [USE88j], pages 347–356. ISBN ???? LCCN ????

Rosenstein:1988:ASM

[RGL88] Mark A. Rosenstein, Daniel E. Geer, Jr., and Peter J. Levine. The Athena Service Management System. In USENIX Association [USE88j], pages 203– 211. ISBN ???? LCCN ????

Rhodes:1985:MAW

[RHH85a] Rocky Rhodes, Paul Haeberli, and Kipp Hickman. Mex — A window manager for the IRIS. In USENIX Association [USE85e], pages 381–392. LCCN QA76.8.U65 U8 1985.

Rhodes:1985:MWM

[RHH85b] Rocky Rhodes, Paul Haeberli, and Kipp Hickman. Mex — A window manager for the IRIS. In USENIX [USE85a], pages 381–392.

Richards:1985:BDA

[Ric85] Robert R. Richards. A basic direct access method for UNIX. In USENIX Association [USE85c], pages 176–182.

Richards:1987:GC

[Ric87] John E. Richards. GKS in C++. EUUG Newsletter, 7 (1):53–64, 1987. CODEN EONLE8. ISSN 1011-4211.

Ritchie:1987:UD

[Rit87] Dennis M. Ritchie. Unix: a dialectic. In USENIX Association [USE87g], pages 29– 34. LCCN QA 76.76 O63 U84 1987. Cover title: Conference proceedings. Spine title: Winter 1987 USENIX Association Conference proceedings, Washington, DC.

Ritchie:1988:GFU

[Rit88] Dennis M. Ritchie. A guest facility for Unicos. In USENIX Association [USE88i], pages 19–24. ISBN ???? LCCN ????

Russo:1988:CIS

[RK88] Vincent F. Russo and Simon M.

Kaplan. A C++ interpreter for Scheme. In USENIX Association [USE88k], pages 95–108.

Ramachandran:1989:IDS

[RK89] Umakishore Ramachandran and M. Yousef A. Khalidi. An implementation of distributed shared memory. In USENIX Association [USE89d], pages 21–38. ISBN ???? LCCN ????

Rodriguez:1986:GFS

[RKH86] R. Rodriguez, M. Koehler, and R. Hyde. The Generic File System. In USENIX Association [USE86c], pages 260–269.

Rodriguez:1988:DUO

[RKPP88] Robert Rodriguez, Matt Koehler, Larry Palmer, and Ricky Palmer. A dynamic UNIX operating system. In Association [Ass88f], pages 305–319.

Rajkumar:1988:TSR

[RL88] Ragunathan Rajkumar and John P. Lehoczky. Task synchronization in real-time operating systems. In Association [Ass88e], pages 18–22.

Rees:1986:EOS

[RLML86a] Jim Rees, Paul H. Levine, Nathaniel Mishkin, and Paul J. Leach. An extensible I/O system. In USENIX Association [USE86c], pages 114–125.

Rees:1986:ES

[RLML86b] Jim Rees, Paul H. Levine, Nathaniel Mishkin, and Paul J. Leach. An extensible I/O system. In USENIX [USE86a], pages 114–125.

Robboy:1984:UBO

[Rob84a] David Robboy. UNIX block I/O optimization on microcomputers. In USENIX Association [USE84c], pages 223–230. ISBN none. LCCN QA76.8.U65 U55 1984.

Robins:1984:AUD

[Rob84b] Charles M. Robins. Adapting UNIX for data communications. In USENIX Association [USE84c], pages 39–46. ISBN none. LCCN QA76.8.U65 U55 1984.

Roberts:1987:WNW

[Rob87]  William Roberts.  What

NeWS? or What light through yonder window breaks? EUUG Newsletter, 7(4):36–41, Winter 1987. CODEN EONLE8. ISSN 1011-4211.

Roberts:1989:IWS

[Rob89] William Roberts. Introduction to window systems. In Association [Ass89e], pages 37–43. CODEN EONLE8. ISSN 1011-4211.

Rochkind:1989:UPI

[Roc89a] Marc J. Rochkind. A unified programming interface for character-based and graphical window systems. In USENIX Association [USE89f], pages 109–117. LCCN QA 76.76 O63 U83 1989.

Rochkind:1989:XAV

[Roc89b] Marc J. Rochkind. XVT:

A virtual toolkit for portability between window systems. In USENIX Association [USE89g], pages 151–163.

Rochkind:1989:XVT

[Roc89c] Marc J. Rochkind. XVT: A virtual toolkit for portability between window systems. In USENIX [USE89c], pages 151– 163.

Rodriguez:1986:SCT

[Rod86] R. Rodriguez. A system call tracer for UNIX. In USENIX Association [USE86c], pages 72–80.

Rodriquez:1987:SDN

[Rod87] Mike Rodriquez. Software distribution in a network environment. In USENIX Association [USE87d], pages 20–?? ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

Rosenthal:1982:SCC

[Ros82] Eric S. Rosenthal. Spelling checkers, compound words, and variant spellings. In Usr Group [Usr82], pages 315–322.

Rees:1987:DES

[ROS87a] Jim Rees, Margaret Olson, and J. Sasidhar. A dynamically extensible streams implementation. In USENIX Association [USE87f], pages 199–207.

Rose:1987:CCL

[Ros87b] John R. Rose. C*: A C++-like language for data-parallel computation. In USENIX Association [USE87a], pages 127–134. ISBN ???? LCCN ????

Rose:1987:ICC

[Ros87c] John R. Rose. Implementing a compiler in C++: Experience and generalizations. In USE-NIX Association [USE87a], pages 135–146. ISBN ???? LCCN ????

Rosenblum:1987:UFS

[Ros87d]  Paul Rosenblum.  Using

the Fortran 8x style of programming. Technical report TMC-148, PL87-7, Thinking Machines Corporation, Cambridge, MA, USA, 1987. 12 pp. Presented at the USENIX C++ Conference, Sante Fe December 1987.

Rosenthal:1988:SXC

[Ros88] David Rosenthal. A simple X11 client program -or- how hard can it really be to write “hello, world”? In USENIX Association [USE88j], pages 229–242. ISBN ???? LCCN ????

Redman:1984:BEB

[RP84] Brian E. Redman and Pat E.

Parseghian. Behind every binary license ins the UNIX heritage. In USENIX Association [USE84c], pages 75–82. ISBN none. LCCN QA76.8.U65 U55 1984.

Rose:1985:MHP

[RR85] Marshall T. Rose and John L.

Romine. MH.5: How to process 200 messages a day and still get some real work done. In USENIX Association [USE85e], pages 455–487. LCCN QA76.8.U65 U8 1985.

Raghavan:1987:CCB

[RRS87] Raghunath Raghavan, Niranjan Ramakrishnan, and Sue Strater. A C++ class browser. In USENIX Association [USE87a], pages 274–281. ISBN ???? LCCN ????

Raeburn:1989:DEC

[RRSZ89] Ken Raeburn, Jon Rochlis, William Sommerfeld, and Stan Zanarotti. Discuss: An electronic conferencing system for a distributed computing environment. In USENIX Association [USE89g], pages 331–342.

Rose:1987:CEC

[RS87] John R. Rose and Guy L.

Steele, Jr. C*: An extended C language for data parallel programming. In USENIX Association [USE87a], pages 361–397. ISBN ???? LCCN ????

Raine:1989:AOF

[RSSW89] Neil Raine, David Seal, William Stoye, and Roger Wilson. The Acorn Outline Font Manager. In USENIX Association [USE89i], pages 25–36. ISBN ???? LCCN ????

Rowe:1985:MPM

[RSV85] P. K. Rowe, S. Sartzetakis, and B. Vishnubhatla. A multiprocessor performance measurement tool. In USENIX Association [USE85e], pages 421– 432. LCCN QA76.8.U65 U8 1985.

Ryan:1983:NPC

[RSW83] Ralph Ryan, Hans Spiller, and Dave Weil. A new portable compiler for XENIX. In Software Tools Users Group [Sof83], pages 229–235. LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Ross:1983:USG

[RT83] Douglas J. Ross and M. Martin Taylor. UNIX support for guaranteed real-time processing. In Software Tools Users Group [Sof83], pages 137–153. LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Raleigh:1988:CAD

[RU88a] T. M. Raleigh and R. W.

Underwood. CRACK: A distributed password advisor. In USENIX Association [USE88g], pages 12– 13. LCCN QA76.8.U65 U55 1988(1)-1990(2)//. Abstract only.

Raleigh:1988:CDP

[RU88b] T. M. Raleigh and R. W. Underwood. CRACK: A distributed password advisor. In USENIX [USE88b], pages 12– 13. Abstract only.

Ruggiero:1982:PAP

[Rug82a] Mario Ruggiero. Ped — A portable editor. In Usr Group [Usr82], pages 35–42.

Ruggiero:1982:PPE

[Rug82b] Mario Ruggiero. Ped — A portable editor. In USENIX [USE82a], pages 35–42.

Rugaber:1983:USU

[Rug83] Spencer Rugaber. A uniform and simple user interface to UNIX. In Association [Ass83a], pages 113–115.

Rieken:1986:HUBb

[RW86a] Bill Rieken and Jim Webb.

HoneyDanBer UUCP — bringing UNIX systems in the information age, part 2: Error handling, administrative aids, and user enhancements. ;login: the USENIX Association newsletter, 11(4):10–35, July/August 1986. CODEN LOGNEM. ISSN 1044-6397.

Rieken:1986:HUBa

[RW86b] Bill Rieken and Jim Webb. HoneyDanBer UUCP — bringing UNIX systems into the information age, part 1: Performance, security, and networking facilities. ;login: the USENIX Association newsletter, 11(3):27–36, May/June 1986. CODEN LOGNEM. ISSN 1044-6397.

Rao:1987:XTS

[RW87] Ram Rao and Smokey Wallace. The X toolkit — the standard toolkit for X version 11. In USENIX Association [USE87f], pages 117–129.

Rosen:1986:NP

[RWFC86] Mordecai B. Rosen, Michael J. Wilde, and Bill Fraser-Campbell. NFS portability. In USENIX Association [USE86c], pages 299–305.

Ramamurthy:1987:PCP

[RWNA87] G. Ramamurthy, Y. T. Wang, Hank Nichols, and Mike Andrews. A prototype capacity planning and configuration modeling tool for UNIX systems. In USENIX Association [USE87g], pages 103–110. LCCN QA 76.76 O63 U84 1987. Cover title: Conference proceedings. Spine title: Winter 1987 USENIX Association Conference proceedings, Washington, DC.

Smith:1986:FOA

[SA86a] Edward T. Smith and David B.

Anderson. Flamingo: Objectoriented abstractions for user interface management. In USENIX Association [USE86e], pages 72–78.

Smith:1986:FOO

[SA86b] Edward T. Smith and David B.

Anderson. Flamingo: Objectoriented abstractions for user interface management. In USE-NIX [USE86b], pages 72–78.

Swick:1988:XTM

[SA88] Ralph R. Swick and Mark S.

Ackerman. The X toolkit: More bricks for building userinterfaces or widgets for hire. In USENIX Association [USE88j], pages 221–228. ISBN ???? LCCN ????

Salwen:1982:RAL

[Sal82] Howard Salwen. On ring architected local networks. In Usr Group [Usr82], pages 187–197.

Salkind:1988:SOS

[Sal88] Lou Salkind. The SAGE operating system. In Association [Ass88e], pages 54–58.

Salus:1989:BRD

[Sal89a] Peter H. Salus. Book review: !%@:: A Directory of Electronic Mail Addressing and Networks. ;login: the USENIX Association newsletter, 14(6):24–??, November/ December 1989. CODEN LOGNEM. ISSN 1044-6397.

Salus:1989:BRL

[Sal89b] Peter H. Salus. Book review: Learning the vi Editor. ;login: the USENIX Association newsletter, 14(3):14–??, May/ June 1989. CODEN LOGNEM. ISSN 1044-6397.

Samadi:1987:KBS

[Sam87] Behrokh Samadi. A knowledgebased system for performance tuning of the UNIX operating system. In USENIX Association [USE87g], pages 110– 123. LCCN QA 76.76 O63 U84 1987. Cover title: Conference proceedings. Spine title: Winter 1987 USENIX Association Conference proceedings, Washington, DC.

Sandel:1983:SVS

[San83] Dave Sandel. System V support offering. In Association [Ass83a], pages 48–49. Abstract only.

Sanger:1989:UC

[San89] Colston Sanger. UNIX clinic. EUUG Newsletter, 9(1):63– 68, Spring 1989. CODEN EONLE8. ISSN 1011-4211.

Sauer:1988:PSS

[Sau88] Charles H. Sauer. Presenting a single system image with fine granularity mounts. ;login: the USENIX Association newsletter, 13(4):13–20, July/August 1988. CODEN LOGNEM. ISSN 1044-6397.

Saxer:1985:IUB

[Sax85a] John Saxer. Interpreting UNIX benchmarks. In USENIX Association [USE85c], pages 78–89.

Saxon:1985:UGA

[Sax85b] Michael S. Saxon. Using gsck — A guide to the UNIX file system check program. ;login: the USENIX Association newsletter, 10(3):13–26, August 1985. CODEN LOGNEM. ISSN 1044-6397.

Saxon:1985:UGG

[Sax85c] Michael S. Saxon. Using gsck — A guide to the UNIX file system check program. ;login: the USENIX Association newsletter, 10(3):13–26, August 1985. CODEN LOGNEM. ISSN 1044-6397.

Shipley:1988:HAS

[SB88a] Peter Shipley and Russell Brand. HACKMAN: A systematic study of real computer security holes. In USENIX Association [USE88g], pages 65– 67. LCCN QA76.8.U65 U55 1988(1)-1990(2)//.

Shipley:1988:HSS

[SB88b] Peter Shipley and Russell Brand. HACKMAN: A systematic study of real computer security holes. In USENIX [USE88b], pages 65–67.

Sparks:1988:BHG

[SC88] Hugh Sparks and Bob Chatham. Butterfly HOSE: Graphical programming for parallel systems. In Association [Ass88e], pages 75–79.

Sznyter:1986:NVI

[SCC86a] Edward W. Sznyter, Patrick Clancy, and James Crossland. A new virtual-memory implementation for Unix. In USENIX Association [USE86c], pages 81–92.

Sznyter:1986:NVM

[SCC86b] Edward W. Sznyter, Patrick Clancy, and James Crossland. A new virtual-memory implementation for Unix. In USE-NIX [USE86a], pages 81–92.

Schaap:1982:PCL

[Sch82] Chaim E. Schaap. Portability of C language programs. In Usr Group [Usr82], pages 43–?? Abstract only.

Scherrer:1983:PTM

[Sch83a] Phil Scherrer. Performance of tools: Minis versus micros. In Software Tools Users Group [Sof83], pages 425–?? LCCN QA76.8.U65 U74 1983. Abstract only.

Scherrer:1983:STC

[Sch83b] Phil Scherrer. Software tools in C? In Association [Ass83b], pages 3–6.

Scheulen:1983:VCU

[Sch83c] Bob Scheulen. Version 7 compatibility under System 3/5. In Software Tools Users Group [Sof83], pages 349–352. LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Schefstrom:1986:RCT

[Sch86a] Dick Schefstrom. Revision control tools and the Ada program library. In USENIX Association [USE86e], pages 241–251.

Schoner:1986:EAC

[Sch86b] James S. Schoner. Ease: A configuration language for Sendmail. ;login: the USENIX Association newsletter, 11(1): 19–31, January/February 1986. CODEN LOGNEM. ISSN 1044-6397.

Schoner:1986:ECL

[Sch86c] James S. Schoner. Ease: A configuration language for sendmail. ;login: the USENIX Association newsletter, 11(1): 19–31, January/February 1986. CODEN LOGNEM. ISSN 1044-6397.

Schaufler:1988:XND

[Sch88a] Robin Schaufler. X11/NeWS design overview. In Association [Ass88f], pages 23–35.

Schwarz:1988:CLI

[Sch88b] Jerry Schwarz. A C++ library for infinite precision floating point. In USENIX Association [USE88k], pages 271–281.

Schaaser:1989:UTN

[Sch89] Horst Schaaser. Using Transputer networks to accelerate communication protocols. In USENIX Association [USE89d], pages 361–372. ISBN ???? LCCN ????

Shein:1989:NAN

[SCW89a] Barry Shein, Mike Callahan, and Paul Woodbury. NFS-STONE: A network file server performance benchmark. In USENIX Association [USE89f], pages 269–275. LCCN QA 76.76 O63 U83 1989.

Shein:1989:NNF

[SCW89b] Barry Shein, Mike Callahan, and Paul Woodbury. NFS-STONE: A network file server performance benchmark. In USENIX [USE89b], pages 269– 275.

Stephens:1987:ESA

[SD87] Lindsey E. Stephens and Lawrence W. Dowdy. Experimental sensitivity analysis of performance in a UNIX system. In USENIX Association [USE87g], pages 60–72. LCCN QA 76.76 O63 U84 1987. Cover title: Conference proceedings. Spine title: Winter 1987 USE-NIX Association Conference proceedings, Washington, DC.

Schulert:1988:ODU

[SE88] Andrew Schulert and Kate Erf. Open dialogue: Using an extensible retained object workspace to support a UIMS. In USENIX Association [USE88k], pages 53–64.

Seeley:1983:VCP

[See83] Donn Seeley. VAX11 compatibility on PDP-11s. In Association [Ass83a], pages 193–198.

Seeley:1989:TW

[See89] Donn Seeley. A tour of the worm. In USENIX Association [USE89g], pages 287–304.

Senft:1987:DDE

[Sen87] Christopher Senft. A distributed design environment for distributed realtime systems. In USENIX Association [USE87b], pages 131–151. ISBN ???? LCCN ????

Sequin:1985:MRM

[Seq85] Carlo H. Sequin. A modular rendering and modeling system. In USENIX Association [USE85d], pages 38–53. ISBN ???? LCCN ????

Sequin:1986:PSI

[Seq86] Carlo H. Sequin. Procedural spline interpolation in UNICU-BIX. In USENIX Association [USE86d], pages 63–83. ISBN ???? LCCN ????

Sequin:1989:MM

[Seq89] Carlo. H. Sequin. Microfabrication on the Macintosh. In USENIX Association [USE89i], pages 1–16. ISBN ???? LCCN ????

Spafford:1986:RAS

[SF86] Eugene H. Spafford and John C. Flaspohler. A report on the accuracy of some floating point math functions on selected computers. ;login: the USENIX Association newsletter, 11(2):31–56, March/April 1986. CODEN LOGNEM. ISSN 1044-6397.

Shapiro:1989:SOO

[SGH^{+}89] Marc Shapiro, Yvon Gourhant, Sabine Habert, Laurence Mosseri, Michel Ruffin, and Celine Valot. SOS: An object-oriented operating system — assessment and perspectives. In Association [Ass89a], pages 287–337.

Sandberg:1985:DIS

[SGK^{+}85] Russel Sandberg, David Goldberg, Steve Kleiman, Dan Walsh, and Bob Lyon. Design and implementation of the Sun Network Filesystem. In USENIX Association [USE85e], pages 119–130. LCCN QA76.8.U65 U8 1985.

Shlag:1989:AM

[SGS89] John F. Shlag, Julian E. Gomez, and Alex D. Seiden. 3D animation on the Macintosh with 3DWorks. In USENIX Association [USE89i], pages 17– 24. ISBN ???? LCCN ????

Shattan:1985:DUB

[SH85] Ariel Shattan and Jenny Hecker. Documenting UNIX: Beyond man pages. In USENIX Association [USE85e], pages 437–454. LCCN QA76.8.U65 U8 1985.

Snider:1989:MK

[SH89] Greg Snider and Jim Hays. The Modix kernel. In USENIX Association [USE89g], pages 377– 392.

Shantz:1983:GSP

[Sha83] Michael Shantz. Graphics standards for personal workstations. In Association [Ass83a], pages 257–259.

Shapiro:1989:PDO

[Sha89]  Marc Shapiro.  Prototyp-

ing a distributed objectoriented operating system on Unix. In USENIX Association [USE89d], pages 311–331. ISBN ???? LCCN ????

Summers-Horton:1985:SUU

[SHH85] Karen Summers-Horton and Mark Horton. Status of the USENIX UUCP project. In USENIX Association [USE85c], pages 183–?? Abstract only.

Shopiro:1987:ECT

[Sho87] Jonathan E. Shopiro. Extending the C++ task system for real-time control. In USENIX Association [USE87a], pages 77–94. ISBN ???? LCCN ????

Shultz:1989:DEO

[Shu89] Roger K. Shultz. A decentralized embedded operating system supporting distributed execution of Ada tasks. In USE-NIX Association [USE89d], pages 391–409. ISBN ???? LCCN ????

Sigmon:1987:ASD

[Sig87] Tim Sigmon. Automatic software distribution. In USENIX Association [USE87d], pages 21–?? ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

Simmons:1988:MLN

[Sim88] Steve Simmons. Making a large network reliable. In USENIX Association [USE88f], pages 47–?? ISBN ???? LCCN ????

Simicich:1989:Y

[Sim89] Nick Simicich. YABS. In USENIX Association [USE89e], pages 109–121. ISBN ???? LCCN ????

Sinkewicz:1988:SSU

[Sin88] Ursula Sinkewicz. A strategy for SMP ULTRIX. In Association [Ass88f], pages 203–212.

Skinner:1983:UNS

[SJ83] Glenn C. Skinner and Bill Jolitz. UNIX on the National Semiconductor NS16032. In Association [Ass83a], pages 291–306.

Sauer:1987:RPD

[SJL^{+}87] Charles H. Sauer, Don W. Johnson, Larry K. Loucks, Amal A. Shaheen-Gouda, and Todd A. Smith. RT PC distributed services: File system. ;login: the USENIX Association newsletter, 12(5):12–22, September/October 1987. CO-DEN LOGNEM. ISSN 1044-6397.

Skubiszewski:1988:SEU

[Sku88] Marcin Skubiszewski. Security of Ethernet under UNIX and Internet protocol. EUUG Newsletter, 8(1):2–10, Spring 1988. CODEN EONLE8. ISSN 1011-4211.

Schragl:1988:PCB

[SL88] R. Schragl and D. Lauber.

A protocol for the communication between objects. In USENIX Association [USE88e], pages 79–87. ISBN ???? LCCN ????

Slezak:1987:MMS

[Sle87]  Tom Slezak.  Managing

modems and serial ports. In USENIX Association [USE87d], pages 45–?? ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

Scott:1989:IIP

[SLM89] Michael L. Scott, Thomas J.

LeBlanc, and Brian D. Marsh. Implementing issues for the Psyche multiprocessor operating system. In USENIX Association [USE89d], pages 227–236. ISBN ???? LCCN ????

Slattery:1984:CDA

[SM84] T. C. Slattery and William McCool. Circuit design aids — CDA: A printed circuit board manufacturing system. In Software Tools Users Group [Sof84], pages 280–286. LCCN QA76.8.U65 U83 1984.

Smith:1988:ECM

[SM88a] Jonathan M. Smith and Gerald Q. Maguire, Jr. Effects of copy-on-write memory management on the response time of UNIX fork operations. In Association [Ass88c], pages 255– 278.

Smith:1988:ECW

[SM88b] Jonathan M. Smith and Gerald Q. Maguire, Jr. Effects of copy-on-write memory management on the response time of UNIX fork operations. In Association [Ass88c], pages 255– 278.

Spencer:1989:FAA

[SM89] Jon F. Spencer and Jackie McAlexander. Factors affecting application portability to a B1 level trusted UNIX. In USENIX Association [USE89f], pages 239–255. LCCN QA 76.76 O63 U83 1989.

Smith:1983:UWW

[Smi83]  Charles R. Smith.  UNIX

writer’s workbench. In Software Tools Users Group [Sof83], pages 321–?? LCCN QA76.8.U65 U74 1983. Abstract only.

Smith:1987:CEN

[Smi87a]  Jeffrey M. Smith.  Creat-

ing an environment for novice users. In USENIX Association [USE87d], pages 37–?? ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

Smith:1987:EA

[Smi87b] Thorn Smith. Excelan administration. In USENIX Association [USE87d], pages 4– ?? ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

Smith:1989:DRA

[Smi89] Griffith G. Smith, Jr. A distributed resource allocator for UNIX systems. In USENIX Association [USE89f], pages 95– 108. LCCN QA 76.76 O63 U83 1989.

Steiner:1988:KAS

[SNS88] Jennifer G. Steiner, Clifford Neuman, and Jeffrey I. Schiller. Kerberos: An authentication service for open network systems. In USENIX Association [USE88j], pages 191–202. ISBN ???? LCCN ????

STUG:1983:PUA

[Sof83] Software Tools Users Group, editor. Proceedings: USE-NIX Association [and] Software Tools Users Group Summer Conference, Toronto 1983, July 1983, Toronto, Ontario, Canada. USENIX, P.O. Box 7, El Cerrito 94530, CA, USA, 1983. LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

STUG:1984:UAS

[Sof84] Software Tools Users Group, editor. USENIX Association [and] Software Tools Users Group Summer Conference, Salt Lake City 1984: proceedings, June 12–15, 1984, Salt Lake City, Utah, USA. USE-NIX, P.O. Box 7, El Cerrito 94530, CA, USA, 1984. LCCN QA76.8.U65 U83 1984.

Somogyi:1988:CFG

[Som88] Zoltan Somogyi. Cake: a fifth generation version of make. EUUG Newsletter, 8(2):13– 20, Summer 1988. CODEN EONLE8. ISSN 1011-4211.

Son:1988:MAD

[Son88a] Sang H. Son. A messagebased approach to distributed database prototyping. In Association [Ass88e], pages 71–74.

Son:1988:MBA

[Son88b] Sang H. Son. A messagebased approach to distributed database prototyping. In Association [Ass88e], pages 71–74.

Spafford:1989:SME

[Spa89] Eugene H. Spafford. Some musings on ethics and computer break-ins. In USENIX Association [USE89g], pages 305–311.

Spence:1987:IDP

[Spe87] Bruce Spence. Intelligent distributed printing/plotting. In USENIX Association [USE87d], pages 44–?? ISBN ????

LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

Spencer:1988:HSC

[Spe88] Henry Spencer. How to steal code -or- inventing the wheel only once. In USENIX Association [USE88j], pages 335–345. ISBN ???? LCCN ???? URL

ftp://ftp.cs.toronto.edu/
doc/programming/steal.ps.

Spence:1989:SUF

[Spe89] Bruce Spence. spy: A Unix file system security monitor. In USENIX Association [USE89e], pages 75–87. ISBN ???? LCCN ????

Scheer:1985:UAR

[SR85a] M. D. Scheer and S. Rajeev. A UNIX-based Ada runtime system. In USENIX Association [USE85c], pages 51–?? Abstract only.

Scheer:1985:UBA

[SR85b] M. D. Scheer and S. Rajeev. A UNIX-based Ada runtime system. In USENIX [USE85b], pages 51–?? Abstract only.

Scott:1988:GEO

[SREC88] Roger Scott, Prakash Reddy, Russel Edwards, and David Campbell. GPIO: Extensible objects for electronic design tools. In USENIX Association [USE88k], pages 109–121.

Stroustrup:1987:SCC

[SS87] Bjarne Stroustrup and Jonathan E.

Shopiro. A set of C++ classes for co-routine style programming. In USENIX Association [USE87a], pages 417–439. ISBN ???? LCCN ????

Sanislo:1988:RLS

[SS88] Jan Sanislo and Mark S. Squillante. An RPC/LWP system for interconnecting heterogeneous systems. In USENIX Association [USE88j], pages 43– 54. ISBN ???? LCCN ????

Sumey:1987:GLN

[SSNU87] Roger A. Sumey, Daniel M. Sunday, David W. Nesbitt, and Kyle M. Upton. A graphics library for Navy Tactical Display Systems. In USENIX Association [USE87c], pages 105–?? ISBN ???? LCCN ???? Abstract only.

Shienbrood:1983:UAC

[SSWW83] Eric R. Shienbrood, Carl A. Soeder, James R. Ward, and Kincade N. Webb. UNIX on Apollo computers (yet another UNIX emulation). In Association [Ass83a], pages 133–142.

Smith-Thomas:1989:SML

[ST89a] Barbara Smith-Thomas. Secure multi-level windowing in a B1 certifiable secure UNIX operating system. In USENIX [USE89c], pages 429–439.

Smith-Thomas:1989:SMW

[ST89b] Barbara Smith-Thomas. Secure multi-level windowing in a B1 certifiable secure UNIX operating system. In USENIX Association [USE89g], pages 429– 439. Describes the architecture of the multi-level version of the AT&T 630 graphics terminal. This terminal was evaluated as part of AT&T System V/MLS, which received a B1 rating.

Straathof:1986:USL

[STA86] Jeffrey H. Straathof, Ashok K.

Thareja, and Ashok K. Agrawala. UNIX scheduling for large systems. In USENIX Association [USE86e], pages 111– 139.

Stay:1987:DRB

[Sta87a] Paul Randal Stay. The definition and ray-tracing of B-spline objects in a combinatorial solid geometric modeling system. In USENIX Association [USE87c], pages 81–85. ISBN ???? LCCN ????

Stay:1987:DRT

[Sta87b] Paul Randal Stay. The definition and ray-tracing of B-spline objects in a combinatorial solid geometric modeling system. In USENIX Association [USE87c], pages 81–85. ISBN ???? LCCN ????

Straathof:1987:MRP

[STA87c] Jeffrey H. Straathof, Ashok K.

Thareja, and Ashok K. Agrawala. Methodology and results of performance measurements for a new UNIX scheduler. In USENIX Association [USE87g], pages 165–180. LCCN QA 76.76 O63 U84 1987.

Cover title: Conference proceedings. Spine title: Winter 1987 USENIX Association Conference proceedings, Washington, DC.

Steffen:1983:CAP

[Ste83a] J. L. Steffen. Ctrace — A portable debugger for C programs. In Association [Ass83a], pages 187–191.

Steffen:1983:CPD

[Ste83b] J. L. Steffen. Ctrace — A portable debugger for C programs. In USENIX [USE83a], pages 187–191.

Steffen:1984:SAC

[Ste84] Joseph L. Steffen. Software administration over computer networks — the exptools experience. In USENIX Association [USE84c], pages 17–22. ISBN none. LCCN QA76.8.U65 U55 1984.

Steffen:1985:IEC

[Ste85] Joseph L. Steffen. Interactive examination of a C program with Cscope. In USENIX Association [USE85c], pages 170– 175.

Sterk:1986:FDS

[Ste86] Don Sterk. Full duplex support on mainframes. In USENIX Association [USE86e], pages 165– 171.

Stewart:1988:NAI

[Ste88a] C. A. Stewart. Numerical applications interprocess communication protocol: RPCODE:

RPC server to solve ODEs. In USENIX Association [USE88i], pages 37–42. ISBN ???? LCCN ????

Stewartson:1988:UVB

[Ste88b] Ian Stewartson. UNIX V.3 and beyond. In USENIX Association [USE88e], pages 161–177. ISBN ???? LCCN ????

Stenning:1989:PH

[Ste89a] Vic Stenning. Project Hygiene. In USENIX Association [USE89j], pages 1–9.

Stevens:1989:HDD

[Ste89b] W. Richard Stevens. Heuristics for disk drive positioning in 4.3BSD. In Association [Ass89c], pages 251–274.

Stitt:1983:RDM

[Sti83]  F. W. Stitt.  Research

database management software for UNIX-based microcomputers. In Association [Ass83a], pages 201–209.

Stokes:1985:UA

[Sto85] Grant Stokes. University application. In USENIX Association [USE85e], pages 191–193. LCCN QA76.8.U65 U8 1985.

Stone:1987:SCH

[Sto87] Ken Stone. System cloning at HP-SDD. In USENIX Association [USE87d], pages 18– ?? ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

Stokes:1988:PDA

[Sto88] Ronan Stokes. Prototyping database applications with a hybrid of C++ and 4GL. In USENIX Association [USE88k], pages 41–52.

Stroustrup:1985:EFC

[Str85a] Bjarne Stroustrup. An extensible I/O facility for C++. In USENIX [USE85a], pages 57– 70.

Stroustrup:1985:EOF

[Str85b]  Bjarne Stroustrup.  An

extensible I/O facility for C++. In USENIX Association [USE85e], pages 57–70. LCCN QA76.8.U65 U8 1985.

Strassman:1987:HB

[Str87a]  Steve Strassman.  Hairy

brushes. In USENIX Association [USE87c], pages 99–?? ISBN ???? LCCN ???? Abstract only.

Stroustrup:1987:EC

[Str87b] Bjarne Stroustrup. The evolution of C++: 1985 to 1987. In USENIX Association [USE87a], pages 1–21. ISBN ???? LCCN ????

Stroustrup:1987:PDC

[Str87c] Bjarne Stroustrup. Possible directions for C++. In USENIX Association [USE87a], pages 399–416. ISBN ???? LCCN ????

Stroustrup:1987:WIP

[Str87d]  Bjarne Stroustrup.  What

is “object-oriented programming”? In USENIX Association [USE87a], pages 159–180. ISBN ???? LCCN ????

Stroustrup:1987:WOO

[Str87e]  Bjarne Stroustrup.  What

is “object-oriented programming”? In USENIX Association [USE87a], pages 159–180. ISBN ???? LCCN ????

Stroustrup:1988:PTC

[Str88a] Bjarne Stroustrup. Parameterized types for C++. In USE-NIX Association [USE88k], pages 1–18.

Stroustrup:1988:TLCa

[Str88b] Bjarne Stroustrup. Type-safe linkage for C++. In USENIX Association [USE88k], pages 193–210.

Stroustrup:1988:TLCb

[Str88c] Bjarne Stroustrup. Type-safe linkage for C++. In Association [Ass88a], pages 371–403.

Stroustrup:1988:TLCc

[Str88d] Bjarne Stroustrup. Type-safe linkage for C++. In USENIX Association [USE88k], pages 193–210.

Stroustrup:1988:TSLa

[Str88e] Bjarne Stroustrup. Type-safe linkage for C++. In Association [Ass88a], pages 371–403.

Stroustrup:1988:TSLb

[Str88f] Bjarne Stroustrup. Type-safe linkage for C++. In USENIX [USE88a], pages 193–210.

Stroustrup:1989:EC

[Str89a] Bjarne Stroustrup. The evolution of C++: 1985 to 1989. In Association [Ass89c], pages 191–250.

Stroustrup:1989:MIC

[Str89b] Bjarne Stroustrup. Multiple inheritance for C++. In Association [Ass89a], pages 367–395.

Stroustrup:1989:PTC

[Str89c] Bjarne Stroustrup. Parameterized types for C++. In Association [Ass89d], pages 55–85.

Suzuki:1986:REC

[STT86a] Tatsuo Suzuki, Hideo Taniguchi, and Hisayasu Takada. A real-time electronic conferencing system based on distributed UNIX. In USENIX Association [USE86c], pages 189–199.

Suzuki:1986:RTE

[STT86b] Tatsuo Suzuki, Hideo Taniguchi, and Hisayasu Takada. A real-time electronic conferencing system based on distributed UNIX. In USENIX [USE86a], pages 189–199.

Spezzano:1987:NED

[STV87] Giandomenico Spezzano, Domenico Talia, and Marco Vanneschi. NERECO: An environment for the development of distributed software. In USE-NIX Association [USE87b], pages 153–167. ISBN ???? LCCN ????

Sullivan:1987:VCG

[Sul87] Michael J. Sullivan. Visualization: Computer graphics in the research laboratory. In USENIX Association [USE87c], pages 104–?? ISBN ???? LCCN ???? Abstract only.

Sunderam:1988:FTO

[Sun88] V. S. Sunderam. A fast transaction oriented protocol for distributed applications. In USENIX Association [USE88j], pages 79–87. ISBN ???? LCCN ????

Sunderam:1989:ENS

[Sun89] V. S. Sunderam. An experiment with network shared libraries. In USENIX Association [USE89f], pages 39–49. LCCN QA 76.76 O63 U83 1989.

Steffen:1983:ESC

[SV83] Joseph L. Steffen and Michael T.

Veach. The edit shell — combining screen editing and the history list. In Software Tools Users Group [Sof83], pages 187–190. LCCN QA76.8.U65 U74 1983. Sponsored by USE-NIX Association in cooperation with Software Tools Users Group.

Sventek:1983:PMS

[Sve83] Joe Sventek. A portable mail system for the software tools virtual operating system. In Association [Ass83b], pages 7– 8. Abstract only.

Sanford:1982:DCA

[SW82] Curtis Sanford and David Walden. Development of a commercial applications system under UNIX. In Usr Group [Usr82], pages 177–?? Abstract only.

Shah:1984:DBB

[SW84a] Bakul Shah and Robert P. Warnock, III. A dynamic badblock forwarding algorithm. In USENIX [USE84a], pages 192– ?? Abstract only.

Shah:1984:DBF

[SW84b] Bakul Shah and Robert P. Warnock, III. A dynamic badblock forwarding algorithm. In Software Tools Users Group [Sof84], pages 192–?? LCCN QA76.8.U65 U83 1984. Abstract only.

Schaffer:1988:LII

[SW88] Mark A. Schaffer and Geoff Walsh. LOCK/ix: An implementation of UNIX for the LOCK TCB. ;login: the USENIX Association newsletter, 13(3):11–24, May/June 1988. CODEN LOGNEM. ISSN 1044-6397.

Swartz:1983:CS

[Swa83] Robert Swartz. Criteria for standards. In Association [Ass83a], pages 349–?? Abstract only.

Toy:1982:RWI

[TA82] Michael C. Toy and Kenneth C.

R. C. Arnold. Rogue: Where it has been, why it was there, and why it shouldn’t have been there in the first place. In Usr Group [Usr82], pages 139– ?? Abstract only.

Tague:1983:USN

[Tag83] Berkley A. Tague. The UNIX system: New directions. In Association [Ass83a], pages 109– ?? Abstract only.

Talati:1989:DLT

[Tal89a] Kirit Talati. Distributed online transaction processing on UNIX. In USENIX Association [USE89h], pages 7–12. ISBN ???? LCCN ????

Talati:1989:DOT

[Tal89b] Kirit Talati. Distributed online transaction processing on UNIX. In USENIX Association [USE89h], pages 7–12. ISBN ???? LCCN ????

Tannenbaum:1984:PHU

[Tan84] Andrew Tannenbaum. Political history of UNIX. In USENIX Association [USE84c], pages 83–88. ISBN none. LCCN QA76.8.U65 U55 1984.

Tannenbaum:1986:R

[Tan86] Ed Tannenbaum. Recollections. In USENIX Association [USE86d], pages 15–16. ISBN ???? LCCN ????

Tanenbaum:1987:MUC

[Tan87a] Andrew S. Tanenbaum. MINIX: A UNIX clone with source code for the IBM PC. ;login: the USENIX Association newsletter, 12(2):3–9, March/April 1987. CODEN LOGNEM. ISSN 1044-6397.

Tannenbaum:1987:UML

[Tan87b] Andrew Tannenbaum. The UNIX marketplace in 1987: Life, the UNIverse, and everything. In USENIX Association [USE87f], pages 419–424.

Taylor:1986:PIO

[Tay86] Dave Taylor. Personalizing the impersonal and other tales of communication in the computer age. ;login: the USE-NIX Association newsletter, 11 (6):5–12, November/December 1986. CODEN LOGNEM. ISSN 1044-6397.

Taylor:1988:PAR

[Tay88] Dave Taylor. The postman always rings twice: Electronic mail in a highly distributed environment. In USENIX Association [USE88j], pages 145–153. ISBN ???? LCCN ????

Thompson:1988:TLB

[TBJW88] Michael Y. Thompson, J. M. Barton, T. A. Jermoluk, and J. C. Wagner. Translation lookaside buffer synchronization in a multiprocessor system. In USENIX Association [USE88j], pages 297–302. ISBN ???? LCCN ????

Tolchin:1987:RWU

[TBS87] Stephen Tolchin, Eric Bergan, and Marvin Schneider. Real world UNIX DBMS applications: Experiences and observations. In USENIX Association [USE87g], pages 210– 222. LCCN QA 76.76 O63 U84 1987. Cover title: Conference proceedings. Spine title: Winter 1987 USENIX Association Conference proceedings, Washington, DC.

Terry:1987:ONL

[Ter87] Michael J. C. Terry. An overview of the Native Language System. EUUG Newsletter, 7(2):25–33, 1987. CODEN EONLE8. ISSN 1011-4211.

Test:1982:NWS

[Tes82] Jack A. Test. NUnix window system description. In Usr Group [Usr82], pages 45–50.

Test:1986:MMC

[Tes86a] Jack A. Test. Multi-processor management in the Concentrix operating system. In USENIX Association [USE86e], pages 172–182.

Test:1986:MPM

[Tes86b] Jack A. Test. Multi-processor management in the concentrix operating system. In USENIX [USE86b], pages 172–182.

Thomas:1989:PDE

[TF89a] Spencer W. Thomas and Martin Friedmann. PEX — A 3-D extension to X Windows. In USENIX [USE89c], pages 139– 149.

Thomas:1989:PEX

[TF89b] Spencer W. Thomas and Martin Friedmann. PEX — A 3-D extension to X Windows. In USENIX Association [USE89g], pages 139–149.

Taylor:1986:SNS

[TG86] Bradley Taylor and David Goldberg. Secure networking in the Sun environment. In USENIX Association [USE86c], pages 28–37.

Teixeira:1988:SUG

[TG88] Thomas J. Teixeira and Robert F. Gurwitz. Stellix: UNIX for a graphics supercomputer. In Association [Ass88f], pages 321– 330.

Tsai:1989:TNA

[TGB^{+}89] Chii-Ren Tsai, Virgil D. Gligor, Wilhelm Burger, Mark E. Carson, Pau-Chen Cheng, Janet A. Cugini, Matthew S. Hecht, Shau-Ping Lo, Sohail Malik, and N. Vasudevan. A trusted network architecture for AIX systems. In USENIX Association [USE89g], pages 457–471.

Tanenbaum:1982:SB

[TH82] Andy Tanenbaum and Teus Hagen. Some benchmarks. EUUG Newsletter, 2(4):29– ??, Winter 1982. CODEN EONLE8. ISSN 1011-4211.

Tanenbaum:1983:TPMa

[TH83a] Andrew S. Tanenbaum and Teus Hagen. Two programs, many UNIX systems. EUUG Newsletter, 3(1):12–13, Spring 1983. CODEN EONLE8. ISSN 1011-4211.

Tanenbaum:1983:TPMb

[TH83b] Andrew S. Tanenbaum and Teus Hagen. Two programs, many UNIX systems (reprint). EUUG Newsletter, 3(4):24– 26, Winter 1983. CODEN EONLE8. ISSN 1011-4211.

Tillman:1986:NUV

[TH86] Irwin Tillman and Peter Honeyman. Netnews under VM/ CMS. ;login: the USENIX Association newsletter, 11(4):36– 38, July/August 1986. CODEN LOGNEM. ISSN 1044-6397.

Taylor:1987:ASA

[TH87] Lloyd W. Taylor and John R.

Hayes. An automated student account system. In USENIX Association [USE87d], pages 29–?? ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

Thomas:1985:ACA

[Tho85a] Spencer W. Thomas. The Alpha 1 computer-aided geometric design system in the UNIX environment. ;login: the USE-NIX Association newsletter, 10 (4):54–64, October/November 1985. CODEN LOGNEM. ISSN 1044-6397.

Thomas:1985:ACG

[Tho85b] Spencer W. Thomas. The Alpha 1 computer-aided geometric design system in the UNIX environment. ;login: the USE-NIX Association newsletter, 10 (4):54–64, October/November 1985. CODEN LOGNEM. ISSN 1044-6397.

Thomas:1985:LCG

[Tho85c]  Spencer W. Thomas.  A

low cost graphics workstation. In USENIX Association [USE85d], pages 13–22. ISBN ???? LCCN ????

Thomas:1986:STB

[Tho86] Spencer W. Thomas. Scattered thoughts on B-splines. In USE-NIX Association [USE86d], pages 49–62. ISBN ???? LCCN ????

Tiemann:1988:SRP

[Tie88] Michael D. Tiemann. Solving the RPC problem in GNU C++. In USENIX Association [USE88k], pages 343–361.

Tilson:1982:HUL

[Til82] Michael C. Tilson. How to use lots of memory. In Usr Group [Usr82], pages 107–?? Abstract only.

Tilson:1983:TCP

[Til83] Michael Tilson. A tutorial on C portability. In Association [Ass83a], pages 315–323.

Tilson:1984:TUS

[Til84] Michael Tilson. Towards a UNIX standard. In Software Tools Users Group [Sof84], pages 1–10. LCCN QA76.8.U65 U83 1984.

Tilson:1987:UAT

[Til87a] Michael Tilson. UNIX at the turn of the century. In USENIX Association [USE87f], pages 425–435.

Tilson:1987:UTC

[Til87b] Michael Tilson. UNIX at the turn of the century. In USENIX Association [USE87f], pages 425–435.

Tilbrook:1988:UTC

[Til88]  David Tilbrook.  USENIX

Technical Conference, Dallas, TX, 1988, 1988. 1 videocassette (1 hr. + 51 min.).

Tilbrook:1989:UFA

[Til89] David Tilbrook. Under 10 flags (not always smooth sailing). In USENIX Association [USE89j], pages 137–146.

Tims:1985:PRO

[Tim85] Tracy Tims. A portable reference optimizer for the System V loader. In USENIX Association [USE85e], pages 591–596. LCCN QA76.8.U65 U8 1985.

Tokuda:1988:SBP

[TK88a] Hideyuki Tokuda and Makoto Kotera. Scheduler 1-2-3: It’s better to be predictable than ad hoc. In Association [Ass88e], pages 1–6.

Tokuda:1988:SIB

[TK88b] Hideyuki Tokuda and Makoto Kotera. Scheduler 1-2-3: It’s better to be predictable than ad hoc. In Association [Ass88e], pages 1–6.

Tanenbaum:1982:DSO

[TM82] Andrew S. Tanenbaum and Sape J. Mullender. Design and structure of an open distributed operating system. EUUG Newsletter, 2(4):20–21, Winter 1982. CODEN EONLE8. ISSN 1011-4211.

Ton:1987:RFD

[Ton87] Patricia J. Ton. RDBMS features and data integrity issues in an Army Budget Database System. In USENIX Association [USE87g], pages 196– 209. LCCN QA 76.76 O63 U84 1987. Cover title: Conference proceedings. Spine title: Winter 1987 USENIX Association Conference proceedings, Washington, DC.

Torcaso:1983:IWV

[Tor83a] William Torcaso. The IS/1 workbench for VAX/VMS. In Association [Ass83a], pages 199–?? Abstract only.

Torcaso:1983:WVV

[Tor83b] William Torcaso. The IS/1 workbench for VAX/VMS. In USENIX [USE83a], pages 199– ?? Abstract only.

Tilbrook:1986:TMI

[TP86] D. M. Tilbrook and P. R. H.

Place. Tools for the maintenance and installation of a large software distribution. In USENIX Association [USE86c], pages 223–237.

Terry:1984:BIN

[TPRZ84] Douglas B. Terry, Mark Painter, David W. Riggle, and Songnian Zhou. The Berkeley Internet name domain server. In Software Tools Users Group [Sof84], pages 23–31. LCCN QA76.8.U65 U83 1984.

Tichy:1984:TDF

[TR84] Walter F. Tichy and Zuwang Ruan. Towards a distributed file system. In Software Tools Users Group [Sof84], pages 87– 97. LCCN QA76.8.U65 U83 1984.

Trammell:1985:CBH

[Tra85] R. D. Trammell. A capability based hierarchic architecture for UNIX window management. In USENIX Association [USE85e], pages 373–379. LCCN QA76.8.U65 U8 1985.

Treese:1988:BUW

[Tre88] G. Winfield Treese. Berkeley UNIX on 1000 workstations: Athena changes to 4.3BSD. In USENIX Association [USE88j], pages 175–182. ISBN ???? LCCN ????

Tevanian:1987:MTU

[TRG^{+}87] Avadis Tevanian, Jr., Richard F. Rashid, David B. Golub, David L. Black, Eric Cooper, and Michael W. Young. Mach threads and the Unix kernel: The battle for control. In USENIX Association [USE87f], pages 185–197.

Trickey:1987:CVL

[Tri87] Howard Trickey. C++ versus Lisp: A case study. In USENIX Association [USE87a], pages 440–449. ISBN ???? LCCN ????

Tristam:1989:CVW

[Tri89] David A. Tristam. Controlling virtual words with the panel library. In USENIX Association [USE89i], pages 83–92. ISBN ???? LCCN ????

Tevanian:1987:UIS

[TRY^{+}87] Avadis Tevanian, Jr., Richard F. Rashid, Michael W. Young, David B. Golub, Mary R. Thompson, William Bolosky, and Richard Sanzi. A Unix interface for shared memory and memory mapped files under Mach. In USENIX Association [USE87f], pages 53–67.

Tilbrook:1987:CUS

[TS87] David Tilbrook and Zalman Stern. Cleaning up UNIX source or bringing discipline to anarchy. In USENIX Association [USE87b], pages 275–286. ISBN ???? LCCN ????

Tuori:1982:UBT

[Tuo82] Martin Tuori. A UNIX benchmarking tool with results from the PDP-11/44, VAX 11/780, and Perkin-Elmer 3242. In Usr Group [Usr82], pages 237–246.

Tuori:1983:TUS

[Tuo83] Martin Tuori. Talking to UNIX — some experience with speech input. In Software Tools Users Group [Sof83], pages 179–185. LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Turner:1987:MAF

[Tur87] David Turner. Miranda — an advanced functional programming system running under UNIX. In USENIX Association [USE87f], pages 459–469.

Turner:1988:OM

[Tur88] David Turner. An overview of Miranda. In USENIX Association [USE88e], pages 59–67. ISBN ???? LCCN ????

Tuthill:1982:TAF

[Tut82] Bill Tuthill. Teaching awk as a first programming language. In Usr Group [Usr82], pages 44–?? Abstract only.

Tuthill:1983:DRB

[Tut83] Bill Tuthill. Development of refer: Bug fixes and enhancements (or (unofficially) “refer madness”). In Association [Ass83a], pages 99–103.

Tanenbaum:1983:UTK

[TvK83] Andrew S. Tanenbaum, Hans van Staveren, and E. G. Keizer. A UNIX tool kit for making portable compilers. In Software Tools Users Group [Sof83], pages 255–261. LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Torek:1984:MWS

[TW84] Chris Torek and Mark Weiser. The Maryland Window System. In Software Tools Users Group [Sof84], pages 166–172. LCCN QA76.8.U65 U83 1984.

Truscott:1986:SUD

[TWM86a] Tom Truscott, Bob Warren, and Kent Moat. A state-wide UNIX distributed computing system. In USENIX Association [USE86c], pages 499–513.

Truscott:1986:SWU

[TWM86b] Tom Truscott, Bob Warren, and Kent Moat. A state-wide UNIX distributed computing system. In USENIX [USE86a], pages 499–513.

Thomas:1982:CU

[TY82] Rebecca Thomas and Jean Yates. The commercialization of UNIX. In Usr Group [Usr82], pages 281–282. Abstract only.

Association:1988:CSJ

[UE88] Computing systems: the journal of the USENIX Association, page various, 1988. ISSN 0895-6340. University of California Press, Berkeley, CA, USA.

Uhler:1987:MWS

[Uhl87] Stephen A. Uhler. MGR — a window system for UNIX. In USENIX Association [USE87c], pages 106–?? ISBN ???? LCCN ???? Abstract only.

Uitti:1987:HAC

[Uit87a] Stephen Uitti. Hacct — A charge back system. In USE-NIX Association [USE87d], pages 34–?? ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

Uitti:1987:HCB

[Uit87b] Stephen Uitti. Hacct — A charge back system. In USE-NIX Association [USE87d], pages 34–?? ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

Upshaw:1983:NTV

[UJ83] Bob Upshaw and Van Jacobson. New tools for the virtual operating system. In Association [Ass83b], pages 9–13.

Upshaw:1982:PVO

[Ups82] Bob Upshaw. Portability in the Virtual Operating System. In Usr Group [Usr82], pages 303– 312.

Uttal:1985:TIU

[URK85] Judi Uttal, Jeff Rothschild, and Charles Kline. Transparent integration of UNIX and MS-DOS. In USENIX Association [USE85c], pages 104–116.

USENIX:1982:UCP

[USE82a] USENIX, editor. USENIX Conference Proceedings, Summer, 1982. Boston, MA. USE-NIX, Berkeley, CA, USA, Summer 1982.

USENIX:1982:UST

[USE82b]  USENIX, editor.  USE-

NIX/Software Tools Conference Proceedings, Summer, 1982. Boston, MA. Software Tools Users Group, Berkeley, CA, USA, Summer 1982.

USENIX:1983:UCPa

[USE83a] USENIX, editor. USENIX Conference Proceedings, Winter, 1983. San Diego, CA. USENIX, Berkeley, CA, USA, Winter 1983.

USENIX:1983:USTa

[USE83b]  USENIX, editor.  USE-

NIX/Software Tools Conference Proceedings, Summer, 1983. Toronto, Ontario. Software Tools Users Group, Berkeley, CA, USA, Summer 1983.

USENIX:1984:UCP

[USE84a] USENIX, editor. USENIX Conference Proceedings, Summer, 1984. Salt Lake City, UT. USENIX, Berkeley, CA, USA, Summer 1984.

USENIX:1984:UST

[USE84b]  USENIX, editor.  USE-

NIX/Software Tools Conference Proceedings, Summer, 1984. Salt Lake City, UT. Software Tools Users Group, Berkeley, CA, USA, Summer 1984.

USENIX:1984:UUC

[USE84c] USENIX Association, editor.

USENIX UniForum Conference Proceedings, January 17– 20, 1984. Washington DC, USA. USENIX, Berkeley, CA, USA, January 17–20, 1984. ISBN none. LCCN QA76.8.U65 U55 1984.

USENIX:1985:UCPa

[USE85a] USENIX, editor. USENIX Conference Proceedings, Summer, 1985. Portland, OR. USE-NIX, Berkeley, CA, USA, Summer 1985.

USENIX:1985:UCPb

[USE85b] USENIX, editor. USENIX Conference Proceedings, Winter, 1985. Dallas, TX. USE-NIX, Berkeley, CA, USA, Winter 1985.

USENIX:1985:PUA

[USE85c] USENIX Association, editor.

Proceedings: USENIX Association Winter Conference, January 23–25, 1985, Dallas, Texas, USA. USENIX, P.O. Box 7, El Cerrito 94530, CA, USA, 1985.

USENIX:1985:SCG

[USE85d] USENIX Association, editor.

Second Computer Graphics Workshop Proceedings, December 12–13, 1985. Monterey, CA. USENIX, Berkeley, CA, USA, December 12–13, 1985. ISBN ???? LCCN ????

USENIX:1985:SCP

[USE85e] USENIX Association, editor. Summer conference proceedings, Portland 1985: June 11– 14, 1985, Portland, Oregon USA. USENIX, P.O. Box 7, El Cerrito 94530, CA, USA, 1985. LCCN QA76.8.U65 U8 1985.

USENIX:1986:UCPa

[USE86a] USENIX, editor. USENIX Conference Proceedings, Summer, 1986. Atlanta, GA. USE-NIX, Berkeley, CA, USA, Summer 1986.

USENIX:1986:UCPb

[USE86b] USENIX, editor. USENIX Conference Proceedings, Winter, 1986. Denver, CO. USE-NIX, Berkeley, CA, USA, Winter 1986.

USENIX:1986:SCP

[USE86c] USENIX Association, editor. Summer conference proceedings, Atlanta 1986: June 9–13, 1986, Atlanta, Georgia, USA. USENIX, P.O. Box 7, El Cerrito 94530, CA, USA, 1986.

USENIX:1986:TCG

[USE86d] USENIX Association, editor. Third Computer Graphics Workshop Proceedings, November 20–21, 1986. Monterey, CA. USENIX, Berkeley, CA, USA, November 20–21, 1986. ISBN ???? LCCN ????

USENIX:1986:UAW

[USE86e] USENIX Association, editor.

USENIX Association Winter Conference proceedings: January 15–17, 1986, Denver, Colorado USA. USENIX, P.O. Box 7, El Cerrito 94530, CA, USA, 1986.

USENIX:1987:CWP

[USE87a] USENIX Association, editor.

C++ Workshop Proceedings, November 9–10, 1987. Santa Fe, NM. USENIX, Berkeley, CA, USA, November 9–10, 1987. ISBN ???? LCCN ????

USENIX:1987:ECP

[USE87b] USENIX Association, editor.

EUUG Conference Proceedings, Autumn, 1987. Dublin, Ireland. EUUG, Buntingford, Herts, UK, Autumn 1987. ISBN ???? LCCN ????

USENIX:1987:FCG

[USE87c] USENIX Association, editor.

Fourth Computer Graphics Workshop Proceedings, October 8–9, 1987. Cambridge, MA. USENIX, Berkeley, CA, USA, October 8–9, 1987. ISBN ???? LCCN ????

USENIX:1987:LIS

[USE87d] USENIX Association, editor.

Large Installation System Administrators Workshop Proceedings, April 9–10, 1987. Philadelphia, PA. USENIX, Berkeley, CA, USA, April 9–10, 1987. ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

USENIX:1987:P

[USE87e] USENIX Association, editor. Proceedings. USENIX, Berkeley, CA, USA, 1987. LCCN T 385 C67 1987. Cover title: USENIX.

USENIX:1987:PSU

[USE87f] USENIX Association, editor. Proceedings of the Summer 1987 USENIX Conference: June 8–12, 1987, Phoenix, Arizona, USA. USENIX, Berkeley, CA, USA, 1987.

USENIX:1987:UAW

[USE87g] USENIX Association, editor. USENIX Association Winter Conference proceedings, Washington, DC, 1987: January 21–23, 1987, Washington, District of Columbia, USA. USE-NIX, P.O. Box 7, El Cerrito 94530, CA, USA, 1987. LCCN QA 76.76 O63 U84 1987. Cover title: Conference proceedings. Spine title: Winter 1987 USENIX Association Conference proceedings, Washington, DC.

USENIX:1988:CCP

[USE88a] USENIX, editor. C++ Conference Proceedings, October 17– 21, 1988. Denver, CO. USE-NIX, Berkeley, CA, USA, October 17–21, 1988.

USENIX:1988:USWa

[USE88b] USENIX, editor. UNIX Security Workshop Proceedings, August 29–30, 1988. Portland, OR. USENIX, Berkeley, CA, USA, August 29–30, 1988.

USENIX:1988:AIU

[USE88c] USENIX Association, editor.

Abstracts of IEEE and USE-NIX of the Fifth Workshop on Real-Time Software and Operating Systems. USENIX, Berkeley, CA, USA, 1988.

USENIX:1988:APS

[USE88d] USENIX Association, editor.

The Andrew Project selected papers from the USENIX technical conference Winter 1988. Information Technology Center, Carnegie Mellon University, Pittsburgh, PA, USA, 1988.

USENIX:1988:ECP

[USE88e] USENIX Association, editor.

EUUG Conference Proceedings, Spring, 1988. London, England. EUUG, Buntingford, Herts, UK, Spring 1988. ISBN ???? LCCN ????

USENIX:1988:LIS

[USE88f] USENIX Association, editor.

Large Installation Systems Administration Workshop Proceedings, November 17–18, 1988. Monterey, CA. USENIX, Berkeley, CA, USA, November 17–18, 1988. ISBN ???? LCCN ????

USENIX:1988:PFU

[USE88g] USENIX Association, editor.

Proceedings of the (First) USE-NIX Security Workshop, August 29–30, 1988, Portland, OR, USA. USENIX, Berkeley, CA, USA, 1988. LCCN QA76.8.U65 U55 1988(1)- 1990(2)//.

USENIX:1988:PSU

[USE88h] USENIX Association, editor. Proceedings of the Summer 1988 USENIX Conference: June 20–24, 1988, San Francisco, California, USA. USE-NIX, Berkeley, CA, USA, 1988.

USENIX:1988:USWb

[USE88i] USENIX Association, editor. UNIX and Supercomputers Workshop Proceedings, September 26–27, 1988. Pittsburgh, PA. USENIX, Berkeley, CA, USA, September 26–27, 1988. ISBN ???? LCCN ????

USENIX:1988:UCPb

[USE88j] USENIX Association, editor. USENIX Conference Proceedings (Dallas, TX, USA). USE-NIX, Berkeley, CA, USA, Winter 1988. ISBN ???? LCCN ????

USENIX:1988:UPC

[USE88k] USENIX Association, editor. USENIX proceedings: C++ Conference, Denver, CO, October 17–21, 1988. USENIX, Berkeley, CA, USA, 1988.

USENIX:1989:SMW

[USE89a]  USENIX, editor.  Software

Management Workshop Proceedings, April 3–4, 1989. New Orleans, LA. USENIX, Berkeley, CA, USA, April 3–4, 1989.

USENIX:1989:UCPb

[USE89b] USENIX, editor. USENIX Conference Proceedings, Summer, 1989. Baltimore, MD. USENIX, Berkeley, CA, USA, Summer 1989.

USENIX:1989:UCPa

[USE89c] USENIX, editor. USENIX Conference Proceedings, Winter, 1989. San Diego, CA. USENIX, Berkeley, CA, USA, Winter 1989.

USENIX:1989:DMS

[USE89d] USENIX Association, editor.

Distributed and Multiprocessor Systems Workshop Proceedings, October 5–6, 1989. Fort Lauderdale, FL. USENIX, Berkeley, CA, USA, October 5– 6, 1989. ISBN ???? LCCN ????

USENIX:1989:LIS

[USE89e] USENIX Association, editor.

Large Installation Systems Administration III Workshop Proceedings, September 7–8, 1989. Austin, TX. USENIX, Berkeley, CA, USA, September 7–8, 1989. ISBN ???? LCCN ????

USENIX:1989:PSU

[USE89f] USENIX Association, editor.

Proceedings of the Summer 1989 USENIX Conference: June 12 — June 16, 1989, Baltimore, Maryland USA. USE-NIX, Berkeley, CA, USA, 1989. LCCN QA 76.76 O63 U83 1989.

USENIX:1989:PWU

[USE89g] USENIX Association, editor. Proceedings of the Winter 1989 USENIX Conference: January 30–February 3, 1989, San Diego, California, USA. USE-NIX, Berkeley, CA, USA, 1989.

USENIX:1989:UTP

[USE89h] USENIX Association, editor. UNIX Transaction Processing Workshop Proceedings, May 1– 2, 1989. Pittsburgh, PA. USE-NIX, Berkeley, CA, USA, May 1–2, 1989. ISBN ???? LCCN ????

USENIX:1989:UFC

[USE89i] USENIX Association, editor. USENIX Fifth Computer Graphics workshop, November 16–17, 1989. Monterey, CA. USENIX, Berkeley, CA, USA, November 16–17, 1989. ISBN ???? LCCN ????

USENIX:1989:UWP

[USE89j] USENIX Association, editor. USENIX workshop proceedings: software management, April 3–4, 1989, New Orleans, Louisiana. USENIX, Berkeley, CA, USA, 1989.

USENIX:1989:WP

[USE89k] USENIX Association, editor. Workshop on Experiences with Building Distributed and Multiprocessor Systems (WEB-DMS). USENIX, Berkeley, CA, USA, 1989.

Upshaw:1983:WCI

[USJ83] Bob Upshaw, Joe Sventek, and Van Jacobson. West Coast Implementors Group proposed standards. In Association [Ass83b], pages 15–16. Abstract only.

USENIX:1982:UAS

[Usr82] Usr Group, editor. /usr/ group, USENIX Association, Software Tools Users Group Joint Conference Proceedings: Boston, July 1982. USENIX, Berkeley, CA, USA, 1982.

Ullman:1984:BTB

[UTC84] Ellen Ullman, Page Thompson, and Jerry Carlin. Building tunnels and bridges: Constructing a commercial application under UNIX. In USENIX Association [USE84c], pages 259–284. ISBN none. LCCN QA76.8.U65 U55 1984.

Veer:1986:UEO

[van86] Gerrit C. van der Veer. UNIX and the electronic office — cognitive ergonomic reflections. EUUG Newsletter, 6(3):41– 49, Winter 1986. CODEN EONLE8. ISSN 1011-4211.

VanBaak:1987:VDN

[Van87] Thomas Van Baak. Virtual disks: A new approach to disk configuration. In USENIX Association [USE87g], pages 145– 146. LCCN QA 76.76 O63 U84 1987. Abstract only.

VanCleef:1988:SAM

[Van88] Robert E. Van Cleef. System administration and maintenance of fully configured workstations. In USENIX Association [USE88f], pages 79– 81. ISBN ???? LCCN ????

Vasilatos:1987:ADA

[Vas87a] Alix Vasilatos. Automated dumping at Project Athena. In USENIX Association [USE87d], pages 7–?? ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

Vasilatos:1987:ADP

[Vas87b] Alix Vasilatos. Automated dumping at project Athena. In USENIX Association [USE87d], pages 7–?? ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

Verbaeten:1983:PU

[VB83] P. Verbaeten and Y. Berbers. Porting UNIX. In Association [Ass83a], pages 285–289.

Huu:1987:ESD

[vH87] Le van Huu. An environment for SGML document preparation. In USENIX Association [USE87f], pages 43–52.

Vlissides:1988:AOD

[VL88a] John M. Vlissides and Mark A. Linton. Applying objectoriented design to structured graphics. In USENIX Association [USE88k], pages 81–94.

Vlissides:1988:AOO

[VL88b] John M. Vlissides and Mark A.

Linton. Applying objectoriented design to structured graphics. In USENIX [USE88a], pages 81–94.

Vaish:1984:TDX

[VM84] Paresh K. Vaish and Jean Marie McNamara. Techniques for debugging XENIX device drivers. In Software Tools Users Group [Sof84], pages 214–223. LCCN QA76.8.U65 U83 1984.

vandeGoor:1988:UMS

[vM88] A. J. van de Goor and A. Moolenaar. UNIX I/O in a multiprocessor system. In USENIX Association [USE88j], pages 251–258. ISBN ???? LCCN ????

vandeGoor:1988:MUS

[vMM88] A. J. van de Goor, A. Moolenaar, and J. M. Mulder. Multiprocessor UNIX: Separate processing of I/O. In USENIX Association [USE88e], pages 123– 134. ISBN ???? LCCN ????

Waidhofer:1982:TID

[Wai82] Gordon W. Waidhofer. Tabstar — information data base management. In Usr Group [Usr82], pages 280–?? Abstract only.

Walsh:1982:UUAa

[Wal82a] Daniel Walsh. UTS: UNIX on the Amdahl 470. In USENIX [USE82a], pages 210–?? Abstract only.

Walsh:1982:UUAb

[Wal82b] Daniel Walsh. UTS: UNIX on the Amdahl 470. In USENIX [USE82a], pages 247–?? Abstract only.

Walsh:1982:UUAc

[Wal82c] Daniel Walsh. UTS: UNIX on the Amdahl 470. In Usr Group [Usr82], pages 210–?? Abstract only.

Waldo:1986:MTA

[Wal86a] James Waldo. Modelling text as A hierarchical object. In USENIX Association [USE86c], pages 270–283.

Waldo:1986:MTH

[Wal86b] James Waldo. Modelling text as A hierarchical object. In USENIX [USE86a], pages 270– 283.

Waldo:1987:UCD

[Wal87] Jim Waldo. Using C++ to develop a WYSISYG hypertext toolkit. In USENIX Association [USE87a], pages 246–255. ISBN ???? LCCN ????

Wambecq:1983:NAD

[Wam83a]  A. Wambecq.  NETIX: A

distributed operating system based on UNIX software and local networking. In Software Tools Users Group [Sof83], pages 295–310. LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Wambecq:1983:NDO

[Wam83b]  A. Wambecq.  NETIX: A

distributed operating system based on UNIX software and local networking. In Association [Ass83a], pages 295–310.

Wareham:1982:SDL

[War82] Elwyn Wareham. Systems designers limited vendor presentation on Angus. In Usr Group [Usr82], pages 210–?? Abstract only.

Ward:1983:DID

[War83] J. Robert Ward. The design and implementation of the DB relational database management system. In Association [Ass83a], pages 211–228.

Warnock:1984:UMD

[War84] Robert P. Warnock, III. Usermode development of hardware and kernel software. In USE-NIX [USE84a], pages 224–226. Abstract only.

Watt:1983:LSU

[Wat83] Alan S. Watt. List of sources for UNIX device drivers. ;login: the USENIX Association newsletter, 8(3):14–39, June 1983. CODEN LOGNEM. ISSN 1044-6397.

Watson:1988:OAD

[Wat88a] Pat H. Watson. An overview of architectural directions for realtime distributed systems. In Association [Ass88e], pages 59– 65.

Watson:1988:CTH

[Wat88b] W. Bruce Watson. Capacity testing a HYPERchannel-Based local area network. In USENIX Association [USE88f], pages 83–85. ISBN ???? LCCN ????

Watson:1988:CAC

[Wat88c] W. Bruce Watson. Computer aided capacity planning of a very large information management system. In USENIX Association [USE88f], pages 63–65. ISBN ???? LCCN ????

Wood:1985:PLE

[WB85a] Jean Wood and Hans-Joachim Brede. Parlez-vous l’UNIX? The European perspective, past and future. In USENIX Association [USE85e], pages 17–23. LCCN QA76.8.U65 U8 1985.

Wood:1985:PVL

[WB85b] Jean Wood and Hans-Joachim Brede. Parlez-vous l’UNIX? The European perspective, past and future. In USENIX [USE85a], pages 17–23.

Waldo:1987:GFF

[WDL87] James Waldo, Marcia Delaney, and John Laporta. A generalized font file format. In USENIX Association [USE87c], pages 107–?? ISBN ???? LCCN ???? Abstract only.

Wehr:1983:UFS

[Weh83] Larry A. Wehr. UNIX file system evolution. In Association [Ass83a], pages 110–?? Abstract only.

Weinberger:1984:VNF

[Wei84a] P. J. Weinberger. The Version 8 Network File System. In Software Tools Users Group [Sof84], pages 86–?? LCCN QA76.8.U65 U83 1984. Abstract only.

Weinstein:1984:BNN

[Wei84b]  Lauren Weinstein.  Broad-

casting of Netnews and network mail via satellite. In Software Tools Users Group [Sof84], pages 18–22. LCCN QA76.8.U65 U83 1984.

Weinstein:1985:NSA

[Wei85a] Lauren Weinstein. Netnews via satellite: A progress report (12/84). In USENIX Association [USE85c], pages 1–3.

Weinstein:1985:NSP

[Wei85b] Lauren Weinstein. Netnews via satellite: A progress report (12/84). In USENIX [USE85b], pages 1–3.

Weinstein:1985:PS

[Wei85c] Lauren Weinstein. Project Stargate. In USENIX Association [USE85e], pages 79–80. LCCN QA76.8.U65 U8 1985.

Walsh:1984:CBT

[WG84] Robert Walsh and Robert Gurwitz. Converting the BBN TCP/IP to 4.2BSD. In Software Tools Users Group [Sof84], pages 52–61. LCCN QA76.8.U65 U83 1984.

Wolfe:1983:PAP

[WH83] Peter Wolfe and Allen Hustler. A powerful accounting package for UNIX-Based systems. In Software Tools Users Group [Sof83], pages 321–?? LCCN QA76.8.U65 U74 1983. Abstract only.

Weiser:1989:LVG

[WHM89] Mark Weiser, Barry Hayes, and Jock Mackinlay. Learning from visualized garbage collector. In USENIX Association [USE89i], pages 93–98. ISBN ???? LCCN ????

Williams:1982:BFM

[Wil82a] Gary Williams. A businessoriented file manager under UNIX, with contention control and ISAM. In Usr Group [Usr82], pages 269–279.

Williams:1982:BOF

[Wil82b] Gary Williams. A businessoriented file manager under UNIX, with contention control and ISAM. In USENIX [USE82a], pages 269–279.

Wilder:1983:GVC

[Wil83a] Henry Wilder. Getting venture capital. In Association [Ass83a], pages 308–?? Abstract only.

Wilens:1983:SAH

[Wil83b] Michael E. Wilens. SERIX — A high performance implementation of UNIX for the IBM Series/1. In Software Tools Users Group [Sof83], pages 89– 99. LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Wilens:1983:SHP

[Wil83c] Michael E. Wilens. SERIX — A high performance implementation of UNIX for the IBM Series/1. In Association [Ass83a], pages 89–99.

Williams:1983:E

[Wil83d]  Ellen Williams.  EUNICE.

In Association [Ass83a], pages 284–?? Abstract only.

Wilhelms:1987:DE

[Wil87a] Jane Wilhelms. Dynamics for everyone. In USENIX Association [USE87c], pages 49–72. ISBN ???? LCCN ????

Williams:1987:GMP

[Wil87b] Alain D. D. Williams. Grouse:

Messages and prompts in programs. EUUG Newsletter, 7(2):35–44, 1987. CODEN EONLE8. ISSN 1011-4211.

Wilhelms:1988:DCG

[Wil88] Jane Wilhelms. Dynamics for computer graphics: A tutorial. In Association [Ass88d], pages 63–93.

Williams:1989:SMS

[Wil89] Tim Williams. Session management in System V Release 4. In USENIX Association [USE89g], pages 365–375.

Winsor:1988:AOR

[Win88] Don Winsor. Access for operators that require root privileges (SUID & SGID). In USENIX Association [USE88g], pages 57–60. LCCN QA76.8.U65 U55 1988(1)-1990(2)//.

Weiner:1982:UPP

[WJ82] James L. Weiner and Brian L. Johnson. UNIX/Prime: Porting the UNIX operating system to Prime machines. In Usr Group [Usr82], pages 247–248. Abstract only.

Wasserman:1983:FUA

[WK83a] Anthony I. Wasserman and Martin Kersten. Focus/USE: A low keystroke database editor and browser. In Association [Ass83a], pages 241–244.

Wasserman:1983:FUL

[WK83b] Anthony I. Wasserman and Martin Kersten. Focus/USE: A low keystroke database editor and browser. In USENIX [USE83a], pages 241–244.

Walsh:1985:OSN

[WLS^{+}85] Dan Walsh, Bob Lyon, Gary Sager, J. M. Chang, D. Goldberg, S. Kleiman, T. Lyon, R. Sandberg, and P. Weiss. Overview of the Sun network file system. In USENIX Association [USE85c], pages 117–124.

Weisman:1982:CCD

[WM82] Robert Weisman and Mike Meissner. C compiler for Data General AOS/VS. In Usr Group [Usr82], pages 199–209.

Welch:1988:PDU

[WO88] Brent B. Welch and John K.

Ousterhout. Pseudo devices: User-level extensions to the Sprite File System. In Association [Ass88f], pages 37–49.

Wohlever:1988:USA

[Woh88] Kevin Wohlever. UNICOS system administration at the Ohio Supercomputer Center — tuning considerations. In USENIX Association [USE88i], pages 135–136. ISBN ???? LCCN ???? Abstract only.

Wood:1983:VCC

[Woo83a] Jean Wood. VMS C compiler.

In Association [Ass83a], pages 330–?? Abstract only.

Woods:1983:FFF

[Woo83b] James A. Woods. Finding files fast. ;login: the USENIX Association newsletter, 8(1):8– 10, February 1983. CODEN LOGNEM. ISSN 1044-6397.

Woznick:1982:MRU

[Woz82] Benjamin J. Woznick. Managing a roomful of UNIX systems. In Usr Group [Usr82], pages 179–185.

Wilkenloh:1989:CEB

[WRM^{+}89] C. J. Wilkenloh, U. Ramachandran, S. Menon, R. J. LeBlanc, M. Y. A. Khalidi, P. W. Hutto, P. Dasgupta, R. C. Chen, J. M. Bernabeu, W. F. Appelbe, and M. Ahamad. The Clouds experience: Building an object-based distributed operating system. In USENIX Association [USE89d], pages 333– 347. ISBN ???? LCCN ????

Wasserman:1983:RAT

[WS83a] Anthony I. Wasserman and David T. Shewmake. RAPID: A tool for building interactive information systems. In Association [Ass83a], pages 105–108.

Wasserman:1983:RTB

[WS83b] Anthony I. Wasserman and David T. Shewmake. RAPID: A tool for building interactive information systems. In USE-NIX [USE83a], pages 105–108.

Warnock:1984:SST

[WS84] Robert P. Warnock, III and Bakul Shah. A simple simulation toolkit in C. In USENIX [USE84a], pages 2–11. Abstract on p. 110, paper in Addendum pp. 2–11.

Yamasaki:1988:SPU

[Yam88] Michael J. Yamasaki. Special purpose user-space network protocols. In USENIX Association [USE88j], pages 63–69. ISBN ???? LCCN ????

Yao:1983:UA

[Yao83] Joseph Yao. UNIX APL. In Association [Ass83a], pages 330– ?? Abstract only.

Yashinovitz:1989:ODB

[YKK89] C. J. Yashinovitz, T. Kovacs, and John Kalucki. An optical disk backup/restore system. In USENIX Association [USE89e], pages 123–131. ISBN ???? LCCN ????

Yost:1985:CTM

[Yos85] David Yost. The cloned tree method of revision control or A rich person’s revision control system or how I adapted the UNIX file system and tools that manipulate it to perform project revision control. In USENIX Association [USE85e], pages 229–245. LCCN QA76.8.U65 U8 1985.

Yeong:1989:UIS

[YSF89] Wengyik Yeong, Martin Lee Schoffstall, and Mark S. Fedor. A UNIX implementation of the simple network management protocol. In USENIX Association [USE89g], pages 209– 217.

Yates:1983:WUD

[YT83] Jean Yates and Rebecca Thomas. Writing user documentation for UNIX systems. In Association [Ass83a], pages 117–?? Abstract only.

Yee:1988:SSS

[YTS88] B. S. Yee, J. D. Tygar, and A. Z. Spector. StrongBox: Support for self-securing programs. In USENIX Association [USE88g], pages 50– ?? LCCN QA76.8.U65 U55 1988(1)-1990(2)//. Abstract only.

Zadrozny:1989:ARS

[Zad89] Peter Zadrozny. Administering remote sites. In USENIX Association [USE89e], pages 45–47. ISBN ???? LCCN ????

Zhou:1985:FST

[ZDS85] Songnian Zhou, Herve Dacosta, and Alan Jay Smith. A file system tracing package for Berkeley UNIX. In USENIX Association [USE85e], pages 407–419. LCCN QA76.8.U65 U8 1985.

Zemon:1983:FTP

[Zem83] Arthur Zemon. A friendly text processing environment. In Association [Ass83a], pages 116– ?? Abstract only.

Zorn:1988:MAP

[ZH88] Benjamin Zorn and Paul Hilfinger. A memory allocation profiler for C and Lisp programs. In Association [Ass88f], pages 223–237.

Zhou:1987:EAR

[Zho87] Songnian Zhou. An experimental assessment of resource queue lengths as load indices. In USENIX Association [USE87g], pages 73–82. LCCN QA 76.76 O63 U84 1987. Cover title: Conference proceedings. Spine title: Winter 1987 USE-NIX Association Conference proceedings, Washington, DC.

Zimmerman:1985:DUK

[Zim85] Steven A. Zimmerman. A debugger for the UNIX kernel. In USENIX Association [USE85e], pages 151–153. LCCN QA76.8.U65 U8 1985.

Zwicky:1989:MLP

[ZP89] Elizabeth D. Zwicky and Paul W. Placeway. Modifying the line printer system for a large networked environment. In USENIX Association [USE89e], pages 53–57. ISBN ???? LCCN ????

Zelitzky:1983:CN

[ZS83] Jay Zelitzky and Sunil Srivastava. Compilers on the NS16000. In Software Tools Users Group [Sof83], pages 237–245. LCCN QA76.8.U65 U74 1983. Sponsored by USE-NIX Association in cooperation with Software Tools Users Group.

Zucker:1983:CLM

[Zuc83a] Steve Zucker. Contiguous load modules for UNIX. In Association [Ass83a], pages 39–?? Abstract only.

Zucker:1983:CEU

[Zuc83b] Steven Zucker. IS/3: A compatible extension of UNIX system III. In USENIX [USE83a], pages 325–329.

Zucker:1983:IAC

[Zuc83c]  Steven Zucker.  IS/3: A

compatible extension of UNIX system III. In Association [Ass83a], pages 325–329.

Zwicky:1988:BAO

[Zwi88a] Elizabeth Zwicky. Backup at Ohio state. In USENIX Association [USE88f], pages 43–44. ISBN ???? LCCN ????

Zwicky:1988:BOS

[Zwi88b] Elizabeth Zwicky. Backup at Ohio state. In USENIX Association [USE88f], pages 43–44. ISBN ???? LCCN ????

Zwicky:1989:DSM

[Zwi89] Elizabeth D. Zwicky. Disk space management without quotas. In USENIX Association [USE89e], pages 41–43. ISBN ???? LCCN ????